A Complete Bibliography of the Journal of Number Theory (2000–2009)

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

16 July 2020 Version 1.00

Title word cross-reference

(0, 1) [622]. (1^{2} + 1) ⋯ (n^{2} + 1) [1431]. (3)^{2} [75]. (α, α^{2}) [490]. (U(1), U(2)) [805, 806]. ({nα + γ} − 12) [39]. (O_{K}/I)^{×} [163]. −2^{2s}qr [140]. 0 [587]. 1 [220, 421, 574, 676, 932, 1094, 1310, 1538, 1688]. 1/2 [1477]. 11 [592]. 12 [355]. 13 [592]. 2 [69, 88, 123, 146, 181, 203, 218, 276, 295, 365, 452, 522, 526, 645, 703, 708, 718, 729, 732, 734, 860, 980, 1111, 1150, 1203, 1220, 1296, 1321, 1326, 1361, 1469, 1494, 1517, 1540, 1581]. 24m [271]. 2^{t} [297]. 2k [108]. 2n [1568]. 2s [1268]. 3 [75, 123, 692, 755, 1157, 1321, 1353]. 3x + 1 [948, 1111]. 4 [123, 208, 466, 536, 1369, 1666]. 40 [1528]. 4ℓ [24]. 5 [56, 323, 1525]. 7 [1580]. 8 [585, 812, 1525]. 9 [1531]. ?(x) [694]. A [1115]. a(modp) [636, 637]. A(q) [179]. a^{3} + b^{3} + c^{3} + d^{3} = 0 [339]. A_{4} [1056]. A_{6} [891]. A_{n} [1543]. abc [50]. α [425]. α^{3} + kα − 1 = 0 [490]. ax(x − 1)/2 + by(y − 1)/2 [1552]. axy + bx + cy + d [512]. B [115, 844]. B_{2}[g] [1089]. B_{h}[g] [406]. F̄_{p}(x) [1149]. β [1259]. mod [758]. modℓ [213, 779, 1629]. modm [1441, 1646]. modp [23, 104, 114, 145, 548, 652, 684, 891, 1210]. C^{n} [735]. C_{2} × C_{2} × C_{2} [1494]. C_{p} [193]. d [786, 1445]. D(4) [1368]. ℓ [26, 168]. ℓ^{j} [1663]. f(x, y) = f(u, v) [206]. F_{2}(X) [15]. F_{l}[[T]] [390]. F_{p} [395, 409]. F_{q} [81]. F_{q}[T] [112, 562]. F_{r}(T) [735]. F_{r}[T] [735]. \frac{x^{3}−1}{x−1} = \frac{y^{n}−1}{y−1} [813]. G [283, 926]. g(modp) [870, 962, 1031]. g − 1 [926]. G_{2} [70]. G_{4} [241]. G_{6} [241]. Γ [318, 744]. Γ_{0}(13) [1098]. Γ_{0}(4) [817]. Γ_{0}(N) [824, 991]. Γ_{0}(N) \ H [399]. Γ_{1}(m) [147]. Γ_{1}(N) [824]. h [443]. H^{1} [534]. Â_{5} [236]. ∞ [855]. ∫_{−1}^{1} ∫_{−1}^{1} ∫_{−1}^{1} Φ((ax + by + cz)R)Ψ(x, y, z) dx dy dz [1435]. J [132, 458, 1149, 1292, 1704]. J_{0}(431) [783, 953]. K [69, 392, 729, 998, 1024, 1025, 1044, 1055, 1240, 1595]. {k − 1, k + 1} [1290]. k − 2^{n} [223]. k2^{n} + 1 [223]. k2^{n} ± 1 [1189]. K3 [98, 756, 1118, 1466]. k = 4 [242]. K[ϵ] [1390]. k^{r} − 2^{n} [461]. k^{r}2^{n} + 1 [461]. K_{2} [530]. K_{2}(Q) [842, 1300]. K_{2}(O) [208]. k ± 2^{n} [1189]. L [23, 26, 36, 48, 70, 72, 100, 103, 108, 129, 137, 152, 169, 171, 184, 197, 307, 430, 446, 452, 459, 510, 608, 621, 650, 677, 722, 769, 780, 791, 848, 888, 924, 1005, 1032, 1066, 1081, 1101, 1130, 1132, 1148, 1167, 1176, 1177, 1196, 1198, 1208, 1291, 1319, 1336, 1362, 1409, 1422, 1429, 1463, 1576, 1579, 1592, 1654, 1665, 1677, 1691, 1698]. L^{∞} [1171]. Λ [500, 990, 1702]. ⌊αp + β⌋ [1651]. log η_{g,h}(z) [495]. m [185, 224, 1038]. F_{1} [1593]. F_{13} [1343]. F_{p} [821]. F_{q} [926]. F_{q}[t] [1445]. K_{p}((X^{−1})) [916]. P^{1} [1600]. P^{n} [1277]. Q [816, 860, 891, 1292, 1365, 1521, 1558]. Q(\sqrt{2}) [749]. Q(ζ_{p}) [1090]. R^{n} [1661]. Z^{2} [1423, 1662]. Z^{3} [1559]. Z^{k} [941]. Z_{p} [990]. Z_{p}^{2} [1625]. Z_{p}^{d} [1083, 1454]. C_{p} [1574]. F_{q}[t] [1657]. Q_{p} [1574]. Z_{m} [1440]. Z_{p} [836]. O_{∞} [1705]. CH^{3}(F, 5) [1146]. GL(n) [1036]. GL_{2} [1130]. GL_{2n} [1213]. GL_{m} [1081]. SL(4, Z) [1418]. SL_{2}(F) [1641]. Cl_{2}(k) ≃ (2, 2, 2) [583]. GL(2) [1374]. GL(m, D) [814]. GL(n, F) [814]. GL(p) [476]. GL_{2} [71, 816]. GL_{2}(F_{q^{2}}) [176]. GL_{2}(F_{ℓ}) [1558]. GL_{3}(Z) [891]. GL_{m}(A_{K}) [1344]. GL_{n} [293, 572, 1457]. GSp(4) [1422]. GSp_{4} × GL_{2} [1576]. M_{m}(D) [1224]. PSL(2, Z) [160]. PSL(2, Z) \ H [605]. PSL_{2}(Z[1/N]) [491]. SL(2) [119]. SL(2, Z)_{+} [371]. SL(n, Z)_{+} [1661]. SL_{2}(Z) [1399]. SL_{2}(Z_{3}) [1157]. SL_{4}(Z) [336]. SL_{n} [64]. Sp(2n) [1522]. Sp(4) [1521]. Sp_{2} [822]. Sp_{2} × (Sp_{1})^{2} [822]. μ [282, 314, 568]. n! [54, 523, 573, 1096, 1159, 1615]. N [168, 679, 730, 786, 804, 915, 1190, 1506, 1552, 1636]. n = 3 [1680]. n = 7 [1555]. n^{2} [804]. n_{1}n_{2} = n_{3}n_{4} [1289]. Nℓ^{m} [168]. n ≤ 6 [1554]. n × (n − 2) [1036]. P [2, 25, 58, 72, 80, 100, 119, 180, 187, 198, 235, 255, 353, 446, 474, 476, 532, 604, 607, 618, 625, 654, 688, 697, 724, 792, 807, 850, 855, 934, 956, 968, 1012, 1017, 1036, 1090, 1135, 1325, 1333, 1353, 1390, 1403, 1404, 1411, 1422, 1543, 1625, 1640, 1656, 1665, 1683, 1686, 1701, 1704]. P (f) = Q(g) [654]. p = k^{2} + l^{2} + 1 [1349]. p > 0 [109]. p > 2 [863]. P^{1} [540]. P^{2} [540, 548, 1176]. p^{5} [1112]. p^{a} [344]. p^{j} [344]. p^{m} [262]. p^{q} [515]. p^{x} − q^{y} = c [646]. ϕ [1260]. ϕ(n) + σ(n) [1636]. π [96, 1707]. ±1 [107, 451]. ±ax ± by = c [992]. pq [463, 1573]. px^{2} + q^{2m} = y^{p} [1378, 1713]. Q [71, 141, 228, 236, 273, 363, 505, 656, 713, 780, 1067, 1172, 1543, 1557, 1610]. Q(−2) [703]. Q(ζ_{p^{r}}) [393]. Q + Qα [55]. q^{k} − 1 [562]. R [46, 283, 284, 498]. r_{s}(n) [1268]. S [121, 153, 608, 697, 1000, 1160, 1611, 1686]. s = 0 [1278, 1382, 1579]. s = 1 [137, 1278]. s_{0} [1463]. S_{3}(Γ(4, 8)) [1196]. S_{n} [1161]. S_{n}(n ≤ 5) [571]. ∑ +l + 6m = nσ(l)σ(m) [1182]. ∑_{2l+3m=n} σ(l)σ(m) [1182].

∑_{i=1}^{9} \frac{1}{x_{i}} = 1 [1255]. ∑_{l=0}^{∞} d^{l}/ ∏_{j=1}^{l}(1 + d^{j}r + d^{2j}s) [1272]. T [963, 1110, 1217, 1442, 1489, 1611, 1686]. τ [434, 524, 574, 769]. Θ [805, 806]. Ã_{5} [236]. U(2, 1) [874, 1062]. φ(n)/n [1022]. w [1003, 1405]. x^{2} + D^{m} = p^{n} [796]. x^{2} + q^{2k+1} = y^{n} [356]. x^{2} + y^{2} + z^{2} + 7s^{2} + 7t^{2} + 7u^{2} [1580]. x^{2} − Dy^{2} = −1, ±2 [856]. x^{3} + y^{9} = z^{2} [798]. x^{4} + ax^{2} + bx [1012]. x^{n} − 1 [1671]. X^{p} + aX^{p−1} + a [1541]. X_{0}(N) [1635]. X_{0}(p)/w_{p} [843]. X_{0}(p^{n}) [764]. X_{0}^{∗}(p) [506]. x_{0}y_{0} + ⋯ + x_{s}y_{s} = 0 [1591]. X_{1}(n) [569]. x_{1}x_{2} = x_{3}x_{4} + λ [1436]. Ξ [1030]. y^{2} = x(x − 2^{m})(x + q − 2^{m}) [1174]. y^{2} = x(x^{2} − n^{2}) [17, 1487, 1533]. y^{2} = x^{3} − n^{2}x [739]. y^{2} = x^{ℓ} + A [322]. Z [78]. Z_{2} [645]. Z_{ℓ} [311, 453]. Z_{p} [153, 382, 500]. ζ(2k + 1) [989]. ζ(s) [946]. ζ_{q}(1) [1230]. ζ_{q}(2) [1230].

-additive [115]. -Adic [2, 25, 58, 69, 80, 96, 100, 119, 168, 180, 187, 235, 255, 353, 446, 452, 476, 510, 604, 607, 625, 654, 688, 697, 724, 734, 792, 807, 850, 968, 1017, 1036, 1090, 1135, 1361, 1403, 1411, 1422, 1469, 1640, 1656, 1702, 1704]. -adics [708]. -adiques [850]. -Analogue [780]. -analytic [390]. -arithmetic [1000]. -ary [185]. -binomial [1067]. -body [26]. -Catalan [515]. -class [522, 980, 1203, 1494, 1666, 1686]. -coatings [283]. -continued [425, 1557]. -core [297, 1442, 1543, 1611]. -cores [1580]. -correspondences [805, 806]. -curves [1292]. -cycles [1111]. -cyclotomic [1325]. -derivations [96]. -dimensional [323, 679, 786]. -equivalence [283, 284, 498]. -espaces [409]. -expansions [1259]. -extension [645]. -extensions [26, 153, 382, 500, 645, 836, 860, 891, 990, 1056, 1083, 1150, 1157, 1296, 1454, 1568, 1625]. -forms [458]. -free [844, 1055]. -function [1030, 1167, 1422]. -Functions [23, 36, 70, 72, 100, 103, 108, 137, 152, 169, 171, 184, 197, 307, 430, 446, 452, 621, 722, 735, 791, 848, 888, 924, 1005, 1032, 1066, 1081, 1101, 1132, 1148, 1177, 1196, 1198, 1208, 1291, 1319, 1336, 1362, 1409, 1429, 1463, 1576, 1579, 1592, 1665, 1677, 1691, 1698]. -group [121, 729]. -Groupe [121, 729]. -groups [505, 1404]. -Hypermonogenic [1240]. -ideals [153]. -ièmes [1353]. -invariant [1149]. -invariants [132, 314, 568, 990, 1292, 1704]. -lattices [78, 1657]. -linéaires [1574]. -linear [81, 1574]. -Mahler [1172]. -modules [574, 1115]. -monomials [318]. -motifs [421, 1489]. -motives [963, 1365]. -multinomial [228]. -orbits [1661]. -orderings [474]. -power [203, 526]. -powers [1506]. -primary [276]. -ramified [1625, 1686]. -rank [208, 311, 453, 466, 1321, 1701]. -ranks [75]. -rational [26, 395, 730, 1517]. -rationnels [26]. -reducing [198]. -réduisants [198]. -regular [344]. -revêtements [283]. -Selmer [1581]. -sequences [622]. -series [48, 129, 459, 608, 677, 713, 769, 780, 1176]. -sextuple [1368]. -sheaves [434, 524, 769]. -space [822]. -split [1686]. -submodules [500]. -sums [1024]. -Sylow [181]. -th [108, 392, 934, 1044, 1445, 1595]. -theorems [50]. -théorie [729]. -theory [69, 729, 998]. -torsion [56, 755, 1190, 1390]. -tower [1683]. -transform [744]. -tuples [224, 1025, 1038]. -twisted [780]. -type [71, 816]. -unit [1160]. -unités [121]. -universality [1110, 1217]. -values [650, 1130, 1654]. -variable [915]. -vector [409]. -zeros [1704]. -zeta [1610].

1 [920]. 102 [589, 629]. 104 [1017]. 107 [1712, 1715]. 109 [126, 1714]. 10th [533]. 110 [953]. 114 [1062]. 120 [263]. 124 [1296]. 125 [1709]. 127 [1391]. 128 [1688, 1713]. 129 [1533, 1710, 1711, 1716]. 12th [701]. 144 [25].

216 [84].

450 [83].

73 [83].

89 [263].

92 [453]. 95 [760]. 99 [628]. 9th [701].

=1 [608, 697].

A. [886, 1133]. ABC [31, 59, 342, 369, 696, 1168, 1266, 1389]. abc-conjecture [342]. ABC-hits [1389]. Abelian [1, 21, 47, 71, 83, 125, 148, 184, 201, 207, 298, 301, 306, 311, 329, 334, 345, 394, 416, 453, 500, 509, 549, 552, 553, 611, 612, 626, 647, 662, 697, 774, 815, 816, 828, 860, 890, 931, 998, 1024, 1041, 1042, 1121, 1163, 1190, 1218, 1238, 1267, 1465, 1514, 1579, 1594, 1620, 1660, 1679, 1716]. abélienne [125]. abéliennes [207]. absolute [75, 224, 1227, 1305]. absolutely [298]. Abu [1713]. abundancy [1369]. action [316]. Actions [1346, 1371, 1414]. Adams [671]. Addendum [453, 485, 1391, 1709]. addition [122, 404, 940, 1646]. additional [1023]. Additive [37, 115, 188, 331, 424, 534, 624, 665, 1058, 1123, 1415, 1515, 1597, 1703]. additives [1123]. Adelic [184, 1545]. Adic [2, 25, 58, 69, 80, 96, 100, 119, 168, 180, 187, 235, 255, 353, 446, 452, 476, 510, 604, 607, 625, 654, 688, 697, 724, 734, 792, 807, 850, 968, 1017, 1036, 1090, 1135, 1361, 1403, 1411, 1422, 1469, 1640, 1656, 1702, 1704]. adics [708]. adiques [850]. adjoint [505, 1422]. admissible [1082]. affine [1346, 1627, 1699]. AGM [1407]. Aiba [886]. algebra [1705]. Algebraic [34, 43, 68, 95, 136, 196, 207, 235, 250, 254, 318, 385, 388, 497, 557, 571, 617, 634, 684, 690, 773, 792, 926, 988, 1070, 1091, 1131, 1195, 1220, 1310, 1411, 1414, 1421, 1568, 1688, 1706, 1712]. algebraicity [652]. algebraischen [68]. algebras [61, 157, 272, 330, 414, 638, 759, 1227, 1460, 1524, 1698]. algébrique [207]. algébriques [43, 95, 385, 1195]. algorithm [1526, 1637]. algorithms [784, 1647]. almost [1025, 1094]. Alon [1682]. along [821]. alternate [619]. alternating [246]. ambiges [516]. ambiguous [140, 516]. among [1537, 1546]. amplitudes [1509]. analogous [217, 1381]. Analogue [30, 33, 100, 126, 482, 533, 535, 780, 1082, 1338, 1440, 1606]. analogues [1009, 1281]. analysis [432, 1446]. analytic [390, 524, 544, 576, 934, 1344, 1419, 1619, 1650]. Anderson [1497, 1532]. angled [1088]. Ankeny [248]. anneau [1370]. Answers [1703]. Antoniadis [86].

Any [261, 662, 1603]. Apollonian [504, 1010]. Apostol [944]. Appendix [1039, 1522]. Application [114, 122, 551, 653, 1510, 1590, 1698]. Applications [149, 159, 379, 769, 827, 856, 1106, 1110, 1281, 1318, 1436, 1560, 1574, 1623, 1642, 1708]. approach [23, 75, 1027, 1669]. Approche [1027, 1669]. approximable [584, 1638, 1699]. Approximating [1348, 1500]. Approximation [19, 34, 55, 95, 107, 130, 334, 385, 441, 490, 499, 778, 792, 974, 1006, 1131, 1180, 1252, 1346, 1406, 1627]. Approximations [43, 408, 557, 1435, 1557]. April [40, 177, 319, 486, 643, 800, 954, 1128, 1314, 1534]. Arabia [1713]. Arakelov [807, 1701]. arbitrarily [480]. arbitrary [298, 690, 1712]. Archimedean [283, 348, 482, 533, 769, 1374, 1419, 1604]. archimédiens [283]. arcs [279]. arctangent [1387]. area [1273]. arguments [681, 1151]. arising [161, 256, 595, 1275, 1387, 1400]. Arithmetic [17, 111, 114, 293, 296, 299, 309, 322, 354, 381, 399, 748, 781, 793, 839, 844, 862, 904, 1000, 1015, 1059, 1090, 1099, 1154, 1219, 1248, 1250, 1293, 1316, 1360, 1379, 1380, 1515, 1516, 1547, 1596, 1674]. Arithmetical [9, 122, 883, 1033, 1085, 1387]. arithmétique [1379]. arithmétiques [748]. Arnold [763]. arrangements [1601]. arrays [1294]. article [83, 84, 284]. Artin [33, 101, 126, 139, 200, 248, 459, 563, 1148, 1463, 1489]. ary [185]. associated [458, 505, 522, 534, 547, 608, 634, 802, 928– 930, 963, 1014, 1069, 1100, 1106, 1118, 1247, 1396, 1401, 1494, 1524, 1544, 1560]. Asymptotic [20, 115, 266, 321, 544, 745, 1442]. Asymptotically [82, 389]. Asymptotics [1161]. Atkin [157, 837, 1426, 1635]. Atkinson [1687]. attached [161, 414, 639, 966, 968, 1090]. attains [669]. August [85, 227, 358, 541, 698, 840, 1007, 1206, 1412, 1612]. Author [27, 52, 73, 92, 117, 143, 166, 190, 215, 237, 264, 289, 312, 326, 349, 374, 402, 435, 469, 501, 527, 554, 577, 600, 630, 657, 685, 714, 736, 761, 788, 809, 829, 851, 877, 907, 936, 970, 994, 1018, 1072]. Automata [5]. Automatic [51]. automorphes [120]. Automorphic [90, 120, 157, 565, 1081, 1105, 1130, 1240, 1291, 1319, 1482, 1586]. Automorphisms [489, 497, 506]. auxiliary [341]. Average [101, 132, 755, 861, 881, 1336, 1409, 1516]. averages [845, 1039]. Ax [1084]. axis [171].

B [1647]. back [852, 1324]. bad [708]. Badly [584, 1638, 1699]. Baladi [1647]. balanced [1126]. balls [786]. Banach [267]. Barban [1336, 1676]. barcode [852]. Barker [1697]. Barnes’ [291, 1633]. Barsotti [467, 1233]. base [46, 119, 476, 496, 730, 874, 1028, 1062, 1312, 1363, 1493]. base- [46]. based [1590]. bases [29, 111, 133, 150, 203, 420, 429, 443, 536, 646, 665, 1046, 1057, 1123, 1536, 1597, 1608, 1703]. basis [188, 242, 382, 1415, 1465]. Bateman [175, 401, 949]. be [480]. Beatty [343, 1138]. behaved [825]. behavior [20, 446]. behaviour [544]. being [1499]. Belyi [377]. Bernoulli [102, 247, 291, 309, 328, 412, 671, 768, 808, 834, 964, 1112, 1153, 1288, 1318, 1469, 1616, 1637, 1673]. Bernstein [814, 1457]. Bessel [176]. Beta [1450]. Beta-expansions [1450]. between [12, 75, 495, 534, 617, 753, 890, 896, 1032, 1052, 1228, 1287, 1303, 1351, 1459, 1594, 1606, 1621]. Beurling [79, 803, 825]. Bézout [114]. bi [180]. bi-URS [180]. Bianchi [1629]. bias [171]. bicyclic [651, 1483]. Binary [99, 161, 221, 304, 320, 371, 384, 664, 804, 833, 941, 1114, 1152, 1158, 1349, 1634]. Binomial [305, 325, 515, 898, 969, 1045, 1067, 1155, 1242, 1274, 1455, 1670]. biquadrates [782]. biquadratic [383, 1052]. Birch [353, 1581]. Birkhoff [1421]. Bloch [58, 1146]. blocks [1120, 1543]. Blumenthal [306]. Board [419, 439, 440, 455, 471, 487, 503, 514, 529, 542, 556, 567, 579, 591, 602, 615, 632, 644, 659, 674, 687, 699, 716, 728, 738, 751, 762, 777, 790, 801, 811, 819, 831, 841, 854, 867, 880, 893, 910, 923, 939, 955, 973, 984, 997, 1008, 1021, 1035, 1050, 1064, 1075, 1093, 1109, 1129, 1145, 1166, 1186, 1207, 1223, 1232, 1245, 1257, 1271, 1284, 1298, 1315, 1341, 1359, 1393, 1413, 1434, 1452, 1462, 1471, 1480, 1496, 1519, 1535, 1551, 1570, 1599, 1613, 1631, 1644, 1668, 1690]. bodies [1197]. body [26, 125]. Bombieri [1642]. borné [385]. Borwein [61]. Bost [1501]. both [1611]. bound [21, 59, 224, 433, 626, 669, 1089, 1100, 1116, 1188, 1453, 1658, 1701]. Bounded [34, 131, 385, 1000, 1131, 1252]. Boundedness [39, 520]. bounding [965]. Bounds [137, 179, 302, 406, 608, 679, 732, 797, 913, 1066, 1070, 1175, 1215, 1226, 1309, 1321, 1389, 1415, 1671]. branched [273]. Brauer [882, 1190, 1476]. breaks [1486]. Breuil [1468]. Brownawell [1532]. Brun [49, 239, 1526]. bundles [849].

Calabi [48]. Calculating [1660]. can [480]. candies [31]. canonical [281, 626, 1287, 1382]. capitulation [125, 153, 1083, 1308]. capturing [1279]. caractéristique [1665]. cardinality [378]. Carlitz [111, 507]. Carmichael [588, 1330]. carrés [1572]. Cartan [1202]. cas [346, 1094, 1262]. case [60, 80, 105, 262, 285, 346, 478, 805, 806, 1094, 1259, 1262, 1356, 1627]. Cassels [448]. Catalan [488, 515, 950, 982]. causality [184]. cells [1601]. Central [36, 129, 272, 725, 848, 1101, 1176, 1227]. Certain [37, 46, 48, 77, 78, 104, 152, 221, 276, 335, 381, 394, 415, 465, 468, 479, 484, 517, 520, 544, 557, 592, 598, 634, 639, 645, 647, 660, 695, 704, 722, 839, 848, 935, 981, 1013, 1082, 1114, 1141, 1199, 1306, 1307, 1326, 1377, 1401, 1602, 1696, 1704, 1706]. certains [46]. Châtelet [785]. chains [795]. champion [1379]. champions [1379]. change [119, 476, 874, 1062, 1363]. changing [218]. Character [35, 274, 300, 864, 1173, 1289, 1354, 1429, 1548, 1634]. Characteristic [5, 20, 72, 109, 170, 202, 772, 834, 863, 928– 930, 933, 956, 961, 1180, 1409, 1545, 1606, 1665, 1694, 1701]. characteristics [548]. characterization [902, 1559]. characterizing [499, 1071, 1258]. Characters [62, 97, 386, 526, 608, 713, 1105, 1134, 1148, 1224, 1342]. Chebotarev [288]. Chebyshev [171]. Chen [1202, 1355]. Chevalley [1084]. chiffres [46, 861]. choosing [1191]. Chow [1146]. Chowla [248, 663]. Circle [1, 83, 269, 279, 504, 988, 1010, 1254, 1391]. circulant [1275]. Circular [214, 651, 903, 1339, 1483]. Class [18, 38, 75, 105, 137, 145, 179, 205, 222, 263, 276, 352, 361, 397, 407, 420, 445, 468, 483, 488, 517, 522, 536, 537, 539, 550, 662, 666, 705, 752, 846, 912, 927, 959, 980, 982, 1015, 1045, 1053, 1124, 1192, 1193, 1203, 1263, 1317, 1321, 1326, 1366, 1388, 1408, 1410, 1428, 1439, 1494, 1525, 1531, 1537, 1566, 1571, 1578, 1601, 1660, 1664, 1666, 1686, 1695]. class-number [1053]. classe [1192, 1709]. classes [24, 121, 140, 306, 462, 604, 709, 729, 812, 870, 962, 967, 1017, 1031, 1132, 1187, 1334, 1362, 1662]. classical [168, 538, 1400]. classification [895, 1636]. classify [835]. classifying [97]. Clifford [576]. Clifford-analytic [576]. closed [832]. closure [863, 1645]. closures [235]. CM [363, 395, 593, 733, 1091, 1101, 1193, 1306, 1328, 1692]. CM-fields [1091, 1193]. coatings [283]. Cochrane [947, 1104, 1188]. cocompact [1398]. Coefficients [107, 228, 305, 325, 357, 465, 481, 508, 515, 562, 628, 660, 661, 690, 839, 887, 898, 918, 969, 975, 1067, 1179, 1209, 1242, 1252, 1274, 1329, 1455, 1516, 1641, 1681, 1693, 1712]. Cohen [292, 675, 1404, 1458]. Cohn [35, 97, 300, 619]. Cohomological [633]. Cohomology [69, 336, 475, 491, 692, 797, 891, 967, 1418, 1490, 1653]. Coleman [1538]. Colmez [1427]. colored [138]. combinatorial [1239, 1404]. Combinatorics [1331, 1430]. Comments [1065]. common [832, 1215]. commutative [351, 642, 1502]. compact [1071]. compactifications [131]. comparison [111]. complement [284, 1277]. complements [1419]. Complete [683, 782, 1347, 1493, 1525]. Completing [1373]. completion [112]. completions [735]. Complex [197, 365, 509, 743, 802, 858, 1444, 1457, 1542, 1578]. components [764]. comporte [46]. composite [101]. compositum [1339]. Compter [189]. Computation [383, 522, 534, 647, 1220, 1494, 1641]. computations [1352]. Computing [133, 380, 393, 890]. concerning [162, 485, 787, 1703]. condition [794]. conditional [1127]. Conditions [240, 656, 719, 1023, 1175, 1706]. conductor [332, 620, 978]. conductors [156]. cones [505]. configurations [1511]. conformal [209, 559, 778]. Congruence [134, 241, 336, 352, 452, 837, 864, 1052, 1066, 1067, 1097, 1099, 1418, 1426]. Congruences [12, 102, 138, 185, 291, 297, 328, 412, 465, 560, 562, 629, 758, 771, 887, 950, 1086, 1235, 1242, 1268, 1288, 1302, 1478, 1592, 1614, 1621]. congruent [75, 1581]. Conjecture [30, 33, 59, 86, 101, 126, 139, 175, 183, 200, 225, 248, 258, 331, 332, 342, 353, 363, 364, 369, 401, 457, 463, 478, 479, 482, 488, 520, 575, 597, 665, 697, 720, 794, 826, 846, 855, 921, 931, 963, 1016, 1056, 1083, 1088, 1117, 1168, 1194, 1202, 1210, 1265, 1367, 1381, 1396, 1402, 1440, 1442, 1457, 1498, 1542, 1554, 1565, 1581, 1587, 1591, 1618, 1682]. conjectures [329, 390, 593, 1555]. conjugate [1543]. Connes [1501]. consecutive [597, 872, 1151, 1375]. consequences [781, 1587]. constancy [1482]. constant [680, 726, 783, 882, 953, 1149, 1179, 1700]. constants [847]. constraints [974]. Constructing [1136]. Construction [29, 664, 802, 990, 1059, 1306, 1608]. constructions [1626]. constructive [75]. contain [46]. containing [280, 442, 1506]. contains [1603]. contenant [280]. Contents [436, 437, 852, 878, 908, 937, 971, 995, 1019, 1048, 1073, 1107, 1143, 1184, 1221, 1243, 1269]. continuation [544]. continue [15, 682]. Continued [130, 172, 248, 268, 425, 634, 702, 740, 958, 1002, 1169, 1234, 1250, 1261, 1264, 1279, 1343, 1366, 1424, 1504, 1547, 1557, 1609, 1705, 1706]. continues [916, 1574]. Continuous [15, 94, 96, 112, 226, 682, 916, 1574]. continuous-fraction [15]. Control [58, 195]. Controlling [332]. Convergence [410, 1526]. Convergent [1596]. convergents [1279]. converse [234, 1036, 1098, 1301]. Convexity [87]. Convolution [1153, 1182]. coordinates [1351]. core [297, 1442, 1543, 1611]. cores [1580]. Corps [26, 121, 125, 198, 283, 516, 729, 1027, 1197]. correlation [113, 605, 1456]. Correspondence [90, 119, 213, 1105, 1176, 1276]. correspondences [805, 806]. corresponding [1407]. Corrigendum [25, 83, 263, 485, 628, 1017, 1296, 1533, 1688, 1709–1712, 1714–1716]. cotangent [1383, 1384]. countable [1258]. Counter [97, 931]. counter-example [931]. Counter-examples [97]. counterparts [834]. Counting [98, 1120, 1239, 1473, 1692]. courbes [1068, 1427, 1572]. cover [618, 632, 644, 659, 674, 687, 699, 716, 728, 738, 751, 762, 777, 790, 801, 811, 819, 831, 841, 852, 854, 867, 880, 893, 910, 923, 939, 955, 1641]. Covering [551, 771, 799, 1488, 1528, 1664]. Coverings [318, 389]. covers [540, 771, 1011, 1059]. Coxeter [1556]. Cramér [164]. crank [1607]. cranks [123]. critére [916, 1192, 1709]. Criteria [726]. criterion [434, 488, 791, 916, 1192, 1532]. critical [1130, 1679, 1716]. cryptography [1590]. cubes [449, 787]. Cubic [287, 361, 422, 511, 560, 561, 589, 629, 691, 720, 1013, 1053, 1122, 1152, 1177, 1292, 1332, 1542, 1588, 1715]. cubiques [1332]. cuboids [820]. Cuntz [1705]. Curve [173, 380, 408, 617, 770, 821, 1425, 1523, 1531, 1590, 1691, 1695]. curves [56, 100, 103, 170, 186, 211, 218, 222, 244, 250, 259, 263, 276, 302, 322, 348, 363, 369, 370, 372, 389, 395, 398, 480, 497, 520, 538, 551, 568, 569, 593, 623, 626, 666, 692, 732, 733, 739, 741, 755, 760, 797, 827, 858, 860, 865, 873, 875, 894, 897, 913, 933, 968, 999, 1005, 1009, 1040, 1068, 1136, 1141, 1149, 1266, 1287, 1292, 1306, 1307, 1311, 1328, 1333, 1388, 1390, 1410, 1419, 1427, 1475, 1481, 1484, 1491, 1505, 1565, 1572, 1581, 1692, 1708]. cusp [91, 824, 1135, 1171, 1397, 1516, 1681]. cuspidal [1077, 1130]. cycle [710]. cycles [6, 255, 871, 1111]. Cyclic [105, 121, 561, 589, 1273, 1311, 1695]. cyclicity [395]. cyclique [121]. Cyclotomic [29, 69, 203, 249, 382, 500, 508, 876, 895, 924, 982, 1046, 1253, 1325, 1327, 1370, 1380, 1529, 1556, 1693]. cyclotomique [1370]. cyclotomy [746, 1714].

D. [594]. dans [43, 125, 207, 273, 346, 516, 916, 1123, 1427, 1574]. Darmon [388]. data [1664]. databases [1123]. Davenport [1336, 1676, 1700]. dead [1350]. December [127, 277, 418, 590, 750, 892, 1063, 1256, 1470, 1689]. decidability [365]. decidable [730]. Decomposable [489, 547, 884]. decomposition [814, 1126, 1347, 1421]. Dedekind [57, 137, 178, 217, 340, 376, 458, 460, 899, 1002, 1173, 1320, 1338, 1624]. defect [1507]. definability [365]. defined [363, 731, 828, 1120, 1473]. definite [157, 208, 585, 1158, 1657]. definition [733]. deformation [775]. Deformations [195, 232, 1041]. Degenerate [1173, 1318]. Degert [1326, 1525]. degré [121, 385, 1353]. degree [24, 34, 121, 152, 365, 385, 428, 430, 692, 703, 752, 812, 882, 926, 1058, 1131, 1292, 1353, 1540, 1686]. Degrees [244, 359, 1038]. Dembowski [294]. Demjanenko [445]. Denef [1411]. denominators [447, 885, 1456]. dense [146, 717, 718]. denses [146]. Densité [280, 667]. Densities [80, 466, 734, 745, 980, 1666]. Density [267, 280, 288, 587, 667, 741, 757, 928, 1013, 1156, 1205, 1277, 1310, 1474, 1688, 1691]. d’entiers [748]. Department [1713]. dependence [901]. dependent [1097]. Depth [874, 1062]. Depth-zero [874, 1062]. derivation [1632]. derivations [96]. derivative [129, 900, 946]. Derivatives [36, 111, 291, 460, 1677]. derived [42]. descent [827]. description [702]. Detecting [901]. determinant [883]. Determination [966, 1130, 1525]. determines [620]. Determining [424, 621]. deux [280]. development [15, 682]. développement [15, 682]. d’hypersurfaces [1466]. Dickson [1347]. Dickson-type [1347]. dictionary [1606]. difference [112, 913, 1287, 1511]. differences [1595]. different [151, 1438, 1608]. differentiability [1446]. Differential [65, 81, 187, 409, 444, 1015, 1078, 1335, 1421]. differentiation [1030]. différentielles [409]. digamma [1127, 1211, 1618]. digit [110, 111, 1577, 1684]. digits [220, 432, 861, 1120, 1366, 1589]. d’Igusa [89]. Dihedral [352, 812, 1317]. dilates [67]. dilogarithm [1653]. Dimension [271, 298, 385, 441, 574, 921]. dimensional [315, 323, 679, 694, 786, 944, 947, 1004, 1137, 1148, 1188, 1664]. Dimensions [296, 824, 1164, 1482]. d’indépendance [1572]. diophantienne [340]. Diophantine [16, 34, 55, 206, 224, 233, 240, 339, 340, 356, 365, 405, 408, 441, 473, 499, 547, 599, 646, 778, 782, 787, 796, 813, 856, 872, 974, 1001, 1038, 1180, 1200, 1290, 1299, 1347, 1378, 1487, 1533, 1557, 1601, 1603, 1627, 1713]. Diophantus [42, 683]. DiPippo [83]. direct [963]. Dirichlet [36, 51, 108, 171, 570, 608, 881, 1032, 1106, 1134, 1346, 1394, 1401, 1592, 1654, 1677]. Discrepancy [296, 622, 1302]. discrete [772, 900, 1639, 1711]. discriminant [140, 287, 422, 1102]. discriminants [393, 468, 494, 1053, 1267]. dispersion [94]. distance [207, 617, 952, 1594]. distances [1351, 1376]. distinct [975, 1121, 1247, 1255]. distinguished [795, 1004, 1213]. distributed [67, 580, 1302]. Distribution [7, 32, 151, 160, 194, 210, 220, 256, 286, 287, 308, 432, 459, 543, 550, 613, 636, 637, 693, 747, 803, 821, 870, 962, 1002, 1022, 1031, 1080, 1094, 1254, 1391, 1405, 1420, 1423, 1512, 1562, 1584, 1628, 1663, 1680]. Distributions [13, 192, 214, 261, 347, 755, 903, 1385]. disturbed [1094]. divide [1636]. diviseurs [146]. Divisibility [407, 660, 974, 979, 1045, 1141, 1660]. divisible [18, 871, 1304, 1428]. division [202, 638, 1328]. Divisor [752, 881, 1408, 1566, 1584, 1698]. Divisors [146, 228, 625, 717, 718, 817, 917, 926, 979, 1138, 1419, 1492, 1566, 1671]. do [618]. does [46, 1368]. dominated [1475]. dont [46, 861]. Double [318, 450, 587, 650, 677, 743, 920, 1032, 1051, 1076, 1125, 1169, 1323, 1633, 1641]. doubling [559]. Drinfeld [132, 200, 250, 261, 282, 405, 467, 569, 596, 616, 627, 669, 781, 918, 928–930, 963, 999, 1079, 1139, 1194, 1364, 1377, 1397, 1485, 1538, 1544, 1545, 1563, 1696]. droite [409]. dual [429, 1608]. Duality [492, 1051]. due [212]. d’un [121, 125, 207, 729]. d’une [125, 189, 333, 1195, 1370]. Dyadic [80, 564, 987]. Dyer [353, 837, 1426, 1581]. dynamical [255, 1197, 1582]. Dynamics [155, 1003, 1405]. dynamiques [1197]. Dyson [1484].

écriture [46]. edges [793, 1286]. Editorial [419, 439, 440, 455, 471, 487, 503, 514, 529, 542, 556, 567, 579, 591, 602, 615, 632, 644, 659, 674, 687, 699, 716, 728, 738, 751, 762, 777, 790, 801, 811, 819, 831, 841, 854, 867, 880, 893, 910, 923, 939, 955, 973, 984, 997, 1008, 1021, 1035, 1050, 1064, 1075, 1093, 1109, 1129, 1145, 1166, 1186, 1207, 1223, 1232, 1245, 1257, 1271, 1284, 1298, 1315, 1341, 1359, 1393, 1413, 1434, 1452, 1462, 1471, 1480, 1496, 1519, 1535, 1551, 1570, 1599, 1613, 1631, 1644, 1668, 1690]. Effective [377, 442, 1201, 1321, 1530]. effectiveness [1168]. effectivité [1168]. Egyptian [99]. Ehrhart [961]. Eichler [700, 1465]. eigenforms [399, 906, 1246, 1335]. Eigenfunctions [847, 1490]. eigenspaces [1410, 1482]. eigenvalues [160, 605, 613, 684, 965, 1530]. eighth [1681]. Einheiten [68]. Einheiten- [68]. Eisenstein [135, 310, 412, 495, 562, 570, 576, 607, 640, 641, 887, 958, 981, 1060, 1312, 1329, 1383, 1384, 1609, 1704]. Elasticity [338]. element [188]. elementary [23, 47, 860, 1054, 1218]. Elements [69, 274, 785, 842, 1300, 1454, 1498]. Eleven [1376]. Elliptic [41, 56, 100, 161, 170, 186, 199, 211, 222, 244, 263, 302, 348, 359, 363, 370, 372, 380, 395, 398, 480, 520, 530, 568, 593, 606, 623, 626, 639, 666, 733, 739, 741, 755, 760, 770, 797, 858, 860, 865, 873, 875, 897, 913, 968, 979, 1009, 1040, 1118, 1136, 1141, 1149, 1205, 1287, 1306, 1307, 1311, 1328, 1333, 1338, 1388, 1390, 1410, 1425, 1475, 1481, 1491, 1505, 1565, 1572, 1590, 1691, 1692, 1695, 1708]. elliptiques [1572]. elucidation [1029]. embed [1427]. embedding [140, 700]. embeddings [1460]. endomorphism [79]. endomorphisms [112, 1115]. ends [1350]. Engel [22, 581]. Enlargement [1483]. enough [1376]. Ensembles [280, 748, 850]. entier [1420]. entières [748]. Entiers [146, 280, 861, 1370, 1420]. entire [1345]. entries [1200]. Entropy [281]. Enumeration [553, 604, 1017]. Ephraim [763]. epsilon [234]. Epstein [1324]. Equal [392, 1038, 1606]. Equality [571, 1683]. equation [16, 136, 285, 339, 340, 356, 388, 473, 595, 654, 668, 741, 749, 782, 796, 813, 982, 992, 1001, 1015, 1174, 1175, 1200, 1255, 1289, 1378, 1436, 1487, 1533, 1713]. equations [68, 81, 112, 164, 187, 206, 381, 519, 558, 638, 646, 691, 711, 856, 872, 976, 977, 1001, 1058, 1160, 1214, 1347, 1360, 1505, 1601, 1653, 1659, 1715]. Equidistribution [399, 1094, 1246, 1530]. equivalence [283, 284, 498, 896, 1332]. equivariant [131, 1056, 1574]. équivariantes [1574]. Erdos [175, 575, 665, 949, 1009, 1016, 1238, 1440]. Erratum [84, 126, 589, 629, 760, 953, 1062]. error [9, 37, 347, 881]. Escott [1137]. espace [1427]. espaces [43, 409]. Essentialité [1123]. Essentiality [1123]. est [861]. estimate [323, 1088, 1188, 1560]. estimates [37, 279, 1084]. eta [465, 1381]. eta-product [465]. étale [69, 475, 1026]. étales [1027]. Étude [1262]. Euclidean [63, 784, 1070, 1376, 1647]. Euler [214, 231, 328, 670, 726, 742, 772, 780, 839, 859, 905, 935, 942, 989, 1044, 1179, 1265, 1288, 1478, 1507, 1592, 1610, 1623, 1637, 1640]. Euler-type [989]. Evaluation [294, 526, 650, 920, 1182]. Evaluations [899, 942, 1437, 1654]. Even [88, 355, 608, 639, 978, 1372, 1648, 1695]. Everett [83]. every [1506]. everywhere [1094]. Exact [82, 188, 1313]. example [931, 1343]. examples [97, 586, 986]. Exceptional [10, 181, 465, 545, 603, 779, 864, 1026, 1291, 1336]. exceptions [1158]. Exchanging [855]. exercises [987]. Exhibiting [993]. exist [1368]. Existence [298, 347, 415, 468, 474, 926, 1111, 1217]. Expander [1590]. expansion [248, 582, 634, 1343, 1589]. Expansions [22, 172, 581, 702, 804, 1234, 1250, 1259, 1261, 1450, 1536, 1705]. Explicit [137, 316, 476, 570, 584, 608, 775, 814, 834, 942, 966, 1040, 1228, 1241, 1260, 1282, 1437, 1498, 1608]. exponent [912]. Exponential [113, 182, 239, 386, 430, 646, 655, 681, 721, 919, 925, 1134, 1602, 1642]. exponentially [1054]. exponentiation [710]. Exponents [222, 260, 263, 373, 388, 523, 1627]. Expressible [3, 84]. expressions [1099]. Extended [899, 942, 1407, 1637]. extensibility [1290]. extension [42, 125, 212, 424, 635, 645, 1370, 1372, 1373, 1502, 1621, 1686]. Extensions [21, 24, 26, 133, 141, 153, 254, 273, 301, 311, 365, 382, 390, 429, 453, 462, 485, 500, 563, 571, 604, 626, 645, 656, 752, 812, 828, 836, 846, 860, 882, 891, 914, 982, 987, 990, 1017, 1041, 1056, 1082, 1083, 1150, 1157, 1190, 1208, 1218, 1267, 1296, 1317, 1334, 1454, 1508, 1568, 1625]. extremal [242].

face [820]. facteur [1420]. facteurs [280]. factor [935, 1356, 1420]. factorial [1318]. factorials [260, 320]. Factoring [362]. factorisatio [1181]. Factorization [54, 174, 260, 523, 573, 1096, 1159, 1260, 1615]. factorizations [32, 286]. factors [32, 234, 280, 754, 857, 1292, 1304, 1439, 1556]. Fadwa [1713]. faiblement [273]. False [1621]. families [302, 415, 551, 664, 756, 915, 1135, 1136, 1450, 1600, 1696]. family [8, 56, 423, 561, 589, 691, 704, 732, 835, 1043, 1428, 1695, 1715]. Farey [447, 693, 885]. Farhi [1703]. Fast [383]. February [14, 154, 303, 454, 614, 776, 922, 1092, 1283, 1495]. Fermat [16, 285, 328, 410, 450, 473, 749, 1357, 1425, 1435, 1464]. Fermat-type [16, 473]. Fermatian [198]. fermatiens [198]. few [485]. fibers [606, 1205]. Fibonacci [1054]. Fibrations [826, 1205]. Field [13, 111, 115, 121, 179, 201, 287, 298, 368, 420, 424, 476, 537, 550, 651, 654, 729, 733, 752, 795, 828, 846, 980, 985, 1036, 1122, 1164, 1180, 1220, 1267, 1325, 1403, 1473, 1483, 1502, 1523, 1526, 1658, 1664, 1686]. Fields [1, 18, 30, 42, 75, 83, 103, 128, 145, 179, 181, 183, 192, 198, 226, 249, 283, 292, 306, 342, 352, 360, 361, 367, 370, 390, 415, 423, 430, 445, 468, 477, 483, 494, 516, 522, 535, 564, 583, 604, 623, 666, 675, 690, 712, 731, 743, 767, 823, 849, 863, 876, 896, 912–914, 924, 926, 976, 980, 1000, 1001, 1017, 1026, 1027, 1039, 1046, 1052, 1070, 1091, 1102, 1103, 1142, 1157, 1178, 1201, 1227, 1229, 1234, 1263, 1307, 1325, 1326, 1361, 1371, 1386, 1402, 1417, 1428, 1436, 1453, 1454, 1475, 1488, 1497, 1501, 1517, 1565, 1586, 1712]. fields [38, 47, 68, 105, 125, 148, 170, 203, 211, 245, 253, 274, 330, 345, 348, 357, 365, 394, 407, 428, 456, 511, 533, 537, 538, 571, 593, 603, 606, 645, 647, 662, 672, 688, 692, 720, 722, 871, 882, 927, 933, 1053, 1056, 1057, 1101, 1113, 1115, 1169, 1193, 1203, 1218, 1247, 1295, 1321, 1328, 1339, 1380, 1385, 1394, 1408, 1410, 1494, 1542, 1567, 1568, 1571, 1578, 1649, 1660, 1666, 1676, 1695]. fini [498]. finis [1197]. Finite [1, 47, 67, 83, 103, 116, 148, 179, 211, 274, 293, 306, 314, 351, 368, 370, 406, 430, 498, 500, 535, 540, 549, 553, 598, 611, 642, 720, 735, 933, 976, 985, 1004, 1115, 1122, 1135, 1163, 1164, 1197, 1215, 1229, 1234, 1342, 1386, 1400, 1415, 1443, 1497, 1523, 1524, 1679, 1716]. Finite-dimensional [1004]. finitely [1158]. Finiteness [64, 104, 335, 390, 475, 484, 537, 695]. first [285, 1167]. Five [711, 1058, 1255, 1428, 1513]. fixed [857, 1225]. fixing [539]. flag [360]. Flat [368, 985, 1253, 1342]. flows [1293]. fluctuations [1523]. Fonction [346, 1022, 1195, 1379]. fonctionnel [346]. fonctions [89, 120, 516, 748, 1665]. Fontaine [30, 1606]. Fontaine-Theory [1606]. form [68, 380, 838, 884, 1154, 1349, 1580, 1651]. Formal [15, 534, 535, 596, 836, 1115, 1197, 1526, 1527, 1563]. formel [1197]. formelles [15]. Formes [409, 747, 932, 1262, 1353]. Forms [65, 80, 91, 109, 140, 157, 161, 168, 169, 221, 223, 257, 262, 304, 316, 371, 397, 409, 413, 414, 444, 458, 461, 472, 481, 489, 547, 628, 639, 676, 678, 684, 690, 703, 719, 734, 747, 766, 817, 824, 833, 918, 932, 958, 966, 1077, 1078, 1097, 1135, 1152, 1158, 1171, 1189, 1201, 1204, 1235, 1240, 1262, 1285, 1353, 1365, 1388, 1397– 1399, 1406, 1407, 1465, 1477, 1482, 1513, 1516, 1540, 1586, 1605, 1621, 1629, 1634, 1648, 1656, 1681, 1696, 1699, 1710, 1712]. formula [79, 100, 217, 266, 321, 445, 467, 495, 512, 570, 594, 613, 627, 705, 838, 839, 848, 989, 1082, 1198, 1212, 1248, 1260, 1285, 1324, 1442, 1469, 1546, 1562, 1660, 1678, 1687]. formulae [834, 1058]. formulas [341, 747, 814, 918, 986, 1110, 1154, 1293, 1295, 1309, 1313, 1624]. formules [747]. four [400, 449, 619, 1235, 1322]. Four-Squares [400]. fourfolds [720]. Fourier [376, 690, 1329, 1516, 1681, 1712]. fourth [610, 723, 1464, 1698]. fourths [275]. Fractal [1446]. fractals [1638]. Fraction [15, 130, 172, 248, 634, 682, 702, 1234, 1250, 1261, 1343, 1547, 1609, 1705]. fractional [39, 151, 1382, 1503]. fractions [99, 268, 396, 425, 447, 740, 916, 958, 1002, 1169, 1264, 1279, 1366, 1424, 1504, 1557, 1706]. free [488, 558, 588, 611, 759, 791, 844, 1044, 1055, 1585]. French [15, 26, 43, 46, 89, 95, 120, 121, 125, 146, 198, 207, 273, 280, 283, 284, 333, 340, 346, 385, 409, 421, 498, 516, 667, 682, 729, 747, 748, 826, 850, 861, 889, 916, 932, 1022, 1027, 1068, 1094, 1123, 1168, 1192, 1195, 1197, 1262, 1308, 1332, 1353, 1370, 1379, 1420, 1427, 1466, 1572, 1574, 1665, 1669]. frequencies [432]. Frequency [1589]. fresh [185]. Frey [1042]. Frobenius [138, 261, 376, 494, 755, 964, 1116, 1350, 1680]. front [632, 644, 659, 674, 687, 699, 716, 728, 738, 751, 762, 777, 790, 801, 811, 819, 831, 841, 854, 867, 880, 893, 910, 923, 939, 955]. full [1339, 1367]. Fun [1593]. Function [11, 30, 111, 128, 149, 179, 183, 185, 192, 219, 220, 226, 253, 256, 282, 324, 330, 342– 346, 348, 360, 424, 428, 431, 445, 456, 479, 516, 538, 559, 588, 623, 624, 640, 641, 662, 675, 677, 694, 731, 743, 744, 753, 765, 823, 846, 882, 900, 912, 924, 926, 927, 944, 956, 957, 1001, 1003, 1029, 1030, 1037, 1044, 1095, 1100, 1103, 1125, 1127, 1142, 1167, 1195, 1201, 1211, 1220, 1303, 1324, 1367, 1379, 1402, 1403, 1405, 1408, 1422, 1442, 1446, 1449, 1454, 1459, 1475, 1566, 1577, 1586, 1618, 1639, 1663, 1687, 1698, 1711]. function [38, 105, 170, 209, 245, 304, 625, 692, 752, 799, 904, 1022, 1113, 1169, 1385, 1394, 1410, 1473, 1521, 1584, 1660, 1666]. Functional [164, 346, 381, 595, 654, 1032, 1653]. Functions [9, 23, 25, 36, 70, 72, 81, 89, 96, 100, 103, 108, 115, 120, 137, 152, 155, 161, 164, 169, 171, 176, 180, 184, 197, 199, 226, 253, 297, 317, 329, 341, 357, 379, 430, 451, 452, 459, 495, 533, 544, 621, 634, 677, 681, 722, 781, 791, 821, 822, 848, 899, 924, 925, 934, 991, 1005, 1032, 1058, 1066, 1081, 1099, 1110, 1132, 1148, 1172, 1196, 1198, 1208, 1217, 1219, 1234, 1278, 1281, 1291, 1312, 1313, 1316, 1319, 1336, 1345, 1377, 1383, 1384, 1409, 1438, 1463, 1477, 1515, 1576, 1579, 1583, 1610, 1633, 1655, 1665, 1672, 1677, 1691, 1698]. functions [2, 137, 307, 446, 642, 735, 748, 888, 1101, 1177, 1362, 1429, 1522, 1592, 1654]. functors [633]. fundamental [239, 468, 561, 589, 1052, 1674]. further [466, 890].

Galois [104, 141, 195, 246, 273, 301, 308, 434, 463, 494, 509, 522, 534, 563, 574, 669, 689, 692, 779, 812, 845, 873, 928–930, 966– 968, 978, 1039, 1041, 1102, 1150, 1203, 1218, 1227, 1275, 1296, 1334, 1371, 1382, 1481, 1494, 1508, 1541, 1544, 1558, 1567, 1568, 1574, 1648]. Galois-equivariant [1574]. Galois-équivariantes [1574]. galoisienne [273]. game [1638]. games [1699]. gamma [743, 1356, 1633, 1639, 1711]. gamma-monomials [743]. Gap [243, 653]. gaps [243, 1303]. Garrett [1285]. Gauss [128, 212, 225, 249, 305, 307, 351, 526, 546, 661, 889, 1228, 1400, 1402, 1407]. Gaussian [115, 355, 784, 1282, 1647]. gcd [832]. gcd-closed [832]. Gekeler [183]. Gelfand [79, 518, 1457]. general [108, 610]. généralisations [340]. Generalised [1626]. généralisée [1669]. generalization [130, 404, 1212, 1346, 1575, 1676]. Generalizations [340, 417, 976]. Generalized [11, 57, 240, 307, 342, 431, 450, 472, 609, 665, 704, 726, 781, 986, 992, 1028, 1100, 1104, 1260, 1276, 1294, 1318, 1338, 1669]. Generalizing [1476]. generated [1345, 1706]. Générateurs [1370]. generating [343, 944, 1577]. Generation [78, 721, 1000]. Generators [1370]. generic [1545]. genres [125]. genus [125, 218, 259, 354, 456, 732, 827, 894, 933, 1011]. genus-changing [218]. geometric [1154, 1286]. Géométrie [333]. Geometry [333, 400, 405, 1650]. German [68, 1435]. Gille [284]. Ginzburg [1238]. given [298, 550, 820, 917, 959, 1415, 1473]. Global [13, 159, 345, 616, 617, 662, 741, 743, 802, 863, 882, 927, 934, 1001, 1410, 1447, 1650, 1664]. GLY [1088]. Goldbach [707, 723, 959, 1349]. Goldbach-type [723]. Göllnitz [317, 1547]. Golod [1683]. Goncharov [1146, 1147]. good [434, 1233]. goodness [1435]. Gordon [317, 1547]. Gorenstein [414]. Goss [11, 1622]. Götzky [400]. Gouvêa [332, 478]. graded [703]. Graham [1016]. Grandes [1379]. grands [682]. Granville [388, 1133]. Graphs [731, 856, 1590]. Greenberg [195, 1083]. GRH [1590]. Gross [705]. Grothendieck [1363]. Group [29, 47, 56, 121, 211, 246, 276, 333, 395, 397, 498, 534, 729, 785, 812, 873, 890, 912, 927, 930, 985, 1124, 1146, 1156, 1197, 1213, 1282, 1339, 1342, 1398, 1414, 1474, 1483, 1494, 1498, 1541, 1544, 1566]. Groupe [121, 333, 498, 729, 1197]. Groups [12, 58, 75, 131, 134, 186, 195, 205, 218, 222, 236, 241, 254, 263, 299, 308, 379, 416, 424, 480, 494, 500, 505, 522, 536, 538, 549, 553, 564, 571, 611, 662, 752, 774, 797, 836, 865, 882, 897, 961, 982, 1000, 1024, 1071, 1121, 1163, 1190, 1203, 1233, 1238, 1258, 1275, 1291, 1316, 1317, 1321, 1371, 1400, 1404, 1408, 1410, 1494, 1505, 1527, 1565, 1568, 1573, 1581, 1674, 1679, 1683, 1716]. grow [1054]. Guinand [164]. Güte [1435].

H [594]. H. [35]. Halberstam [1336, 1676]. half [116, 149, 481, 628, 676, 690, 1712]. half-integral [149, 481, 628, 676]. Halton [1512]. Hans [126]. happy [1375]. Hardy [495, 1104, 1687]. harmonic [275, 357, 670, 1294, 1324]. Hasse [394, 538, 666]. Hausdorff [385, 441]. Hauteur [89, 667]. hauteurs [346]. having [280, 550, 741, 1234]. Hawkins [1014]. Hecke [90, 129, 134, 241, 272, 316, 399, 414, 505, 521, 572, 613, 684, 722, 725, 839, 847, 848, 906, 965, 991, 1246, 1324, 1397, 1427, 1482, 1524, 1530, 1605, 1710]. Heegner [623, 1247, 1498]. Height [21, 89, 131, 281, 369, 626, 667, 913, 1091, 1266, 1287, 1403, 1473, 1671]. Heights [79, 346, 612, 638, 1414, 1443]. Heilbronn [182, 1210]. Herglotz [677]. Hermitian [161, 304, 379, 703, 849, 1158]. Heron [1061, 1118, 1225]. heuristic [1458]. heuristics [82, 292, 675]. high [36, 947, 1149, 1188]. high-dimensional [947, 1188]. Higher [102, 296, 623, 944, 1146, 1313, 1485, 1566, 1616, 1664]. higher-dimensional [944]. Higher-order [1616]. Hilbert [6, 60, 76, 179, 262, 306, 354, 420, 533, 678, 690, 966, 980, 1078, 1135, 1246, 1325, 1465, 1477, 1712]. Hilbertian [1039]. Hirzebruch [346]. Hisashi [84]. hits [1389]. holomorphic [676, 1171, 1463]. Homogeneous [505, 771, 822, 1353, 1694]. homogènes [1353]. Homological [1249]. homologie [1068]. homology [1068]. homomorphisms [652]. Hooley [49]. Hopf [1395]. Horn [401]. Howe [83]. Hurwitz [976, 1029, 1388]. hyperbolas [1351, 1448]. hyperbolic [160, 605, 1655]. hypercomplex [210]. Hyperdifferential [226]. Hyperelliptic [276, 897, 933, 943, 993, 1523, 1707]. hypergeometric [479, 545]. hypergraphs [116]. Hypermonogenic [1240]. hyperplane [1601]. hyperplanes [1277]. hypersurface [1419]. hypersurfaces [89, 114, 1332, 1466]. hypothèse [1192, 1669, 1709]. hypothesis [72, 625, 981, 1192, 1521, 1669].

icosahedral [978]. ideal [75, 310, 912, 1124, 1260, 1382]. Ideals [38, 153, 172, 516, 550, 688]. Idéaux [516]. Identities [22, 77, 135, 247, 949, 1040, 1069, 1153, 1155, 1372, 1400, 1583, 1637, 1673]. Identity [61, 70, 417, 496, 768, 1560, 1575, 1609]. ièmes [1353]. IFSs [921]. Igusa [89, 219, 1623]. Ihara [1417]. II [8, 131, 160, 270, 278, 293, 295, 322, 382, 484, 609, 637, 641, 660, 668, 795, 806, 929, 949, 962, 987, 1038, 1096, 1149, 1178, 1179, 1355, 1418, 1459, 1585]. III [119, 396, 483, 930, 1031, 1094, 1291, 1549]. Ikeda [1401]. Image [574, 730, 930, 1544]. images [779, 966]. imaginaire [125]. Imaginary [105, 125, 367, 394, 415, 583, 722, 767, 1157, 1203, 1247, 1263, 1417, 1428, 1494, 1649]. implies [369, 1266]. improvement [1280]. Improvements [1084]. Improving [1238]. Incomplete [748, 919, 1226]. incongruent [1012]. increasing [1366]. Indecomposability [688]. indefinite [257]. independence [254, 634, 1247, 1438, 1572, 1618, 1706]. Index [27, 52, 68, 73, 92, 117, 143, 166, 190, 215, 237, 264, 289, 312, 326, 349, 372, 374, 402, 435, 469, 501, 527, 550, 554, 577, 600, 630, 657, 676, 685, 714, 736, 760, 761, 788, 809, 829, 851, 865, 870, 877, 885, 907, 936, 962, 970, 994, 1018, 1031, 1047, 1072, 1339]. indexed [384]. Indexformgleichungen [68]. indices [1499]. inductive [1111]. inequalities [8, 233, 422, 431, 547, 599, 884, 1043, 1661]. inequality [324, 369, 875, 1079, 1139, 1266, 1299, 1509, 1701]. Inertia [774]. Infinite [209, 270, 301, 415, 429, 559, 873, 921, 980, 1199, 1261, 1381, 1450, 1697]. infinitely [1119, 1506]. Infinitesimal [252]. infinity [927]. Inose [756]. Integer [3, 20, 67, 84, 230, 681, 884, 951, 952, 1420, 1509, 1527, 1536, 1684, 1694]. integer-valued [20, 1694]. Integers [3, 32, 69, 80, 84, 88, 136, 210, 223, 257, 280, 286, 366, 396, 461, 592, 595, 598, 708, 717, 718, 746, 748, 759, 803, 825, 857, 861, 940, 964, 988, 1100, 1189, 1215, 1218, 1239, 1255, 1282, 1304, 1350, 1370, 1420, 1467, 1580, 1636, 1714]. Integral [17, 133, 149, 221, 269, 302, 323, 371, 382, 420, 429, 481, 525, 536, 564, 628, 676, 690, 888, 1010, 1046, 1057, 1088, 1179, 1277, 1329, 1491, 1576, 1608, 1645, 1661, 1712]. Integrality [863]. integrals [90, 174, 981, 1069, 1640, 1655, 1707]. integration [348, 769]. intermediate [510]. Interpolation [341]. interpretation [1249, 1404]. interval [94, 1499, 1548]. intervals [10, 113, 288, 442, 1322]. Invariant [847, 1149, 1251]. Invariants [132, 250, 314, 367, 568, 647, 666, 744, 990, 1040, 1263, 1292, 1371, 1481, 1621, 1649, 1702, 1704]. Inverse [1069, 1529, 1608]. inverses [543]. inversion [1110, 1293]. involutions [539]. involving [247, 328, 782, 925, 946, 1067, 1114, 1288, 1583, 1706]. Irrational [413]. Irrationality [391, 609, 726, 1127, 1230, 1272]. Irreducibility [60, 424, 517, 532]. Irregularities [7, 1385]. Isogeny [306, 890, 963, 1154, 1311]. isometries [78]. isomorphism [604, 1017]. Isospectral [1657]. isotropic [1000]. Issai [597]. iterated [209, 559]. itself [1574]. IV [7]. Iwasawa [201, 225, 314, 568, 647, 744, 990, 1444, 1621, 1625, 1702].

J [589, 628, 629, 760, 953, 1017, 1062, 1296, 1391, 1533, 1688, 1709–1716]. Jacobi [65, 159, 212, 496, 676, 747, 1251, 1605, 1710]. Jacobian [492, 890]. Jacobians [173, 259, 510, 943, 993, 1363]. January [4, 144, 290, 438, 601, 909, 1074, 1270, 1479]. Jarden [1042]. Jarvis’ [1562]. joint [1110, 1217]. Journal [453]. Julia [155, 850]. July [74, 216, 350, 528, 686, 830, 996, 1185, 1392, 1598]. June [66, 204, 337, 513, 673, 818, 983, 1165, 1358, 1569].

K. [1133, 1381]. Kac [1009]. Kakeya [1164]. Kalmár [1379]. Kato [58].

Katz [1084]. Katz-type [1084]. kernel [181, 208, 310, 397, 603, 729, 1319]. kernels [466, 712, 1026, 1308]. Kida [100, 1082]. kind [1469]. Klingen [570]. Kloosterman [108, 433, 477, 580, 610, 679, 960, 1226]. knapsack [362]. Kneser [404]. Knopp [417]. known [1255]. Kodaira [1015]. Koecher [570]. Kolyvagin [1498]. König [665]. Kronecker [627, 1295, 1323]. Kubota [452]. Kudla [1198]. Kummer [102, 258, 291, 456, 485, 758, 1626]. Kummer-type [291]. Kurihara [773]. Kuznietsov [613].

L. [1249]. lacunaires [748]. Lacunary [894, 1153]. Laguerre [517, 704]. Lambert [495]. Lang [252, 520, 1582]. Langlands [213, 1291]. Laplacian [160, 605]. Laplacians [1251]. Large [186, 279, 365, 480, 664, 682, 753, 1011, 1229, 1280, 1303, 1379, 1459, 1475, 1509, 1515]. Largest [3, 84, 1611]. Lattice [240, 539, 786, 1080, 1236, 1624]. Lattices [78, 205, 271, 355, 379, 493, 496, 525, 585, 869, 876, 1657]. Laurent [535]. law [458, 460, 531, 944]. laws [1197, 1320]. least [141, 218, 832, 1215, 1321]. Legendre [370, 1005]. Lehmer [594, 1079, 1117, 1139, 1492, 1614]. Lehner [157, 1635]. lemma [239, 665, 1023, 1305, 1417]. length [1504]. lengths [1279, 1553]. Lenstra [292, 675, 1404, 1458]. Leopoldt [452, 794, 855]. Lerch [986, 1110, 1217]. Level [272, 775, 1080, 1176, 1313, 1531, 1691]. levels [168]. Levinson [706]. Lévy [1156, 1474]. Li [791, 1192, 1709]. Lift [359]. Lifting [548, 1115, 1401, 1567]. liftings [678, 1704]. lifts [655]. like [384, 1057]. like-indexed [384]. Limit [256, 627, 1122, 1295, 1384]. Limiting [347, 464, 1584, 1680]. Limits [1694]. Lindelöf [1100]. line [409, 1346]. linéaire [1572]. linéaires [1262, 1574]. Linear [81, 112, 187, 254, 413, 452, 558, 638, 719, 901, 902, 1006, 1038, 1085, 1114, 1201, 1204, 1219, 1234, 1237, 1262, 1274, 1438, 1561, 1572, 1574, 1618, 1659, 1661, 1706]. linearizing [1582]. lines [1564]. Links [534]. Liouvillian [1094]. liouvillien [1094]. lisse [333]. Littlewood [945, 1327]. Ljunggren [1175]. LLL [696]. Local [80, 159, 213, 219, 249, 272, 283, 286, 534, 537, 546, 564, 616, 734, 795, 871, 882, 889, 896, 914, 985, 1036, 1139, 1180, 1218, 1356, 1447, 1482, 1641, 1658]. locales [889]. locally [1041]. locaux [283]. loci [444]. locus [1701]. log [1639, 1711]. logarithm [1037]. logarithmes [1262]. logarithmic [409, 1027]. logarithmique [1027]. logarithmiques [409]. logarithms [1204, 1262]. lois [1197]. long [493]. look [185, 1562]. Lösen [68]. Lower [21, 59, 179, 406, 626, 1100, 1116, 1215, 1309, 1367, 1389, 1453, 1536, 1691]. Lubin [1527]. Lucas’ [325, 701, 872, 1112, 1183, 1499, 1520]. lui [1574]. lui-même [1574]. Lüroth [1589]. Lyapunov [464].

même [1574]. Maass [570, 690, 745, 1430, 1712]. Mackey [633]. Mahler [136, 165, 512, 586, 634, 915, 1069, 1172, 1352, 1416, 1466, 1604, 1675]. Main [225]. manifolds [441]. Manin [783, 953, 1591]. many [389, 581, 1119, 1158, 1160, 1439, 1506]. maps [836, 901, 998, 1582, 1600]. March [28, 167, 313, 470, 631, 789, 938, 1108, 1297, 1518]. mark [1446]. Markoff [172, 340, 976]. Markov [835]. mass [869]. Massy [1039].

Mathematics [1713]. matrices [149, 832, 1275, 1638]. matrix [431, 445, 753, 1459]. mauvaise [1332]. Maximal [645, 1157, 1181, 1209, 1671]. May [53, 191, 327, 502, 658, 810, 972, 1144, 1340, 1550]. Mazur [30]. McKay [1161]. mean [9, 62, 108, 178, 217, 483, 511, 594, 610, 752, 888, 947, 1029, 1037, 1134, 1154, 1267, 1289, 1429]. means [898, 1655]. measure [165, 512, 814, 915, 1404, 1416, 1604]. measurement [1466]. Measures [245, 334, 391, 559, 586, 778, 1069, 1127, 1172, 1272, 1352, 1474, 1572, 1675]. mechanics [256, 1142]. medians [411]. members [1499]. meromorphic [180, 533]. meromorphy [1403]. Mersenne [770]. Mertens’ [1248]. Mesure [1466]. Mesures [1572]. metaplectic [1641]. method [386, 911, 925, 1239, 1291, 1510]. methods [267]. Metric [165, 425, 474, 1180, 1250, 1604]. Metrical [408, 1366]. metrics [1251]. mid [898]. mid-coefficients [898]. middle [1536]. Miki [768]. mild [974]. Minima [271, 849, 1070, 1305]. Minimal [64, 254, 494, 741, 917, 1163, 1635]. Minimum [1573]. Minkowski [694, 1446, 1555]. minoration [667]. minus [1565]. Mixed [519, 1549]. Möbius [149, 765, 1293]. Mod [463]. model [816]. Models [827, 999]. moderate [1136]. modulaires [932, 1068]. Modular [6, 64, 76, 100, 109, 161, 168, 169, 205, 250, 262, 282, 316, 317, 354, 414, 428, 444, 464, 472, 481, 496, 519, 569, 599, 628, 639, 676, 678, 684, 689, 703, 709, 775, 779, 817, 876, 888, 918, 932, 958, 966, 967, 969, 999, 1058, 1068, 1077, 1078, 1097, 1124, 1154, 1235, 1246, 1281, 1282, 1285, 1299, 1313, 1351, 1362, 1365, 1377, 1388, 1399, 1407, 1448, 1465, 1477, 1485, 1531, 1540, 1575, 1605, 1621, 1629, 1648, 1696, 1710]. Modularity [71, 733, 1307, 1328, 1708]. Module [201, 405, 563, 812, 928–930, 1068, 1334, 1364, 1563]. modules [132, 200, 261, 467, 507, 534, 574, 596, 616, 627, 781, 963, 1079, 1115, 1139, 1194, 1468, 1538, 1544, 1545, 1625]. Moduli [101, 123, 960, 1086, 1280, 1313, 1372, 1432, 1515, 1600]. modulo [75, 220, 262, 344, 532, 536, 543, 560, 618, 629, 710, 889, 905, 1012, 1094, 1268, 1274, 1310, 1333, 1414, 1478, 1525, 1663, 1685, 1688]. modulus [1528]. moment [725, 1167, 1367]. Moments [307, 640, 641, 900, 1607, 1681, 1698]. monochromatic [1170]. monogeneity [423]. monomials [318, 743]. Mordell [252, 398, 520, 977, 1032, 1079, 1410, 1582]. morphic [268]. most [280, 585]. motifs [421, 1489]. motives [963, 1365, 1497]. Motivic [1653]. Motzkin [950]. moyenne [861]. multi [724]. multi-zeta [724]. multidimensional [1094, 1448]. multidimensionnel [1094]. multifractal [432]. multinomial [228, 660, 661, 975]. Multiperfect [1369, 1564]. Multiple [44, 291, 544, 586, 670, 832, 847, 899, 1051, 1172, 1215, 1226, 1383, 1384, 1421, 1437, 1537, 1546, 1632, 1654, 1678]. Multiples [1491, 1684]. multiplication [197, 576, 595, 802, 858, 943, 1331, 1386, 1444, 1594]. multiplications [509]. Multiplicative [231, 386, 451, 1228, 1414]. multiplicity [565, 873, 1237, 1344, 1619]. multiquadratic [1571]. multivariable [1623]. Mumford [653]. Muriefah [1713]. my [485].

Nagell [1175]. Näherungen [1435]. narrow [1525]. Nathanson [1443]. natural [1156, 1672]. near [82, 171]. nearest [952]. Nearly [46]. necessarily [1077]. necessity [1486]. negative [287, 422, 1053]. newforms [159, 465, 745, 824, 1276, 1374]. Newman [1209]. Newton [42, 219, 430, 450, 532, 859, 1140]. Nicol [1337]. Nicolas [364]. Niebur [1329]. Niederreiter [1512]. Nochka [1202]. Nombres [26, 46, 95, 121, 385, 729, 1027, 1379]. Non [2, 25, 36, 80, 283, 348, 353, 395, 414, 423, 474, 533, 597, 606, 669, 769, 871, 926, 933, 1111, 1163, 1201, 1217, 1238, 1259, 1364, 1419, 1446, 1498, 1536, 1604]. non-abelian [1163, 1238]. Non-Archimedean [348, 1419]. non-differentiability [1446]. Non-divisible [871]. non-dyadic [80]. non-existence [474, 1111, 1217]. Non-hyperelliptic [933]. non-integer [1536]. non-linear [1201]. Non-monogeneity [423]. non-perfect [606]. non-residues [597]. non-semistable [353]. non-special [926]. non-torsion [1364]. non-trivial [1498]. non-unit [1259]. Non-Urs [25]. Non-vanishing [36]. nonabelian [1334, 1573]. nonanalytic [90]. Noncanonical [1467]. noncongruence [1399]. nonic [1281]. nonnegative [1381]. Nonrational [76]. nonresidues [1119]. Nonsingularity [832]. nonsplit [141]. Nonstandard [267]. Nontrivial [1215]. nonunits [1646]. Nörlund [1469]. Norm [156, 285, 564, 571, 836, 1568]. Normal [162, 380, 382, 429, 536, 1193, 1608]. Normality [115, 1345]. Normalized [898]. Normalizer [147]. Normalizers [241]. norms [132, 345, 678, 1171]. normset [338]. Note [187, 243, 245, 249, 318, 407, 415, 447, 473, 886, 934, 951, 977, 1133, 1188, 1209, 1248, 1445, 1517, 1532, 1610, 1614, 1647, 1713]. November [118, 265, 403, 578, 737, 879, 1049, 1244, 1461, 1667]. noyau [729]. Noyaux [1027, 1308]. nuclei [1027]. nulle [280]. Number [10, 25, 30, 32, 83, 84, 115, 121, 126, 136, 181, 201, 255, 263, 292, 321, 323, 331, 347, 363, 373, 390, 445, 453, 477, 488, 504, 538, 589, 592, 628, 629, 633, 639, 688, 690, 695, 705, 712, 729, 732, 760, 828, 849, 857, 913, 917, 941, 953, 959, 975, 1000, 1017, 1026, 1027, 1053, 1056, 1062, 1070, 1102, 1200, 1204, 1210, 1214, 1218, 1220, 1263, 1267, 1296, 1307, 1325, 1326, 1371, 1391, 1428, 1441, 1453, 1488, 1501, 1523, 1525, 1531, 1533, 1567, 1571, 1597, 1649, 1679, 1688, 1695, 1709–1716]. number [68, 105, 256, 357, 365, 396, 571, 593, 603, 692, 786, 1012, 1057, 1081, 1193, 1330, 1351, 1467, 1542, 1552, 1568, 1578, 1581, 1624, 1676]. number-theoretic [347]. Numbers [18, 19, 26, 38, 46, 95, 102, 137, 146, 151, 162, 172, 196, 229, 247, 278, 309, 328, 361, 385, 400, 407, 410, 412, 427, 468, 483, 515, 557, 582, 649, 666, 670, 671, 768, 780, 792, 808, 834, 844, 895, 905, 935, 945, 950, 964, 1033, 1055, 1065, 1091, 1112, 1131, 1153, 1156, 1161, 1181, 1239, 1265, 1288, 1294, 1310, 1330, 1345, 1369, 1375, 1379, 1388, 1389, 1396, 1414, 1439, 1450, 1469, 1478, 1549, 1564, 1592, 1601, 1616, 1628, 1640, 1660, 1672, 1680, 1688]. numerorum [1181].

objects [1004]. obstruction [372, 760]. obtained [1415]. occupy [786]. octahedral [987]. Octic [1361, 1520]. October [106, 251, 387, 566, 727, 866, 1034, 1231, 1451, 1643]. Odd [173, 271, 447, 856, 935, 1255, 1369]. one [39, 105, 188, 200, 259, 315, 354, 565, 756, 827, 894, 959, 1053, 1060, 1136, 1148, 1193, 1344, 1349, 1508, 1531, 1578, 1619, 1691]. one-dimensional [315, 1148]. one-level [1691]. one-parameter [756, 1136]. one-sided [39]. ones [75]. only [1255, 1508]. Ono [158, 367, 1263, 1649]. open [1140]. Openness [574, 929, 1545]. operator [61]. operators [65, 226, 316, 505, 521, 572, 847, 991, 1078, 1397, 1430, 1490, 1530, 1605, 1710]. Oppenheim [1250]. Optimal [230, 783, 875, 953, 1116, 1460]. Optimally [549]. orbit [1506]. orbits [371, 1661, 1685]. order [102, 150, 443, 526, 636, 637, 662, 722, 785, 842, 870, 947, 962, 1024, 1031, 1181, 1253, 1300, 1367, 1415, 1436, 1514, 1573, 1616, 1691]. ordered [32, 286]. orderings [474, 823]. orders [361, 700, 752, 802, 881, 1224, 1536]. Ordinary [720, 1149, 1287, 1390]. origin [207]. origine [207]. orthogonal [1000]. oscillantes [1094]. oscillating [1094]. oscillations [451]. Ostrom [294]. other [343, 1381]. Overpartitions [672, 1372, 1607].

P. [284]. packings [504, 1010]. Padé [309]. Pages [4, 14, 28, 40, 53, 66, 74, 85, 93, 106, 118, 127, 144, 154, 167, 177, 191, 204, 216, 227, 238, 251, 265, 277, 290, 303, 313, 319, 327, 337, 350, 358, 375, 387, 403, 418, 438, 454, 470, 486, 502, 513, 528, 541, 555, 566, 578, 590, 601, 614, 631, 643, 658, 673, 686, 698, 715, 727, 737, 750, 776, 789, 800, 810, 818, 830, 840, 853, 866, 879, 892, 909, 922, 938, 954, 972, 983, 996, 1007, 1020, 1034, 1049, 1063, 1074, 1092, 1128, 1165, 1206, 1256]. Pages [25, 83, 84, 126, 263, 1108, 1144, 1185, 1222, 1231, 1244, 1270, 1283, 1297, 1314, 1340, 1358, 1392, 1412, 1433, 1451, 1461, 1470, 1479, 1495, 1518, 1534, 1550, 1569, 1598, 1612, 1630, 1643, 1667, 1689]. Pair [113, 605, 1456]. paired [483]. pairing [119, 448, 1363, 1390]. pairs [304, 371, 398, 518, 584, 859, 1121, 1290]. Papanikolas [1532]. paper [182, 485, 1039, 1133]. parameter [756, 1136]. parameterized [1695]. parameters [169, 1438]. parametric [691, 1715]. parametrizable [1416]. Parametrization [18]. parametrizations [244, 428]. Parity [260, 523, 885, 1095, 1311, 1444, 1565, 1571]. Part [805, 806, 920]. partial [278, 682, 1234, 1278, 1424]. partiels [682]. Partition [175, 185, 297, 344, 949, 1095, 1442, 1611, 1663]. Partitions [123, 138, 243, 364, 1281, 1543, 1672]. partout [1094]. parts [39, 151, 1503]. Pascal [1564]. past [185]. Patterns [543, 1119, 1159]. peculiar [1513]. PEL [1432]. Pell [668, 1214, 1360]. perfect [606, 863, 935, 1506]. period [174, 372, 760, 865, 991, 1279, 1504, 1655]. period-index [372, 760, 865]. Périodes [421, 932, 1572]. Periodic [196, 835, 850, 1259, 1264, 1504, 1582]. periodicity [969]. périodiques [850]. Periods [334, 421, 932, 1572, 1639, 1685, 1711]. permutation [642]. permutations [862]. permutative [1705]. perturbées [1094]. Pesenti [875]. Petersson [678, 1077]. petit [1420]. phase [386]. Picard [218, 709, 1154, 1195, 1282, 1362, 1407, 1506]. pictures [1147]. Pierce [1492]. Pillai [457, 992]. Pisot [19, 427, 649, 695, 895, 945, 1065, 1259, 1450]. Pisot-cyclotomic [895]. places [735, 855]. Plancherel [814]. Plane [497, 999, 1652]. Platonic [448]. plongent [1427]. plus [280, 1320, 1565]. plus/minus [1565]. poids [932]. Poincaré [1329, 1411]. Point [36, 207, 333, 1364, 1514, 1692]. Points [17, 34, 43, 98, 124, 131, 189, 218, 240, 269, 302, 323, 333, 360, 389, 395, 408, 540, 581, 617, 623, 655, 667, 732, 786, 835, 843, 850, 858, 1013, 1080, 1088, 1195, 1236, 1247, 1254, 1277, 1351, 1391, 1432, 1448, 1491, 1498, 1523, 1582, 1624, 1695]. poles [70, 1132, 1291]. Polignac [1396]. polygons [430, 1273]. polyhedra [219]. polylogarithms [586]. polynômes [850]. Polynomial [115, 321, 426, 531, 642, 666, 668, 683, 915, 1324, 1472, 1509, 1511, 1577]. Polynomials [1, 2, 20, 25, 42, 83, 107, 180, 202, 291, 294, 309, 335, 362, 430, 484, 508, 517, 532, 561, 589, 598, 642, 702, 704, 839, 850, 902, 961, 1038, 1080, 1106, 1126, 1140, 1209, 1252, 1253, 1275, 1318, 1327, 1347, 1416, 1485, 1493, 1503, 1527, 1529, 1556, 1585, 1637, 1640, 1673, 1693, 1694]. polytopes [376, 532]. poor [1332]. Porosity [209]. posed [1703]. Positive [5, 208, 221, 299, 371, 585, 729, 834, 1088, 1158, 1180, 1187, 1205, 1661]. positives [729]. possédant [280]. Postage [242]. Potential [1708]. Potentially [1233]. pour [1192, 1709]. Power [54, 105, 108, 133, 203, 228, 235, 368, 526, 573, 610, 842, 898, 914, 952, 1046, 1057, 1090, 1096, 1151, 1216, 1300, 1343, 1445, 1464, 1499, 1526, 1538, 1585, 1615]. power-free [1585]. Powers [45, 88, 140, 151, 194, 229, 231, 274, 295, 392, 568, 720, 723, 758, 905, 934, 951, 964, 1033, 1044, 1304, 1353, 1396, 1455, 1478, 1506, 1561, 1595, 1663]. premier [121, 1420]. premiers [46, 280, 1420]. prescribed [990]. Prescribing [1514]. presence [982]. preserved [338]. presque [46, 1094]. primality [1425]. primary [276]. Prime [46, 54, 105, 121, 140–142, 151, 228, 260, 280, 413, 543, 550, 560, 573, 629, 646, 710, 711, 758, 842, 857, 863, 927, 959, 978, 1009, 1025, 1046, 1081, 1096, 1138, 1274, 1292, 1300, 1304, 1322, 1325, 1349, 1420, 1436, 1439, 1456, 1508, 1615, 1659, 1686]. prime-power [1046]. Primes [10, 37, 41, 88, 113, 239, 295, 352, 410, 442, 449, 770, 775, 825, 843, 1003, 1014, 1119, 1170, 1385, 1405, 1425, 1513, 1602, 1651, 1685]. Primitive [33, 82, 126, 194, 274, 315, 507, 608, 757, 798, 858, 979, 1178, 1191, 1492]. Principal [134, 172, 1224, 1260, 1641]. principle [110, 538, 653]. Pro [26, 645]. Pro- [26, 645]. probabilities [1404]. probability [1191]. problem [35, 97, 105, 128, 140, 162, 184, 185, 242, 256, 300, 362, 376, 382, 394, 510, 533, 552, 594, 616, 683, 707, 723, 787, 864, 865, 881, 959, 964, 998, 1053, 1111, 1116, 1137, 1164, 1181, 1193, 1326, 1337, 1349, 1350, 1432, 1465, 1472, 1531, 1578, 1582, 1620, 1693]. problems [37, 44, 267, 775]. product [6, 231, 384, 465, 483, 709, 839, 1362, 1502, 1573, 1623]. Products [872, 906, 986, 1067, 1077, 1199, 1261, 1323, 1381, 1475, 1499, 1620]. Progression [17, 611, 793, 1286]. Progression-free [611]. progressions [296, 844, 904, 1248, 1360, 1515, 1516, 1596]. projectifs [43]. projective [43, 114, 124, 409, 1566]. Proof [86, 175, 331, 400, 433, 619, 838, 1054, 1111, 1122, 1202, 1554]. properties [44, 309, 338, 381, 425, 607, 660, 859, 883, 935, 1045, 1085, 1097, 1250, 1366, 1387, 1512]. property [19, 427, 559, 636, 637, 695, 1033]. Proportionally [599, 1299]. Prouhet [1137]. pseudomagic [957]. Pseudomoments [957]. pseudorandom [664]. Pseudorepresentations [232]. puissances [1353]. pull [1324]. pull-back [1324]. pullback [1285]. Punctured [192]. Pure [1497]. Purely [1259]. Pythagorean [198, 275]. pythagoriciens [198].

Q2 [713]. QM [816]. QM-abelian [816]. Quadratic [15, 18, 33, 42, 75, 80, 119, 125, 126, 140, 145, 221, 233, 257, 307, 351, 352, 367, 371, 397, 398, 407, 415, 468, 477, 483, 522, 550, 583, 597, 645, 666, 672, 702, 707, 711, 719, 722, 734, 752, 759, 766, 767, 833, 838, 912, 919, 940, 977, 980, 1113, 1119, 1152, 1157, 1169, 1178, 1203, 1247, 1263, 1292, 1295, 1302, 1321, 1326, 1331, 1339, 1354, 1406, 1408, 1417, 1428, 1494, 1513, 1580, 1634, 1649, 1666]. quadratique [125]. quadratiques [15]. quadrics [1006, 1406]. quadruple [1603]. Quantum [50, 595, 746, 940, 1142, 1714]. Quartic [8, 133, 199, 333, 560, 629, 833, 1043, 1343, 1520, 1578]. quartique [333]. Quasi [334, 961, 1264, 1380, 1504, 1572, 1632, 1696]. quasi-cyclotomic [1380]. quasi-derivation [1632]. quasi-modular [1696]. quasi-périodes [1572]. Quasi-periodic [1264, 1504]. quasi-periods [334, 1572]. quasi-polynomials [961]. quasimodular [1398]. Quasirandom [862]. quaternary [838]. quaternion [24, 157, 330, 1460, 1588, 1698]. quaternionic [838, 1132, 1362]. Quelques [1427]. question [1042, 1446, 1622]. questions [288, 1402, 1703]. quinary [221]. quintic [206, 1493, 1695]. quintuple [1603]. quotient [201, 783, 953, 1464]. Quotients [173, 278, 328, 682, 887, 1234, 1357, 1424, 1635].

R [1039]. Rabinowitsch [335, 484]. races [821]. Rademacher [1338]. radicalized [1266]. Rallis [1198]. Ramanujan [22, 77, 130, 138, 199, 519, 740, 1281, 1372, 1547, 1557, 1609, 1617]. Ramanujan-Type [138, 1617]. Ramification [311, 424, 453, 656, 985, 1342, 1486, 1658]. ramified [806, 1011, 1508, 1625, 1686]. ramifiées [273]. random [32, 431, 753, 1014, 1054, 1459]. range [799]. Rank [200, 208, 261, 301, 311, 355, 453, 466, 585, 620, 739, 927, 977, 1042, 1136, 1149, 1205, 1220, 1321, 1386, 1475, 1485, 1538, 1566, 1607, 1701]. Rankin [1066, 1401]. ranks [75, 123, 398, 520, 1311, 1666]. Rational [26, 42, 98, 124, 130, 155, 192, 218, 246, 360, 376, 380, 389, 395, 411, 424, 540, 730, 732, 744, 793, 820, 821, 925, 952, 1013, 1061, 1234, 1262, 1273, 1286, 1332, 1517, 1600, 1645, 1695]. rational-derived [42]. Rationality [572, 681, 1411]. rationals [299, 1348, 1456, 1500]. rationnel [1262]. rationnelle [1332]. rationnels [26]. ray [145, 536]. Raynaud [1468]. Real [1, 62, 83, 171, 352, 365, 407, 468, 477, 509, 550, 582, 645, 647, 672, 943, 980, 1053, 1088, 1113, 1131, 1169, 1295, 1307, 1326, 1346, 1565, 1594, 1655]. Realizable [24, 462, 812, 1334]. Realization [236, 1553]. realizations [1558]. reals [1348, 1500]. rearrangements [1597]. Reciprocal [3, 84]. reciprocals [391, 410]. reciprocity [458, 460, 531, 944, 1320, 1538]. Rectilinearization [1411]. recurrence [320, 391, 587, 1219, 1237, 1274]. recurrences [384, 964, 1114, 1153, 1455, 1616, 1706]. recurrent [1085]. recursion [445]. recursively [731]. Reducibility [114]. reducible [1493]. reducing [198]. Reduction [155, 353, 434, 652, 667, 767, 901, 998, 1233, 1332, 1364, 1629]. reductions [1514, 1692]. réduisants [198]. refined [705, 931, 1532]. refinement [1542]. refinements [37]. Reflection [1408]. regarding [1693]. regions [791]. regular [344, 582, 847, 1559, 1603]. Regularity [1159]. regularized [986]. Regularly [748]. regulator [1453]. regulators [1352]. régulièrement [748]. related [102, 116, 184, 250, 410, 646, 670, 950, 1069, 1347, 1402, 1425, 1601, 1640, 1670]. relation [1147, 1632]. Relations [75, 231, 317, 452, 495, 837, 1002, 1032, 1051, 1052, 1146, 1228, 1274, 1281, 1426, 1537]. Relative [38, 133, 137, 420, 445, 1043, 1190, 1408, 1660]. Remark [71, 182, 1424]. remarks [1039, 1076]. Removing [188, 1415]. renormalization [1421]. répartition [1022, 1420]. repeated [710, 1030]. representable [396]. Representation [88, 140, 624, 689, 838, 873, 1130, 1513, 1544, 1576, 1672, 1702]. Representations [58, 80, 104, 221, 230, 257, 366, 463, 509, 510, 548, 565, 592, 754, 779, 928– 930, 959, 966– 968, 978, 1028, 1041, 1213, 1374, 1406, 1524, 1552, 1567, 1580, 1641, 1658, 1705]. represented [702]. Representing [229]. residual [636, 637]. residue [156, 285, 383, 550, 606, 765, 821, 870, 959, 962, 1031, 1445]. residues [137, 597, 618, 833, 1012, 1119, 1151, 1152, 1331, 1520]. resolution [1635]. restricted [311, 378, 408, 453, 868, 1510, 1546]. Restricting [1605, 1710]. restriction [754]. restrictions [426, 1626]. result [9, 635, 903, 1642, 1700]. results [104, 475, 609, 949, 1055, 1162, 1163, 1381]. revêtements [283]. revisited [457, 671]. Rham [1481]. Richard [1476]. Richaud [1326, 1525]. Riemann [72, 324, 431, 625, 640, 641, 753, 900, 904, 957, 981, 1030, 1037, 1192, 1303, 1367, 1435, 1449, 1459, 1521, 1669, 1709]. Riemannscher [1435]. Riesel [1396]. right [1088]. right-angled [1088]. rigid [48, 934]. Ring [69, 703, 812, 930, 1316, 1370, 1398, 1410, 1544]. rings [79, 351, 642, 1330, 1472, 1540, 1588]. Riordan [1294]. Riyadh [1713]. Robbins’ [1343]. Rodriguez [479]. Rodriguez-Villegas [479]. Rogers [22, 130, 740, 1281, 1372, 1575, 1609]. Rohrlich [1618]. Romanoff [680]. root [194, 1102, 1608]. Roots [1, 33, 82, 83, 126, 315, 485, 507, 561, 589, 873, 1178, 1302, 1458, 1561]. Roskam [126]. Ross [763]. rounded [1423, 1662]. RSA [373]. Ru [1202]. Rubin [846]. rules [940].

S [1713]. Saito [1381]. Salem [1033, 1450]. same [581, 1672]. sans [1420]. Sárközy [364]. Satake [169]. satisfied [1015]. saturated [795]. Saudi [1713]. sauvage [729]. sauvages [1027]. scalars [754]. Schanuel [1587]. scheme [985]. schemes [1059, 1342, 1468, 1674]. Schinzel [1091]. Schläfli [519]. Schläfli-type [519]. Schmidt [1201, 1638, 1699]. Scholz [1103]. Schreier [563]. Schur [597]. Schweiger [162]. se [1427]. second [1469].

Selberg [253, 1066, 1099, 1192, 1212, 1401, 1709]. Self [429, 505, 778, 1543, 1608]. self-adjoint [505]. self-conformal [778]. Self-conjugate [1543]. Self-dual [429, 1608]. Selmer [12, 58, 195, 480, 500, 890, 1057, 1565, 1581]. Selmer-like [1057]. semi [620, 889, 1411, 1658]. semi-algebraic [1411]. Semi-local [889]. semi-locales [889]. semi-stable [620, 1658]. semigroup [948, 1463]. semigroups [1293]. semistable [353]. separability [1645]. Separable [759]. September [93, 238, 375, 555, 715, 853, 1020, 1222, 1433, 1630]. septic [496, 1312]. sequence [580, 693, 730, 1179, 1387]. Sequences [20, 67, 115, 196, 320, 325, 343, 391, 406, 450, 499, 587, 622, 664, 701, 872, 950, 979, 1054, 1071, 1085, 1094, 1138, 1141, 1183, 1215, 1219, 1237, 1258, 1395, 1424, 1492, 1499, 1506, 1512, 1520, 1539, 1596, 1597, 1694]. sequential [1669]. séquentielle [1669]. series [15, 48, 51, 91, 129, 135, 235, 270, 330, 368, 391, 410, 412, 459, 495, 535, 545, 562, 570, 576, 607–609, 639–641, 670, 677, 713, 725, 769, 780, 887, 914, 981, 1060, 1090, 1106, 1114, 1120, 1176, 1312, 1323, 1329, 1343, 1383, 1384, 1394, 1401, 1411, 1526, 1538, 1609, 1641, 1697, 1704, 1706]. Serre [288, 463, 1457]. set [10, 864, 1299, 1573]. Sets [155, 229, 230, 280, 545, 558, 598, 611, 748, 757, 832, 850, 921, 1087, 1089, 1121, 1156, 1191, 1229, 1239, 1411, 1446, 1511, 1539, 1553, 1694]. Seven [496]. Seven-modular [496]. seventh [45]. several [65, 381, 919]. sextenary [1580]. sextic [206, 423, 1076]. sextuple [1368]. SHA [993]. shadow [493]. Shafarevich [56, 186, 416, 797, 897, 1498, 1683]. Shahidi [1291]. sharp [323]. sheaves [434, 524, 769]. shifts [835]. Shimura [173, 244, 1132, 1176, 1276, 1362, 1430, 1432]. Shintani [304]. Shioda [756]. Short [10, 113, 288, 442, 844, 1322, 1548]. Shrinking [1279, 1504]. shuffle [1051]. side [1317]. sided [39]. sides [1225, 1273]. Siegel [70, 120, 169, 316, 607, 676, 684, 965, 1023, 1168, 1198, 1251, 1285, 1305, 1324, 1540, 1605, 1710]. Sierpiński [635, 1396]. sieve [49, 279, 911, 1014, 1280, 1509, 1596]. sign [1331]. similarity [1662]. Simple [272, 298, 900, 946, 1111, 1151, 1224, 1227, 1554]. simplices [1088]. simplified [1122]. simultanée [284]. Simultaneous [55, 151, 284, 334, 441, 474, 490, 499, 804, 823, 1513, 1543, 1627]. sine [1125, 1169]. Singer [497]. singular [103, 282, 960, 1013, 1086, 1260, 1313]. six [430]. sixth [1681]. size [224, 1309, 1611]. sizes [1163, 1573]. Skew [676]. Skew-holomorphic [676]. Skinner [158]. Slope [91, 1135]. Slopes [525, 815, 849]. Small [194, 229, 278, 494, 549, 719, 1091, 1102, 1420, 1503, 1539]. smallest [1528]. Smallness [1674]. smooth [321, 333, 1239]. sociable [1628]. solids [448, 1076]. Solomon [903]. Solution [558, 683, 782, 982]. Solution-free [558]. solutions [136, 381, 798, 884, 1160, 1200, 1214, 1255, 1299, 1661]. solvability [206]. solvable [553]. Solving [68, 1487, 1533]. Some [9, 37, 135, 218, 247, 250, 288, 347, 414, 425, 446, 475, 519, 522, 584, 586, 607, 642, 781, 797, 827, 895, 915, 918, 935, 980, 1032, 1039, 1053, 1055, 1069, 1086, 1090, 1162, 1163, 1203, 1215, 1238, 1261, 1399, 1402, 1427, 1432, 1437, 1492, 1494, 1578, 1583, 1586, 1587, 1614, 1639, 1708, 1711]. somme [861]. Sommes [340, 889].

Soundararajan [1133]. sous [498]. space [55, 63, 116, 124, 161, 304, 786, 822, 1251, 1427]. spaces [43, 114, 409, 414, 753, 824, 847, 1443, 1459, 1502, 1561, 1600, 1655]. spacing [324]. sparse [373]. Special [60, 693, 712, 742, 846, 926, 928–930, 967, 1169, 1198, 1563, 1579, 1659, 1665]. spéciales [1665]. specializability [1261]. Spectra [598, 649, 1065]. Spectral [1655]. spectrum [94, 464, 772, 1597]. Speiser [1325]. Spencer [1015, 1016]. sphere [1236]. spheres [1080]. Spherical [176, 822]. Spiegelung [75]. Spiegelungssatz [292, 675]. spine [1282]. spinor [564]. split [1686]. splitting [1227]. Square [9, 588, 594, 701, 1029, 1167, 1187, 1280, 1608, 1655]. Square-free [588]. square-root [1608]. Squared [149]. squarefree [515, 775]. Squares [88, 142, 210, 295, 366, 400, 592, 619, 957, 1170, 1183, 1209, 1309, 1322, 1431, 1447, 1549, 1572]. Stability [1356]. Stabilized [11]. stable [620, 1658]. Stamp [242]. standard [523]. Stanton [1442]. star [1437]. Stark [329, 593, 697, 846, 1169, 1542]. state [1027]. stationary [386]. stationary-phase [386]. statistical [256, 1142]. Steinitz [1334]. step [343]. Stepanov [925]. Stephen [83]. steps [669]. Stickelberger [38, 1241, 1454]. still [1279]. Stirling [247]. stratification [1140]. stratum [1140]. strict [536, 1472]. Strong [499, 565, 1071, 1258, 1344, 1406, 1619]. strongly [1082]. Structural [1029]. Structure [163, 211, 273, 524, 563, 649, 710, 812, 1087, 1218, 1481, 1565, 1648]. structures [47, 756, 1316]. study [42, 1069, 1262]. Sturmian [268]. Subconvexity [1066]. subextensions [145, 571]. subfields [393]. subgroup [181, 771]. subgroups [134, 241, 293, 336, 662, 860, 882, 1000, 1066, 1071, 1099, 1395, 1399, 1418]. sublattices [1423, 1662]. submodules [314, 500]. subrings [365]. subset [229, 230, 618, 1162, 1415]. subset-sums [618]. subsets [426, 693]. subspace [1201]. subspaces [847, 1473, 1627]. subsum [1539]. Subvariety [612, 815]. subword [1126]. Successive [849, 1305]. sufficient [794]. suites [1094]. Sum [3, 44, 84, 142, 182, 217, 220, 266, 300, 610, 648, 861, 946, 947, 1336, 1387, 1546, 1642, 1670, 1677, 1678]. sum-of-digits-function [220]. Summation [808, 1120]. Sums [35, 39, 45, 88, 108, 113, 178, 225, 229, 239, 249, 274, 278, 294, 305, 307, 320, 340, 351, 366, 376, 386, 392, 396, 426, 430, 433, 449, 460, 477, 495, 526, 546, 592, 618, 655, 661, 679, 721, 857, 864, 889, 919, 920, 925, 942, 944, 960, 963, 964, 1002, 1024, 1045, 1104, 1121, 1134, 1162, 1170, 1173, 1182, 1188, 1216, 1226, 1228, 1242, 1289, 1304, 1309, 1318, 1322, 1323, 1338, 1348, 1388, 1400, 1429, 1449, 1455, 1500, 1548, 1549, 1595, 1602, 1624, 1634, 1684]. sums-of-squares [1309]. sumset [1163]. sumsets [63, 87, 378, 549, 868, 951, 1229, 1510]. supercongruence [479]. supercongruences [1617]. Supercuspidal [518]. Supersequences [1597]. Supersingular [47, 148, 666, 843, 1068, 1307, 1704]. supersingulier [1068]. Superspecial [1465, 1594]. Support [510, 552, 931, 998, 1620]. surds [702]. surface [399, 1013, 1118, 1362]. surfaces [6, 98, 346, 709, 756, 774, 816, 871, 1080, 1132, 1475]. surjectivity [689]. Swan [1395]. Swinnerton [353, 837, 1426, 1581]. Swinnerton-Dyer [353, 837, 1426, 1581]. Sylow [181]. Sylvester [581]. symbol [212, 285, 383, 1445]. symbols [57, 64, 156, 458, 464, 1320]. Symmetric [568, 721, 1167, 1200, 1565, 1654, 1673]. symmetric-square [1167]. symmetries [78]. Symmetry [1261]. symplectic [1213]. system [1528]. Systèmes [1197]. systems [209, 214, 255, 559, 561, 589, 695, 742, 941, 1006, 1197, 1214, 1467, 1488, 1501, 1507]. Szpiro [875]. Szüsz [535].

Table [437, 852]. taking [451]. Tamagawa [363, 1056, 1507]. Tame [208, 397, 466, 712, 1342, 1558]. Tao [1510]. Tarry [1137]. Tasoev [740]. Tate [6, 56, 186, 380, 416, 448, 482, 492, 709, 720, 797, 826, 897, 1132, 1194, 1233, 1308, 1362, 1498, 1527, 1621]. technique [343]. techniques [551]. Teichmüller [359]. tensor [1386]. term [646, 701, 881, 1002, 1134]. terms [9, 37, 208, 347, 384, 397, 872, 1154, 1367, 1691]. ternary [257, 508, 766, 959, 1657]. test [770]. tests [1425]. tetrahedra [323, 793, 1061, 1286, 1559, 1624]. tetrahedral [462]. Texan [921]. th [108, 392, 934, 1044, 1445, 1595]. their [42, 243, 275, 365, 458, 460, 568, 834, 856, 1304, 1318]. theorem [58, 60, 90, 122, 124, 158, 164, 195, 234, 270, 288, 377, 400, 404, 511, 535, 619, 663, 671, 700, 706, 752, 765, 773, 815, 886, 1009, 1036, 1037, 1079, 1081, 1091, 1098, 1103, 1112, 1168, 1201, 1238, 1241, 1249, 1260, 1267, 1301, 1344, 1346, 1355, 1373, 1408, 1476, 1484, 1553, 1619, 1670, 1676]. théorème [1168]. theorems [50, 408, 537, 572, 1513]. theoretic [256, 347]. théorie [729, 1262]. Theory [69, 157, 159, 199, 243, 331, 431, 453, 504, 537, 589, 628, 629, 633, 729, 753, 760, 767, 807, 953, 998, 1017, 1062, 1262, 1296, 1391, 1444, 1459, 1533, 1597, 1606, 1664, 1688, 1709–1716]. There [1368]. Theta [91, 119, 330, 357, 495, 781, 847, 1058, 1105, 1312, 1477, 1583]. theta-functions [495]. Thin [150, 443]. three [18, 138, 142, 411, 646, 857, 933, 1002, 1069, 1154, 1253, 1263, 1595]. Three-term [1002]. three-variable [1069]. threefolds [48, 76, 354]. Thue [8, 136, 422, 691, 1001, 1043, 1505, 1715]. tiling [598]. top [1536]. topological [1087]. tori [315, 1514, 1620]. torique [189]. Tornheim [920, 1032]. Toroidal [1586]. torseurs [284, 498]. Torsion [56, 380, 405, 491, 755, 860, 882, 1190, 1364, 1390, 1505, 1625, 1658]. torsors [284, 498]. torus [961]. total [152]. totally [365, 1187, 1307, 1565, 1578]. tower [669, 1122, 1683]. Towers [179, 522, 731, 980, 1203, 1475, 1494, 1686]. Trace [193, 613]. Traces [202, 521, 960, 1086, 1313]. Transcendance [120, 421, 916]. Transcendence [5, 120, 268, 270, 421, 582, 916, 1114, 1199, 1639, 1711]. Transcendental [1211]. transfer [1490]. transform [744, 1155]. transformation [495, 1179]. Transition [62]. translates [598]. treatment [243]. trees [157]. triangle [1564]. Triangles [411, 1004, 1118, 1225]. triangular [1549]. trigonometric [266, 648, 1069]. trilogarithm [1147]. Trinomial [656]. trinomials [246]. triples [275]. trivial [1498]. truncated [479, 914]. Tschakaloff [609, 1438]. Tunnell [1373]. tuples [224, 1025, 1038]. Turán [575, 665, 911, 1440]. twelfth [725]. twin [1014]. Twisted [57, 197, 329, 780, 1005, 1132, 1208, 1278, 1432]. twists [398, 621, 977, 1130]. Two [6, 44, 139, 150, 161, 170, 202, 280, 392, 446, 617, 694, 709, 785, 905, 912, 933, 1076, 1134, 1137, 1160, 1164, 1225, 1235, 1348, 1409, 1478, 1525, 1585, 1620, 1703]. two-dimensional [694, 1137]. two-term [1134]. two-variable [139, 446]. Type [16, 71, 79, 138, 206, 291, 324, 473, 519, 570, 723, 816, 989, 1037, 1084, 1326, 1347, 1401, 1432, 1501, 1525, 1557, 1617, 1623, 1642, 1659]. types [272, 598].

unification [243]. unified [1554]. Uniform [220, 302, 732]. uniformly [67, 580, 1302]. Unimodular [271, 355, 493, 869]. uniqueness [124, 180]. Unit [1, 68, 83, 269, 396, 988, 1160, 1259, 1447]. unitary [379, 1213, 1356]. Unités [121, 889]. units [29, 121, 249, 287, 530, 546, 550, 561, 589, 651, 689, 742, 889, 968, 1052, 1124, 1169, 1187, 1339, 1354, 1483, 1646]. unity [485, 873, 1458]. Universal [102, 345, 758, 1563]. universality [1110, 1217]. Universally [708]. University [1713]. univoque [1087]. unknowns [711]. unobstructed [478, 775]. unramified [201, 262, 429, 476, 645, 805, 874, 1041, 1062, 1150, 1157, 1296]. upon [1084]. Upper [116, 267, 323, 406, 608, 679, 1070, 1088, 1089, 1188, 1415, 1701]. Urs [2, 25, 180]. Ursim [2, 25]. Using [380, 890, 925].

V [1647]. Valeurs [120, 1195, 1379, 1665]. Vallée [1647]. valuations [1514]. value [178, 217, 459, 483, 511, 594, 752, 848, 947, 1037, 1134, 1267, 1289, 1429, 1516]. valued [20, 533, 1285, 1527, 1694]. Values [11, 62, 120, 137, 282, 321, 371, 451, 588, 608, 609, 634, 650, 677, 724, 725, 888, 899, 1044, 1051, 1101, 1125, 1127, 1130, 1169, 1176, 1195, 1198, 1211, 1216, 1260, 1345, 1379, 1384, 1409, 1421, 1437, 1438, 1537, 1546, 1579, 1585, 1632, 1654, 1665, 1678]. Vanishing [36, 722, 797]. variable [139, 446, 915, 1069]. variables [65, 161, 381, 413, 919, 1585, 1659]. variant [1117, 1510, 1618]. variants [1620]. variation [897]. variations [1022, 1112]. variété [189]. variétés [207]. Varieties [1, 47, 71, 83, 148, 207, 298, 301, 306, 334, 345, 360, 416, 492, 500, 509, 537, 552, 828, 890, 931, 998, 1042, 1246, 1432, 1465, 1514, 1566, 1594, 1620]. variety [612]. various [868]. vector [55, 131, 409, 849, 1285, 1443, 1502]. vectoriels [409]. vectors [1357]. Verifying [697]. version [377, 683, 773, 1241, 1532]. vertex [61]. via [243, 432, 532, 1457, 1654]. viewpoint [474]. Villegas [479]. Vinogradov [266, 648, 1642]. Visibility [416, 1652]. Visible [1236, 1448]. Visualizing [785]. Vladut [669]. Vojta [369, 1266]. Voloch [1194]. Volume [25, 27, 52, 73, 83, 84, 92, 117, 126, 143, 166, 190, 215, 237, 263, 264, 289, 312, 326, 349, 374, 402, 435, 436, 469, 501, 527, 554, 577, 600, 657, 685, 714, 736, 761, 788, 809, 829, 851, 877, 878, 907, 908, 936, 937, 970, 971, 994, 995, 1018, 1019, 1047, 1048, 1072, 1073, 1107, 1143, 1184, 1221, 1243, 1269]. Volumes [547].

W [83]. Waring [707, 723, 1210, 1441, 1472]. Warning [1084]. Washington [1249]. wave [690, 1655, 1712]. ways [786]. WC [865]. WC-groups [865].

Weak [562, 1006]. weakly [273]. Weierstrass [333, 741]. Weight [108, 481, 562, 628, 676, 690, 932, 1060, 1097, 1329, 1429, 1477, 1629, 1648, 1712]. Weight-dependent [1097]. Weighted [9, 307, 613, 911, 1678]. Weights [1235, 1457]. Weil [70, 294, 398, 433, 467, 520, 785, 1079, 1091, 1198, 1390, 1410, 1524, 1560, 1626]. Well [825, 1423, 1662]. Well-behaved [825]. well-rounded [1423, 1662]. Wendt [883]. Weng [1220, 1521]. whether [730]. which [669, 1525]. while [1279]. Whittaker [1374]. whole [146, 748]. whose [18, 46, 561, 589, 701, 861, 1428, 1528]. Wieferich [41]. wild [181, 603, 712, 729, 1026, 1027]. Wildly [1011]. Wiles’ [1538]. Wirtinger [324, 431]. without [1420]. Witt [1208, 1357]. Wolstenholme [1112, 1301, 1670]. Women [1713]. Wong [1202]. Woods [1554, 1555]. writing [46]. Wronskians [894, 1060].

Yau [48]. Yokota [84]. Yu [182].

zêta [89, 346]. Zagier [677, 944]. Zahlkörpern [68]. Zariski [928, 1277]. Zariski-density [1277]. Zeilberger [1637]. Zero [44, 272, 280, 444, 739, 752, 791, 874, 1062, 1336, 1701]. zero-free [791]. zero-sum [44]. Zeros [113, 171, 324, 431, 459, 719, 753, 900, 904, 946, 956, 1177, 1303, 1377, 1449, 1459, 1656, 1677, 1704]. Zeta [11, 89, 137, 161, 219, 245, 253, 304, 324, 329, 346, 431, 459, 544, 625, 640, 641, 677, 724, 753, 899, 900, 904, 956, 957, 986, 1029, 1037, 1051, 1099, 1110, 1172, 1217, 1220, 1278, 1303, 1324, 1367, 1403, 1421, 1437, 1449, 1459, 1521, 1522, 1537, 1546, 1566, 1610, 1632, 1633, 1675, 1678]. zeta-function [324, 431, 753, 900, 957, 1029, 1459]. zetafunctions [329, 459, 544, 1278]. zeta-regularized [986]. zeta-star [1437]. Zhang [1337]. Ziv [1238]. Zonal [379].

References

DiPippo:1998:RPA

[1] Stephen A. DiPippo and Everett W. Howe. Real polynomials with all roots on the unit circle and abelian varieties over finite fields. Journal of Number Theory, 73(2):426–450, December 1998. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

      sciencedirect.com/science/article/pii/S0022314X98923022. See

corrigendum [83].

Escassut:1999:UUN

[2] Alain Escassut, Labib Haddad, and Robert Vidal. Urs, Ursim, and Non–Urs for p-adic Functions and Polynomials. Journal of Number Theory, 75(1):133–144, March 1999. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

      science/article/pii/S0022314X98923241. See corrigendum [25].

Yokota:1999:LIE

[3] Hisashi Yokota. The largest integer expressible as a sum of reciprocal of integers. Journal of Number Theory, 76(2):206–216, June 1999. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X98923599. See erratum [84].

Anonymous:2000:PJa

[4] Anonymous. Pages 1–167 (January 2000). Journal of Number Theory, 80(1):??, January 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Fresnel:2000:ATP

[5] Jean Fresnel, Michel Koskas, and Bernard de Mathan. Automata and transcendence in positive characteristic. Journal of Number Theory, 80(1):1–24, January 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X99924447.

Murty:2000:TCP

[6] V. Kumar Murty and Dipendra Prasad. Tate cycles on a product of two Hilbert modular surfaces. Journal of Number Theory, 80(1):25– 43, January 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X99924460.

Chen:2000:IDI

[7] W. W. L. Chen. On irregularities of distribution IV. Journal of Number Theory, 80(1):44–59, January 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X99924423.

Wakabayashi:2000:FQT

[8] Isao Wakabayashi. On a family of quartic Thue inequalities, II. Journal of Number Theory, 80(1):60–88, January 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X99924393.

Lau:2000:WMS

[9] Yuk-Kam Lau. On a weighted mean square result of the error terms of some arithmetical functions. Journal of Number Theory, 80(1):89– 108, January 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X99924368.

Bazzanella:2000:ESN

[10] D. Bazzanella and A. Perelli. The exceptional set for the number of primes in short intervals. Journal of Number Theory, 80(1):109–124, January 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X99924290.

Brumley:2000:SVG

[11] Farrell Brumley. Stabilized values of the generalized Goss zeta function. Journal of Number Theory, 80(1):125–134, January 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X99924319.

Guo:2000:CBS

[12] Li Guo. Congruences between Selmer groups. Journal of Number Theory, 80(1):135–153, January 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X99924459.

Yin:2000:DGF

[13] Linsheng Yin. Distributions on a Global Field. Journal of Number Theory, 80(1):154–167, January 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X9992440X.

Anonymous:2000:PF

[14] Anonymous. Pages 169–320 (February 2000). Journal of Number Theory, 80(2):??, February 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Mkaouar:2000:DFC

[15] Mohamed Mkaouar. Sur le développement en fraction continue des séries formelles quadratiques sur F_{2}(X). (French) [On the continuous-fraction development of the quadratic formal series on F_{2}(X)]. Journal of Number Theory, 80(2):169–173, February 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X99924228.

Sitaraman:2000:FTD

[16] Sankar Sitaraman. On a Fermat-type diophantine equation. Journal of Number Theory, 80(2):174–186, February 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X99924381.

Bremner:2000:IPA

[17] A. Bremner, J. H. Silverman, and N. Tzanakis. Integral points in arithmetic progression on y^{2} = x(x^{2} − n^{2}). Journal of Number Theory, 80(2):187–208, February 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X99924307.

Kishi:2000:PQF

[18] Yasuhiro Kishi and Katsuya Miyake. Parametrization of the quadratic fields whose class numbers are divisible by three. Journal of Number Theory, 80(2):209–217, February 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X99924551.

Komornik:2000:APP

[19] Vilmos Komornik, Paola Loreti, and Marco Pedicini. An approximation property of Pisot numbers. Journal of Number Theory, 80 (2):218–237, February 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X99924563.

Boulanger:2000:ABC

[20] Jacques Boulanger and Jean-Luc Chabert. Asymptotic behavior of characteristic sequences of integer-valued polynomials. Journal of Number Theory, 80(2):238–259, February 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X99924320.

Amoroso:2000:LBH

[21] Francesco Amoroso and Roberto Dvornicich. A lower bound for the height in abelian extensions. Journal of Number Theory, 80(2):260– 272, February 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X99924514.

Andrews:2000:EER

[22] George E. Andrews, Arnold Knopfmacher, and John Knopfmacher. Engel expansions and the Rogers–Ramanujan identities. Journal of Number Theory, 80(2):273–290, February 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X99924538.

Anderson:2000:EAF

[23] G. W. Anderson. An elementary approach to L-functions modp. Journal of Number Theory, 80(2):291–303, February 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X99924526.

Sodaigui:2000:RCQ

[24] Bouchaïb Sodaïgui. Realizable classes of quaternion extensions of degree 4ℓ. Journal of Number Theory, 80(2):304–315, February 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X99924484.

Escassut:2000:CUU

[25] Alain Escassut, Labib Haddad, and Robert Vidal. Corrigendum: Urs, Ursim, and Non-Urs for p-adic Functions and Polynomials: Volume 75, Number 1 (1999), pages 133–144. Journal of Number Theory, 80(2):316–317, February 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X99924666. See [2].

Jaulent:2000:PEC

[26] J.-F. Jaulent and O. Sauzet. Pro-ℓ-extensions de corps de nombres lrationnels. (French) [pro-ℓ-body extensions of l-rational numbers]. Journal of Number Theory, 80(2):318–319, February 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X99924356.

Anonymous:2000:AIVa

[27] Anonymous. Author index for volume 80. Journal of Number Theory, 80(2):320, February 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X99925167.

Anonymous:2000:PMa

[28] Anonymous. Pages 1–204 (March 2000). Journal of Number Theory, 81 (1):??, March 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Conrad:2000:CBG

[29] Marc Conrad. Construction of bases for the group of cyclotomic units. Journal of Number Theory, 81(1):1–15, March 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X99924435.

Holden:2000:FMC

[30] Joshua Brandon Holden. On the Fontaine–Mazur Conjecture for number fields and an analogue for function fields. Journal of Number Theory, 81(1):16–47, March 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X99924344.

Ribenboim:2000:AC

[31] Paulo Ribenboim. ABC candies. Journal of Number Theory, 81(1): 48–60, March 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X99900008.

Hwang:2000:DNF

[32] Hsien-Kuei Hwang. Distribution of the number of factors in random ordered factorizations of integers. Journal of Number Theory, 81(1): 61–92, March 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X99924678.

Roskam:2000:QAA

[33] Hans Roskam. A quadratic analogue of Artin’s Conjecture on primitive roots. Journal of Number Theory, 81(1):93–109, March 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X99924708. See erratum [126].

Ru:2000:DAA

[34] Min Ru and Julie Tzu-Yueh Wang. Diophantine approximation with algebraic points of bounded degree. Journal of Number Theory, 81(1):

110–119, March 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X99924629.

Cochrane:2000:PHC

[35] Todd Cochrane, David Garth, and Zhiyong Zheng. On a problem of H. Cohn for character sums. Journal of Number Theory, 81(1):120– 129, March 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X99924575.

Michel:2000:NVH

[36] Philippe Michel and Jeffrey VanderKam. Non-vanishing of high derivatives of Dirichlet L-functions at the central point. Journal of Number Theory, 81(1):130–148, March 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X99924605.

Languasco:2000:SRE

[37] A. Languasco. Some refinements of error terms estimates for certain additive problems with primes. Journal of Number Theory, 81(1):149– 161, March 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X9992468X.

Yin:2000:SIR

[38] Linsheng Yin. Stickelberger ideals and relative class numbers in function fields. Journal of Number Theory, 81(1):162–169, March 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X99924721.

Pinner:2000:OSB

[39] Christopher G. Pinner. On the one-sided boundedness of sums of fractional parts ({nα + γ} − 12). Journal of Number Theory, 81(1):170– 204, March 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X99924691.

Anonymous:2000:PAa

[40] Anonymous. Pages 205–374 (April 2000). Journal of Number Theory, 81(2):??, April 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Voloch:2000:EWP

[41] José Felipe Voloch. Elliptic Wieferich primes. Journal of Number Theory, 81(2):205–209, April 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X9992471X.

Buchholz:2000:WNM

[42] Ralph H. Buchholz and James A. MacDougall. When Newton met Diophantus: a study of rational-derived polynomials and their extension to quadratic fields. Journal of Number Theory, 81(2):210– 233, April 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X99924733.

Philippon:2000:AAP

[43] Patrice Philippon. Approximations algébriques des points dans les espaces projectifs I. (French) [Algebraic approximations of points in projective spaces I]. Journal of Number Theory, 81(2):234–253, April 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X99924617.

Gao:2000:TZS

[44] W. D. Gao. Two zero-sum problems and multiple properties. Journal of Number Theory, 81(2):254–265, April 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X99924599.

Choudhry:2000:SSP

[45] Ajai Choudhry. On sums of seventh powers. Journal of Number Theory, 81(2):266–269, April 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X99924654.

Dartyge:2000:NPP

[46] Cécile Dartyge and Christian Mauduit. Nombres presque premiers dont l’écriture en base r ne comporte pas certains chiffres. (French) [Nearly prime numbers whose base-r writing does not contain certain numbers]. Journal of Number Theory, 81(2):270–291, April 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X99924587.

Zhu:2000:GSE

[47] Hui June Zhu. Group structures of elementary supersingular abelian varieties over finite fields. Journal of Number Theory, 81(2):292– 309, April 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X99924630.

Verrill:2000:SCR

[48] H. A. Verrill. The L-series of certain rigid Calabi–Yau threefolds. Journal of Number Theory, 81(2):310–334, April 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X99924496.

Ford:2000:BHS

[49] Kevin Ford and H. Halberstam. The Brun–Hooley sieve. Journal of Number Theory, 81(2):335–350, April 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X99924794.

Vaserstein:2000:QT

[50] L. N. Vaserstein. Quantum (abc)-theorems. Journal of Number Theory, 81(2):351–358, April 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X9992485X.

Allouche:2000:ADS

[51] J.-P Allouche, M. Mendès France, and J. Peyrière. Automatic Dirichlet series. Journal of Number Theory, 81(2):359–373, April 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X99924873.

Anonymous:2000:AIVb

[52] Anonymous. Author index for volume 81. Journal of Number Theory, 81 (2):374, April 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X00925368.

Anonymous:2000:PMb

[53] Anonymous. Pages 1–163 (May 2000). Journal of Number Theory, 82(1): ??, May 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Chen:2000:PPF

[54] Yong-Gao Chen and Yao-Chen Zhu. On the prime power factorization of n!. Journal of Number Theory, 82(1):1–11, May 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X99924770.

Burger:2000:SDA

[55] Edward B. Burger. On simultaneous Diophantine approximation in the vector space Q + Qα. Journal of Number Theory, 82(1):12– 24, May 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X99924885.

Beaver:2000:TST

[56] Cheryl DeLorme Beaver. 5-torsion in the Shafarevich–Tate group of a family of elliptic curves. Journal of Number Theory, 82(1):25– 46, May 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X99924939.

Fukuhara:2000:TGD

[57] Shinji Fukuhara. Twisted generalized Dedekind symbols. Journal of Number Theory, 82(1):47–68, May 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X99924940.

Ochiai:2000:CTB

[58] Tadashi Ochiai. Control theorem for Bloch–Kato’s Selmer groups of p-adic representations. Journal of Number Theory, 82(1):69–90, May 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X99924836.

Frankenhuysen:2000:LBA

[59] Machiel Van Frankenhuysen. A lower bound in the abc conjecture. Journal of Number Theory, 82(1):91–95, May 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X99924848.

Cavachi:2000:SCH

[60] Marius Cavachi. On a special case of Hilbert’s irreducibility theorem. Journal of Number Theory, 82(1):96–99, May 2000. CODEN JNUTA9.

ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X99924769.

Kitazume:2000:BIV

[61] Masaaki Kitazume, Masahiko Miyamoto, and Hiromichi Yamada. Borwein identity and vertex operator algebras. Journal of Number Theory, 82(1):100–108, May 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X99924824.

Conrey:2000:TMV

[62] J. B. Conrey, D. W. Farmer, and K. Soundararajan. Transition mean values of real characters. Journal of Number Theory, 82(1):109– 120, May 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X99924927.

Fainleib:2000:SES

[63] A. S. Fainleib. On sumsets in Euclidean space. Journal of Number Theory, 82(1):121–133, May 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X99924782.

Gunnells:2000:FMM

[64] Paul E. Gunnells. Finiteness of minimal modular symbols for SL_{n}. Journal of Number Theory, 82(1):134–139, May 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X99924812.

Choie:2000:DOJ

[65] YoungJu Choie and Haesuk Kim. Differential operators on Jacobi forms of several variables. Journal of Number Theory, 82(1):140– 163, May 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X99924915.

Anonymous:2000:PJb

[66] Anonymous. Pages 165–323 (June 2000). Journal of Number Theory, 82 (2):??, June 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Konyagin:2000:UDD

[67] S. V. Konyagin, I. Z. Ruzsa, and W. Schlag. On uniformly distributed dilates of finite integer sequences. Journal of Number Theory, 82(2):165–187, June 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X99924204.

Wildanger:2000:LEI

[68] K. Wildanger. Über das Lösen von Einheiten- und Indexformgleichungen in algebraischen Zahlkörpern. (German) [On solving unit and index form equations in algebraic number fields]. Journal of Number Theory, 82(2):188–224, June 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X99924149.

Arlettaz:2000:ACE

[69] Dominique Arlettaz, Grzegorz Banaszak, and Wojciech Gajda. On 2adic cyclotomic elements in K-theory and étale cohomology of the ring of integers. Journal of Number Theory, 82(2):225–255, June 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X99924757.

Ginzburg:2000:SWI

[70] David Ginzburg and Dihua Jiang. A Siegel–Weil identity for G_{2} and poles of L-functions. Journal of Number Theory, 82(2):256– 287, June 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X99924800.

Murabayashi:2000:RMA

[71] Naoki Murabayashi. A remark on the modularity of abelian varieties of GL_{2}-type over Q. Journal of Number Theory, 82(2):288– 298, June 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X99925039.

Goss:2000:RHC

[72] David Goss. A Riemann Hypothesis for characteristic p L-functions. Journal of Number Theory, 82(2):299–322, June 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X99924952.

Anonymous:2000:AIVc

[73] Anonymous. Author index for volume 82. Journal of Number Theory, 82(2):323, June 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X00925642.

Anonymous:2000:PJc

[74] Anonymous. Pages 1–184 (July 2000). Journal of Number Theory, 83(1): ??, July 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Kishi:2000:CAS

[75] Yasuhiro Kishi. A constructive approach to Spiegelung relations between 3-ranks of absolute ideal class groups and congruent ones modulo (3)^{2} in quadratic fields. Journal of Number Theory, 83(1):1– 49, July 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X99924988.

Grundman:2000:NHM

[76] H. G. Grundman. Nonrational Hilbert modular threefolds. Journal of Number Theory, 83(1):50–58, July 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X99924964.

Liu:2000:CIR

[77] Zhi-Guo Liu. On certain identities of Ramanujan. Journal of Number Theory, 83(1):59–75, July 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X00925137.

Kim:2000:GIC

[78] Myung-Hwan Kim and Byeong-Kweon Oh. Generation of isometries of certain Z-lattices by symmetries. Journal of Number Theory, 83(1): 76–90, July 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X99924976.

Talamanca:2000:GBT

[79] Valerio Talamanca. A Gelfand–Beurling type formula for heights on endomorphism rings. Journal of Number Theory, 83(1):91–105, July 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X99925064.

Sato:2000:LDR

[80] Fumihiro Sato and Yumiko Hironaka. Local densities of representations of quadratic forms over p-adic integers (the non-dyadic case). Journal of Number Theory, 83(1):106–136, July 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X99925052.

Kochubei:2000:DEL

[81] Anatoly N. Kochubei. Differential equations for F_{q}-linear functions. Journal of Number Theory, 83(1):137–154, July 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X9992499X.

Moree:2000:AEH

[82] Pieter Moree. Asymptotically exact heuristics for (near) primitive roots. Journal of Number Theory, 83(1):155–181, July 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X99925027.

Anonymous:2000:CVN

[83] Anonymous. Corrigendum: Volume 73, number 2 (1998), in the article “Real Polynomials with All Roots on the Unit Circle and Abelian Varieties over Finite Fields,” by Stephen A. DiPippo and Everett W. Howe, pages 426–450. Journal of Number Theory, 83(1): 182, July 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X00925654. See [1].

Anonymous:2000:EVNa

[84] Anonymous. Erratum: Volume 76, number 2 (1999), in the article “The Largest Integer Expressible as a Sum of Reciprocal of Integers,” by Hisashi Yokota, pages 206–216. Journal of Number Theory, 83(1):183–184, July 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X00925666. See [3].

Anonymous:2000:PAb

[85] Anonymous. Pages 185–297 (August 2000). Journal of Number Theory, 83(2):??, August 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Cao:2000:NPC

[86] Zhenfu Cao, Shanzhi Mu, and Xiaolei Dong. A new proof of a conjecture of Antoniadis. Journal of Number Theory, 83(2):185–193, August 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X99924897.

Elekes:2000:CS

[87] György Elekes, Melvyn B. Nathanson, and Imre Z. Ruzsa. Convexity and Sumsets. Journal of Number Theory, 83(2):194–201, August 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X99923867.

Liu:2000:REI

[88] Jianya Liu and Ming-Chit Liu. Representation of even integers as sums of squares of primes and powers of 2. Journal of Number Theory, 83(2):202–225, August 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X99925003.

Cassaigne:2000:HHF

[89] Julien Cassaigne and Vincent Maillot. Hauteur des hypersurfaces et fonctions Zêta d’Igusa. (French) [Height of hypersurfaces and Zeta functions of Igusa]. Journal of Number Theory, 83(2):226–255, August 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X99924903.

Pasles:2000:HCT

[90] Paul C. Pasles. A Hecke correspondence theorem for nonanalytic automorphic integrals. Journal of Number Theory, 83(2):256–281, August 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X99925040.

Manni:2000:SCF

[91] R. Salvati Manni. Slope of cusp forms and theta series. Journal of Number Theory, 83(2):282–296, August 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X99925015.

Anonymous:2000:AIVd

[92] Anonymous. Author index for volume 83. Journal of Number Theory, 83(2):297, August 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X00925733.

Anonymous:2000:PS

[93] Anonymous. Pages 1–183 (September 2000). Journal of Number Theory, 84(1):??, September 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Gupta:2000:CID

[94] Soma Gupta. Continuous interval in the dispersion spectrum. Journal of Number Theory, 84(1):1–14, September 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X99924745.

Bugeaud:2000:APN

[95] Yann Bugeaud. Approximation par des nombres algébriques. (French) [Approximation by algebraic numbers]. Journal of Number Theory, 84(1):15–33, September 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X00925174.

Buium:2000:CAF

[96] Alexandru Buium. Continuous π-adic functions and π-derivations. Journal of Number Theory, 84(1):34–39, September 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X00925290.

Choi:2000:CEP

[97] Kwok-Kwong Choi and Man-Keung Siu. Counter-examples to a problem of Cohn on classifying characters. Journal of Number Theory, 84(1):40–48, September 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X00925319.

McKinnon:2000:CRP

[98] David McKinnon. Counting rational points on K3 surfaces. Journal of Number Theory, 84(1):49–62, September 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X00925332.

Croot:2000:BEF

[99] Ernest S. Croot, David E. Dobbs, John B. Friedlander, Andrew J. Hetzel, and Francesco Pappalardi. Binary Egyptian fractions. Journal of Number Theory, 84(1):63–79, September 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X00925083.

Matsuno:2000:AKF

[100] Kazuo Matsuno. An analogue of Kida’s formula for the p-adic L-functions of modular elliptic curves. Journal of Number Theory, 84(1):80–92, September 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X00925101.

Li:2000:ACA

[101] Shuguang Li. Artin’s Conjecture on average for composite moduli. Journal of Number Theory, 84(1):93–118, September 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X00925125.

Adelberg:2000:UHO

[102] Arnold Adelberg. Universal higher order Bernoulli numbers and Kummer and related congruences. Journal of Number Theory, 84(1):119–135, September 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X00925265.

Castro:2000:FSC

[103] Francis N. Castro and Carlos J. Moreno. L-functions of singular curves over finite fields. Journal of Number Theory, 84(1):136–155, September 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X00925307.

Moon:2000:FRC

[104] Hyunsuk Moon. Finiteness results on certain modp Galois representations. Journal of Number Theory, 84(1):156–165, September 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X00925344.

Semirat:2000:CNO

[105] Stéphan Sémirat. Class number one problem for imaginary function fields: the cyclic prime power case. Journal of Number Theory, 84 (1):166–183, September 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X00925356.

Anonymous:2000:PO

[106] Anonymous. Pages 185–361 (October 2000). Journal of Number Theory, 84(2):??, October 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Peres:2000:APC

[107] Yuval Peres and Boris Solomyak. Approximation by polynomials with coefficients ±1. Journal of Number Theory, 84(2):185–198, October 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X00925149.

Wenpeng:2000:PMD

[108] Zhang Wenpeng, Yi Yuan, and He Xiali. On the 2k-th power mean of Dirichlet L-functions with the weight of general Kloosterman sums. Journal of Number Theory, 84(2):199–213, October 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X00925150.

Reversat:2000:MFC

[109] Marc Reversat. On modular forms of characteristic p > 0. Journal of Number Theory, 84(2):214–229, October 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X00925253.

Conrad:2000:DP

[110] Keith Conrad. The digit principle. Journal of Number Theory, 84(2):230– 257, October 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X00925071.

Jeong:2000:CCD

[111] Sangtae Jeong. A comparison of the Carlitz and digit derivatives bases in function field arithmetic. Journal of Number Theory, 84(2):258– 275, October 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X00925277.

Jeong:2000:CLE

[112] Sangtae Jeong. Continuous linear endomorphisms and difference equations over the completion of F_{q}[T]. Journal of Number Theory, 84(2):276– 291, October 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X00925320.

Languasco:2000:PCZ

[113] Alessandro Languasco and Alberto Perelli. Pair correlation of zeros, primes in short intervals and exponential sums over primes. Journal of Number Theory, 84(2):292–304, October 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X00925113.

Erne:2000:RHP

[114] Reinie Erné. Reducibility modp of hypersurfaces in projective spaces — an application of arithmetic Bézout. Journal of Number Theory, 84(2):305–316, October 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X00925204.

Gittenberger:2000:ANA

[115] Bernhard Gittenberger and Jörg M. Thuswaldner. Asymptotic normality of b-additive functions on polynomial sequences in the Gaussian number field. Journal of Number Theory, 84(2):317–341, October 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X00925228.

Martinez:2000:FUH

[116] María G. Martínez. The finite upper half space and related hypergraphs. Journal of Number Theory, 84(2):342–360, October 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X00925241.

Anonymous:2000:AIVe

[117] Anonymous. Author index for volume 84. Journal of Number Theory, 84(2):361, October 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X00925988.

Anonymous:2000:PN

[118] Anonymous. Pages 1–108 (November 2000). Journal of Number Theory, 85(1):??, November 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Manderscheid:2000:QBC

[119] David Manderscheid. Quadratic base change for p-adic SL(2) as a theta correspondence III: pairing. Journal of Number Theory, 85(1):1–17, November 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X00925095.

Derome:2000:TVF

[120] Geoffroy Derome. Transcendance des valeurs des fonctions automorphes de Siegel. (French) [Transcendence of values of the Siegel automorphic functions]. Journal of Number Theory, 85(1):18–34, November 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X00925186.

Dubois:2000:UGC

[121] I. Dubois. S-unités et S-groupe de classes d’un corps de nombres cyclique de degré premier. (French) [S- units and S-group of classes of a cyclic number field of prime degree]. Journal of Number Theory, 85(1):35–58, November 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X00925198.

Freiman:2000:ATA

[122] Gregory Freiman and Alfred Geroldinger. An addition theorem and its arithmetical application. Journal of Number Theory, 85(1):59–73, November 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X00925216.

Andrews:2000:RCP

[123] George A. Andrews and Richard Lewis. The ranks and cranks of partitions moduli 2, 3, and 4. Journal of Number Theory, 85(1):74–84, November 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X0092537X.

Ru:2000:UTR

[124] Min Ru. A uniqueness theorem for rational points in projective space. Journal of Number Theory, 85(1):85–91, November 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X00925484.

Thiebaud:2000:CCG

[125] Caroline Thiebaud. Sur la capitulation dans les corps de genres d’une extension abélienne d’un corps quadratique imaginaire. (French) [On the capitulation in genus fields of an abelian extension of an imaginary quadratic body]. Journal of Number Theory, 85(1):92–107, November 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X00925502.

Anonymous:2000:EVNb

[126] Anonymous. Erratum: Volume 81, number 1 (2000), “A Quadratic Analogue of Artin’s Conjecture on Primitive Roots,” by Hans Roskam, pages 93–109. Journal of Number Theory, 85(1):108, November 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X0092585X. See [33].

Anonymous:2000:PD

[127] Anonymous. Pages 109–360 (December 2000). Journal of Number Theory, 85(2):??, December 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Lachaud:2000:GPF

[128] Gilles Lachaud and Serge Vlăduţ. Gauss problem for function fields. Journal of Number Theory, 85(2):109–129, December 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X0092523X.

Yang:2000:CDH

[129] Tonghai Yang. On the central derivative of Hecke L-series. Journal of Number Theory, 85(2):130–157, December 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X00925289.

Bowman:2000:RAN

[130] Douglas Bowman and Geumlan Choi. Rational approximation to a new generalization of the Rogers–Ramanujan continued fraction. Journal of Number Theory, 85(2):158–171, December 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X00925381.

Chambert-Loir:2000:PBH

[131] Antoine Chambert-Loir and Yuri Tschinkel. Points of bounded height on equivariant compactifications of vector groups, II. Journal of Number Theory, 85(2):172–188, December 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X00925393.

Chen:2000:ANI

[132] Zesen Chen and David R. Hayes. Average norms of j-invariants of Drinfeld modules. Journal of Number Theory, 85(2):189–200, December 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X0092540X.

Gaal:2000:CPI

[133] István Gaál and Michael Pohst. Computing power integral bases in quartic relative extensions. Journal of Number Theory, 85(2):201–219, December 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X00925411.

Lang:2000:PCS

[134] Mong-Lung Lang, Chong-Hai Lim, and Ser-Peow Tan. Principal congruence subgroups of the Hecke groups. Journal of Number Theory, 85(2):220–230, December 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X00925423.

Liu:2000:SES

[135] Zhi-Guo Liu. Some Eisenstein series identities. Journal of Number Theory, 85(2):231–252, December 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X00925435.

Locher:2000:NSA

[136] Helmut Locher. On the number of solutions in algebraic integers of the Thue–Mahler equation. Journal of Number Theory, 85(2):253–262, December 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X00925447.

Louboutin:2000:EBR

[137] Stéphane Louboutin. Explicit bounds for residues of Dedekind zeta functions, values of L-Functions at s = 1, and relative class numbers. Journal of Number Theory, 85(2):263–282, December 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X00925459.

Lovejoy:2000:RTC

[138] Jeremy Lovejoy. Ramanujan-type congruences for three colored Frobenius partitions. Journal of Number Theory, 85(2):283–290, December 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X00925460.

Moree:2000:TVA

[139] Pieter Moree and Peter Stevenhagen. A two-variable Artin conjecture. Journal of Number Theory, 85(2):291–304, December 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X00925472.

Siligardos:2000:EPR

[140] Giorgos Siligardos. The embedding problem and representation of prime powers by ambiguous classes of quadratic forms of discriminant −2^{2s}qr. Journal of Number Theory, 85(2):305–319, December 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X00925496.

Vaaler:2000:LNP

[141] Jeffrey D. Vaaler and José Felipe Voloch. The least nonsplit prime in Galois extensions of Q. Journal of Number Theory, 85(2):320–335, December 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X00925514.

Bauer:2000:STP

[142] Claus Bauer, Ming-Chit Liu, and Tao Zhan. On a sum of three prime squares. Journal of Number Theory, 85(2):336–359, December 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X00925526.

Anonymous:2000:AIVf

[143] Anonymous. Author index for volume 85. Journal of Number Theory, 85(2):360, December 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X00926362.

Anonymous:2001:PJa

[144] Anonymous. Pages 1–174 (January 2001). Journal of Number Theory, 86(1):??, January 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Kawamoto:2001:QSR

[145] Fuminori Kawamoto. On quadratic subextensions of ray class fields of quadratic fields modp. Journal of Number Theory, 86(1):1–38, January 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X00925538.

Saias:2001:EDD

[146] Eric Saias. Entiers à diviseurs denses 2. (French) [Whole numbers with dense divisors 2]. Journal of Number Theory, 86(1):39–49, January 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X00925599.

Lang:2001:N

[147] Mong-Lung Lang. Normalizer of Γ_{1}(m). Journal of Number Theory, 86(1):50–60, January 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X00925551.

Zhu:2001:SAV

[148] Hui June Zhu. Supersingular abelian varieties over finite fields. Journal of Number Theory, 86(1):61–77, January 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X00925629.

Ibukiyama:2001:SMF

[149] Tomoyoshi Ibukiyama and Hidenori Katsurada. Squared Möbius function for half-integral matrices and its applications. Journal of Number Theory, 86(1):78–117, January 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X0092568X.

Hofmeister:2001:TBO

[150] Gerd Hofmeister. Thin bases of order two. Journal of Number Theory, 86(1):118–132, January 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X00925678.

Zhai:2001:SDF

[151] Wenguang Zhai. On the simultaneous distribution of the fractional parts of different powers of prime numbers. Journal of Number Theory, 86(1):133–155, January 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X00925630.

Lopez:2001:TDC

[152] Ricardo García López. On the total degree of certain L-functions. Journal of Number Theory, 86(1):156–162, January 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X00925563.

Sumida:2001:CIE

[153] Hiroki Sumida. On capitulation of S-ideals in Z_{p}-extensions. Journal of Number Theory, 86(1):163–174, January 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X00925617.

Anonymous:2001:PF

[154] Anonymous. Pages 175–388 (February 2001). Journal of Number Theory, 86(2):??, February 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Benedetto:2001:RDJ

[155] Robert L. Benedetto. Reduction, dynamics, and Julia sets of rational functions. Journal of Number Theory, 86(2):175–195, February 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X00925770.

Sharifi:2001:NRS

[156] Romyar T. Sharifi. On norm residue symbols and conductors. Journal of Number Theory, 86(2):196–209, February 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X00925605.

Rhodes:2001:AFD

[157] John A. Rhodes. Automorphic forms, definite quaternion algebras, and Atkin–Lehner theory on trees. Journal of Number Theory, 86 (2):210–243, February 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X00925708.

McGraw:2001:TOS

[158] William J. McGraw. On a Theorem of Ono and Skinner. Journal of Number Theory, 86(2):244–252, February 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X00925575.

Schmidt:2001:LNG

[159] Ralf Schmidt. Local newforms with global applications in the Jacobi theory. Journal of Number Theory, 86(2):253–283, February 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X00925757.

Mozzochi:2001:DEH

[160] C. J. Mozzochi. On the distribution of the eigenvalues of the hyperbolic Laplacian for PSL(2, Z), II. Journal of Number Theory, 86 (2):284–301, February 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X00925587.

Ueno:2001:EMF

[161] Takahiko Ueno. Elliptic modular forms arising from zeta functions in two variables attached to the space of binary Hermitian forms. Journal of Number Theory, 86(2):302–329, February 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X00925721.

Kraaikamp:2001:PSC

[162] Cor Kraaikamp and Hitoshi Nakada. On a problem of Schweiger concerning normal numbers. Journal of Number Theory, 86(2):330–340, February 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X0092554X.

Sorensen:2001:S

[163] Claus Mazanti Sorensen. On the structure of (O_{K}/I)^{×}. Journal of Number Theory, 86(2):341–350, February 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X0092571X.

Jorgenson:2001:GTF

[164] Jay Jorgenson and Serge Lang. Guinand’s Theorem and functional equations for the Cramér functions. Journal of Number Theory, 86 (2):351–367, February 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X00925812.

Dubickas:2001:MMM

[165] A. Dubickas and C. J. Smyth. On the metric Mahler measure. Journal of Number Theory, 86(2):368–387, February 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X00925794.

Anonymous:2001:AIVa

[166] Anonymous. Author index for volume 86. Journal of Number Theory, 86(2):388, February 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X01926617.

Anonymous:2001:PMa

[167] Anonymous. Pages 1–171 (March 2001). Journal of Number Theory, 87 (1):??, March 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Hatada:2001:CAM

[168] Kazuyuki Hatada. On classical and ℓ-adic modular forms of levels Nℓ^{m} and N. Journal of Number Theory, 87(1):1–14, March 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X00925836.

Veenstra:2001:SMF

[169] Tamara B. Veenstra. Siegel modular forms, L-functions, and Satake parameters. Journal of Number Theory, 87(1):15–30, March 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X00925861.

Schweizer:2001:ECF

[170] Andreas Schweizer. On elliptic curves over function fields of characteristic two. Journal of Number Theory, 87(1):31–53, March 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X00925824.

Bays:2001:ZDF

[171] Carter Bays, Kevin Ford, Richard H. Hudson, and Michael Rubinstein. Zeros of Dirichlet L-functions near the real axis and Chebyshev’s bias. Journal of Number Theory, 87(1):54–76, March 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X00926015.

Button:2001:MNP

[172] J. O. Button. Markoff numbers, principal ideals and continued fraction expansions. Journal of Number Theory, 87(1):77–95, March 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X00925782.

Baba:2001:SCQ

[173] Srinath Baba. Shimura curve quotients with odd Jacobians. Journal of Number Theory, 87(1):96–108, March 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X00926003.

Jacquet:2001:FPI

[174] Hervé Jacquet. Factorization of period integrals. Journal of Number Theory, 87(1):109–143, March 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X00925848.

Bell:2001:PPC

[175] Jason P. Bell. A proof of a partition conjecture of Bateman and Erdős. Journal of Number Theory, 87(1):144–153, March 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X00925897.

Mao:2001:SBF

[176] Zhengyu Mao. Spherical Bessel functions of GL_{2}(F_{q^{2}}). Journal of Number Theory, 87(1):154–171, March 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X00925769.

Anonymous:2001:PAa

[177] Anonymous. Pages 173–332 (April 2001). Journal of Number Theory, 87(2):??, April 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Jia:2001:MVD

[178] Chaohua Jia. On the mean value of Dedekind sums. Journal of Number Theory, 87(2):173–188, April 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X00925800.

Temkine:2001:HCF

[179] Alexandre Temkine. Hilbert class field towers of function fields over finite fields and lower bounds for A(q). Journal of Number Theory, 87(2):189–210, April 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X00925964.

Khoai:2001:UPB

[180] Ha Huy Khoai and Ta Thi Hoai An. On uniqueness polynomials and bi-URS for p-adic meromorphic functions. Journal of Number Theory, 87(2):211–221, April 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X00925915.

Hutchinson:2001:SSW

[181] Kevin Hutchinson. The 2-Sylow subgroup of the wild kernel of exceptional number fields. Journal of Number Theory, 87(2):222–238, April 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X00925903.

Puchta:2001:RPY

[182] Jan-Christoph Puchta. Remark on a paper of Yu on Heilbronn’s exponential sum. Journal of Number Theory, 87(2):239–241, April 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X00925939.

Angles:2001:GCF

[183] Bruno Anglès. On Gekeler’s Conjecture for function fields. Journal of Number Theory, 87(2):242–252, April 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X00925873.

Burnol:2001:ACP

[184] Jean-François Burnol. An adelic causality problem related to abelian L-functions. Journal of Number Theory, 87(2):253–269, April 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X00926167.

Rodseth:2001:APF

[185] Øystein J. Rødseth and James A. Sellers. On m-ary partition function congruences: a fresh look at a past problem. Journal of Number Theory, 87(2):270–281, April 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X00925940.

Atake:2001:ECL

[186] Daisuke Atake. On elliptic curves with large Tate–Shafarevich groups. Journal of Number Theory, 87(2):282–300, April 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X0092599X.

Boutabaa:2001:NAL

[187] Abdelbaki Boutabaa. A note on p-adic linear differential equations. Journal of Number Theory, 87(2):301–305, April 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X00926027.

Plagne:2001:ROE

[188] Alain Plagne. Removing one element from an exact additive basis. Journal of Number Theory, 87(2):306–314, April 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X00925927.

delaBreteche:2001:CPV

[189] Régis de la Bretèche. Compter des points d’une variété torique. Journal of Number Theory, 87(2):315–331, April 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X00926052.

Anonymous:2001:AIVb

[190] Anonymous. Author index for volume 87. Journal of Number Theory, 87 (2):332, April 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X01926769.

Anonymous:2001:PMb

[191] Anonymous. Pages 1–209 (May 2001). Journal of Number Theory, 88(1): ??, May 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Bae:2001:PDR

[192] Sunghan Bae. Punctured distributions in the rational function fields. Journal of Number Theory, 88(1):1–12, May 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X00925885.

Alexandru:2001:T

[193] Victor Alexandru, Nicolae Popescu, and Alexandru Zaharescu. Trace on C_{p}. Journal of Number Theory, 88(1):13–48, May 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X00926106.

Cobeli:2001:DSP

[194] C. I. Cobeli, S. M. Gonek, and A. Zaharescu. On the distribution of small powers of a primitive root. Journal of Number Theory, 88(1): 49–58, May 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X00926040.

Ochiai:2001:CTG

[195] Tadashi Ochiai. Control theorem for Greenberg’s Selmer groups of Galois deformations. Journal of Number Theory, 88(1):59–85, May 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X00926118.

Garrity:2001:PSA

[196] Thomas Garrity. On periodic sequences for algebraic numbers. Journal of Number Theory, 88(1):86–103, May 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X00926088.

Sebbar:2001:TFC

[197] Abdellah Sebbar. Twisted L-functions and complex multiplication. Journal of Number Theory, 88(1):104–113, May 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X00926131.

Deschamps:2001:CPF

[198] Bruno Deschamps. Corps pythagoriciens, fermatiens et P -réduisants. (French) [Pythagorean, Fermatian, and P -reducing fields]. Journal of Number Theory, 88(1):114–128, May 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X00926076.

Berndt:2001:RQT

[199] Bruce C. Berndt, Heng Huat Chan, and Wen-Chin Liaw. On Ramanujan’s quartic theory of elliptic functions. Journal of Number Theory, 88(1):129–156, May 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X00926155.

Hsu:2001:ACR

[200] Chih-Nung Hsu and Jing Yu. On Artin’s Conjecture for rank one Drinfeld modules. Journal of Number Theory, 88(1):157–174, May 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X00926179.

Ichimura:2001:QUI

[201] Humio Ichimura. On a quotient of the unramified Iwasawa module over an abelian number field. Journal of Number Theory, 88(1):175– 190, May 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X00926180.

Bluher:2001:TDP

[202] Antonia W. Bluher. Traces of division polynomials in characteristic two. Journal of Number Theory, 88(1):191–195, May 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X00926258.

Robertson:2001:PBP

[203] Leanne Robertson. Power bases for 2-power cyclotomic fields. Journal of Number Theory, 88(1):196–209, May 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X00926210.

Anonymous:2001:PJb

[204] Anonymous. Pages 211–380 (June 2001). Journal of Number Theory, 88 (2):??, June 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Rains:2001:CGM

[205] E. M. Rains. Class groups and modular lattices. Journal of Number Theory, 88(2):211–224, June 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X00926337.

Choudhry:2001:SQS

[206] Ajai Choudhry. On the solvability of quintic and sextic diophantine equations of the type f(x, y) = f(u, v). Journal of Number Theory, 88(2):225–240, June 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X00926039.

Pellarin:2001:DPA

[207] Federico Pellarin. Sur la distance d’un point algébrique à l’origine dans les variétés abéliennes. (French) [On the distance of an algebraic point at the origin in the abelian varieties]. Journal of Number Theory, 88(2):241–262, June 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X0092612X.

Conner:2001:RTK

[208] P. E. Conner and Jurgen Hurrelbrink. On the 4-rank of the tame kernel K_{2}(O) in positive definite terms. Journal of Number Theory, 88(2):263–282, June 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X0092626X.

Urbanski:2001:PCI

[209] Mariusz Urbański. Porosity in conformal infinite iterated function systems. Journal of Number Theory, 88(2):283–312, June 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X00926453.

Kuba:2001:DSH

[210] G. Kuba. On the distribution of squares of hypercomplex integers. Journal of Number Theory, 88(2):313–334, June 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X00926441.

Wittmann:2001:GSE

[211] Christian Wittmann. Group structure of elliptic curves over finite fields. Journal of Number Theory, 88(2):335–344, June 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X00926222.

Groger:2001:EJS

[212] Detlef Gröger. An extension of the Jacobi symbol due to Gauss. Journal of Number Theory, 88(2):345–356, June 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X0092643X.

Khare:2001:LLC

[213] Chandrashekhar Khare. On the local Langlands correspondence modℓ. Journal of Number Theory, 88(2):357–365, June 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X00925745.

Seo:2001:CDE

[214] Soogil Seo. Circular distributions and Euler systems. Journal of Number Theory, 88(2):366–379, June 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X00926349.

Anonymous:2001:AIVc

[215] Anonymous. Author index for volume 88. Journal of Number Theory, 88(2):380, June 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X01927052.

Anonymous:2001:PJc

[216] Anonymous. Pages 1–178 (July 2001). Journal of Number Theory, 89(1): ??, July 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Zhang:2001:SAD

[217] Wenpeng Zhang. A sum analogous to the Dedekind sum and its mean value formula. Journal of Number Theory, 89(1):1–13, July 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X00926246.

Jeong:2001:RPP

[218] Sangtae Jeong. Rational points on Picard groups of some genuschanging curves of genus at least 2. Journal of Number Theory, 89(1): 14–30, July 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X0092609X.

Denef:2001:NPI

[219] Jan Denef and Kathleen Hoornaert. Newton polyhedra and Igusa’s local zeta function. Journal of Number Theory, 89(1):31–64, July 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X00926064.

Drmota:2001:SDF

[220] Michael Drmota and Gerhard Larcher. The sum-of-digits-function and uniform distribution modulo 1. Journal of Number Theory, 89(1):65– 96, July 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X00926283.

Kim:2001:RBF

[221] Myung-Hwan Kim, Ja Kyung Koo, and Byeong-Kweon Oh. Representations of binary forms by certain quinary positive integral quadratic forms. Journal of Number Theory, 89(1):97–113, July 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X00926313.

Wong:2001:ECG

[222] Siman Wong. Exponents of class groups and elliptic curves. Journal of Number Theory, 89(1):114–120, July 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X00926234. See

corrigendum [263].

Chen:2001:IF

[223] Yong-Gao Chen. On integers of the forms k − 2^{n} and k2^{n} + 1. Journal of Number Theory, 89(1):121–125, July 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X0192640X.

Dujella:2001:ABS

[224] Andrej Dujella. An absolute bound for the size of diophantine mtuples. Journal of Number Theory, 89(1):126–150, July 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X00926271.

Aoki:2001:IMC

[225] Miho Aoki. The Iwasawa Main Conjecture and Gauss sums. Journal of Number Theory, 89(1):151–164, July 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X0192637X.

Jeong:2001:HOC

[226] Sangtae Jeong. Hyperdifferential operators and continuous functions on function fields. Journal of Number Theory, 89(1):165–178, July 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X00926295.

Anonymous:2001:PAb

[227] Anonymous. Pages 179–369 (August 2001). Journal of Number Theory, 89(2):??, August 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Kolesnik:2001:PPD

[228] Grigori Kolesnik. Prime power divisors of multinomial and qmultinomial coefficients. Journal of Number Theory, 89(2):179–192, August 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X01926575.

Moulton:2001:RPN

[229] David Petrie Moulton. Representing powers of numbers as subset sums of small sets. Journal of Number Theory, 89(2):193–211, August 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X00926465.

Develin:2001:OSR

[230] Mike Develin. On optimal subset representations of integer sets. Journal of Number Theory, 89(2):212–221, August 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X00926477.

Ahlgren:2001:MRP

[231] Scott Ahlgren. Multiplicative Relations in Powers of Euler’s Product. Journal of Number Theory, 89(2):222–233, August 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X01926484.

Gomez:2001:DP

[232] Emiliano Gómez. Deformations of Pseudorepresentations. Journal of Number Theory, 89(2):234–267, August 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X01926514.

Freeman:2001:QDI

[233] D. Eric Freeman. Quadratic diophantine inequalities. Journal of Number Theory, 89(2):268–307, August 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X01926496.

Kameswari:2001:CTE

[234] P. Anuradha Kameswari and Rajat Tandon. A converse theorem for epsilon factors. Journal of Number Theory, 89(2):308–323, August 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X00926192.

Kedlaya:2001:PSA

[235] Kiran S. Kedlaya. Power series and p-adic algebraic closures. Journal of Number Theory, 89(2):324–339, August 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X00926301.

Jehanne:2001:RG

[236] A. Jehanne. Realization over Q of the groups Ã_{5} and Â_{5}. Journal of Number Theory, 89(2):340–368, August 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X01926563.

Anonymous:2001:AIVd

[237] Anonymous. Author index for volume 89. Journal of Number Theory, 89(2):369, August 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X0192726X.

Anonymous:2001:PS

[238] Anonymous. Pages 1–183 (September 2001). Journal of Number Theory, 90(1):??, September 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Daboussi:2001:BFL

[239] H. Daboussi. Brun’s fundamental lemma and exponential sums over primes. Journal of Number Theory, 90(1):1–18, September 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X01926393.

Iosevich:2001:LPG

[240] Alex Iosevich. Lattice points and generalized diophantine conditions. Journal of Number Theory, 90(1):19–30, September 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X01926551.

Lang:2001:NCS

[241] Mong-Lung Lang. Normalizers of the congruence subgroups of the Hecke groups G_{4} and G_{6}. Journal of Number Theory, 90(1):31–43, September 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X00925691.

Mossige:2001:PSP

[242] Svein Mossige. The Postage Stamp Problem: the extremal basis for k = 4. Journal of Number Theory, 90(1):44–61, September 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X00926209.

Eichhorn:2001:GUT

[243] Dennis Eichhorn. Gap unification theory: a note on the treatment of partitions via their gaps. Journal of Number Theory, 90(1):62–73, September 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X00926416.

Takahashi:2001:DPE

[244] Shuzo Takahashi. Degrees of parametrizations of elliptic curves by Shimura curves. Journal of Number Theory, 90(1):74–88, September 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X00926143.

Yang:2001:NZM

[245] Zifeng Yang. A note on zeta measures over function fields. Journal of Number Theory, 90(1):89–112, September 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X00925976.

Hermez:2001:RTA

[246] Alain Hermez and Alain Salinier. Rational trinomials with the alternating group as Galois group. Journal of Number Theory, 90(1):113–129, September 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X01926538.

Shirai:2001:SII

[247] Susumu Shirai and Ken ichi Sato. Some identities involving Bernoulli and Stirling numbers. Journal of Number Theory, 90(1):130–142, September 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X01926599.

Hashimoto:2001:AAC

[248] Ryūta Hashimoto. Ankeny–Artin–Chowla Conjecture and continued fraction expansion. Journal of Number Theory, 90(1):143–153, September 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X01926526.

Ichimura:2001:NCU

[249] Humio Ichimura. Note on cyclotomic units and Gauss sums in local cyclotomic fields. Journal of Number Theory, 90(1):154–165, September 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X0192654X.

Gekeler:2001:ISA

[250] Ernst-Ulrich Gekeler. Invariants of some algebraic curves related to Drinfeld modular curves. Journal of Number Theory, 90(1):166–183, September 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X00926428.

Anonymous:2001:PO

[251] Anonymous. Pages 185–378 (October 2001). Journal of Number Theory, 90(2):??, October 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Buium:2001:IML

[252] Alexandru Buium. Infinitesimal Mordell–Lang. Journal of Number Theory, 90(2):185–206, October 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X01926630.

Nagoshi:2001:SZF

[253] Hirofumi Nagoshi. Selberg zeta functions over function fields. Journal of Number Theory, 90(2):207–238, October 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X01926587.

Brownawell:2001:MEA

[254] W. Dale Brownawell. Minimal extensions of algebraic groups and linear independence. Journal of Number Theory, 90(2):239–254, October 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X01926381.

Khrennikov:2001:NCA

[255] Andrei Khrennikov and Marcus Nilsson. On the number of cycles of p-adic dynamical systems. Journal of Number Theory, 90(2):255–264, October 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X01926654.

Peter:2001:LDN

[256] Manfred Peter. The limit distribution of a number theoretic function arising from a problem in statistical mechanics. Journal of Number Theory, 90(2):265–280, October 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X01926666.

Borovoi:2001:RII

[257] Mikhail Borovoi. On representations of integers by indefinite ternary quadratic forms. Journal of Number Theory, 90(2):281–293, October 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X01926629.

Murty:2001:KC

[258] M. Ram Murty and Yiannis N. Petridis. On Kummer’s Conjecture. Journal of Number Theory, 90(2):294–303, October 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X01926678.

An:2001:JGO

[259] Sang Yook An, Seog Young Kim, David C. Marshall, Susan H. Marshall, William G. McCallum, and Alexander R. Perlis. Jacobians of genus one curves. Journal of Number Theory, 90(2):304–315, October 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X00926325.

Sander:2001:PEP

[260] J. W. Sander. On the parity of exponents in the prime factorization of factorials. Journal of Number Theory, 90(2):316–328, October 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X00926684.

David:2001:FDD

[261] Chantal David. Frobenius distributions of Drinfeld modules of any rank. Journal of Number Theory, 90(2):329–340, October 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X00926647.

Goren:2001:HMF

[262] Eyal Z. Goren. Hilbert modular forms modulo p^{m}: the unramified case. Journal of Number Theory, 90(2):341–375, October 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X01926605.

Wong:2001:CEC

[263] Siman Wong. Corrigendum: Exponents of class groups and elliptic curves: Volume 89, Number 1 (2001), pages 114–120. Journal of Number Theory, 90(2):376–377, October 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X01927386. See [222].

Anonymous:2001:AIVe

[264] Anonymous. Author index for volume 90. Journal of Number Theory, 90(2):378, October 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X01927465.

Anonymous:2001:PN

[265] Anonymous. Pages 1–185 (November 2001). Journal of Number Theory, 91(1):??, November 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Cochrane:2001:AFT

[266] Todd Cochrane and J. C. Peral. An asymptotic formula for a trigonometric sum of Vinogradov. Journal of Number Theory, 91(1):1–19, November 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X01926794.

Jin:2001:NMU

[267] Renling Jin. Nonstandard methods for upper Banach density problems. Journal of Number Theory, 91(1):20–38, November 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X0192671X.

Allouche:2001:TSM

[268] J.-P. Allouche, J. L. Davison, M. Queffélec, and L. Q. Zamboni. Transcendence of Sturmian or morphic continued fractions. Journal of Number Theory, 91(1):39–66, November 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X01926691.

Shastri:2001:IPU

[269] Parvati Shastri. Integral points on the unit circle. Journal of Number Theory, 91(1):67–70, November 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X00926350.

Nyblom:2001:TTI

[270] M. A. Nyblom. A theorem on transcendence of infinite series II. Journal of Number Theory, 91(1):71–80, November 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X01926721.

Gaulter:2001:MOU

[271] Mark Gaulter. Minima of odd unimodular lattices in dimension 24m. Journal of Number Theory, 91(1):81–91, November 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X01926708.

Grabitz:2001:LZT

[272] Martin Grabitz, Allan J. Silberger, and Ernst-Wilhelm Zink. Level zero types and Hecke algebras for local central simple algebras. Journal of Number Theory, 91(1):92–125, November 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X01926848.

Vinatier:2001:SGE

[273] Stéphane Vinatier. Structure galoisienne dans les extensions faiblement ramifiées de Q. (French) [Galois structure in the weakly branched extensions of Q]. Journal of Number Theory, 91(1):126–152, November 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X01926745.

Winterhof:2001:CSP

[274] Arne Winterhof. Character sums, primitive elements, and powers in finite fields. Journal of Number Theory, 91(1):153–163, November 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X01926757.

Rieger:2001:PTT

[275] G. J. Rieger. On Pythagorean triples and their harmonic fourths. Journal of Number Theory, 91(1):164–173, November 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X01926733.

Cornelissen:2001:PCG

[276] Gunther Cornelissen. The 2-primary class group of certain hyperelliptic curves. Journal of Number Theory, 91(1):174–185, November 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X01926800.

Anonymous:2001:PD

[277] Anonymous. Pages 187–372 (December 2001). Journal of Number Theory, 91(2):??, December 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Astels:2001:SNS

[278] S. Astels. Sums of numbers with small partial quotients II. Journal of Number Theory, 91(2):187–205, December 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X01926782.

Golinskii:2001:LSE

[279] Leonid Golinskii, Doron S. Lubinsky, and Paul Nevai. Large sieve estimates on arcs of a circle. Journal of Number Theory, 91(2):206–229, December 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X01926836.

Dartyge:2001:EDN

[280] Cécile Dartyge and Christian Mauduit. Ensembles de densité nulle contenant des entiers possédant au plus deux facteurs premiers. (French) [Zero density sets containing integers having at most two prime factors].

Journal of Number Theory, 91(2):230–255, December 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X01926812.

Einsiedler:2001:ECH

[281] M. Einsiedler, G. Everest, and T. Ward. Entropy and the canonical height. Journal of Number Theory, 91(2):256–273, December 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X01926824.

Jeon:2001:SVD

[282] Daeyeol Jeon and Chang Heon Kim. On the singular values of the Drinfeld modular function μ. Journal of Number Theory, 91(2):274–283, December 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X0192685X.

Gille:2001:ERC

[283] Philippe Gille. R-équivalence sur les G-revêtements sur les corps locaux non archimédiens. (French) [R-equivalence on G-coatings on local non–Archimedean fields]. Journal of Number Theory, 91(2):284–292, December 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X01926502.

Moret-Bailly:2001:EST

[284] Laurent Moret-Bailly. R-équivalence simultanée de torseurs: un complément à l’article de P. Gille. (French) [simultaneous R-equivalence of torsors: a complement to the article by P. Gille]. Journal of Number Theory, 91(2):293–296, December 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X01926770.

Angles:2001:NRS

[285] Bruno Anglès. Norm residue symbol and the first case of Fermat’s equation. Journal of Number Theory, 91(2):297–311, December 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X01926939.

Lau:2001:LDO

[286] Yuk-Kam Lau. Local distribution of ordered factorizations of integers. Journal of Number Theory, 91(2):312–317, December 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X01926873.

Kitaoka:2001:DUC

[287] Yoshiyuki Kitaoka. Distribution of units of a cubic field with negative discriminant. Journal of Number Theory, 91(2):318–355, December 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X01926861.

Balog:2001:CDT

[288] Antal Balog and Ken Ono. The Chebotarev density theorem in short intervals and some questions of Serre. Journal of Number Theory, 91(2):356–371, December 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X01926940.

Anonymous:2001:AIVf

[289] Anonymous. Author index for volume 91. Journal of Number Theory, 91(2):372, December 2001. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X01927660.

Anonymous:2002:PJa

[290] Anonymous. Pages 1–203 (January 2002). Journal of Number Theory, 92(1):??, January 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Ota:2002:KTC

[291] Kaori Ota. On Kummer-type congruences for derivatives of Barnes’ multiple Bernoulli polynomials. Journal of Number Theory, 92(1):1– 36, January 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X01927027.

Lee:2002:CLH

[292] Yoonjin Lee. Cohen–Lenstra heuristics and the Spiegelungssatz: number fields. Journal of Number Theory, 92(1):37–66, January 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X0192699X.

Mazur:2002:FAS

[293] Marcin Mazur. Finite arithmetic subgroups of GL_{n}, II. Journal of Number Theory, 92(1):67–86, January 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X01926903.

Mills:2002:EWS

[294] Donald Mills. On the evaluation of Weil sums of Dembowski– Ostrom polynomials. Journal of Number Theory, 92(1):87–98, January 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X01927015.

Liu:2002:SPP

[295] Jianya Liu, Ming-Chit Liu, and Tao Zhan. Squares of primes and powers of 2, II. Journal of Number Theory, 92(1):99–116, January 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X01926897.

Valko:2002:DAP

[296] Benedek Valkó. Discrepancy of arithmetic progressions in higher dimensions. Journal of Number Theory, 92(1):117–130, January 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X01927040.

Boylan:2002:CCP

[297] Matthew Boylan. Congruences for 2^{t}-core partition functions. Journal of Number Theory, 92(1):131–138, January 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X01926952.

Howe:2002:EAS

[298] Everett W. Howe and Hui June Zhu. On the existence of absolutely simple abelian varieties of a given dimension over an arbitrary field. Journal of Number Theory, 92(1):139–163, January 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X01926976.

Queen:2002:AGP

[299] Clifford S. Queen. Arithmetic on groups of positive rationals. Journal of Number Theory, 92(1):164–173, January 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X01927039.

Kurlberg:2002:CSP

[300] Pär Kurlberg. On a character sum problem of Cohn. Journal of Number Theory, 92(1):174–181, January 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X01926988.

Rosen:2002:RAV

[301] Michael Rosen and Siman Wong. The rank of abelian varieties over infinite Galois extensions. Journal of Number Theory, 92(1):182–196, January 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X01926927.

Fastenberg:2002:UBI

[302] Lisa A. Fastenberg. Uniform bounds for integral points on families of elliptic curves. Journal of Number Theory, 92(1):197–203, January 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X01927106.

Anonymous:2002:PF

[303] Anonymous. Pages 205–405 (February 2002). Journal of Number Theory, 92(2):??, February 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Yukie:2002:SZF

[304] Akihiko Yukie. On the Shintani zeta function for the space of pairs of binary Hermitian forms. Journal of Number Theory, 92(2):205–256, February 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X01927076.

Lee:2002:GSB

[305] Dong Hoon Lee and Sang Geun Hahn. Gauss sums and binomial coefficients. Journal of Number Theory, 92(2):257–271, February 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X01926885.

Achter:2002:ICH

[306] Jeffrey D. Achter and Clifton L. R. Cunningham. Isogeny classes of Hilbert–Blumenthal abelian varieties over finite fields. Journal of Number Theory, 92(2):272–303, February 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X01927167.

Wenpeng:2002:MGQ

[307] Zhang Wenpeng. Moments of generalized quadratic Gauss sums weighted by L-functions. Journal of Number Theory, 92(2):304–314, February 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X01927155.

Malle:2002:DGG

[308] Gunter Malle. On the distribution of Galois groups. Journal of Number Theory, 92(2):315–329, February 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X01927131.

Dilcher:2002:APB

[309] Karl Dilcher and Louise Louise. Arithmetic properties of Bernoulli–Padé numbers and polynomials. Journal of Number Theory, 92(2):330–347, February 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X01926964.

Csirik:2002:KEI

[310] János A. Csirik. The kernel of the Eisenstein ideal. Journal of Number Theory, 92(2):348–375, February 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X0192709X.

Maire:2002:RAE

[311] Christian Maire. On the Z_{ℓ}-rank of abelian extensions with restricted ramification. Journal of Number Theory, 92(2):376–404, February 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X0192712X. See addendum [453].

Anonymous:2002:AIVa

[312] Anonymous. Author index for volume 92. Journal of Number Theory, 92(2):405, February 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X02927745.

Anonymous:2002:PMa

[313] Anonymous. Pages 1–85 (March 2002). Journal of Number Theory, 93 (1):??, March 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Drinen:2002:FSI

[314] Michael J. Drinen. Finite submodules and Iwasawa μ-invariants. Journal of Number Theory, 93(1):1–22, March 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X01927179.

Chen:2002:PRO

[315] Yen-Mei J. Chen. On primitive roots of one-dimensional tori. Journal of Number Theory, 93(1):23–33, March 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X01927064.

Hafner:2002:EAH

[316] James Lee Hafner and Lynne H. Walling. Explicit action of Hecke operators on Siegel modular forms. Journal of Number Theory, 93(1):34– 57, March 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X01927118.

Chen:2002:NMR

[317] Shu-Ling Chen and Sen-Shan Huang. New modular relations for the Göllnitz–Gordon functions. Journal of Number Theory, 93(1):58– 75, March 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X01927088.

Seo:2002:NAM

[318] Soogil Seo. A note on algebraic Γ-monomials and double coverings. Journal of Number Theory, 93(1):76–85, March 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X01927143.

Anonymous:2002:PAa

[319] Anonymous. Pages 87–285 (April 2002). Journal of Number Theory, 93 (2):??, April 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Grossman:2002:SFB

[320] George Grossman and Florian Luca. Sums of factorials in binary recurrence sequences. Journal of Number Theory, 93(2):87–107, April 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X01927180.

Martin:2002:AFN

[321] Greg Martin. An asymptotic formula for the number of smooth values of a polynomial. Journal of Number Theory, 93(2):108–182, April 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X01927222.

Stoll:2002:ACI

[322] Michael Stoll. On the arithmetic of the curves y^{2} = x^{ℓ} + A, II. Journal of Number Theory, 93(2):183–206, April 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X01927271.

Lin:2002:SUE

[323] Ke-Pao Lin and Stephen S.-T. Yau. A sharp upper estimate of the number of integral points in a 5-dimensional tetrahedra. Journal of Number Theory, 93(2):207–234, April 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X01927209.

Hall:2002:WTI

[324] R. R. Hall. A Wirtinger type inequality and the spacing of the zeros of the Riemann zeta-function. Journal of Number Theory, 93(2):235– 245, April 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X01927192.

Flammenkamp:2002:BCL

[325] Achim Flammenkamp and Florian Luca. Binomial coefficients and Lucas sequences. Journal of Number Theory, 93(2):246–284, April 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X01927210.

Anonymous:2002:AIVb

[326] Anonymous. Author index for volume 93. Journal of Number Theory, 93 (2):285, April 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X0292799X.

Anonymous:2002:PMb

[327] Anonymous. Pages 1–212 (May 2002). Journal of Number Theory, 94(1): ??, May 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Agoh:2002:CIB

[328] Takashi Agoh. Congruences involving Bernoulli numbers and Fermat– Euler quotients. Journal of Number Theory, 94(1):1–9, May 2002. CO-DEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL

http://www.sciencedirect.com/science/article/pii/S0022314X01927283.

Solomon:2002:TZF

[329] David Solomon. Twisted zeta-functions and abelian Stark conjectures. Journal of Number Theory, 94(1):10–48, May 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X01927246.

Rosson:2002:TSQ

[330] Holly J. Rosson. Theta series of quaternion algebras over function fields. Journal of Number Theory, 94(1):49–79, May 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X01927234.

Gica:2002:PCA

[331] Alexandru Gica. The proof of a conjecture of additive number theory. Journal of Number Theory, 94(1):80–89, May 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X01927313.

Yamagami:2002:GCC

[332] Atsushi Yamagami. On Gouvêa’s Conjecture on Controlling the Conductor. Journal of Number Theory, 94(1):90–102, May 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X01927258.

Girard:2002:GDG

[333] Martine Girard. Géométrie du groupe des points de Weierstrass d’une quartique lisse. (French) [Weierstrass point group geometry of a smooth quartic]. Journal of Number Theory, 94(1):103–135, May 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X01927325.

Grinspan:2002:MSA

[334] Pierre Grinspan. Measures of simultaneous approximation for quasiperiods of abelian varieties. Journal of Number Theory, 94(1):136– 176, May 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X01927337.

Byeon:2002:FCR

[335] Dongho Byeon and H. M. Stark. On the finiteness of certain Rabinowitsch polynomials. Journal of Number Theory, 94(1):177–180, May 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X01927295.

Ash:2002:CCS

[336] Avner Ash, Paul E. Gunnells, and Mark McConnell. Cohomology of congruence subgroups of SL_{4}(Z). Journal of Number Theory, 94(1): 181–212, May 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X01927301.

Anonymous:2002:PJb

[337] Anonymous. Pages 213–405 (June 2002). Journal of Number Theory, 94 (2):??, June 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Coykendall:2002:EPP

[338] Jim Coykendall. Elasticity properties preserved in the normset. Journal of Number Theory, 94(2):213–218, June 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X01927404.

Gar-el:2002:DE

[339] Rachel Gar-el and Leonid Vaserstein. On the diophantine equation a^{3} + b^{3} + c^{3} + d^{3} = 0. Journal of Number Theory, 94(2):219– 223, June 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X01927374.

Perrine:2002:SDG

[340] Serge Perrine. Sommes de Dedekind et généralisations de l’équation diophantienne de Markoff. (French) [Dedekind sums and generalizations of the Markoff diophantine equation]. Journal of Number Theory, 94(2):224–247, June 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X01927453.

Roy:2002:IFA

[341] Damien Roy. Interpolation formulas and auxiliary functions. Journal of Number Theory, 94(2):248–285, June 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X00925952.

Hu:2002:GAC

[342] Pei-Chu Hu and Chung-Chun Yang. A generalized abc-conjecture over function fields. Journal of Number Theory, 94(2):286–298, June 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X01927428.

OBryant:2002:GFT

[343] Kevin O’Bryant. A generating function technique for Beatty sequences and other step sequences. Journal of Number Theory, 94(2):299– 319, June 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X0192743X.

Penniston:2002:RPF

[344] David Penniston. The p^{a}-regular partition function modulo p^{j}. Journal of Number Theory, 94(2):320–325, June 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X01927362.

Papanikolas:2002:UNA

[345] Matthew A. Papanikolas. Universal norms on abelian varieties over global function fields. Journal of Number Theory, 94(2):326–342, June 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X01927441.

Bourqui:2002:FZH

[346] David Bourqui. Fonction zêta des hauteurs des surfaces de Hirzebruch dans le cas fonctionnel. (French) [Zeta function of the heights of the surfaces of Hirzebruch in the functional case]. Journal of Number Theory, 94(2):343–358, June 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X01927398.

Lau:2002:ELD

[347] Yuk-Kam Lau. On the existence of limiting distributions of some number-theoretic error terms. Journal of Number Theory, 94(2):359– 374, June 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X01927349.

Longhi:2002:NAI

[348] Ignazio Longhi. Non-Archimedean integration and elliptic curves over function fields. Journal of Number Theory, 94(2):375–404, June 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X01927350.

Anonymous:2002:AIVc

[349] Anonymous. Author index for volume 94. Journal of Number Theory, 94(2):405, June 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X02928210.

Anonymous:2002:PJc

[350] Anonymous. Pages 1–121 (July 2002). Journal of Number Theory, 95(1): ??, July 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Szechtman:2002:QGS

[351] Fernando Szechtman. Quadratic Gauss sums over finite commutative rings. Journal of Number Theory, 95(1):1–13, July 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X01927520.

Brown:2002:DCP

[352] Alexander F. Brown and Eknath P. Ghate. Dihedral congruence primes and class fields of real quadratic fields. Journal of Number Theory, 95(1):14–37, July 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X01927532.

Delbourgo:2002:ABS

[353] Daniel Delbourgo. On the p-adic Birch, Swinnerton-Dyer Conjecture for non-semistable reduction. Journal of Number Theory, 95(1):38– 71, July 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X01927556.

Grundman:2002:HMT

[354] H. G. Grundman and L. E. Lippincott. Hilbert modular threefolds of arithmetic genus one. Journal of Number Theory, 95(1):72– 76, July 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X01927477.

Kitazume:2002:EUG

[355] Masaaki Kitazume and Akihiro Munemasa. Even unimodular Gaussian lattices of rank 12. Journal of Number Theory, 95(1):77– 94, July 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X01927490.

Arif:2002:DE

[356] S. A. Arif and Fadwa S. Abu Muriefah. On the diophantine equation x^{2} + q^{2k+1} = y^{n}. Journal of Number Theory, 95(1):95–100, July 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X01927507.

Richter:2002:TFH

[357] Olav K. Richter. Theta functions with harmonic coefficients over number fields. Journal of Number Theory, 95(1):101–121, July 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X01927581.

Anonymous:2002:PAb

[358] Anonymous. Pages 123–351 (August 2002). Journal of Number Theory, 95(2):??, August 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Finotti:2002:DET

[359] Luís R. A. Finotti. Degrees of the Elliptic Teichmüller Lift. Journal of Number Theory, 95(2):123–141, August 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X01927568.

Lai:2002:RPF

[360] King Fai Lai and Kit Ming Yeung. Rational points in flag varieties over function fields. Journal of Number Theory, 95(2):142–149, August 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X0192757X.

Deitmar:2002:CNO

[361] Anton Deitmar. Class numbers of orders in cubic fields. Journal of Number Theory, 95(2):150–166, August 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X01927544.

vanHoeij:2002:FPK

[362] Mark van Hoeij. Factoring polynomials and the knapsack problem. Journal of Number Theory, 95(2):167–189, August 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X01927635.

Bars:2002:TNC

[363] Francesc Bars. On the Tamagawa Number Conjecture for CM elliptic curves defined over Q. Journal of Number Theory, 95(2):190– 208, August 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X02927769.

BenSaid:2002:CNS

[364] F. Ben Saïd. On a conjecture of Nicolas–Sárközy about partitions. Journal of Number Theory, 95(2):209–226, August 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X01927714.

Shlapentokh:2002:DDD

[365] Alexandra Shlapentokh. Diophantine definability and decidability in large subrings of totally real number fields and their totally complex extensions of degree 2. Journal of Number Theory, 95(2):227– 252, August 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X01927593.

Ono:2002:RIS

[366] Ken Ono. Representations of Integers as Sums of Squares. Journal of Number Theory, 95(2):253–258, August 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X01927659.

Cohen:2002:OII

[367] Joseph Cohen and Jack Sonn. On the Ono invariants of imaginary quadratic fields. Journal of Number Theory, 95(2):259–267, August 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X01927684.

Lasjaunias:2002:FPS

[368] A. Lasjaunias and J.-J. Ruch. Flat power series over a finite field. Journal of Number Theory, 95(2):268–288, August 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X01927647.

Frankenhuysen:2002:ACI

[369] Machiel Van Frankenhuysen. The ABC Conjecture implies Vojta’s height inequality for curves. Journal of Number Theory, 95(2):289– 302, August 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X01927696.

Auer:2002:LEC

[370] Roland Auer and Jaap Top. Legendre elliptic curves over finite fields. Journal of Number Theory, 95(2):303–312, August 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X0192760X.

Dani:2002:OVB

[371] S. G. Dani and Arnaldo Nogueira. On orbits of SL(2, Z)_{+} and values of binary quadratic forms on positive integral pairs. Journal of Number Theory, 95(2):313–328, August 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X01927611.

ONeil:2002:PIO

[372] Catherine O’Neil. The period-index obstruction for elliptic curves. Journal of Number Theory, 95(2):329–339, August 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X01927702. See

[760].

Banks:2002:NSR

[373] William D. Banks and Igor E. Shparlinski. On the number of sparse RSA exponents. Journal of Number Theory, 95(2):340–350, August 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X01927751.

Anonymous:2002:AIVd

[374] Anonymous. Author index for volume 95. Journal of Number Theory, 95(2):351, August 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X02928404.

Anonymous:2002:PS

[375] Anonymous. Pages 1–224 (September 2002). Journal of Number Theory, 96(1):??, September 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Beck:2002:FPR

[376] Matthias Beck, Ricardo Diaz, and Sinai Robins. The Frobenius Problem, rational polytopes, and Fourier–Dedekind sums. Journal of Number Theory, 96(1):1–21, September 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X02927861.

Khadjavi:2002:EVB

[377] Lily S. Khadjavi. An effective version of Belyi’s Theorem. Journal of Number Theory, 96(1):22–47, September 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X01927489.

Schoen:2002:CRS

[378] Tomasz Schoen. The cardinality of restricted sumsets. Journal of Number Theory, 96(1):48–54, September 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X01927726.

Bachoc:2002:ZFU

[379] Christine Bachoc and Gabriele Nebe. Zonal functions for the unitary groups and applications to Hermitian lattices. Journal of Number Theory, 96(1):55–75, September 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X0292785X.

Garcia-Selfa:2002:CRT

[380] Irene García-Selfa, Miguel A. Olalla, and José M. Tornero. Computing the rational torsion of an elliptic curve using Tate Normal Form. Journal of Number Theory, 96(1):76–88, September 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X02927800.

Amou:2002:APS

[381] Masaaki Amou and Keijo Väänänen. Arithmetic properties of solutions of certain functional equations in several variables. Journal of Number Theory, 96(1):89–104, September 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X02927848.

Ichimura:2002:NIB

[382] Humio Ichimura. On a normal integral basis problem over cyclotomic Z_{p}-extensions, II. Journal of Number Theory, 96(1):105–132, September 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X02927915.

Weilert:2002:FCB

[383] André Weilert. Fast computation of the biquadratic residue symbol. Journal of Number Theory, 96(1):133–151, September 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X02927836.

Luca:2002:PLI

[384] F. Luca and P. G. Walsh. The product of like-indexed terms in binary recurrences. Journal of Number Theory, 96(1):152–173, September 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X02927940.

Bugeaud:2002:APN

[385] Yann Bugeaud. Approximation par des nombres algébriques de degré borné et dimension de Hausdorff. (French) [Approximation by algebraic numbers of bounded degree and Hausdorff dimension ]. Journal of Number Theory, 96(1):174–200, September 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X02927782.

Fisher:2002:SPM

[386] Benji Fisher. The stationary-phase method for exponential sums with multiplicative characters. Journal of Number Theory, 96(1):201–224, September 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X02927903.

Anonymous:2002:PO

[387] Anonymous. Pages 225–449 (October 2002). Journal of Number Theory, 96(2):??, October 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Zuehlke:2002:DGE

[388] John A. Zuehlke. The Darmon–Granville equation with algebraic exponents. Journal of Number Theory, 96(2):225–231, October 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X0292809X.

Li:2002:CCA

[389] Wen-Ching W. Li and Hiren Maharaj. Coverings of curves with asymptotically many rational points. Journal of Number Theory, 96(2):232– 256, October 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X02928040.

Bockle:2002:FCA

[390] Gebhard Böckle. Finiteness conjectures for F_{l}[[T]]-analytic extensions of number fields. Journal of Number Theory, 96(2):257–274, October 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X02927988.

Matala-aho:2002:IMS

[391] Tapani Matala-aho and Marc Prévost. Irrationality measures for the series of reciprocals from recurrence sequences. Journal of Number Theory, 96(2):275–292, October 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X02927952.

Browning:2002:EST

[392] T. D. Browning. Equal sums of two k-th powers. Journal of Number Theory, 96(2):293–318, October 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X02928003.

Neto:2002:CDS

[393] Trajano Pires da Nóbrega Neto, J. Carmelo Interlando, and José Othon Dantas Lopes. On computing discriminants of subfields of Q(ζ_{p^{r}}). Journal of Number Theory, 96(2):319–325, October 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X02927964.

Motoda:2002:PHC

[394] Yasuo Motoda, Toru Nakahara, and Syed Inayat Ali Shah. On a problem of Hasse for certain imaginary abelian fields. Journal of Number Theory, 96(2):326–334, October 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X02928052.

Cojocaru:2002:CGR

[395] Alina Carmen Cojocaru. On the cyclicity of the group of F_{p}-rational points of non–CM elliptic curves. Journal of Number Theory, 96(2):335– 350, October 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X02927897.

Yokota:2002:NIR

[396] Hisashi Yokota. On the number of integers representable as sums of unit fractions, III. Journal of Number Theory, 96(2):351–372, October 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X02927976.

Yue:2002:TKC

[397] Qin Yue. On tame kernel and class group in terms of quadratic forms. Journal of Number Theory, 96(2):373–387, October 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X02928088.

Coogan:2002:MWR

[398] Gwynneth Coogan and Jorge Jimenéz-Urroz. Mordell–Weil ranks of quadratic twists of pairs of elliptic curves. Journal of Number Theory, 96(2):388–399, October 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X02927885.

Lau:2002:EHE

[399] Yuk-Kam Lau. Equidistribution of Hecke eigenforms on the arithmetic surface Γ_{0}(N) \ H. Journal of Number Theory, 96(2):400–416, October 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X02927927.

Deutsch:2002:GNP

[400] Jesse Ira Deutsch. Geometry of numbers proof of Götzky’s Four-Squares Theorem. Journal of Number Theory, 96(2):417–431, October 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X02927794.

Baier:2002:BHC

[401] Stephan Baier. On the Bateman–Horn Conjecture. Journal of Number Theory, 96(2):432–448, October 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X02928118.

Anonymous:2002:AIVe

[402] Anonymous. Author index for volume 96. Journal of Number Theory, 96(2):449, October 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X02928623.

Anonymous:2002:PN

[403] Anonymous. Pages 1–198 (November 2002). Journal of Number Theory, 97(1):??, November 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Hou:2002:GAT

[404] Xiang-Dong Hou, Ka Hin Leung, and Qing Xiang. A generalization of an addition theorem of Kneser. Journal of Number Theory, 97(1): 1–9, November 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X02927939.

Scanlon:2002:DGT

[405] Thomas Scanlon. Diophantine geometry of the torsion of a Drinfeld module. Journal of Number Theory, 97(1):10–25, November 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X01927519.

Cilleruelo:2002:ULB

[406] Javier Cilleruelo, Imre Z. Ruzsa, and Carlos Trujillo. Upper and lower bounds for finite B_{h}[g] sequences. Journal of Number Theory, 97(1):26–34, November 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X01927672.

Yu:2002:NDC

[407] Gang Yu. A note on the divisibility of class numbers of real quadratic fields. Journal of Number Theory, 97(1):35–44, November 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X01927738.

Harman:2002:MTR

[408] Glyn Harman and Huw Jones. Metrical theorems on restricted diophantine approximations to points on a curve. Journal of Number Theory, 97(1):45–57, November 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X02927812.

Pagot:2002:EVF

[409] Guillaume Pagot. F_{p}-espaces vectoriels de formes différentielles logarithmiques sur la droite projective. (French) [F_{p}-vector spaces of logarithmic differential forms on the projective line]. Journal of Number Theory, 97(1):58–94, November 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X02928064.

Krizek:2002:CSR

[410] Michal Křížek, Florian Luca, and Lawrence Somer. On the convergence of series of reciprocals of primes related to the Fermat numbers. Journal of Number Theory, 97(1):95–112, November 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X02927824.

Buchholz:2002:TTR

[411] Ralph H. Buchholz. Triangles with three rational medians. Journal of Number Theory, 97(1):113–131, November 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X02927770.

Gekeler:2002:SNC

[412] Ernst-Ulrich Gekeler. A series of new congruences for Bernoulli numbers and Eisenstein series. Journal of Number Theory, 97(1):132–143, November 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X02928027.

Parsell:2002:ILF

[413] Scott T. Parsell. Irrational linear forms in prime variables. Journal of Number Theory, 97(1):144–156, November 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X02928076.

Kilford:2002:SNG

[414] L. J. P. Kilford. Some non–Gorenstein Hecke algebras attached to spaces of modular forms. Journal of Number Theory, 97(1):157–164, November 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X02928039.

Byeon:2002:NEC

[415] Dongho Byeon. A note on the existence of certain infinite families of imaginary quadratic fields. Journal of Number Theory, 97(1):165–170, November 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X02927873.

Agashe:2002:VST

[416] Amod Agashe and William Stein. Visibility of Shafarevich–Tate groups of abelian varieties. Journal of Number Theory, 97(1):171–185, November 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X02928106.

Chen:2002:GKI

[417] Beifang Chen and Zhi-Wei Sun. Generalizations of Knopp’s Identity. Journal of Number Theory, 97(1):186–198, November 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X02928015.

Anonymous:2002:PD

[418] Anonymous. Pages 199–472 (December 2002). Journal of Number Theory, 97(2):??, December 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2002:EB

[419] Anonymous. Editorial Board. Journal of Number Theory, 97(2):ifc, December 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X02000835.

Soverchia:2002:RIB

[420] Elena Soverchia. Relative integral bases over a Hilbert class field. Journal of Number Theory, 97(2):199–203, December 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X02000094.

Bertolin:2002:PMT

[421] Cristiana Bertolin. Périodes de 1-motifs et transcendance. (French) [Periods of 1-motifs and transcendence]. Journal of Number Theory, 97(2):204–221, December 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X02000021.

Wakabayashi:2002:CTI

[422] Isao Wakabayashi. Cubic Thue inequalities with negative discriminant. Journal of Number Theory, 97(2):222–251, December 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X02000100.

Chang:2002:NMF

[423] Mu-Ling Chang. Non-monogeneity in a family of sextic fields. Journal of Number Theory, 97(2):252–268, December 2002. CODEN JNUTA9.

ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X02000045.

Deolalikar:2002:DIR

[424] Vinay Deolalikar. Determining irreducibility and ramification groups for an additive extension of the rational function field. Journal of Number Theory, 97(2):269–286, December 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X02000124.

Nakada:2002:SMP

[425] Hitoshi Nakada and Rie Natsui. Some metric properties of αcontinued fractions. Journal of Number Theory, 97(2):287–300, December 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X02000082.

Liu:2002:SSP

[426] Jian-Xin Liu and Zhi-Wei Sun. Sums of subsets with polynomial restrictions. Journal of Number Theory, 97(2):301–304, December 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X02000070.

Feng:2002:PPN

[427] De-Jun Feng and Zhi-Ying Wen. A property of Pisot numbers. Journal of Number Theory, 97(2):305–316, December 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X02000136.

Papikian:2002:DMP

[428] Mihran Papikian. On the degree of modular parametrizations over function fields. Journal of Number Theory, 97(2):317–349, December 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X02000161.

Lundstrom:2002:SDN

[429] Patrik Lundström. Self-dual normal integral bases for infinite unramified extensions. Journal of Number Theory, 97(2):350–367, December 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X0200015X.

Hong:2002:NPF

[430] Shaofang Hong. Newton polygons for L-functions of exponential sums of polynomials of degree six over finite fields. Journal of Number Theory, 97(2):368–396, December 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X02000069.

Hall:2002:GWI

[431] R. R. Hall. Generalized Wirtinger inequalities, random matrix theory, and the zeros of the Riemann zeta-function. Journal of Number Theory, 97(2):397–409, December 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X02000057.

Barreira:2002:DFD

[432] L. Barreira, B. Saussol, and J. Schmeling. Distribution of frequencies of digits via multifractal analysis. Journal of Number Theory, 97(2):410–438, December 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X02000033.

Conrad:2002:WPB

[433] Keith Conrad. On Weil’s proof of the bound for Kloosterman sums. Journal of Number Theory, 97(2):439–446, December 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X02000112.

Gardeyn:2002:GCG

[434] Francis Gardeyn. A Galois criterion for good reduction of τ-sheaves. Journal of Number Theory, 97(2):447–471, December 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X02000148.

Anonymous:2002:VAI

[435] Anonymous. Volume author index. Journal of Number Theory, 97(2): 472, December 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X02000860.

Anonymous:2002:VC

[436] Anonymous. Volume Contents. Journal of Number Theory, 97(2):III– IV, December 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X02000896.

Anonymous:2002:TC

[437] Anonymous. Table of Contents. Journal of Number Theory, 97(2):obc, December 2002. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X02000914.

Anonymous:2003:PJa

[438] Anonymous. Pages 1–220 (January 2003). Journal of Number Theory, 98(1):??, January 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2003:EBa

[439] Anonymous. Editorial Board. Journal of Number Theory, 98(1):ifc, January 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X0200094X.

Anonymous:2003:EBb

[440] Anonymous. Editorial Board. Journal of Number Theory, 98(1):iii– iv, January 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X02001038.

Rynne:2003:SDA

[441] Bryan P. Rynne. Simultaneous Diophantine approximation on manifolds and Hausdorff dimension. Journal of Number Theory, 98(1):1– 9, January 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X02000355.

Ramare:2003:SEI

[442] Olivier Ramaré and Yannick Saouter. Short effective intervals containing primes. Journal of Number Theory, 98(1):10–33, January 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X0200029X.

Blomer:2003:TBO

[443] Valentin Blomer. Thin bases of order h. Journal of Number Theory, 98(1):34–46, January 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X02000367.

Hurlburt:2003:ZLD

[444] Chris Hurlburt. Zero loci of differential modular forms. Journal of Number Theory, 98(1):47–54, January 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X02000276.

Jung:2003:DMR

[445] Hwanyup Jung and Jaehyun Ahn. Demjanenko matrix and recursion formula for relative class number over function fields. Journal of Number Theory, 98(1):55–66, January 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X02000239.

Young:2003:BST

[446] Paul Thomas Young. On the behavior of some two-variable padic L-functions. Journal of Number Theory, 98(1):67–88, January 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X02000318.

Haynes:2003:NFF

[447] Alan Haynes. A note on Farey fractions with odd denominators. Journal of Number Theory, 98(1):89–104, January 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X02000264.

Fisher:2003:CTP

[448] Tom A. Fisher. The Cassels–Tate pairing and the Platonic solids. Journal of Number Theory, 98(1):105–155, January 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X02000380.

Ren:2003:SFC

[449] Xiumin Ren. Sums of four cubes of primes. Journal of Number Theory, 98(1):156–171, January 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X02000227.

Du:2003:GFD

[450] Bau-Sen Du, Sen-Shan Huang, and Ming-Chia Li. Generalized Fermat, double Fermat and Newton sequences. Journal of Number Theory, 98(1):172–183, January 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X02000252.

Croot:2003:OMF

[451] Ernest S. Croot. On the oscillations of multiplicative functions taking values ±1. Journal of Number Theory, 98(1):184–194, January 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X02000240.

Urbanowicz:2003:LCR

[452] Jerzy Urbanowicz and Paul van Wamelen. On linear congruence relations for Kubota–Leopoldt 2-adic L-functions. Journal of Number Theory, 98(1):195–216, January 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X02000306.

Maire:2003:ARA

[453] Christian Maire. Addendum: On the Z_{ℓ}-rank of abelian extensions with restricted ramification [Journal of Number Theory 92 (2002) 376–404]. Journal of Number Theory, 98(1):217–220, January 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X02000288. See [311].

Anonymous:2003:PF

[454] Anonymous. Pages 221–438 (February 2003). Journal of Number Theory, 98(2):??, February 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2003:EBc

[455] Anonymous. Editorial Board. Journal of Number Theory, 98(2):ifc, February 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X02001087.

Peng:2003:GFK

[456] Guohua Peng. The genus fields of Kummer function fields. Journal of Number Theory, 98(2):221–227, February 2003. CODEN JNUTA9.

ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X02000422.

Bennett:2003:PCR

[457] Michael A. Bennett. Pillai’s conjecture revisited. Journal of Number Theory, 98(2):228–235, February 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X02000495.

Fukuhara:2003:DSA

[458] Shinji Fukuhara. Dedekind symbols associated with J-forms and their reciprocity law. Journal of Number Theory, 98(2):236–253, February 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X02000392.

Bauer:2003:VDA

[459] Hartmut Bauer. The value distribution of Artin L-series and zeros of Zeta-functions. Journal of Number Theory, 98(2):254–279, February 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X02000483.

Ota:2003:DDS

[460] Kaori Ota. Derivatives of Dedekind sums and their reciprocity law. Journal of Number Theory, 98(2):280–309, February 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X0200046X.

Chen:2003:IF

[461] Yong-Gao Chen. On integers of the forms k^{r} − 2^{n} and k^{r}2^{n} + 1. Journal of Number Theory, 98(2):310–319, February 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X02000513.

Godin:2003:RCT

[462] M. Godin and B. Sodaïgui. Realizable classes of tetrahedral extensions. Journal of Number Theory, 98(2):320–328, February 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X02000409.

Khare:2003:MGR

[463] Chandrashekhar Khare and Ian Kiming. Mod pq Galois representations and Serre’s conjecture. Journal of Number Theory, 98(2):329– 347, February 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X02000434.

Marcolli:2003:LMS

[464] Matilde Marcolli. Limiting modular symbols and the Lyapunov spectrum. Journal of Number Theory, 98(2):348–376, February 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X02000446.

Boylan:2003:ECC

[465] Matthew Boylan. Exceptional congruences for the coefficients of certain eta-product newforms. Journal of Number Theory, 98(2):377–389, February 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X02000379.

Osburn:2003:TKF

[466] Robert Osburn and Brian Murray. Tame kernels and further 4rank densities. Journal of Number Theory, 98(2):390–406, February 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X02000458.

Papanikolas:2003:WBF

[467] Matthew A. Papanikolas and Niranjan Ramachandran. A Weil–Barsotti formula for Drinfeld modules. Journal of Number Theory, 98(2):407– 431, February 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X02000471.

Byeon:2003:ECF

[468] Dongho Byeon. Existence of certain fundamental discriminants and class numbers of real quadratic fields. Journal of Number Theory, 98(2):432–437, February 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X02000501.

Anonymous:2003:VAIa

[469] Anonymous. Volume author index. Journal of Number Theory, 98(2): 438, February 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X02001117.

Anonymous:2003:PMa

[470] Anonymous. Pages 1–224 (March 2003). Journal of Number Theory, 99 (1):??, March 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2003:EBd

[471] Anonymous. Editorial Board. Journal of Number Theory, 99(1): ifc, March 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X03000039.

Knopp:2003:GMF

[472] Marvin Knopp and Geoffrey Mason. Generalized modular forms. Journal of Number Theory, 99(1):1–28, March 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X02000653.

Sitaraman:2003:NFT

[473] Sankar Sitaraman. Note on a Fermat-type diophantine equation. Journal of Number Theory, 99(1):29–35, March 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X02000550.

Wood:2003:OMV

[474] Melanie Wood. P -orderings: a metric viewpoint and the non-existence of simultaneous orderings. Journal of Number Theory, 99(1):36–56, March 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X02000562.

Kahn:2003:SFR

[475] Bruno Kahn. Some finiteness results for étale cohomology. Journal of Number Theory, 99(1):57–73, March 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X02000641.

Bushnell:2003:EUB

[476] Colin J. Bushnell and Guy Henniart. Explicit unramified base change: GL(p) of a p-adic field. Journal of Number Theory, 99(1):74–89, March 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X02000628.

Bruggeman:2003:SKS

[477] R. W. Bruggeman, R. J. Miatello, and I. Pacharoni. Sums of Kloosterman sums for real quadratic number fields. Journal of Number Theory, 99(1):90–119, March 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X02000604.

Yamagami:2003:GCU

[478] Atsushi Yamagami. On Gouvêa’s conjecture in the unobstructed case. Journal of Number Theory, 99(1):120–138, March 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X02000574.

Mortenson:2003:SCR

[479] Eric Mortenson. A supercongruence conjecture of Rodriguez-Villegas for a certain truncated hypergeometric function. Journal of Number Theory, 99(1):139–147, March 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X02000525.

Kloosterman:2003:SGE

[480] Remke Kloosterman and Edward F. Schaefer. Selmer groups of elliptic curves that can be arbitrarily large. Journal of Number Theory, 99(1):148–163, March 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X02000549.

Bruinier:2003:CHI

[481] Jan H. Bruinier and Ken Ono. Coefficients of half-integral weight modular forms. Journal of Number Theory, 99(1):164–179, March 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X02000616. See corrigendum [628].

Prasad:2003:AAT

[482] Dipendra Prasad and C. S. Rajan. On an Archimedean analogue of Tate’s conjecture. Journal of Number Theory, 99(1):180–184, March 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X02000537.

Kable:2003:MVP

[483] Anthony C. Kable and Akihiko Yukie. The mean value of the product of class numbers of paired quadratic fields III. Journal of Number Theory, 99(1):185–218, March 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X02000586.

Byeon:2003:FCR

[484] Dongho Byeon and H. M. Stark. On the finiteness of certain Rabinowitsch polynomials II. Journal of Number Theory, 99(1):219– 221, March 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X0200063X.

Albu:2003:CAM

[485] Toma Albu. Corrigendum and addendum to my paper concerning Kummer extensions with few roots of unity. Journal of Number Theory, 99(1):222–224, March 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X02000598.

Anonymous:2003:PAa

[486] Anonymous. Pages 225–444 (April 2003). Journal of Number Theory, 99(2):??, April 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2003:EBe

[487] Anonymous. Editorial Board. Journal of Number Theory, 99(2): ifc, April 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X03000246.

Mihailescu:2003:CNF

[488] Preda Mihăilescu. A class number free criterion for Catalan’s conjecture. Journal of Number Theory, 99(2):225–231, April 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X02001014.

Summerer:2003:DFA

[489] Leonhard Summerer. Decomposable forms and automorphisms. Journal of Number Theory, 99(2):232–254, April 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X02000732.

Ito:2003:SA

[490] Shunji Ito, Junko Fujii, Hiroko Higashino, and Shin ichi Yasutomi. On simultaneous approximation to (α, α^{2}) with α^{3} + kα − 1 = 0. Journal of Number Theory, 99(2):255–283, April 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X02000768.

Kuhnlein:2003:TC

[491] Stefan Kühnlein. On torsion in the cohomology of PSL_{2}(Z[1/N]). Journal of Number Theory, 99(2):284–297, April 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X0200077X.

Yamazaki:2003:TDJ

[492] Takao Yamazaki. On Tate duality for Jacobian varieties. Journal of Number Theory, 99(2):298–306, April 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X02000677.

Nebe:2003:ULL

[493] Gabriele Nebe and Boris Venkov. Unimodular lattices with long shadow. Journal of Number Theory, 99(2):307–317, April 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X02000793.

Fieker:2003:MDF

[494] Claus Fieker and Jürgen Klüners. Minimal discriminants for fields with small Frobenius groups as Galois groups. Journal of Number Theory, 99(2):318–337, April 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X02000719.

Simsek:2003:RBT

[495] Yilmaz Simsek. Relations between theta-functions Hardy sums Eisenstein and Lambert series in the transformation formula of log η_{g,h}(z). Journal of Number Theory, 99(2):338–360, April 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X02000720.

Chan:2003:SML

[496] Heng Huat Chan, Kok Seng Chua, and Patrick Solé. Seven-modular lattices and a septic base Jacobi identity. Journal of Number Theory, 99(2):361–372, April 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X02000690.

Cossidente:2003:PAC

[497] A. Cossidente and A. Siciliano. Plane algebraic curves with Singer automorphisms. Journal of Number Theory, 99(2):373–382, April 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X02000707.

Moret-Bailly:2003:ETS

[498] Laurent Moret-Bailly. Sur la R-équivalence de torseurs sous un groupe fini. (French) [On the R-equivalence of torsors under a finite group]. Journal of Number Theory, 99(2):383–404, April 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X02000665.

Biro:2003:SCS

[499] András Biró and Vera T. Sós. Strong characterizing sequences in simultaneous diophantine approximation. Journal of Number Theory, 99(2): 405–414, April 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X02000689.

Matsuno:2003:FSS

[500] Kazuo Matsuno. Finite Λ-submodules of Selmer groups of abelian varieties over cyclotomic Z_{p}-extensions. Journal of Number Theory, 99(2): 415–443, April 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X02000781.

Anonymous:2003:VAIb

[501] Anonymous. Volume author index. Journal of Number Theory, 99(2): 444, April 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X03000271.

Anonymous:2003:PMb

[502] Anonymous. Pages 1–202 (May 2003). Journal of Number Theory, 100 (1):??, May 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2003:EBf

[503] Anonymous. Editorial Board. Journal of Number Theory, 100(1): ifc, May 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X03000465.

Graham:2003:ACP

[504] Ronald L. Graham, Jeffrey C. Lagarias, Colin L. Mallows, Allan R. Wilks, and Catherine H. Yan. Apollonian circle packings: number theory. Journal of Number Theory, 100(1):1–45, May 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X03000155.

Gunnells:2003:HOG

[505] Paul E. Gunnells and Mark McConnell. Hecke operators and Q-groups associated to self-adjoint homogeneous cones. Journal of Number Theory, 100(1):46–71, May 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X02000410.

Baker:2003:A

[506] Matthew Baker and Yuji Hasegawa. Automorphisms of X_{0}^{∗}(p). Journal of Number Theory, 100(1):72–87, May 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X02001208.

Yao:2003:PRC

[507] Wei-Chen Yao and Jing Yu. On primitive roots for Carlitz modules. Journal of Number Theory, 100(1):88–103, May 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X0200118X.

Bachman:2003:CTC

[508] Gennady Bachman. On the coefficients of ternary cyclotomic polynomials. Journal of Number Theory, 100(1):104–116, May 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X02001191.

Banaszak:2003:GRA

[509] G. Banaszak, W. Gajda, and P. Krasoń. On Galois representations for abelian varieties with complex and real multiplications. Journal of Number Theory, 100(1):117–132, May 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X0200121X.

Banaszak:2003:SPI

[510] G. Banaszak, W. Gajda, and P. Krasoń. Support problem for the intermediate Jacobians of l-adic representations. Journal of Number Theory, 100(1):133–168, May 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X02001221.

Vaughan:2003:MVT

[511] R. C. Vaughan. A mean value theorem for cubic fields. Journal of Number Theory, 100(1):169–183, May 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X02000756.

Vandervelde:2003:FMM

[512] S. Vandervelde. A formula for the Mahler measure of axy + bx + cy + d. Journal of Number Theory, 100(1):184–202, May 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X02000744.

Anonymous:2003:PJb

[513] Anonymous. Pages 203–396 (June 2003). Journal of Number Theory, 100 (2):??, June 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2003:EBg

[514] Anonymous. Editorial Board. Journal of Number Theory, 100(2): ifc, June 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X03000702.

Stanica:2003:CNS

[515] Pantelimon Stănică. p^{q}-Catalan numbers and squarefree binomial coefficients. Journal of Number Theory, 100(2):203–216, June 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X02001257.

Fleckinger:2003:IAC

[516] Vincent Fleckinger and Caroline Thiébaud. Idéaux ambiges dans les corps de fonctions. (French) [Ambiguous ideals in function fields]. Journal of Number Theory, 100(2):217–228, June 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X02001026.

Filaseta:2003:ICC

[517] Michael Filaseta and Richard L. Williams. On the irreducibility of a certain class of Laguerre polynomials. Journal of Number Theory, 100(2):229–250, June 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X02001294.

Hakim:2003:SGP

[518] Jeffrey Hakim. Supercuspidal Gelfand pairs. Journal of Number Theory, 100(2):251–269, June 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X02001312.

Baruah:2003:SRS

[519] Nayandeep Deka Baruah. On some of Ramanujan’s Schläfli-type “mixed” modular equations. Journal of Number Theory, 100(2):270– 294, June 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X02001270.

Yamagishi:2003:BMW

[520] Hizuru Yamagishi. Boundedness of Mordell–Weil ranks of certain elliptic curves and Lang’s conjecture. Journal of Number Theory, 100(2): 295–306, June 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X02001269.

Arenas:2003:THO

[521] A. Arenas. On the traces of Hecke operators. Journal of Number Theory, 100(2):307–312, June 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X03000143.

Bush:2003:CGG

[522] M. R. Bush. Computation of Galois groups associated to the 2-class towers of some quadratic fields. Journal of Number Theory, 100(2): 313–325, June 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X02001282.

Chen:2003:PES

[523] Yong-Gao Chen. On the parity of exponents in the standard factorization of n!. Journal of Number Theory, 100(2):326–331, June 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X03000131.

Gardeyn:2003:SAS

[524] Francis Gardeyn. The structure of analytic τ-sheaves. Journal of Number Theory, 100(2):332–362, June 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X02001300.

Poor:2003:SIL

[525] Cris Poor and David S. Yuen. Slopes of integral lattices. Journal of Number Theory, 100(2):363–380, June 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X02001245.

Meijer:2003:EGS

[526] P. Meijer and M. van der Vlugt. The evaluation of Gauss sums for characters of 2-power order. Journal of Number Theory, 100(2):381– 395, June 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X02001233.

Anonymous:2003:VAIc

[527] Anonymous. Volume author index. Journal of Number Theory, 100 (2):396, June 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X03000738.

Anonymous:2003:PJc

[528] Anonymous. Pages 1–208 (July 2003). Journal of Number Theory, 101 (1):??, July 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2003:EBh

[529] Anonymous. Editorial Board. Journal of Number Theory, 101(1): ifc, July 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X03000878.

Kimura:2003:EU

[530] Kenichiro Kimura. Elliptic units in K_{2}. Journal of Number Theory, 101(1):1–12, July 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X03000210.

Hsu:2003:PRL

[531] Chih-Nung Hsu. On polynomial reciprocity law. Journal of Number Theory, 101(1):13–31, July 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X03000209.

Gao:2003:IPM

[532] Shuhong Gao and Virgínia M. Rodrigues. Irreducibility of polynomials modulo p via Newton polytopes. Journal of Number Theory, 101(1): 32–47, July 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X03000441.

Vidaux:2003:AHP

[533] X. Vidaux. An analogue of Hilbert’s 10th problem for fields of meromorphic functions over non–Archimedean valued fields. Journal of Number Theory, 101(1):48–73, July 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X03000167.

Bondarko:2003:LBA

[534] M. V. Bondarko. Links between associated additive Galois modules and computation of H^{1} for local formal group modules. Journal of Number Theory, 101(1):74–104, July 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X03000192.

Fuchs:2003:ATS

[535] Michael Fuchs. An analogue of a theorem of Szüsz for formal Laurent series over finite fields. Journal of Number Theory, 101(1):105– 130, July 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X0300012X.

Kawamoto:2003:NIB

[536] Fuminori Kawamoto. Normal integral bases and strict ray class groups modulo 4. Journal of Number Theory, 101(1):131–137, July 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X03000106.

Yoshida:2003:FTC

[537] Teruyoshi Yoshida. Finiteness theorems in the class field theory of varieties over local fields. Journal of Number Theory, 101(1):138– 150, July 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X03000180.

Parimala:2003:HPC

[538] R. Parimala and R. Preeti. Hasse principle for classical groups over function fields of curves over number fields. Journal of Number Theory, 101(1):151–184, July 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X0300009X.

Quebbemann:2003:IFC

[539] H.-G. Quebbemann and E. M. Rains. On the involutions fixing the class of a lattice. Journal of Number Theory, 101(1):185–194, July 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X03000222.

Broberg:2003:RPF

[540] Niklas Broberg. Rational points on finite covers of P^{1} and P^{2}. Journal of Number Theory, 101(1):195–207, July 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X0300043X.

Anonymous:2003:PAb

[541] Anonymous. Pages 209–424 (August 2003). Journal of Number Theory, 101(2):??, August 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2003:EBi

[542] Anonymous. Editorial Board. Journal of Number Theory, 101(2): ifc, August 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X0300115X.

Cobeli:2003:DPI

[543] C. I. Cobeli, S. M. Gonek, and A. Zaharescu. The distribution of patterns of inverses modulo a prime. Journal of Number Theory, 101(2):209–222, August 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X03000593.

Matsumoto:2003:ACA

[544] Kohji Matsumoto. The analytic continuation and the asymptotic behaviour of certain multiple zeta-functions I. Journal of Number Theory, 101(2):223–243, August 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X03000416.

Archinard:2003:ESH

[545] Natália Archinard. Exceptional sets of hypergeometric series. Journal of Number Theory, 101(2):244–269, August 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X03000428.

Aoki:2003:LUG

[546] Miho Aoki. Local units and Gauss sums. Journal of Number Theory, 101(2):270–293, August 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X03000568.

Thunder:2003:VDI

[547] Jeffrey Lin Thunder. Volumes and diophantine inequalities associated with decomposable forms. Journal of Number Theory, 101(2):294– 309, August 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X03000398.

Bockle:2003:LRC

[548] Gebhard Böckle. Lifting modp representations to characteristics p^{2}. Journal of Number Theory, 101(2):310–337, August 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X03000581.

Eliahou:2003:OSS

[549] Shalom Eliahou, Michel Kervaire, and Alain Plagne. Optimally small sumsets in finite abelian groups. Journal of Number Theory, 101(2):338– 348, August 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X0300060X.

Kataoka:2003:DPI

[550] Norisato Kataoka. The distribution of prime ideals in a real quadratic field with units having a given index in the residue class field. Journal of Number Theory, 101(2):349–375, August 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X03000623.

Flynn:2003:ACT

[551] E. V. Flynn and J. Redmond. Application of covering techniques to families of curves. Journal of Number Theory, 101(2):376–397, August 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X03000611.

Larsen:2003:SPA

[552] Michael Larsen. The support problem for abelian varieties. Journal of Number Theory, 101(2):398–403, August 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X03000404.

Alkan:2003:EFA

[553] Emre Alkan. On the enumeration of finite Abelian and solvable groups. Journal of Number Theory, 101(2):404–423, August 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X03000556.

Anonymous:2003:VAId

[554] Anonymous. Volume author index. Journal of Number Theory, 101(2): 424, August 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X03001185.

Anonymous:2003:PS

[555] Anonymous. Pages 1–190 (September 2003). Journal of Number Theory, 102(1):??, September 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2003:EBj

[556] Anonymous. Editorial Board. Journal of Number Theory, 102(1):ifc, September 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X03001264.

Yuan:2003:AAC

[557] Pingzhi Yuan. On algebraic approximations of certain algebraic numbers. Journal of Number Theory, 102(1):1–10, September 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X03000684.

Luczak:2003:SFS

[558] Tomasz Łuczak and Tomasz Schoen. Solution-free sets for linear equations. Journal of Number Theory, 102(1):11–22, September 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X03000647.

Mauldin:2003:DPC

[559] R. Daniel Mauldin and Mariusz Urbański. The doubling property of conformal measures of infinite iterated function systems. Journal of Number Theory, 102(1):23–40, September 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X03000659.

Sun:2003:CQC

[560] Zhi-Hong Sun. Cubic and quartic congruences modulo a prime. Journal of Number Theory, 102(1):41–89, September 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X03000672. See

erratum [629].

Kishi:2003:FCC

[561] Yasuhiro Kishi. A family of cyclic cubic polynomials whose roots are systems of fundamental units. Journal of Number Theory, 102 (1):90–106, September 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X03000854. See erratum [589].

Gallardo:2003:WCC

[562] José Gallardo and Bartolomé López. “Weak” congruences for coefficients of the Eisenstein series for F_{q}[T] of weight q^{k} − 1. Journal of Number Theory, 102(1):107–117, September 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X03000635.

Aiba:2003:ASE

[563] Akira Aiba. Artin–Schreier extensions and Galois module structure. Journal of Number Theory, 102(1):118–124, September 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X03000830.

Beli:2003:ISN

[564] Constantin N. Beli. Integral spinor norm groups over dyadic local fields. Journal of Number Theory, 102(1):125–182, September 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X0300057X.

Rajan:2003:SMO

[565] C. S. Rajan. On strong multiplicity one for automorphic representations. Journal of Number Theory, 102(1):183–190, September 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X03000660.

Anonymous:2003:PO

[566] Anonymous. Pages 191–384 (October 2003). Journal of Number Theory, 102(2):??, October 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2003:EBk

[567] Anonymous. Editorial Board. Journal of Number Theory, 102(2): ifc, October 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X03001331.

Drinen:2003:IIE

[568] Michael J. Drinen. Iwasawa μ-invariants of elliptic curves and their symmetric powers. Journal of Number Theory, 102(2):191–213, October 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X03001057.

Jeon:2003:DMC

[569] Daeyeol Jeon and Chang Heon Kim. On the Drinfeld modular curves X_{1}(n). Journal of Number Theory, 102(2):214–222, October 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X03001008.

Ibukiyama:2003:EFK

[570] Tomoyoshi Ibukiyama and Hidenori Katsurada. An explicit formula for Koecher–Maaß Dirichlet series for Eisenstein series of Klingen type. Journal of Number Theory, 102(2):223–256, October 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X03000994.

Stern:2003:ENG

[571] Leonid Stern. Equality of norm groups of subextensions of S_{n}(n ≤ 5) extensions of algebraic number fields. Journal of Number Theory, 102(2):257–277, October 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X03000829.

Rhodes:2003:RTH

[572] John A. Rhodes and Thomas R. Shemanske. Rationality theorems for Hecke operators on GL_{n}. Journal of Number Theory, 102(2):278– 297, October 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X03000817.

Luca:2003:PPF

[573] Florian Luca and Pantelimon Stănică. On the prime power factorization of n!. Journal of Number Theory, 102(2):298–305, October 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X03001021.

Gardeyn:2003:OGI

[574] Francis Gardeyn. Openness of the Galois image for τ-modules of dimension 1. Journal of Number Theory, 102(2):306–338, October 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X03001069.

Grekos:2003:ETC

[575] G. Grekos, L. Haddad, C. Helou, and J. Pihko. On the Erdős– Turán conjecture. Journal of Number Theory, 102(2):339–352, October 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X03001082.

Krausshar:2003:MCA

[576] Rolf Sören Kraußhar. The multiplication of the Clifford-analytic Eisenstein series. Journal of Number Theory, 102(2):353–382, October 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X03000805.

Anonymous:2003:VAIe

[577] Anonymous. Volume author index. Journal of Number Theory, 102(2): 383, October 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X03001367.

Anonymous:2003:PN

[578] Anonymous. Pages 1–134 (November 2003). Journal of Number Theory, 103(1):??, November 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2003:EBl

[579] Anonymous. Editorial Board. Journal of Number Theory, 103(1):ifc, November 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X03001756.

Strauch:2003:KUD

[580] Oto Strauch, Milan Paštéka, and Georges Grekos. Kloosterman’s uniformly distributed sequence. Journal of Number Theory, 103(1):1–15, November 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X03000118.

Wu:2003:HMP

[581] Jun Wu. How many points have the same Engel and Sylvester expansions? Journal of Number Theory, 103(1):16–26, November 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X03000179.

Adamczewski:2003:TRN

[582] Boris Adamczewski and Julien Cassaigne. On the transcendence of real numbers with a regular expansion. Journal of Number Theory, 103(1):27–37, November 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X03000544.

Benjamin:2003:IQF

[583] E. Benjamin, F. Lemmermeyer, and C. Snyder. Imaginary quadratic fields with Cl_{2}(k) ≃ (2, 2, 2). Journal of Number Theory, 103(1):38–70, November 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X03000842.

Briggs:2003:SEB

[584] Keith Briggs. Some explicit badly approximable pairs. Journal of Number Theory, 103(1):71–76, November 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X03001045.

Griess:2003:PDL

[585] Robert L. Griess. Positive definite lattices of rank at most 8. Journal of Number Theory, 103(1):77–84, November 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X03001070.

Lalin:2003:SEM

[586] Matilde N. Lalín. Some examples of Mahler measures as multiple polylogarithms. Journal of Number Theory, 103(1):85–108, November 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X0300101X.

Moshe:2003:DRD

[587] Yossi Moshe. The density of 0’s in recurrence double sequences. Journal of Number Theory, 103(1):109–121, November 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X03001033.

Pappalardi:2003:SFV

[588] Francesco Pappalardi, Filip Saidak, and Igor E. Shparlinski. Squarefree values of the Carmichael function. Journal of Number Theory, 103(1):122–131, November 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X03001100.

Kishi:2003:EFC

[589] Yasuhiro Kishi. Erratum to “A family of cyclic cubic polynomials whose roots are systems of fundamental units” [J. Number Theory 102 (2003) 90–106]. Journal of Number Theory, 103(1):132–133, November 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X03001641. See [561].

Anonymous:2003:PD

[590] Anonymous. Pages 135–296 (December 2003). Journal of Number Theory, 103(2):??, December 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2003:EBm

[591] Anonymous. Editorial Board. Journal of Number Theory, 103(2):ifc, December 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X03001938.

Cooper:2003:NRC

[592] Shaun Cooper. On the number of representations of certain integers as sums of 11 or 13 squares. Journal of Number Theory, 103 (2):135–162, December 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X03001446.

Stopple:2003:SCC

[593] Jeffrey Stopple. Stark conjectures for CM elliptic curves over number fields. Journal of Number Theory, 103(2):163–196, December 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X03001124.

Zhang:2003:PDH

[594] Wenpeng Zhang, Xu Zongben, and Yi Yuan. A problem of D. H. Lehmer and its mean square value formula. Journal of Number Theory, 103(2):197–213, December 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X03001136.

Nathanson:2003:FEA

[595] Melvyn B. Nathanson. A functional equation arising from multiplication of quantum integers. Journal of Number Theory, 103(2):214–233, December 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X03001094.

Rosen:2003:FDM

[596] Michael Rosen. Formal Drinfeld modules. Journal of Number Theory, 103(2):234–256, December 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X03001112.

Hummel:2003:CQN

[597] Patrick Hummel. On consecutive quadratic non-residues: a conjecture of Issai Schur. Journal of Number Theory, 103(2):257–266, December 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X0300146X.

Konyagin:2003:SCT

[598] Sergei Konyagin and Izabella Łaba. Spectra of certain types of polynomials and tiling of integers with translates of finite sets. Journal of Number Theory, 103(2):267–280, December 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X03001471.

Rosales:2003:PMD

[599] J. C. Rosales, P. A. García-Sánchez, J. I. García-García, and J. M. Urbano-Blanco. Proportionally modular diophantine inequalities. Journal of Number Theory, 103(2):281–294, December 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X03001458.

Anonymous:2003:VAIf

[600] Anonymous. Volume author index. Journal of Number Theory, 103(2): 295, December 2003. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X03001963.

Anonymous:2004:PJa

[601] Anonymous. Pages 1–192 (January 2004). Journal of Number Theory, 104(1):??, January 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2004:EBa

[602] Anonymous. Editorial Board. Journal of Number Theory, 104(1): ifc, January 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X03002063.

Ryan:2004:WKE

[603] Dermot Ryan. The wild kernel of exceptional number fields. Journal of Number Theory, 104(1):1–13, January 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X03001616.

Hou:2004:EIC

[604] Xiang-Dong Hou and Kevin Keating. Enumeration of isomorphism classes of extensions of p-adic fields. Journal of Number Theory, 104(1): 14–61, January 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X03001550. See corrigendum [1017].

Mozzochi:2004:PCE

[605] C. J. Mozzochi. On the pair correlation of the eigenvalues of the hyperbolic Laplacian for PSL(2, Z) \ H. Journal of Number Theory, 104(1):62–74, January 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X03001501.

Szydlo:2004:EFN

[606] Michael Szydlo. Elliptic fibers over non-perfect residue fields. Journal of Number Theory, 104(1):75–99, January 2004. CODEN JNUTA9.

ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X03001537.

Katsurada:2004:SAP

[607] H. Katsurada and S. Nagaoka. On some p-adic properties of Siegel– Eisenstein series. Journal of Number Theory, 104(1):100–117, January 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X03001604.

Louboutin:2004:EUB

[608] Stéphane R. Louboutin. Explicit upper bounds for values at s =1 of Dirichlet L-series associated with primitive even characters. Journal of Number Theory, 104(1):118–131, January 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X03001598.

Amou:2004:IRV

[609] Masaaki Amou and Masanori Katsurada. Irrationality results for values of generalized Tschakaloff series II. Journal of Number Theory, 104(1):132–155, January 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X03001434.

Wenpeng:2004:GKS

[610] Zhang Wenpeng. On the general Kloosterman sum and its fourth power mean. Journal of Number Theory, 104(1):156–161, January 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X03001549.

Lev:2004:PFS

[611] Vsevolod F. Lev. Progression-free sets in finite abelian groups. Journal of Number Theory, 104(1):162–169, January 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X03001483.

Ichikawa:2004:HSA

[612] Takashi Ichikawa. Heights on a subvariety of an abelian variety. Journal of Number Theory, 104(1):170–176, January 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X03001562.

Li:2004:KTF

[613] Charles C. C. Li. Kuznietsov trace formula and weighted distribution of Hecke eigenvalues. Journal of Number Theory, 104(1):177– 192, January 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X03001495.

Anonymous:2004:PF

[614] Anonymous. Pages 193–382 (February 2004). Journal of Number Theory, 104(2):??, February 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2004:EBb

[615] Anonymous. Editorial Board. Journal of Number Theory, 104(2):ifc, February 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X03002324.

vanderHeiden:2004:LGP

[616] Gert-Jan van der Heiden. Local–global problem for Drinfeld modules. Journal of Number Theory, 104(2):193–209, February 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X0300163X.

Laurent:2004:GDB

[617] Michel Laurent and Dimitrios Poulakis. On the global distance between two algebraic points on a curve. Journal of Number Theory, 104(2):210–254, February 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X03001653.

Deshouillers:2004:WSS

[618] Jean-Marc Deshouillers and Gregory A. Freiman. When subset-sums do not cover all the residues modulo p. Journal of Number Theory, 104(2):255–262, February 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X03001677.

Deutsch:2004:APC

[619] Jesse Ira Deutsch. An alternate proof of Cohn’s four squares theorem. Journal of Number Theory, 104(2):263–278, February 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X03001689.

Kisilevsky:2004:RDS

[620] H. Kisilevsky. Rank determines semi-stable conductor. Journal of Number Theory, 104(2):279–286, February 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X03001574.

Kisilevsky:2004:DFT

[621] H. Kisilevsky. Determining L-functions by twists. Journal of Number Theory, 104(2):287–300, February 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X03001586.

Pillichshammer:2004:DS

[622] Friedrich Pillichshammer. On the discrepancy of (0, 1)-sequences. Journal of Number Theory, 104(2):301–314, February 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X03001665.

Breuer:2004:HHP

[623] Florian Breuer. Higher Heegner points on elliptic curves over function fields. Journal of Number Theory, 104(2):315–326, February 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X03001847.

Balasubramanian:2004:ARF

[624] R. Balasubramanian and Gyan Prakash. On an additive representation function. Journal of Number Theory, 104(2):327–334, February 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X0300180X.

Wan:2004:ARH

[625] Daqing Wan and C. Douglas Haessig. On the p-adic Riemann hypothesis for the zeta function of divisors. Journal of Number Theory, 104(2):335–352, February 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X03001690.

Silverman:2004:LBC

[626] Joseph H. Silverman. A lower bound for the canonical height on elliptic curves over abelian extensions. Journal of Number Theory, 104(2):353–372, February 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X03001628.

Kondo:2004:KLF

[627] Satoshi Kondo. Kronecker limit formula for Drinfeld modules. Journal of Number Theory, 104(2):373–377, February 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X03001707.

Bruinier:2004:CCH

[628] Jan H. Bruinier and Ken Ono. Corrigendum to “Coefficients of half-integral weight modular forms”: [J. Number Theory 99 (2003) 164–179]. Journal of Number Theory, 104(2):378–379, February 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X03001720. See [481].

Sun:2004:ECQ

[629] Zhi-Hong Sun. Erratum to “Cubic and quartic congruences modulo a prime”: [J. Number Theory 102 (2003) 41–89]. Journal of Number Theory, 104(2):380, February 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X03001719. See [560].

Anonymous:2004:AI

[630] Anonymous. Author Index. Journal of Number Theory, 104(2):381, February 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X0300235X.

Anonymous:2004:PMa

[631] Anonymous. Pages 1–202 (March 2004). Journal of Number Theory, 105 (1):??, March 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2004:IFCa

[632] Anonymous. Inside front cover: Editorial Board. Journal of Number Theory, 105(1):ifc, March 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X04000034.

Bley:2004:CMF

[633] Werner Bley and Robert Boltje. Cohomological Mackey functors in number theory. Journal of Number Theory, 105(1):1–37, March 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X03001835.

Tanaka:2004:AIV

[634] Taka aki Tanaka. Algebraic independence of the values of Mahler functions associated with a certain continued fraction expansion. Journal of Number Theory, 105(1):38–48, March 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X03001902.

Nyblom:2004:ERS

[635] M. A. Nyblom. An extension of a result of Sierpiński. Journal of Number Theory, 105(1):49–59, March 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X03001896.

Chinen:2004:DPR

[636] Koji Chinen and Leo Murata. On a distribution property of the residual order of a(modp). Journal of Number Theory, 105(1):60– 81, March 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X03001513.

Murata:2004:DPR

[637] Leo Murata and Koji Chinen. On a distribution property of the residual order of a(modp) — II. Journal of Number Theory, 105(1):82– 100, March 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X03001525.

Liebendorfer:2004:LEH

[638] Christine Liebendörfer. Linear equations and heights over division algebras. Journal of Number Theory, 105(1):101–133, March 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X03001872.

Mizumoto:2004:CSA

[639] Shin ichiro Mizumoto. Certain series attached to an even number of elliptic modular forms. Journal of Number Theory, 105(1):134– 149, March 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X03001884.

Beineke:2004:MRZa

[640] Jennifer Beineke and Daniel Bump. Moments of the Riemann zeta function and Eisenstein series — I. Journal of Number Theory, 105(1):150– 174, March 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X03001811.

Beineke:2004:MRZb

[641] Jennifer Beineke and Daniel Bump. Moments of the Riemann zeta function and Eisenstein series — II. Journal of Number Theory, 105(1): 175–191, March 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X03001823.

Zhang:2004:PFP

[642] Qifan Zhang. Polynomial functions and permutation polynomials over some finite commutative rings. Journal of Number Theory, 105(1):192– 202, March 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X03001914.

Anonymous:2004:PAa

[643] Anonymous. Pages 203–402 (April 2004). Journal of Number Theory, 105(2):??, April 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2004:IFCb

[644] Anonymous. Inside front cover: Editorial Board. Journal of Number Theory, 105(2):ifc, April 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X04000320.

Mizusawa:2004:MUP

[645] Yasushi Mizusawa. On the maximal unramified pro-2-extension of Z_{2}- extensions of certain real quadratic fields. Journal of Number Theory, 105(2):203–211, April 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X03002178.

Scott:2004:RTT

[646] Reese Scott and Robert Styer. On p^{x} − q^{y} = c and related three term exponential Diophantine equations with prime bases. Journal of Number Theory, 105(2):212–234, April 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X03002233.

Sumida-Takahashi:2004:CII

[647] Hiroki Sumida-Takahashi. Computation of the Iwasawa invariants of certain real abelian fields. Journal of Number Theory, 105(2):235– 250, April 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X03002245.

Alzer:2004:TSV

[648] Horst Alzer and Stamatis Koumandos. On a trigonometric sum of Vinogradov. Journal of Number Theory, 105(2):251–261, April 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X03002117.

Hare:2004:SSP

[649] Kevin G. Hare. The structure of the spectra of Pisot numbers. Journal of Number Theory, 105(2):262–274, April 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X03002142.

Terhune:2004:EDV

[650] David Terhune. Evaluation of double L-values. Journal of Number Theory, 105(2):275–301, April 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X03002257.

Kraemer:2004:CUB

[651] Pavel Kraemer. Circular units in a bicyclic field. Journal of Number Theory, 105(2):302–321, April 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X03002166.

Khare:2004:RHA

[652] Chandrashekhar Khare and Dipendra Prasad. Reduction of homomorphisms modp and algebraicity. Journal of Number Theory, 105(2):322– 332, April 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X03002154.

Lee:2004:AMG

[653] Jung-Jo Lee and M. Ram Murty. An application of Mumford’s gap principle. Journal of Number Theory, 105(2):333–343, April 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X0300221X.

Escassut:2004:FEA

[654] Alain Escassut and Chung-Chun Yang. The functional equation P (f) = Q(g) in a p-adic field. Journal of Number Theory, 105(2):344– 360, April 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X03002208.

Blache:2004:ESL

[655] Régis Blache. Exponential sums over lifts of points. Journal of Number Theory, 105(2):361–386, April 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X03002129.

Plans:2004:TER

[656] Bernat Plans and Núria Vila. Trinomial extensions of Q with ramification conditions. Journal of Number Theory, 105(2):387–400, April 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X03002221.

Anonymous:2004:VAIa

[657] Anonymous. Volume author index. Journal of Number Theory, 105(2): 401, April 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X04000356.

Anonymous:2004:PMb

[658] Anonymous. Pages 1–186 (May 2004). Journal of Number Theory, 106 (1):??, May 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2004:IFCc

[659] Anonymous. Inside front cover: Editorial Board. Journal of Number Theory, 106(1):ifc, May 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X04000435.

Bachman:2004:DPC

[660] Gennady Bachman and Troy Kessler. On divisibility properties of certain multinomial coefficients — II. Journal of Number Theory, 106(1): 1–12, May 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X04000083.

Young:2004:GSM

[661] Paul Thomas Young. Gauss sums and multinomial coefficients. Journal of Number Theory, 106(1):13–25, May 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X03002270.

Pacelli:2004:ASO

[662] Allison M. Pacelli. Abelian subgroups of any order in class groups of global function fields. Journal of Number Theory, 106(1):26– 49, May 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X04000174.

Yang:2004:TC

[663] Jin Yang and Weizhe Zheng. On a theorem of Chowla. Journal of Number Theory, 106(1):50–55, May 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X03002269.

Goubin:2004:CLF

[664] Louis Goubin, Christian Mauduit, and András Sárközy. Construction of large families of pseudorandom binary sequences. Journal of Number Theory, 106(1):56–69, May 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X04000137.

Nathanson:2004:GAB

[665] Melvyn B. Nathanson. Generalized additive bases, König’s lemma, and the Erdős–Turán conjecture. Journal of Number Theory, 106(1):

70–78, May 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X04000162.

Brillhart:2004:CNQ

[666] John Brillhart and Patrick Morton. Class numbers of quadratic fields, Hasse invariants of elliptic curves, and the supersingular polynomial. Journal of Number Theory, 106(1):79–111, May 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X0400023X.

Ratazzi:2004:DPM

[667] Nicolas Ratazzi. Densité de points et minoration de hauteur. (French) [Density of points and reduction of height]. Journal of Number Theory, 106(1):112–127, May 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X04000186.

Webb:2004:PPE

[668] W. A. Webb and H. Yokota. Polynomial Pell’s equation — II. Journal of Number Theory, 106(1):128–141, May 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X04000228.

Bezerra:2004:TNG

[669] Juscelino Bezerra and Arnaldo Garcia. A tower with non–Galois steps which attains the Drinfeld–Vladut bound. Journal of Number Theory, 106(1):142–154, May 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X03002282.

Tsumura:2004:MHS

[670] Hirofumi Tsumura. Multiple harmonic series related to multiple Euler numbers. Journal of Number Theory, 106(1):155–168, May 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X04000216.

Thangadurai:2004:ATB

[671] R. Thangadurai. Adams theorem on Bernoulli numbers revisited. Journal of Number Theory, 106(1):169–177, May 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X04000204.

Lovejoy:2004:ORQ

[672] Jeremy Lovejoy. Overpartitions and real quadratic fields. Journal of Number Theory, 106(1):178–186, May 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X04000150.

Anonymous:2004:PJb

[673] Anonymous. Pages 187–386 (June 2004). Journal of Number Theory, 106 (2):??, June 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2004:IFCd

[674] Anonymous. Inside front cover: Editorial Board. Journal of Number Theory, 106(2):ifc, June 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X04000642.

Lee:2004:CLH

[675] Yoonjin Lee. Cohen–lenstra heuristics and the Spiegelungssatz; function fields. Journal of Number Theory, 106(2):187–199, June 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X04000241.

Hayashida:2004:SHJ

[676] Shuichi Hayashida. Skew-holomorphic Jacobi forms of index 1 and Siegel modular forms of half-integral weight. Journal of Number Theory, 106(2):200–218, June 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X04000149.

Masri:2004:HZF

[677] Riad Masri. The Herglotz–Zagier function, double zeta functions, and values of L-series. Journal of Number Theory, 106(2):219–237, June 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X04000307.

Lanphier:2004:PNL

[678] Dominic Lanphier. Petersson norms and liftings of Hilbert modular forms. Journal of Number Theory, 106(2):238–258, June 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X04000290.

Cochrane:2004:UBD

[679] Todd Cochrane, Ming-Chit Liu, and Zhiyong Zheng. Upper bounds on n-dimensional Kloosterman sums. Journal of Number Theory, 106(2): 259–274, June 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X03001859.

Chen:2004:RC

[680] Yong-Gao Chen and Xue-Gong Sun. On Romanoff’s constant. Journal of Number Theory, 106(2):275–284, June 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X03002191.

Schmidt:2004:REF

[681] Wolfgang M. Schmidt. Rationality of exponential functions at integer arguments. Journal of Number Theory, 106(2):285–298, June 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X04000253.

Barbolosi:2004:GQP

[682] D. Barbolosi and C. Faivre. Sur les grands quotients partiels du développement en fraction continue. (French) [On the large partial quotients in development as a continuous fraction]. Journal of Number Theory, 106(2):299–325, June 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X04000095.

Dujella:2004:CSP

[683] Andrej Dujella and Clemens Fuchs. Complete solution of the polynomial version of a problem of Diophantus. Journal of Number Theory, 106(2):326–344, June 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X04000101.

Ghitza:2004:HES

[684] Alexandru Ghitza. Hecke eigenvalues of Siegel modular forms (modp) and of algebraic modular forms. Journal of Number Theory, 106(2): 345–384, June 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X04000125.

Anonymous:2004:VAIb

[685] Anonymous. Volume author index. Journal of Number Theory, 106 (2):385, June 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X04000678.

Anonymous:2004:PJc

[686] Anonymous. Pages 1–206 (July 2004). Journal of Number Theory, 107 (1):??, July 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2004:IFCe

[687] Anonymous. Inside front cover: Editorial Board. Journal of Number Theory, 107(1):ifc, July 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X04000861.

Miyata:2004:IIA

[688] Yoshimasa Miyata. Indecomposability of ideals of p-adic number fields. Journal of Number Theory, 107(1):1–7, July 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X04000587.

Rohrlich:2004:MUS

[689] David E. Rohrlich. Modular units and the surjectivity of a Galois representation. Journal of Number Theory, 107(1):8–24, July 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X04000629.

Kojima:2004:FCH

[690] Hisashi Kojima. On the Fourier coefficients of Hilbert–Maass wave forms of half integral weight over arbitrary algebraic number fields. Journal of Number Theory, 107(1):25–62, July 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X04000575. See

corrigendum [1712].

Togbe:2004:PFC

[691] Alain Togbé. A parametric family of cubic Thue equations. Journal of Number Theory, 107(1):63–79, July 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X04000496. See

corrigendum [1715].

Suresh:2004:GCD

[692] V. Suresh. Galois cohomology in degree 3 of function fields of curves over number fields. Journal of Number Theory, 107(1):80– 94, July 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X04000289.

Haynes:2004:DSS

[693] Alan Haynes. The distribution of special subsets of the Farey sequence. Journal of Number Theory, 107(1):95–104, July 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X04000563.

Beaver:2004:TDM

[694] Olga R. Beaver and Thomas Garrity. A two-dimensional Minkowski ?(x) function. Journal of Number Theory, 107(1):105–134, July 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X04000265.

Akiyama:2004:CFP

[695] Shigeki Akiyama, Hui Rao, and Wolfgang Steiner. A certain finiteness property of Pisot number systems. Journal of Number Theory, 107(1):135–160, July 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X04000502.

Dokchitser:2004:LA

[696] Tim Dokchitser. LLL & ABC. Journal of Number Theory, 107(1): 161–167, July 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X04000526.

Roblot:2004:VAA

[697] X.-F. Roblot and D. Solomon. Verifying a p-adic abelian Stark conjecture at s =1. Journal of Number Theory, 107(1):168–206, July 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X04000198.

Anonymous:2004:PAb

[698] Anonymous. Pages 207–406 (August 2004). Journal of Number Theory, 107(2):??, August 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2004:IFCf

[699] Anonymous. Inside front cover: Editorial Board. Journal of Number Theory, 107(2):ifc, August 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X04001052.

Guo:2004:ETE

[700] Xuejun Guo and Hourong Qin. An embedding theorem for Eichler orders. Journal of Number Theory, 107(2):207–214, August 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X04000812.

Bremner:2004:LSW

[701] A. Bremner and N. Tzanakis. Lucas sequences whose 12th or 9th term is a square. Journal of Number Theory, 107(2):215–227, August 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X04000800.

Mollin:2004:DCF

[702] R. A. Mollin and B. Goddard. A description of continued fraction expansions of quadratic surds represented by polynomials. Journal of Number Theory, 107(2):228–240, August 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X04000538.

Dern:2004:GRH

[703] Tobias Dern and Aloys Krieg. The graded ring of Hermitian modular forms of degree 2 over Q(−2). Journal of Number Theory, 107(2):241– 265, August 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X03002130.

Sell:2004:CFG

[704] Elizabeth A. Sell. On a certain family of generalized Laguerre polynomials. Journal of Number Theory, 107(2):266–281, August 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X04000599.

Burns:2004:RCN

[705] David Burns and Joongul Lee. On the refined class number formula of Gross. Journal of Number Theory, 107(2):282–286, August 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X04000514.

Ki:2004:TL

[706] Haseo Ki. On a theorem of Levinson. Journal of Number Theory, 107(2):287–297, August 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X04000782.

Liu:2004:QWG

[707] Jianya Liu, Trevor D. Wooley, and Gang Yu. The quadratic Waring– Goldbach problem. Journal of Number Theory, 107(2):298–321, August 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X04000836.

Eigen:2004:UBI

[708] S. J. Eigen, Y. Ito, and V. S. Prasad. Universally bad integers and the 2adics. Journal of Number Theory, 107(2):322–334, August 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X04000769.

Knightly:2004:TCP

[709] Andrew H. Knightly. Tate classes on a product of two Picard modular surfaces. Journal of Number Theory, 107(2):335–344, August 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X04000757.

Chou:2004:CSR

[710] Wun-Seng Chou and Igor E. Shparlinski. On the cycle structure of repeated exponentiation modulo a prime. Journal of Number Theory, 107(2):345–356, August 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X04000745.

Choi:2004:QEF

[711] Stephen Kwok-Kwong Choi and Angel V. Kumchev. Quadratic equations with five prime unknowns. Journal of Number Theory, 107(2):357– 367, August 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X0400054X.

Hutchinson:2004:TWK

[712] Kevin Hutchinson. On tame and wild kernels of special number fields. Journal of Number Theory, 107(2):368–391, August 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X04000824.

Corson:2004:CSQ

[713] Daniel Corson, David Favero, Kate Liesinger, and Sarah Zubairy. Characters and q-series in Q2. Journal of Number Theory, 107(2):392– 405, August 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X04000551.

Anonymous:2004:VAIc

[714] Anonymous. Volume author index. Journal of Number Theory, 107(2): 406, August 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X04001088.

Anonymous:2004:PS

[715] Anonymous. Pages 1–186 (September 2004). Journal of Number Theory, 108(1):??, September 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2004:IFCg

[716] Anonymous. Inside front cover: Editorial Board. Journal of Number Theory, 108(1):ifc, September 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X04001325.

Weingartner:2004:IDDa

[717] Andreas Weingartner. Integers with dense divisors. Journal of Number Theory, 108(1):1–17, September 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X04001015.

Weingartner:2004:IDDb

[718] Andreas Weingartner. Integers with dense divisors 2. Journal of Number Theory, 108(1):18–28, September 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X04001027.

Fukshansky:2004:SZQ

[719] Lenny Fukshansky. Small zeros of quadratic forms with linear conditions. Journal of Number Theory, 108(1):29–43, September 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X04000927.

Zarhin:2004:TCP

[720] Yuri G. Zarhin. The Tate conjecture for powers of ordinary cubic fourfolds over finite fields. Journal of Number Theory, 108(1):44–59, September 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X04000848.

Sergeyev:2004:GSE

[721] Yaroslav D. Sergeyev. Generation of symmetric exponential sums. Journal of Number Theory, 108(1):60–75, September 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X04000794.

Liu:2004:VOC

[722] Chunlei Liu and Lanju Xu. The vanishing order of certain Hecke Lfunctions of imaginary quadratic fields. Journal of Number Theory, 108(1):76–89, September 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X04000770.

Ren:2004:WGT

[723] Xiumin Ren and Kai-Man Tsang. On a Waring–Goldbach-type problem for fourth powers. Journal of Number Theory, 108(1):90–110, September 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X0400099X.

Unver:2004:AMZ

[724] Sinan Ünver. p-adic multi-zeta values. Journal of Number Theory, 108(1):111–156, September 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X04001003.

Jutila:2004:TMC

[725] Matti Jutila. The twelfth moment of central values of Hecke series. Journal of Number Theory, 108(1):157–168, September 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X04000940.

Pilehrood:2004:CIG

[726] T. Hessami Pilehrood and Kh Hessami Pilehrood. Criteria for irrationality of generalized Euler’s constant. Journal of Number Theory, 108(1):169–185, September 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X04000988.

Anonymous:2004:PO

[727] Anonymous. Pages 187–376 (October 2004). Journal of Number Theory, 108(2):??, October 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2004:IFCh

[728] Anonymous. Inside front cover: Editorial Board. Journal of Number Theory, 108(2):ifc, October 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X04001519.

Jaulent:2004:GCP

[729] Jean-François Jaulent and Florence Soriano-Gafiuk. 2-groupe des classes positives d’un corps de nombres et noyau sauvage de la K-théorie. (French) [2-group of positive classes of a number field and wild kernel of the K-theory]. Journal of Number Theory, 108(2):187–208, October 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X03001860.

Honkala:2004:IDW

[730] Juha Honkala. It is decidable whether the image of an N-rational sequence has a base. Journal of Number Theory, 108(2):209–216, October 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X04001192.

Beelen:2004:GRD

[731] Peter Beelen. Graphs and recursively defined towers of function fields. Journal of Number Theory, 108(2):217–240, October 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X04000915.

Kulesz:2004:UBN

[732] L. Kulesz, G. Matera, and E. Schost. Uniform bounds on the number of rational points of a family of curves of genus 2. Journal of Number Theory, 108(2):241–267, October 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X04000964.

Murabayashi:2004:FDM

[733] Naoki Murabayashi. On the field of definition for modularity of CM elliptic curves. Journal of Number Theory, 108(2):268–286, October 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X04000976.

Yang:2004:LDA

[734] Tonghai Yang. Local densities of 2-adic quadratic forms. Journal of Number Theory, 108(2):287–345, October 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X04000939.

Yang:2004:FCF

[735] Zifeng Yang. C^{n}-functions over completions of F_{r}[T] at finite places of F_{r}(T). Journal of Number Theory, 108(2):346–374, October 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X04001039.

Anonymous:2004:VAId

[736] Anonymous. Volume author index. Journal of Number Theory, 108(2): 375, October 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X04001544.

Anonymous:2004:PN

[737] Anonymous. Pages 1–196 (November 2004). Journal of Number Theory, 109(1):??, November 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2004:IFCi

[738] Anonymous. Inside front cover: Editorial Board. Journal of Number Theory, 109(1):ifc, November 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X04001933.

Feng:2004:ECR

[739] Keqin Feng and Maosheng Xiong. On elliptic curves y^{2} = x^{3} − n^{2}x with rank zero. Journal of Number Theory, 109(1):1–26, November 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X04000113.

Komatsu:2004:TCF

[740] Takao Komatsu. Tasoev’s continued fractions and Rogers–Ramanujan continued fractions. Journal of Number Theory, 109(1):27–40, November 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X04001209.

Bekyel:2004:DEC

[741] Ebru Bekyel. The density of elliptic curves having a global minimal Weierstrass equation. Journal of Number Theory, 109(1):41–58, November 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X04001179.

Seo:2004:ESS

[742] Soogil Seo. Euler systems and special units. Journal of Number Theory, 109(1):59–68, November 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X04001271.

Bae:2004:DCG

[743] Sunghan Bae and Pyung-Lyun Kang. Double complex and gammamonomials in global function fields. Journal of Number Theory, 109 (1):69–88, November 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X04001167.

Rosenberg:2004:IIT

[744] Steven J. Rosenberg. On the Iwasawa invariants of the Γ-transform of a rational function. Journal of Number Theory, 109(1):89–95, November 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X0400126X.

Risager:2004:ADM

[745] Morten S. Risager. Asymptotic densities of Maass newforms. Journal of Number Theory, 109(1):96–119, November 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X04001258.

Borisov:2004:QIC

[746] Alexander Borisov, Melvyn B. Nathanson, and Yang Wang. Quantum integers and cyclotomy. Journal of Number Theory, 109(1):120–135, November 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X04001180. See corrigendum [1714].

Bayad:2004:FJF

[747] Abdelmejid Bayad and Jesús Gómez Ayala. Formes de Jacobi et formules de distribution. (French) [Jacobi forms and distribution formulas]. Journal of Number Theory, 109(1):136–162, November 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X04001155.

Welter:2004:ERL

[748] Michael Welter. Ensembles régulièrement lacunaires d’entiers et fonctions entières arithmétiques. (French) [Regularly incomplete sets of integers and whole arithmetic functions]. Journal of Number Theory, 109(1):163–181, November 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X04001295.

Jarvis:2004:FE

[749] Frazer Jarvis and Paul Meekin. The Fermat equation over Q(\sqrt{2}). Journal of Number Theory, 109(1):182–196, November 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X04001234.

Anonymous:2004:PD

[750] Anonymous. Pages 197–392 (December 2004). Journal of Number Theory, 109(2):??, December 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2004:IFCj

[751] Anonymous. Inside front cover: Editorial Board. Journal of Number Theory, 109(2):ifc, December 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X04002227.

Taniguchi:2004:MVT

[752] Takashi Taniguchi. A mean value theorem for orders of degree zero divisor class groups of quadratic extensions over a function field. Journal of Number Theory, 109(2):197–239, December 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X04001283.

Hall:2004:LSB

[753] R. R. Hall. Large spaces between the zeros of the Riemann zetafunction and random matrix theory. Journal of Number Theory, 109 (2):240–265, December 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X04000277.

Yu:2004:RFR

[754] Hoseog Yu. Representations and factors of restriction of scalars. Journal of Number Theory, 109(2):266–277, December 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X04001301.

James:2004:AFD

[755] Kevin James. Average Frobenius distributions for elliptic curves with 3-torsion. Journal of Number Theory, 109(2):278–298, December 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X04001222.

Long:2004:SIS

[756] Ling Long. On Shioda–Inose structures of one-parameter families of K3 surfaces. Journal of Number Theory, 109(2):299–318, December 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X04001246.

Ahlswede:2004:DPS

[757] R. Ahlswede, L. Khachatrian, and A. Sárközy. On the density of primitive sets. Journal of Number Theory, 109(2):319–361, December 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X04001817.

Adelberg:2004:UKC

[758] Arnold Adelberg. Universal Kummer congruences mod prime powers. Journal of Number Theory, 109(2):362–378, December 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X04001672.

Browkin:2004:SFQ

[759] J. Browkin and J. Brzeziński. Separable free quadratic algebras over quadratic integers. Journal of Number Theory, 109(2):379–389, December 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X04001829.

ONeil:2004:EPI

[760] Catherine O’Neil. Erratum to “The period-index obstruction for elliptic curves” [J. Number Theory 95 (2002) 329–339]. Journal of Number Theory, 109(2):390, December 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X04001660. See [372].

Anonymous:2004:VAIe

[761] Anonymous. Volume author index. Journal of Number Theory, 109(2): 391, December 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X04002252.

Anonymous:2005:IFCa

[762] Anonymous. Inside front cover: Editorial Board. Journal of Number Theory, 110(1):ifc, January 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X04002331.

David:2005:AER

[763] Goss David. Arnold Ephraim Ross (1906–2002). Journal of Number Theory, 110(1):1–2, January 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X04001210.

Coleman:2005:C

[764] Robert F. Coleman. On the components of X_{0}(p^{n}). Journal of Number Theory, 110(1):3–21, January 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X04001398.

Conrad:2005:MFR

[765] Brian Conrad and Keith Conrad. The Möbius function and the residue theorem. Journal of Number Theory, 110(1):22–36, January 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X04001386.

Duke:2005:TQF

[766] W. Duke. On ternary quadratic forms. Journal of Number Theory, 110(1):37–43, January 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X04001404.

Fried:2005:RTQ

[767] David Fried. Reduction theory over quadratic imaginary fields. Journal of Number Theory, 110(1):44–74, January 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X04001490.

Gessel:2005:MIB

[768] Ira M. Gessel. On Miki’s identity for Bernoulli numbers. Journal of Number Theory, 110(1):75–82, January 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X04001416.

Goss:2005:ANA

[769] David Goss. Applications of non–Archimedean integration to the L-series of τ-sheaves. Journal of Number Theory, 110(1):83–113, January 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X04001428.

Gross:2005:ECT

[770] Benedict H. Gross. An elliptic curve test for Mersenne primes. Journal of Number Theory, 110(1):114–119, January 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X0400143X.

Jin:2005:HCC

[771] Boping Jin and Gerry Myerson. Homogeneous covering congruences and subgroup covers. Journal of Number Theory, 110(1):120–135, January 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X04001453.

Gross:2005:ECD

[772] Benedict H. Gross and David Pollack. On the Euler characteristic of the discrete spectrum. Journal of Number Theory, 110(1):136–163, January 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X04001441.

Pollack:2005:AVT

[773] Robert Pollack. An algebraic version of a theorem of Kurihara. Journal of Number Theory, 110(1):164–177, January 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X04001465.

Silverberg:2005:IGA

[774] A. Silverberg and Yu. G. Zarhin. Inertia groups and abelian surfaces. Journal of Number Theory, 110(1):178–198, January 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X04001477.

Weston:2005:EUP

[775] Tom Weston. Explicit unobstructed primes for modular deformation problems of squarefree level. Journal of Number Theory, 110(1):199– 218, January 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X04001489.

Anonymous:2005:PF

[776] Anonymous. Pages 219–414 (February 2005). Journal of Number Theory, 110(2):??, February 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2005:IFCb

[777] Anonymous. Inside front cover: Editorial Board. Journal of Number Theory, 110(2):ifc, February 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X05000260.

Urbanski:2005:DAS

[778] Mariusz Urbański. Diophantine approximation and self-conformal measures. Journal of Number Theory, 110(2):219–235, February 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X04001659.

Kiming:2005:MGR

[779] Ian Kiming and Helena A. Verrill. On modular modℓ Galois representations with exceptional images. Journal of Number Theory, 110 (2):236–266, February 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X04000952.

Simsek:2005:ATS

[780] Yilmaz Simsek. q-Analogue of twisted l-series and q-twisted Euler numbers. Journal of Number Theory, 110(2):267–278, February 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X04001647.

Alvarez:2005:GTF

[781] A. Álvarez. Generalized theta functions, Drinfeld modules and some arithmetic consequences. Journal of Number Theory, 110(2):279–316, February 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X04001684.

Choudhry:2005:CSD

[782] Ajai Choudhry. The complete solution of a diophantine equation involving biquadrates. Journal of Number Theory, 110(2):317–324, February 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X04001854.

Joyce:2005:MCO

[783] Adam Joyce. The Manin constant of an optimal quotient of J_{0}(431). Journal of Number Theory, 110(2):325–330, February 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X0400188X.

See erratum [953].

Baladi:2005:EAG

[784] Viviane Baladi and Brigitte Vallée. Euclidean algorithms are Gaussian. Journal of Number Theory, 110(2):331–386, February 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X04001696.

See note [1647].

Klenke:2005:VEO

[785] Tomas Antonius Klenke. Visualizing elements of order two in the Weil–Châtelet group. Journal of Number Theory, 110(2):387–395, February 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X04001891.

Zunic:2005:NWO

[786] Joviša Žunić. On the number of ways to occupy n lattice points by balls in d-dimensional space. Journal of Number Theory, 110(2):396– 402, February 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X04001799.

Choudhry:2005:DPC

[787] Ajai Choudhry. A diophantine problem concerning cubes. Journal of Number Theory, 110(2):403–412, February 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X04001866.

Anonymous:2005:VAIa

[788] Anonymous. Volume author index. Journal of Number Theory, 110(2): 413, February 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X05000351.

Anonymous:2005:PMa

[789] Anonymous. Pages 1–196 (March 2005). Journal of Number Theory, 111 (1):??, March 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2005:IFCc

[790] Anonymous. Inside front cover: Editorial Board. Journal of Number Theory, 111(1):ifc, March 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X05000466.

Brown:2005:LCZ

[791] Francis C. S. Brown. Li’s criterion and zero-free regions of L-functions. Journal of Number Theory, 111(1):1–32, March 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X04001830.

Beresnevich:2005:AAN

[792] V. V. Beresnevich, V. I. Bernik, and E. I. Kovalevskaya. On approximation of p-adic numbers by p-adic algebraic numbers. Journal of Number Theory, 111(1):33–56, March 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X04001994.

Chisholm:2005:RTE

[793] C. Chisholm and J. A. MacDougall. Rational tetrahedra with edges in arithmetic progression. Journal of Number Theory, 111(1):57–80, March 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X04001842.

Kubotera:2005:SCL

[794] Norikazu Kubotera. A sufficient condition for Leopoldt’s conjecture. Journal of Number Theory, 111(1):81–85, March 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X04001908.

Ota:2005:SDC

[795] Kaori Ota. On saturated distinguished chains over a local field II. Journal of Number Theory, 111(1):86–143, March 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X04002689.

Yuan:2005:DEa

[796] Pingzhi Yuan and Yongzhong Hu. On the diophantine equation x^{2} + D^{m} = p^{n}. Journal of Number Theory, 111(1):144–153, March 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X04002598.

Cha:2005:VSC

[797] Byungchul Cha. Vanishing of some cohomology groups and bounds for the Shafarevich–Tate groups of elliptic curves. Journal of Number Theory, 111(1):154–178, March 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X04001702.

Bruin:2005:PS

[798] Nils Bruin. The primitive solutions to x^{3} + y^{9} = z^{2}. Journal of Number Theory, 111(1):179–189, March 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X04002471.

Sun:2005:RCF

[799] Zhi-Wei Sun. On the range of a covering function. Journal of Number Theory, 111(1):190–196, March 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X04002586.

Anonymous:2005:PAa

[800] Anonymous. Pages 197–392 (April 2005). Journal of Number Theory, 111(2):??, April 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2005:IFCd

[801] Anonymous. Inside front cover: Editorial Board. Journal of Number Theory, 111(2):ifc, April 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X05000612.

Schertz:2005:GCA

[802] Reinhard Schertz. Global construction of associated orders in complex multiplication. Journal of Number Theory, 111(2):197–226, April 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X04002185.

Fainleib:2005:DBI

[803] A. S. Fainleib. On the distribution of Beurling integers. Journal of Number Theory, 111(2):227–247, April 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X04002628.

Melfi:2005:SBE

[804] Giuseppe Melfi. On simultaneous binary expansions of n and n^{2}. Journal of Number Theory, 111(2):248–256, April 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X04002641.

Misaghian:2005:CPU

[805] Manouchehr Misaghian. Θ-correspondences (U(1), U(2)), Part I: unramified case. Journal of Number Theory, 111(2):257–286, April 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X04002501.

Misaghian:2005:CPI

[806] Manouchehr Misaghian. Θ-correspondences (U(1), U(2)), Part II: ramified case. Journal of Number Theory, 111(2):287–317, April 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X04002513.

Besser:2005:AAT

[807] Amnon Besser. p-adic Arakelov theory. Journal of Number Theory, 111(2):318–371, April 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X04002458.

Chen:2005:SBN

[808] Kwang-Wu Chen. A summation on Bernoulli numbers. Journal of Number Theory, 111(2):372–391, April 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X04001714.

Anonymous:2005:VAIb

[809] Anonymous. Volume author index. Journal of Number Theory, 111(2): 392, April 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X05000648.

Anonymous:2005:PMb

[810] Anonymous. Pages 1–204 (May 2005). Journal of Number Theory, 112 (1):??, May 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2005:IFCe

[811] Anonymous. Inside front cover: Editorial Board. Journal of Number Theory, 112(1):ifc, May 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X05000727.

Byott:2005:GMS

[812] Nigel P. Byott and Bouchaïb Sodaïgui. Galois module structure for dihedral extensions of degree 8: realizable classes over the group ring. Journal of Number Theory, 112(1):1–19, May 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X05000387.

Yuan:2005:DEb

[813] Pingzhi Yuan. On the diophantine equation \frac{x^{3}−1}{x−1} = \frac{y^{n}−1}{y−1}. Journal of Number Theory, 112(1):20–25, May 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X04002665.

Aubert:2005:PME

[814] Anne-Marie Aubert and Roger Plymen. Plancherel measure for GL(n, F) and GL(m, D): explicit formulas and Bernstein decomposition. Journal of Number Theory, 112(1):26–66, May 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X05000363.

Viada:2005:SAS

[815] Evelina Viada. Slopes and Abelian Subvariety Theorem. Journal of Number Theory, 112(1):67–115, May 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X0400191X.

Murabayashi:2005:QAS

[816] Naoki Murabayashi. On QM-abelian surfaces with model of GL_{2}-type over Q. Journal of Number Theory, 112(1):116–188, May 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X04002082.

Atkinson:2005:DMF

[817] James R. Atkinson. Divisors of modular forms on Γ_{0}(4). Journal of Number Theory, 112(1):189–204, May 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X04001805.

Anonymous:2005:PJa

[818] Anonymous. Pages 205–402 (June 2005). Journal of Number Theory, 112 (2):??, June 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2005:IFCf

[819] Anonymous. Inside front cover: Editorial Board. Journal of Number Theory, 112(2):CO2, June 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X05000879.

Chen:2005:RCG

[820] Yong-Gao Chen and Shu-Guang Guo. On the rational cuboids with a given face. Journal of Number Theory, 112(2):205–215, June 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X05000521.

Granville:2005:DRF

[821] Andrew Granville, Igor E. Shparlinski, and Alexandru Zaharescu. On the distribution of rational functions along a curve over F_{p} and residue races. Journal of Number Theory, 112(2):216–237, June 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X05000570.

Hironaka:2005:SFS

[822] Yumiko Hironaka. Spherical functions on Sp_{2} as a spherical homogeneous Sp_{2} × (Sp_{1})^{2}-space. Journal of Number Theory, 112(2):238– 286, June 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X04001726.

Adam:2005:SOF

[823] David Adam. Simultaneous orderings in function fields. Journal of Number Theory, 112(2):287–297, June 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X04002409.

Martin:2005:DSC

[824] Greg Martin. Dimensions of the spaces of cusp forms and newforms on Γ_{0}(N) and Γ_{1}(N). Journal of Number Theory, 112(2):298– 331, June 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X04002161.

Hilberdink:2005:WBB

[825] Titus W. Hilberdink. Well-behaved Beurling primes and integers. Journal of Number Theory, 112(2):332–344, June 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X04002069.

Hindry:2005:FCT

[826] Marc Hindry, Amílcar Pacheco, and Rania Wazir. Fibrations et conjecture de Tate. (French) [Fibrations and Tate conjecture]. Journal of Number Theory, 112(2):345–368, June 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X04002495.

ONeil:2005:MSG

[827] Catherine O’Neil. Models of some genus one curves with applications to descent. Journal of Number Theory, 112(2):369–385, June 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X04002045.

Papanikolas:2005:EAV

[828] Matthew A. Papanikolas and Niranjan Ramachandran. Extensions of abelian varieties defined over a number field. Journal of Number Theory, 112(2):386–400, June 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X04001787.

Anonymous:2005:VAIc

[829] Anonymous. Volume author index. Journal of Number Theory, 112 (2):401, June 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X05000909.

Anonymous:2005:PJb

[830] Anonymous. Pages 1–200 (July 2005). Journal of Number Theory, 113 (1):??, July 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2005:IFCg

[831] Anonymous. Inside front cover: Editorial Board. Journal of Number Theory, 113(1):CO2, July 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X05001083.

Hong:2005:NLC

[832] Shaofang Hong. Nonsingularity of least common multiple matrices on gcd-closed sets. Journal of Number Theory, 113(1):1–9, July 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X0500082X.

Sun:2005:QRB

[833] Zhi-Hong Sun. Quartic residues and binary quadratic forms. Journal of Number Theory, 113(1):10–52, July 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X04002574.

Jeong:2005:EFB

[834] Sangtae Jeong, Min-Soo Kim, and Jin-Woo Son. On explicit formulae for Bernoulli numbers and their counterparts in positive characteristic. Journal of Number Theory, 113(1):53–68, July 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X04001738.

Roettger:2005:PPC

[835] C. G. J. Roettger. Periodic points classify a family of Markov shifts. Journal of Number Theory, 113(1):69–83, July 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X04002549.

Kuriya:2005:NMF

[836] Takeshi Kuriya. On the norm maps of formal groups for Z_{p}-extensions. Journal of Number Theory, 113(1):84–116, July 2005. CODEN JNUTA9.

ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X0400174X.

Li:2005:ASD

[837] Wen-Ching Winnie Li, Ling Long, and Zifeng Yang. On Atkin– Swinnerton-Dyer congruence relations. Journal of Number Theory, 113(1):117–148, July 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X04001775.

Deutsch:2005:QPR

[838] Jesse Ira Deutsch. A quaternionic proof of the representation formula of a quaternary quadratic form. Journal of Number Theory, 113(1): 149–174, July 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X04002070.

Li:2005:AFC

[839] Xian-Jin Li. An arithmetic formula for certain coefficients of the Euler product of Hecke polynomials. Journal of Number Theory, 113(1): 175–200, July 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X04001763.

Anonymous:2005:PAb

[840] Anonymous. Pages 201–408 (August 2005). Journal of Number Theory, 113(2):??, August 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2005:IFCh

[841] Anonymous. Inside front cover: Editorial Board. Journal of Number Theory, 113(2):CO2, August 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X05001356.

Zhu:2005:EPP

[842] Qunsheng Zhu. Elements of prime power order in K_{2}(Q). Journal of Number Theory, 113(2):201–207, August 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X04002057.

Jao:2005:SPP

[843] David Jao. Supersingular primes for points on X_{0}(p)/w_{p}. Journal of Number Theory, 113(2):208–225, August 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X0400201X.

Alkan:2005:FNS

[844] Emre Alkan and Alexandru Zaharescu. B-free numbers in short arithmetic progressions. Journal of Number Theory, 113(2):226–243, August 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X04002094.

Massy:2005:GA

[845] Richard Massy. Galois averages. Journal of Number Theory, 113(2):244– 275, August 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X04002021. See remarks [1039].

Popescu:2005:RSC

[846] Cristian D. Popescu. The Rubin–Stark conjecture for a special class of function field extensions. Journal of Number Theory, 113(2):276– 307, August 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X04002173.

Andrianov:2005:ISE

[847] Anatoli Andrianov and Fedor Andrianov. On invariant subspaces and eigenfunctions of regular Hecke operators on spaces of multiple theta constants. Journal of Number Theory, 113(2):308–338, August 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X04002422.

Pacetti:2005:FCV

[848] Ariel Pacetti. A formula for the central value of certain Hecke L-functions. Journal of Number Theory, 113(2):339–379, August 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X04002653.

Borek:2005:SMS

[849] Thomas Borek. Successive minima and slopes of Hermitian vector bundles over number fields. Journal of Number Theory, 113(2):380– 388, August 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X04002677.

Bezivin:2005:EJP

[850] Jean-Paul Bézivin. Ensembles de Julia de polynômes p-adiques et points périodiques. (French) [Julia sets of p-adic polynomials and periodic points]. Journal of Number Theory, 113(2):389–407, August 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X0400246X.

Anonymous:2005:VAId

[851] Anonymous. Volume author index. Journal of Number Theory, 113(2): 408, August 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X05001381.

Anonymous:2005:OBC

[852] Anonymous. Outside back cover: table of contents; barcode. Journal of Number Theory, 113(2):CO4, August 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X05001435.

Anonymous:2005:PS

[853] Anonymous. Pages 1–208 (September 2005). Journal of Number Theory, 114(1):??, September 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2005:IFCi

[854] Anonymous. Inside front cover: Editorial Board. Journal of Number Theory, 114(1):CO2, September 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X05001617.

Deninger:2005:EPL

[855] Christopher Deninger. Exchanging the places p and ∞ in the Leopoldt conjecture. Journal of Number Theory, 114(1):1–17, September 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X0500079X.

Ji:2005:DEO

[856] Chun-Gang Ji. Diophantine equations x^{2} − Dy^{2} = −1, ±2, odd graphs, and their applications. Journal of Number Theory, 114(1):18–36, September 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X0500123X.

Meng:2005:STI

[857] Xianmeng Meng. On sums of three integers with a fixed number of prime factors. Journal of Number Theory, 114(1):37–65, September 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X0500106X.

Chen:2005:PPE

[858] Yen-Mei J. Chen and Jing Yu. On primitive points of elliptic curves with complex multiplication. Journal of Number Theory, 114(1):66–87, September 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X05000533.

Sun:2005:PNE

[859] Zhi-Hong Sun. On the properties of Newton–Euler pairs. Journal of Number Theory, 114(1):88–123, September 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X05001319.

Fujita:2005:TSE

[860] Yasutsugu Fujita. Torsion subgroups of elliptic curves in elementary abelian 2-extensions of Q. Journal of Number Theory, 114(1):124–134, September 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X05000818.

Fouvry:2005:EDS

[861] E. Fouvry and C. Mauduit. Sur les entiers dont la somme des chiffres est moyenne. (French) [On integers whose sum of digits is average]. Journal of Number Theory, 114(1):135–152, September 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X05000806.

Cooper:2005:QAP

[862] Joshua N. Cooper. Quasirandom arithmetic permutations. Journal of Number Theory, 114(1):153–169, September 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X05001198.

Eisentrager:2005:IPG

[863] Kirsten Eisenträger. Integrality at a prime for global fields and the perfect closure of global fields of characteristic p > 2. Journal of Number Theory, 114(1):170–181, September 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X05000557.

Garaev:2005:CSE

[864] M. Z. Garaev and A. A. Karatsuba. On character sums and the exceptional set of a congruence problem. Journal of Number Theory, 114(1):182–192, September 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X04002203.

Clark:2005:PIP

[865] Pete L. Clark. The period-index problem in WC-groups I: elliptic curves. Journal of Number Theory, 114(1):193–208, September 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X04002124.

Anonymous:2005:PO

[866] Anonymous. Pages 209–412 (October 2005). Journal of Number Theory, 114(2):??, October 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2005:IFCj

[867] Anonymous. Inside front cover: Editorial Board. Journal of Number Theory, 114(2):CO2, October 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X05001733.

Sun:2005:VRS

[868] Zhi-Wei Sun and Yeong-Nan Yeh. On various restricted sumsets. Journal of Number Theory, 114(2):209–220, October 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X05001599.

Belolipetsky:2005:MUL

[869] Mikhail Belolipetsky and Wee Teck Gan. The mass of unimodular lattices. Journal of Number Theory, 114(2):221–237, October 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X05000569.

Moree:2005:DOI

[870] Pieter Moree. On the distribution of the order and index of g(modp) over residue classes — I. Journal of Number Theory, 114(2):238–271, October 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X04002033.

Sato:2005:NDC

[871] Kanetomo Sato. Non-divisible cycles on surfaces over local fields. Journal of Number Theory, 114(2):272–297, October 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X04002550.

Luca:2005:DEP

[872] F. Luca and T. N. Shorey. Diophantine equations with products of consecutive terms in Lucas sequences. Journal of Number Theory, 114(2):298–311, October 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X04001751.

Im:2005:IMR

[873] Bo-Hae Im. Infinite multiplicity of roots of unity of the Galois group in the representation on elliptic curves. Journal of Number Theory, 114(2):312–323, October 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X05001514.

Adler:2005:DZB

[874] Jeffrey D. Adler and Joshua M. Lansky. Depth-zero base change for unramified U(2, 1). Journal of Number Theory, 114(2):324–360, October 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X0500096X. See erratum [1062].

Papikian:2005:PSI

[875] Mihran Papikian. Pesenti–Szpiro inequality for optimal elliptic curves. Journal of Number Theory, 114(2):361–393, October 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X04002537.

Bayer-Fluckiger:2005:MLC

[876] Eva Bayer-Fluckiger and Ivan Suarez. Modular lattices over cyclotomic fields. Journal of Number Theory, 114(2):394–411, October 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X04002100.

Anonymous:2005:VAIe

[877] Anonymous. Volume author index. Journal of Number Theory, 114(2): 412, October 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X05001769.

Anonymous:2005:VCa

[878] Anonymous. Volume Contents. Journal of Number Theory, 114(2):III– IV, October 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X05001794.

Anonymous:2005:PN

[879] Anonymous. Pages 1–196 (November 2005). Journal of Number Theory, 115(1):??, November 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2005:IFCk

[880] Anonymous. Inside front cover: Editorial Board. Journal of Number Theory, 115(1):CO2, November 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X05001848.

Furuya:2005:AOE

[881] Jun Furuya. On the average orders of the error term in the Dirichlet divisor problem. Journal of Number Theory, 115(1):1–26, November 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X04001878.

Popescu:2005:TSB

[882] Cristian D. Popescu. Torsion subgroups of Brauer groups and extensions of constant local degree for global function fields. Journal of Number Theory, 115(1):27–44, November 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X04001635.

Helou:2005:APW

[883] Charles Helou and Guy Terjanian. Arithmetical properties of Wendt’s determinant. Journal of Number Theory, 115(1):45–57, November 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X04002008.

Chen:2005:ISD

[884] Zhihua Chen and Min Ru. Integer solutions to decomposable form inequalities. Journal of Number Theory, 115(1):58–70, November 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X04002112.

Hall:2005:PFD

[885] R. R. Hall. The parity of Farey denominators and the Farey index. Journal of Number Theory, 115(1):71–86, November 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X04002136.

Lettl:2005:NTA

[886] Günter Lettl. Note on a theorem of A. Aiba. Journal of Number Theory, 115(1):87–88, November 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X04002148.

Mahlburg:2005:MCC

[887] Karl Mahlburg. More congruences for the coefficients of quotients of Eisenstein series. Journal of Number Theory, 115(1):89–99, November 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X0400215X.

Zhang:2005:IMV

[888] Qiao Zhang. Integral mean values of modular L-functions. Journal of Number Theory, 115(1):100–122, November 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X04002197.

Beliaeva:2005:USL

[889] Tatiana Beliaeva. Unités semi-locales modulo sommes de Gauss. (French) [Semi-local units modulo Gauss sums]. Journal of Number Theory, 115(1):123–157, November 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X04002434.

Schaefer:2005:CSG

[890] Edward F. Schaefer and Joseph L. Wetherell. Computing the Selmer group of an isogeny between Abelian varieties using a further isogeny to a Jacobian. Journal of Number Theory, 115(1):158–175, November 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X04002562.

Ash:2005:EMC

[891] A. Ash, D. Pollack, and W. Sinnott. A_{6}-extensions of Q and the modp cohomology of GL_{3}(Z). Journal of Number Theory, 115(1):176–196, November 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X04002616.

Anonymous:2005:PD

[892] Anonymous. Pages 197–394 (December 2005). Journal of Number Theory, 115(2):??, December 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2005:IFCl

[893] Anonymous. Inside front cover: Editorial Board. Journal of Number Theory, 115(2):CO2, December 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X05002180.

Anderson:2005:LWG

[894] Greg W. Anderson. Lacunary Wronskians on genus one curves. Journal of Number Theory, 115(2):197–214, December 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X04002410.

Bell:2005:CSP

[895] J. P. Bell and K. G. Hare. A classification of (some) Pisotcyclotomic numbers. Journal of Number Theory, 115(2):215–229, December 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X04002446.

Corso:2005:EBL

[896] Ilaria Del Corso and Roberto Dvornicich. An equivalence between local fields. Journal of Number Theory, 115(2):230–248, December 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X04002483.

Papikian:2005:VTS

[897] Mihran Papikian. On the variation of Tate–Shafarevich groups of elliptic curves over hyperelliptic curves. Journal of Number Theory, 115 (2):249–283, December 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X04002525.

Alzer:2005:NBM

[898] Horst Alzer and Bent Fuglede. Normalized binomial mid-coefficients and power means. Journal of Number Theory, 115(2):284–294, December 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X04002604.

Masri:2005:MDZ

[899] Riad Masri. Multiple Dedekind zeta functions and evaluations of extended multiple zeta values. Journal of Number Theory, 115(2):295–309, December 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X0400263X.

Garunkstis:2005:SZD

[900] Ramūnas Garunks̆tis and Jörn Steuding. Simple zeros and discrete moments of the derivative of the Riemann zeta-function. Journal of Number Theory, 115(2):310–321, December 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X04002690.

Banaszak:2005:DLD

[901] G. Banaszak, W. Gajda, and P. Krasoń. Detecting linear dependence by reduction maps. Journal of Number Theory, 115(2):322–342, December 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X05000375.

Gusic:2005:CLP

[902] Ivica Gusić. A characterization of linear polynomials. Journal of Number Theory, 115(2):343–347, December 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X05000405.

Seo:2005:CDR

[903] Soogil Seo. Circular distributions and a result of Solomon. Journal of Number Theory, 115(2):348–359, December 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X05000417.

vanFrankenhuijsen:2005:APZ

[904] Machiel van Frankenhuijsen. Arithmetic progressions of zeros of the Riemann zeta function. Journal of Number Theory, 115(2):360–370, December 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X05000429.

Sun:2005:ENM

[905] Zhi-Wei Sun. On Euler numbers modulo powers of two. Journal of Number Theory, 115(2):371–380, December 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X05000430.

Emmons:2005:PHE

[906] Brad A. Emmons. Products of Hecke eigenforms. Journal of Number Theory, 115(2):381–393, December 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X05000545.

Anonymous:2005:VAIf

[907] Anonymous. Volume author index. Journal of Number Theory, 115(2): 394, December 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X05002210.

Anonymous:2005:VCb

[908] Anonymous. Volume Contents. Journal of Number Theory, 115(2):III– IV, December 2005. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X05002246.

Anonymous:2006:PJa

[909] Anonymous. Pages 1–246 (January 2006). Journal of Number Theory, 116(1):??, January 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2006:IFCa

[910] Anonymous. Inside front cover: Editorial Board. Journal of Number Theory, 116(1):CO2, January 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X05002544.

Liu:2006:WTS

[911] Yu-Ru Liu and M. Ram Murty. A weighted Turán sieve method. Journal of Number Theory, 116(1):1–20, January 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X05000582.

Bautista-Ancona:2006:QFF

[912] Victor Bautista-Ancona and Javier Diaz-Vargas. Quadratic function fields with exponent two ideal class group. Journal of Number Theory, 116(1):21–41, January 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X05000594.

Cremona:2006:HDB

[913] J. E. Cremona, M. Prickett, and Samir Siksek. Height difference bounds for elliptic curves over number fields. Journal of Number Theory, 116(1):42–68, January 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X05000788.

Keating:2006:ELF

[914] Kevin Keating. Extensions of local fields and truncated power series. Journal of Number Theory, 116(1):69–101, January 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X05000831.

Lalin:2006:MMS

[915] Matilde N. Lalín. Mahler measure of some n-variable polynomial families. Journal of Number Theory, 116(1):102–139, January 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X05000843.

Hbaib:2006:CTF

[916] M. Hbaib, M. Mkaouar, and K. Tounsi. Un critére de transcendance de fractions continues dans K_{p}((X^{−1})). (French) [A transcendence criterion of continuous fractions in K_{p}((X^{−1}))]. Journal of Number Theory, 116(1):140–149, January 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X05000855.

Brown:2006:MNG

[917] Ron Brown. The minimal number with a given number of divisors. Journal of Number Theory, 116(1):150–158, January 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X05001009.

Choi:2006:SFC

[918] So Young Choi. Some formulas for the coefficients of Drinfeld modular forms. Journal of Number Theory, 116(1):159–167, January 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X05001010.

Duenez:2006:IQE

[919] Eduardo Dueñez, Steven J. Miller, Amitabha Roy, and Howard Straubing. Incomplete quadratic exponential sums in several variables. Journal of Number Theory, 116(1):168–199, January 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X05001022.

Espinosa:2006:ETD

[920] Olivier Espinosa and Victor H. Moll. The evaluation of Tornheim double sums, Part 1. Journal of Number Theory, 116(1):200–229, January 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X05001034.

Kessebohmer:2006:DSI

[921] Marc Kesseböhmer and Sanguo Zhu. Dimension sets for infinite IFSs: the Texan Conjecture. Journal of Number Theory, 116(1):230–246, January 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X05001046.

Anonymous:2006:PF

[922] Anonymous. Pages 247–488 (February 2006). Journal of Number Theory, 116(2):??, February 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2006:IFCb

[923] Anonymous. Inside front cover: Editorial Board. Journal of Number Theory, 116(2):CO2, February 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X06000035.

Angles:2006:FCF

[924] Bruno Anglès. On L-functions of cyclotomic function fields. Journal of Number Theory, 116(2):247–269, February 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X05000971.

Cochrane:2006:USM

[925] Todd Cochrane and Christopher Pinner. Using Stepanov’s method for exponential sums involving rational functions. Journal of Number Theory, 116(2):270–292, February 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X05000983.

Ballet:2006:ENS

[926] S. Ballet and D. Le Brigand. On the existence of non-special divisors of degree g and g−1 in algebraic function fields over F_{q}. Journal of Number Theory, 116(2):293–310, February 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X05000995.

Pacelli:2006:PIR

[927] Allison M. Pacelli. The prime at infinity and the rank of the class group in global function fields. Journal of Number Theory, 116(2):311–323, February 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X05001149.

Pink:2006:GRAa

[928] Richard Pink. The Galois representations associated to a Drinfeld module in special characteristic — I: Zariski density. Journal of Number Theory, 116(2):324–347, February 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X05001150.

Pink:2006:GRAb

[929] Richard Pink. The Galois representations associated to a Drinfeld module in special characteristic — II: Openness. Journal of Number Theory, 116(2):348–372, February 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X05001162.

Pink:2006:GRAc

[930] Richard Pink and Matthias Traulsen. The Galois representations associated to a Drinfeld module in special characteristic — III: Image of the group ring. Journal of Number Theory, 116(2):373–395, February 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X05001174.

Larsen:2006:RCE

[931] Michael Larsen and René Schoof. A refined counter-example to the support conjecture for abelian varieties. Journal of Number Theory, 116(2):396–398, February 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X05001241.

Martin:2006:PFM

[932] François Martin. Périodes de formes modulaires de poids 1. (French) [Periods of modular forms of weight 1]. Journal of Number Theory, 116(2):399–442, February 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X05001253.

Nart:2006:NHC

[933] Enric Nart and Christophe Ritzenthaler. Non-hyperelliptic curves of genus three over finite fields of characteristic two. Journal of Number Theory, 116(2):443–473, February 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X05001277.

Robinson:2006:NGP

[934] Z. Robinson. A note on global p-th powers of rigid analytic functions. Journal of Number Theory, 116(2):474–482, February 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X05001289.

Starni:2006:SPE

[935] Paolo Starni. On some properties of the Euler’s factor of certain odd perfect numbers. Journal of Number Theory, 116(2):483–486, February 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X05001307.

Anonymous:2006:VAIa

[936] Anonymous. Volume author index. Journal of Number Theory, 116(2): 487, February 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X06000060.

Anonymous:2006:VCa

[937] Anonymous. Volume Contents. Journal of Number Theory, 116(2):III– IV, February 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X06000096.

Anonymous:2006:PMa

[938] Anonymous. Pages 1–240 (March 2006). Journal of Number Theory, 117 (1):??, March 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2006:IFCc

[939] Anonymous. Inside front cover: Editorial Board. Journal of Number Theory, 117(1):CO2, March 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X0600031X.

Kontorovich:2006:QAR

[940] Alex V. Kontorovich and Melvyn B. Nathanson. Quadratic addition rules for quantum integers. Journal of Number Theory, 117(1):

1–13, March 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X05001058.

Barbe:2006:BNS

[941] A. Barbé and F. von Haeseler. Binary number systems for Z^{k}. Journal of Number Theory, 117(1):14–30, March 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X05001186.

Chen:2006:EEE

[942] Kwang-Wu Chen and Minking Eie. Explicit evaluations of extended Euler sums. Journal of Number Theory, 117(1):31–52, March 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X05001204.

Elkin:2006:HJR

[943] Arsen Elkin. Hyperelliptic Jacobians with real multiplication. Journal of Number Theory, 117(1):53–86, March 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X05001216.

Fukuhara:2006:GFH

[944] Shinji Fukuhara and Noriko Yui. A generating function of higherdimensional Apostol–Zagier sums and its reciprocity law. Journal of Number Theory, 117(1):87–105, March 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X05001228.

Mukunda:2006:LPN

[945] Keshav Mukunda. Littlewood Pisot numbers. Journal of Number Theory, 117(1):106–121, March 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X05001265.

Garaev:2006:SID

[946] M. Z. Garaev and A. Sankaranarayanan. The sum involving derivative of ζ(s) over simple zeros. Journal of Number Theory, 117(1):122– 130, March 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X05001290.

Zhefeng:2006:OHD

[947] Xu Zhefeng and Zhang Wenpeng. On the order of the high-dimensional Cochrane sum and its mean value. Journal of Number Theory, 117(1): 131–145, March 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X05001320.

Applegate:2006:S

[948] David Applegate and Jeffrey C. Lagarias. The 3x + 1 semigroup. Journal of Number Theory, 117(1):146–159, March 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X05001459.

Bell:2006:PII

[949] Jason P. Bell and Stanley N. Burris. Partition identities II. The results of Bateman and Erdős. Journal of Number Theory, 117(1):160– 190, March 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X05001460.

Deutsch:2006:CCM

[950] Emeric Deutsch and Bruce E. Sagan. Congruences for Catalan and Motzkin numbers and related sequences. Journal of Number Theory, 117(1):191–215, March 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X05001472.

Pan:2006:NIP

[951] Hao Pan. Note on integer powers in sumsets. Journal of Number Theory, 117(1):216–221, March 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X05001551.

Dubickas:2006:DRP

[952] Artūras Dubickas. On the distance from a rational power to the nearest integer. Journal of Number Theory, 117(1):222–239, March 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X0500168X.

Joyce:2006:EMC

[953] Adam Joyce. Erratum to “The Manin constant of an optimal quotient of J_{0}(431)” [J. Number Theory 110 (2005) 325–330]. Journal of Number Theory, 117(1):240, March 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X05001332. See [783].

Anonymous:2006:PAa

[954] Anonymous. Pages 241–482 (April 2006). Journal of Number Theory, 117(2):??, April 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2006:IFCd

[955] Anonymous. Inside front cover: Editorial Board. Journal of Number Theory, 117(2):CO2, April 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X06000497.

Diaz-Vargas:2006:ZCZ

[956] Javier Diaz-Vargas. On zeros of characteristic p zeta function. Journal of Number Theory, 117(2):241–262, April 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X05001484.

Conrey:2006:PRZ

[957] Brian Conrey and Alex Gamburd. Pseudomoments of the Riemann zetafunction and pseudomagic squares. Journal of Number Theory, 117(2): 263–278, April 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X05000399.

Folsom:2006:MFE

[958] Amanda Folsom. Modular forms and Eisenstein’s continued fractions. Journal of Number Theory, 117(2):279–291, April 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X05001496.

Halupczok:2006:NRT

[959] Karin Halupczok. On the number of representations in the ternary Goldbach problem with one prime number in a given residue class. Journal of Number Theory, 117(2):292–300, April 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X05001502.

Jenkins:2006:KST

[960] Paul Jenkins. Kloosterman sums and traces of singular moduli. Journal of Number Theory, 117(2):301–314, April 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X05001526.

Lawrence:2006:CPE

[961] Jim Lawrence. Characteristic polynomials, Ehrhart quasi-polynomials, and torus groups. Journal of Number Theory, 117(2):315–329, April 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X05001538.

Moree:2006:DOIa

[962] Pieter Moree. On the distribution of the order and index of g(mod p) over residue classes II. Journal of Number Theory, 117(2):330– 354, April 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X0500154X.

Pink:2006:ICM

[963] Richard Pink and Matthias Traulsen. The isogeny conjecture for tmotives associated to direct sums of Drinfeld modules. Journal of Number Theory, 117(2):355–375, April 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X05001563.

Tuenter:2006:FPS

[964] Hans J. H. Tuenter. The Frobenius problem, sums of powers of integers, and recurrences for the Bernoulli numbers. Journal of Number Theory, 117(2):376–386, April 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X05001575.

Walling:2006:BHS

[965] Lynne H. Walling. On bounding Hecke–Siegel eigenvalues. Journal of Number Theory, 117(2):387–396, April 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X05001587.

Dieulefait:2006:EDI

[966] Luis Dieulefait and Mladen Dimitrov. Explicit determination of images of Galois representations attached to Hilbert modular forms. Journal of Number Theory, 117(2):397–405, April 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X05001678.

Howard:2006:SCC

[967] Benjamin Howard. Special cohomology classes for modular Galois representations. Journal of Number Theory, 117(2):406–438, April 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X05001691.

Lozano-Robledo:2006:EUA

[968] Álvaro Lozano-Robledo. On elliptic units and p-adic Galois representations attached to elliptic curves. Journal of Number Theory, 117(2): 439–470, April 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X05001708.

Mattarei:2006:MPB

[969] Sandro Mattarei. Modular periodicity of binomial coefficients. Journal of Number Theory, 117(2):471–481, April 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X0500171X.

Anonymous:2006:VAIb

[970] Anonymous. Volume author index. Journal of Number Theory, 117(2): 482, April 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X06000527.

Anonymous:2006:VCb

[971] Anonymous. Volume Contents. Journal of Number Theory, 117(2): III–IV, April 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X06000552.

Anonymous:2006:PMb

[972] Anonymous. Pages 1–144 (May 2006). Journal of Number Theory, 118 (1):??, May 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2006:EBa

[973] Anonymous. Editorial Board. Journal of Number Theory, 118(1): CO2, May 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X06000722.

Alkan:2006:DAM

[974] Emre Alkan, Glyn Harman, and Alexandru Zaharescu. Diophantine approximation with mild divisibility constraints. Journal of Number Theory, 118(1):1–14, May 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X05001903.

Andrews:2006:NDM

[975] George E. Andrews, Arnold Knopfmacher, and Burkhard Zimmermann. On the number of distinct multinomial coefficients. Journal of Number Theory, 118(1):15–30, May 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X05001915.

Baoulina:2006:GMH

[976] Ioulia Baoulina. Generalizations of the Markoff–Hurwitz equations over finite fields. Journal of Number Theory, 118(1):31–52, May 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X05001927.

Chang:2006:NRQ

[977] Sungkon Chang. Note on the rank of quadratic twists of Mordell equations. Journal of Number Theory, 118(1):53–61, May 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X05001939.

Doud:2006:EIG

[978] Darrin Doud and Michael W. Moore. Even icosahedral Galois representations of prime conductor. Journal of Number Theory, 118(1): 62–70, May 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X05001940.

Everest:2006:PDE

[979] Graham Everest, Gerard Mclaren, and Thomas Ward. Primitive divisors of elliptic divisibility sequences. Journal of Number Theory, 118(1):

71–89, May 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X05001952.

Gerth:2006:DSR

[980] Frank Gerth. Densities for some real quadratic fields with infinite Hilbert 2-class field towers. Journal of Number Theory, 118(1):90– 97, May 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X05001964.

Lagarias:2006:RHC

[981] Jeffrey C. Lagarias and Masatoshi Suzuki. The Riemann hypothesis for certain integrals of Eisenstein series. Journal of Number Theory, 118(1):98–122, May 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X05001988.

Mihailescu:2006:CGC

[982] Preda Mihăilescu. On the class groups of cyclotomic extensions in presence of a solution to Catalan’s equation. Journal of Number Theory, 118(1):123–144, May 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X0500199X.

Anonymous:2006:PJb

[983] Anonymous. Pages 145–292 (June 2006). Journal of Number Theory, 118 (2):??, June 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2006:EBb

[984] Anonymous. Editorial Board. Journal of Number Theory, 118(2): CO2, June 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X06000898.

Hattori:2006:RFF

[985] Shin Hattori. Ramification of a finite flat group scheme over a local field. Journal of Number Theory, 118(2):145–154, June 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X05001976.

Mizuno:2006:GLF

[986] Yoshinori Mizuno. Generalized Lerch formulas: Examples of zetaregularized products. Journal of Number Theory, 118(2):155–171, June 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X05002003.

Rio:2006:DEO

[987] Anna Rio. Dyadic exercises for octahedral extensions II. Journal of Number Theory, 118(2):172–188, June 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X05002015.

Daileda:2006:AIU

[988] Ryan C. Daileda. Algebraic integers on the unit circle. Journal of Number Theory, 118(2):189–191, June 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X05002027.

Dancs:2006:ETF

[989] Michael J. Dancs and Tian-Xiao He. An Euler-type formula for ζ(2k + 1). Journal of Number Theory, 118(2):192–199, June 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X05002039.

Fujii:2006:CEP

[990] Satoshi Fujii, Yoshihiro Ohgi, and Manabu Ozaki. Construction of Z_{p}- extensions with prescribed Iwasawa λ-invariants. Journal of Number Theory, 118(2):200–207, June 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X05002040.

Muhlenbruch:2006:HOP

[991] T. Mühlenbruch. Hecke operators on period functions for Γ_{0}(n). Journal of Number Theory, 118(2):208–235, June 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X05002052.

Scott:2006:GPE

[992] Reese Scott and Robert Styer. On the generalized Pillai equation ±ax ± by = c. Journal of Number Theory, 118(2):236–265, June 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X05002064.

Bruin:2006:ESH

[993] N. Bruin and E. V. Flynn. Exhibiting SHA[2] on hyperelliptic Jacobians. Journal of Number Theory, 118(2):266–291, June 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X05002088.

Anonymous:2006:VAIc

[994] Anonymous. Volume author index. Journal of Number Theory, 118 (2):292, June 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X06000928.

Anonymous:2006:VCc

[995] Anonymous. Volume Contents. Journal of Number Theory, 118(2): III, June 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X06000953.

Anonymous:2006:PJc

[996] Anonymous. Pages 1–154 (July 2006). Journal of Number Theory, 119 (1):??, July 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2006:EBc

[997] Anonymous. Editorial Board. Journal of Number Theory, 119(1): CO2, July 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X06001004.

Baranczuk:2006:RMS

[998] Stefan Barańczuk. On reduction maps and support problem in Ktheory and abelian varieties. Journal of Number Theory, 119(1):1– 17, July 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X05002076.

Choi:2006:PMD

[999] So Young Choi, Kuk Jin Hong, and Daeyeol Jeon. On plane models for Drinfeld modular curves. Journal of Number Theory, 119(1):

18–27, July 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X0500209X.

Erovenko:2006:BGA

[1000] Igor V. Erovenko and Andrei S. Rapinchuk. Bounded generation of Sarithmetic subgroups of isotropic orthogonal groups over number fields. Journal of Number Theory, 119(1):28–48, July 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X05002106.

Gaal:2006:DEG

[1001] István Gaál and Michael Pohst. Diophantine equations over global function fields I: the Thue equation. Journal of Number Theory, 119(1): 49–65, July 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X05002118.

Girstmair:2006:CFD

[1002] Kurt Girstmair. Continued fractions and Dedekind sums: Threeterm relations and distribution. Journal of Number Theory, 119(1): 66–85, July 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X0500212X.

Goldring:2006:DFP

[1003] Wushi Goldring. Dynamics of the w function and primes. Journal of Number Theory, 119(1):86–98, July 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X05002131.

Guletskiii:2006:FDO

[1004] Vladimir Guletskiıĭ. Finite-dimensional objects in distinguished triangles. Journal of Number Theory, 119(1):99–127, July 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X05002143.

Hall:2006:FTL

[1005] Chris Hall. L-functions of twisted Legendre curves. Journal of Number Theory, 119(1):128–147, July 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X05002155.

Im:2006:WAL

[1006] Bo-Hae Im and Michael Larsen. Weak approximation for linear systems of quadrics. Journal of Number Theory, 119(1):148–154, July 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X05002167.

Anonymous:2006:PAb

[1007] Anonymous. Pages 155–318 (August 2006). Journal of Number Theory, 119(2):??, August 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2006:EBd

[1008] Anonymous. Editorial Board. Journal of Number Theory, 119(2): CO2, August 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X06001405.

Liu:2006:PAE

[1009] Yu-Ru Liu. Prime analogues of the Erdős–Kac theorem for elliptic curves. Journal of Number Theory, 119(2):155–170, August 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X05002313.

Northshield:2006:IAC

[1010] S. Northshield. On integral Apollonian circle packings. Journal of Number Theory, 119(2):171–193, August 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X05002301.

Pries:2006:WRC

[1011] Rachel J. Pries. Wildly ramified covers with large genus. Journal of Number Theory, 119(2):194–209, August 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X05002325.

Sun:2006:NIR

[1012] Zhi-Hong Sun. On the number of incongruent residues of x^{4} + ax^{2} + bx modulo p. Journal of Number Theory, 119(2):210–241, August 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X05002349.

Browning:2006:DRP

[1013] T. D. Browning. The density of rational points on a certain singular cubic surface. Journal of Number Theory, 119(2):242–283, August 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X05002374.

Bui:2006:TPA

[1014] H. M. Bui and J. P. Keating. On twin primes associated with the Hawkins random sieve. Journal of Number Theory, 119(2):284–296, August 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X05002386.

Buium:2006:DES

[1015] Alexandru Buium. A differential equation satisfied by the arithmetic Kodaira–Spencer class. Journal of Number Theory, 119(2):297–306, August 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X05002398.

Chen:2006:CEG

[1016] Yong-Gao Chen. On a conjecture of Erdős, Graham and Spencer. Journal of Number Theory, 119(2):307–314, August 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X05002404.

Hou:2006:CEI

[1017] Xiang-Dong Hou and Kevin Keating. Corrigendum to “Enumeration of isomorphism classes of extensions of p-adic fields” [J. Number Theory 104 (2004) 14–61]. Journal of Number Theory, 119(2):315–316, August 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X05002362. See [604].

Anonymous:2006:AIVa

[1018] Anonymous. Author index for volume 119. Journal of Number Theory, 119(2):317, August 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X06001491.

Anonymous:2006:CVa

[1019] Anonymous. Contents of volume 119. Journal of Number Theory, 119 (2):III, August 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X06001521.

Anonymous:2006:PS

[1020] Anonymous. Pages 1–192 (September 2006). Journal of Number Theory, 120(1):??, September 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2006:EBe

[1021] Anonymous. Editorial Board. Journal of Number Theory, 120(1):CO2, September 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X06001715.

Toulmonde:2006:VFR

[1022] Vincent Toulmonde. Sur les variations de la fonction de répartition de φ(n)/n. (French) [On the variations of the distribution function of φ(n)/n]. Journal of Number Theory, 120(1):1–12, September 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X05002337.

Fukshansky:2006:SLA

[1023] Lenny Fukshansky. Siegel’s lemma with additional conditions. Journal of Number Theory, 120(1):13–25, September 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X05002416.

Gao:2006:SSA

[1024] Weidong Gao and Imre Leader. Sums and k-sums in abelian groups of order k. Journal of Number Theory, 120(1):26–32, September 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X05002428.

Ho:2006:APT

[1025] Kwan-Hung Ho and Kai-Man Tsang. On almost prime k-tuples. Journal of Number Theory, 120(1):33–46, September 2006. CODEN JNUTA9.

ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X0500243X.

Hutchinson:2006:EWK

[1026] Kevin Hutchinson. Étale wild kernels of exceptional number fields. Journal of Number Theory, 120(1):47–71, September 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X05002441.

Jaulent:2006:ALN

[1027] Jean-François Jaulent and Alexis Michel. Approche logarithmique des noyaux étales sauvages des corps de nombres. (French) [Logarithmic approach of wild state nuclei of number fields]. Journal of Number Theory, 120(1):72–91, September 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X05002453.

Kane:2006:GBR

[1028] Daniel M. Kane. Generalized base representations. Journal of Number Theory, 120(1):92–100, September 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X05002465.

Kanemitsu:2006:SEM

[1029] Shigeru Kanemitsu, Yoshio Tanigawa, and Masami Yoshimoto. Structural elucidation of the mean square of the Hurwitz zeta-function. Journal of Number Theory, 120(1):101–119, September 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X05002477.

Ki:2006:RFU

[1030] Haseo Ki. The Riemann Ξ-function under repeated differentiation. Journal of Number Theory, 120(1):120–131, September 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X05002489.

Moree:2006:DOIb

[1031] Pieter Moree. On the distribution of the order and index of g(modp) over residue classes III. Journal of Number Theory, 120(1):132–160, September 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X05002490.

Tsumura:2006:SFR

[1032] Hirofumi Tsumura. On some functional relations between Mordell– Tornheim double L-functions and Dirichlet L-functions. Journal of Number Theory, 120(1):161–178, September 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X05002519.

Zaimi:2006:APP

[1033] Toufik Zaïmi. An arithmetical property of powers of Salem numbers. Journal of Number Theory, 120(1):179–191, September 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X05002520.

Anonymous:2006:PO

[1034] Anonymous. Pages 193–386 (October 2006). Journal of Number Theory, 120(2):??, October 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2006:EBf

[1035] Anonymous. Editorial Board. Journal of Number Theory, 120(2): CO2, October 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X06001788.

Chen:2006:LCT

[1036] Jiang-Ping Jeff Chen. The n × (n − 2) local converse theorem for GL(n) over a p-adic field. Journal of Number Theory, 120(2):193– 205, October 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X05002623.

Ding:2006:LTM

[1037] Xia-Qi Ding and Shao-Ji Feng. A logarithm type mean value theorem of the Riemann zeta function. Journal of Number Theory, 120(2):206– 212, October 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X05002635.

Dujella:2006:DTL

[1038] Andrej Dujella, Clemens Fuchs, and Gary Walsh. Diophantine m-tuples for linear polynomials II. Equal degrees. Journal of Number Theory, 120(2):213–228, October 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X05002647.

Jensen:2006:SRH

[1039] C. U. Jensen and Richard Massy. Some remarks on Hilbertian fields (an appendix to the paper “Galois averages” by R. Massy). Journal of Number Theory, 120(2):229–233, October 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X05002659. See

[845].

Morton:2006:EII

[1040] Patrick Morton. Explicit identities for invariants of elliptic curves. Journal of Number Theory, 120(2):234–271, October 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X05002660.

Ohtani:2006:DLA

[1041] Sachiko Ohtani. Deformations of locally abelian Galois representations and unramified extensions. Journal of Number Theory, 120(2):272– 286, October 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X05002672.

Petersen:2006:QFJ

[1042] Sebastian Petersen. On a question of Frey and Jarden about the rank of Abelian varieties. Journal of Number Theory, 120(2):287–302, October 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X05002684.

Ziegler:2006:FRQ

[1043] Volker Ziegler. On a family of relative quartic Thue inequalities. Journal of Number Theory, 120(2):303–325, October 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X05002696.

Banks:2006:VEF

[1044] William D. Banks and Francesco Pappalardi. Values of the Euler function free of k-th powers. Journal of Number Theory, 120(2):326–348, October 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X06000151.

Chamberland:2006:DPC

[1045] Marc Chamberland and Karl Dilcher. Divisibility properties of a class of binomial sums. Journal of Number Theory, 120(2):349–371, October 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X06000163.

Gaal:2006:PIB

[1046] István Gaál and Leanne Robertson. Power integral bases in primepower cyclotomic fields. Journal of Number Theory, 120(2):372–384, October 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X06000175.

Anonymous:2006:IV

[1047] Anonymous. Index for volume 120. Journal of Number Theory, 120(2): 385, October 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X06001818.

Anonymous:2006:CVb

[1048] Anonymous. Contents of volume 120. Journal of Number Theory, 120(2):III–IV, October 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X06001843.

Anonymous:2006:PN

[1049] Anonymous. Pages 1–186 (November 2006). Journal of Number Theory, 121(1):??, November 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2006:EBg

[1050] Anonymous. Editorial Board. Journal of Number Theory, 121(1):CO2, November 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X06001983.

Kajikawa:2006:DDS

[1051] Jun Kajikawa. Duality and double shuffle relations of multiple zeta values. Journal of Number Theory, 121(1):1–6, November 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X06000199.

Kim:2006:CRB

[1052] Sey Kim. On congruence relations between the fundamental units of biquadratic fields. Journal of Number Theory, 121(1):7–29, November 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X06000205.

Louboutin:2006:CNO

[1053] Stéphane R. Louboutin. The class-number one problem for some real cubic number fields with negative discriminants. Journal of Number Theory, 121(1):30–39, November 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X06000217.

Makover:2006:EPR

[1054] Eran Makover and Jeffrey McGowan. An elementary proof that random Fibonacci sequences grow exponentially. Journal of Number Theory, 121(1):40–44, November 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X06000229.

Meng:2006:SNR

[1055] Zaizhao Meng. Some new results on k-free numbers. Journal of Number Theory, 121(1):45–66, November 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X06000230.

Navilarekallu:2006:ETN

[1056] Tejaswi Navilarekallu. On the equivariant Tamagawa number conjecture for A_{4}-extensions of number fields. Journal of Number Theory, 121(1):67–89, November 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X06000242.

Ratliff:2006:PIB

[1057] Louis J. Ratliff, David E. Rush, and Kishor Shah. Power integral bases for Selmer-like number fields. Journal of Number Theory, 121 (1):90–113, November 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X06000254.

Son:2006:AFT

[1058] Seung H. Son. Additive formulae of theta functions and modular equations of degree five. Journal of Number Theory, 121(1):114–117, November 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X06000266.

Wiesend:2006:CCA

[1059] Götz Wiesend. A construction of covers of arithmetic schemes. Journal of Number Theory, 121(1):118–131, November 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X06000278.

Borisov:2006:WWO

[1060] Lev A. Borisov. On Wronskians of weight one Eisenstein series. Journal of Number Theory, 121(1):132–152, November 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X06000400.

Chisholm:2006:RHT

[1061] C. Chisholm and J. A. MacDougall. Rational and Heron tetrahedra. Journal of Number Theory, 121(1):153–185, November 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X06000412.

Adler:2006:EDZ

[1062] Jeffrey D. Adler and Joshua M. Lansky. Erratum to “Depthzero base change for unramified U(2, 1)” [J. Number Theory 114 (2005) 324–360]. Journal of Number Theory, 121(1):186, November 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X06000394. See [874].

Anonymous:2006:PD

[1063] Anonymous. Pages 187–356 (December 2006). Journal of Number Theory, 121(2):??, December 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2006:EBh

[1064] Anonymous. Editorial Board. Journal of Number Theory, 121(2):CO2, December 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X06002113.

Garth:2006:CSP

[1065] David Garth and Kevin G. Hare. Comments on the spectra of Pisot numbers. Journal of Number Theory, 121(2):187–203, December 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X06000424.

Lau:2006:SBR

[1066] Yuk-Kam Lau, Jianya Liu, and Yangbo Ye. Subconvexity bounds for Rankin–Selberg L-functions for congruence subgroups. Journal of Number Theory, 121(2):204–223, December 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X06000436.

Pan:2006:CIP

[1067] Hao Pan and Hui-Qin Cao. A congruence involving products of qbinomial coefficients. Journal of Number Theory, 121(2):224–233, December 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X06000448.

Rebolledo:2006:MSH

[1068] Marusia Rebolledo. Module supersingulier et homologie des courbes modulaires. (French) [Supersingular module and homology of modular curves]. Journal of Number Theory, 121(2):234–264, December 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X0600045X.

Rogers:2006:SIT

[1069] Mathew D. Rogers. A study of inverse trigonometric integrals associated with three-variable Mahler measures, and some related identities. Journal of Number Theory, 121(2):265–304, December 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X06000461.

Fluckiger:2006:UBE

[1070] Eva Bayer Fluckiger. Upper bounds for Euclidean minima of algebraic number fields. Journal of Number Theory, 121(2):305–323, December 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X0600062X.

Biro:2006:SCS

[1071] András Biró. Strong characterizing sequences for subgroups of compact groups. Journal of Number Theory, 121(2):324–354, December 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X06000631.

Anonymous:2006:AIVb

[1072] Anonymous. Author index for volume 121. Journal of Number Theory, 121(2):355, December 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X06002149.

Anonymous:2006:CVc

[1073] Anonymous. Contents of volume 121. Journal of Number Theory, 121(2):III–IV, December 2006. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X06002174.

Anonymous:2007:PJa

[1074] Anonymous. Pages 1–260 (January 2007). Journal of Number Theory, 122(1):??, January 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2007:EBa

[1075] Anonymous. Editorial Board. Journal of Number Theory, 122(1): CO2, January 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X0600254X.

Cheltsov:2007:TRS

[1076] Ivan Cheltsov and Jihun Park. Two remarks on sextic double solids. Journal of Number Theory, 122(1):1–12, January 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X06000643.

Chiera:2007:PPN

[1077] F. L. Chiera. On Petersson products of not necessarily cuspidal modular forms. Journal of Number Theory, 122(1):13–24, January 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X06000655.

Choie:2007:DOH

[1078] YoungJu Choie, Haesuk Kim, and Olav K. Richter. Differential operators on Hilbert modular forms. Journal of Number Theory, 122(1): 25–36, January 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X06000667.

Ghioca:2007:LIM

[1079] Dragos Ghioca. The Lehmer inequality and the Mordell–Weil theorem for Drinfeld modules. Journal of Number Theory, 122(1):37–68, January 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X06000680.

Magyar:2007:DLP

[1080] Akos Magyar. On the distribution of lattice points on spheres and level surfaces of polynomials. Journal of Number Theory, 122(1):69– 83, January 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X06000692.

Qu:2007:PNT

[1081] Yan Qu. The prime number theorem for automorphic L-functions for GL_{m}. Journal of Number Theory, 122(1):84–99, January 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X06000709.

Bhave:2007:AKF

[1082] Amala Bhave. Analogue of Kida’s formula for certain strongly admissible extensions. Journal of Number Theory, 122(1):100–120, January 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X06000473.

Bandini:2007:GCC

[1083] Andrea Bandini. Greenberg’s conjecture and capitulation in Z_{p}^{d}- extensions. Journal of Number Theory, 122(1):121–134, January 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X06000783.

Cao:2007:IUC

[1084] Wei Cao and Qi Sun. Improvements upon the Chevalley–Warning– Ax–Katz-type estimates. Journal of Number Theory, 122(1):135–141, January 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X06000795.

Dubickas:2007:APL

[1085] Artūras Dubickas. Arithmetical properties of linear recurrent sequences. Journal of Number Theory, 122(1):142–150, January 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X06000801.

Guerzhoy:2007:SCT

[1086] P. Guerzhoy. Some congruences for traces of singular moduli. Journal of Number Theory, 122(1):151–156, January 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X06000813.

Komornik:2007:TSU

[1087] Vilmos Komornik and Paola Loreti. On the topological structure of univoque sets. Journal of Number Theory, 122(1):157–183, January 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X06000837.

Wang:2007:GCU

[1088] Xuejun Wang and Stephen Yau. On the GLY Conjecture of upper estimate of positive integral points in real right-angled simplices. Journal of Number Theory, 122(1):184–210, January 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X06000849.

Yu:2007:UBS

[1089] Gang Yu. An upper bound for B_{2}[g] sets. Journal of Number Theory, 122(1):211–220, January 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X06000850.

Angles:2007:SAP

[1090] Bruno Anglès. On some p-adic power series attached to the arithmetic of Q(ζ_{p}). Journal of Number Theory, 122(1):221–246, January 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X06000862.

Amoroso:2007:ANS

[1091] Francesco Amoroso and Filippo A. E. Nuccio. Algebraic numbers of small Weil’s height in CM-fields: on a theorem of Schinzel. Journal of Number Theory, 122(1):247–260, January 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X06000874.

Anonymous:2007:PF

[1092] Anonymous. Pages 261–518 (February 2007). Journal of Number Theory, 122(2):??, February 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2007:EBb

[1093] Anonymous. Editorial Board. Journal of Number Theory, 122(2):CO2, February 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X06002629.

Rittaud:2007:EPP

[1094] Benoît Rittaud. Équidistribution presque partout modulo 1 de suites oscillantes perturbées III : cas liouvillien multidimensionnel. (French) [Almost everywhere modulo 1 distribution of disturbed oscillating sequences III: multidimensional Liouvillian case]. Journal of Number Theory, 122(2):261–282, February 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X06001090.

Dai:2007:PPF

[1095] Li-Xia Dai and Yong-Gao Chen. On the parity of the partition function. Journal of Number Theory, 122(2):283–289, February 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X06001107.

Chen:2007:PPF

[1096] Yong-Gao Chen and Wei Liu. On the prime power factorization of n!, II. Journal of Number Theory, 122(2):290–300, February 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X06001119.

Choi:2007:WDC

[1097] D. Choi and Y. Choie. Weight-dependent congruence properties of modular forms. Journal of Number Theory, 122(2):301–313, February 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X06001120.

Conrey:2007:CT

[1098] J. B. Conrey, David W. Farmer, B. E. Odgers, and N. C. Snaith. A converse theorem for Γ_{0}(13). Journal of Number Theory, 122(2):314– 323, February 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X06001132.

Hashimoto:2007:AES

[1099] Yasufumi Hashimoto. Arithmetic expressions of Selberg’s zeta functions for congruence subgroups. Journal of Number Theory, 122(2):324–335, February 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X06001144.

Hilberdink:2007:LBL

[1100] Titus W. Hilberdink. A lower bound for the Lindelöf function associated to generalized integers. Journal of Number Theory, 122(2):336– 341, February 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X06001156.

Xue:2007:CVF

[1101] Hui Xue. Central values of L-functions over CM fields. Journal of Number Theory, 122(2):342–378, February 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X06001168.

Jones:2007:GNF

[1102] John W. Jones and David P. Roberts. Galois number fields with small root discriminant. Journal of Number Theory, 122(2):379–407, February 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X0600117X.

Lee:2007:STF

[1103] Yoonjin Lee. The Scholz theorem in function fields. Journal of Number Theory, 122(2):408–414, February 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X06001181.

Liu:2007:GCS

[1104] Huaning Liu and Wenpeng Zhang. Generalized Cochrane sums and Cochrane–Hardy sums. Journal of Number Theory, 122(2):415–428, February 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X0600120X.

Snitz:2007:TCA

[1105] Kobi Snitz. Theta correspondence of automorphic characters. Journal of Number Theory, 122(2):429–465, February 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X06001235.

Tanigawa:2007:DSA

[1106] Yoshio Tanigawa and Wenguang Zhai. Dirichlet series associated with polynomials and applications. Journal of Number Theory, 122(2):466– 518, February 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X06001247.

Anonymous:2007:CVa

[1107] Anonymous. Contents of volume 122. Journal of Number Theory, 122(2):III–IV, February 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X06002678.

Anonymous:2007:PMa

[1108] Anonymous. Pages 1–254 (March 2007). Journal of Number Theory, 123 (1):??, March 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2007:EBc

[1109] Anonymous. Editorial Board. Journal of Number Theory, 123(1): CO2, March 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X07000030.

Nakamura:2007:AIF

[1110] Takashi Nakamura. Applications of inversion formulas to the joint tuniversality of Lerch zeta functions. Journal of Number Theory, 123(1): 1–9, March 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X06001211.

Simons:2007:SIP

[1111] John L. Simons. A simple (inductive) proof for the non-existence of 2-cycles of the 3x + 1 problem. Journal of Number Theory, 123(1):10– 17, March 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X06001223.

Zhao:2007:BNW

[1112] Jianqiang Zhao. Bernoulli numbers, Wolstenholme’s theorem, and p^{5} variations of Lucas’ theorem. Journal of Number Theory, 123(1):18– 26, March 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X06001259.

Yaouanc:2007:RQF

[1113] Erwan Le Yaouanc. On real quadratic function fields. Journal of Number Theory, 123(1):27–34, March 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X06001272.

Kurosawa:2007:TCS

[1114] Takeshi Kurosawa. Transcendence of certain series involving binary linear recurrences. Journal of Number Theory, 123(1):35–58, March 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X06001284.

Li:2007:LEF

[1115] Hua-Chieh Li. Lifting endomorphisms of formal A-modules over finite fields. Journal of Number Theory, 123(1):59–70, March 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X06001296.

Aliev:2007:OLB

[1116] Iskander M. Aliev and Peter M. Gruber. An optimal lower bound for the Frobenius problem. Journal of Number Theory, 123(1):71– 79, March 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X06001302.

Murty:2007:VLC

[1117] V. Kumar Murty. A variant of Lehmer’s conjecture. Journal of Number Theory, 123(1):80–91, March 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X06001314.

vanLuijk:2007:ESA

[1118] Ronald van Luijk. An elliptic K3 surface associated to Heron triangles. Journal of Number Theory, 123(1):92–119, March 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X06001326.

Wright:2007:PQR

[1119] Steve Wright. Patterns of quadratic residues and nonresidues for infinitely many primes. Journal of Number Theory, 123(1):120–132, March 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X06001338.

Allouche:2007:SSD

[1120] J.-P. Allouche, J. Shallit, and J. Sondow. Summation of series defined by counting blocks of digits. Journal of Number Theory, 123(1):133– 143, March 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X0600134X.

Haanpaa:2007:SAG

[1121] Harri Haanpää and Patric R. J. Östergård. Sets in Abelian groups with distinct sums of pairs. Journal of Number Theory, 123(1):144– 153, March 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X06001351.

Bassa:2007:SPL

[1122] Alp Bassa and Henning Stichtenoth. A simplified proof for the limit of a tower over a cubic finite field. Journal of Number Theory, 123(1): 154–169, March 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X06001363.

Deschamps:2007:EBA

[1123] Bruno Deschamps and Bakir Farhi. Essentialité dans les bases additives. (French) [Essentiality in additive databases]. Journal of Number Theory, 123(1):170–192, March 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X06001375.

Itoh:2007:GMU

[1124] Tsuyoshi Itoh and Keiichi Komatsu. On the group of modular units and the ideal class group. Journal of Number Theory, 123(1):193– 203, March 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X06001557.

Koyama:2007:VDS

[1125] Shin ya Koyama and Nobushige Kurokawa. Values of the double sine function. Journal of Number Theory, 123(1):204–223, March 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X06001569.

Moshe:2007:SDB

[1126] Yossi Moshe. On subword decomposition and balanced polynomials. Journal of Number Theory, 123(1):224–240, March 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X06001582.

Pilehrood:2007:CIM

[1127] Kh. Hessami Pilehrood and T. Hessami Pilehrood. On conditional irrationality measures for values of the digamma function. Journal of Number Theory, 123(1):241–253, March 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X06001594.

Anonymous:2007:PAa

[1128] Anonymous. Pages 255–528 (April 2007). Journal of Number Theory, 123(2):??, April 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2007:EBd

[1129] Anonymous. Editorial Board. Journal of Number Theory, 123(2): ifc, April 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X07000200.

Li:2007:DAC

[1130] Ji Li. Determination of a GL_{2} automorphic cuspidal representation by twists of critical L-values. Journal of Number Theory, 123(2):255– 289, April 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X06001570.

Tsishchanka:2007:ARN

[1131] Kiryl I. Tsishchanka. On approximation of real numbers by algebraic numbers of bounded degree. Journal of Number Theory, 123(2):290– 314, April 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X06001600.

Virdol:2007:TCP

[1132] Cristian Virdol. Tate classes and poles of L-functions of twisted quaternionic Shimura surfaces. Journal of Number Theory, 123(2):315– 328, April 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X06001612.

Wu:2007:NPG

[1133] J. Wu. Note on a paper by A. Granville and K. Soundararajan. Journal of Number Theory, 123(2):329–351, April 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X06001624.

Xu:2007:MVT

[1134] Zhefeng Xu, Tianping Zhang, and Wenpeng Zhang. On the mean value of the two-term exponential sums with Dirichlet characters. Journal of Number Theory, 123(2):352–362, April 2007. CODEN JNUTA9.

ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X06001636.

Yamagami:2007:AFH

[1135] Atsushi Yamagami. On p-adic families of Hilbert cusp forms of finite slope. Journal of Number Theory, 123(2):363–387, April 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X06001648.

Arms:2007:COP

[1136] Scott Arms, Álvaro Lozano-Robledo, and Steven J. Miller. Constructing one-parameter families of elliptic curves with moderate rank. Journal of Number Theory, 123(2):388–402, April 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X0600165X.

Alpers:2007:TDP

[1137] Andreas Alpers and Rob Tijdeman. The two-dimensional Prouhet– Tarry–Escott problem. Journal of Number Theory, 123(2):403–412, April 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X06001661.

Banks:2007:PDB

[1138] William D. Banks and Igor E. Shparlinski. Prime divisors in Beatty sequences. Journal of Number Theory, 123(2):413–425, April 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X06001673.

Ghioca:2007:LLI

[1139] Dragos Ghioca. The local Lehmer inequality for Drinfeld modules. Journal of Number Theory, 123(2):426–455, April 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X06001685.

Blache:2007:NSP

[1140] Régis Blache and Éric Férard. Newton stratification for polynomials: the open stratum. Journal of Number Theory, 123(2):456– 472, April 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X06001880.

Ingram:2007:EDS

[1141] Patrick Ingram. Elliptic divisibility sequences over certain curves. Journal of Number Theory, 123(2):473–486, April 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X06001910.

Consani:2007:QSM

[1142] Caterina Consani and Matilde Marcolli. Quantum statistical mechanics over function fields. Journal of Number Theory, 123(2):487– 528, April 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X06002873.

Anonymous:2007:CVb

[1143] Anonymous. Contents of volume 123. Journal of Number Theory, 123(2): III–IV, April 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X0700025X.

Anonymous:2007:PMb

[1144] Anonymous. Pages 1–258 (May 2007). Journal of Number Theory, 124 (1):??, May 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2007:EBe

[1145] Anonymous. Editorial Board. Journal of Number Theory, 124(1): ifc, May 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X07000352.

Zhao:2007:GRB

[1146] Jianqiang Zhao and With appendix by Herbert Gangl. Goncharov’s relations in Bloch’s higher Chow group CH^{3}(F, 5). Journal of Number Theory, 124(1):1–25, May 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X06001260.

Gangl:2007:GTR

[1147] Herbert Gangl. Goncharov’s trilogarithm relation on pictures. Journal of Number Theory, 124(1):17–25, May 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X0700008X.

Nicolae:2007:AFO

[1148] Florin Nicolae. Artin’s L-functions and one-dimensional characters. Journal of Number Theory, 124(1):26–30, May 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X06001922.

Diem:2007:OEC

[1149] Claus Diem and Jasper Scholten. Ordinary elliptic curves of high rank over F̄_{p}(x) with constant j-invariant II. Journal of Number Theory, 124(1):31–41, May 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X06001934.

Jossey:2007:GEU

[1150] John Jossey. Galois 2-extensions unramified outside 2. Journal of Number Theory, 124(1):42–56, May 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X06001946. See corrigendum [1296].

Sun:2007:SAC

[1151] Zhi-Wei Sun. Simple arguments on consecutive power residues. Journal of Number Theory, 124(1):57–61, May 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X06001958.

Sun:2007:CRB

[1152] Zhi-Hong Sun. Cubic residues and binary quadratic forms. Journal of Number Theory, 124(1):62–104, May 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X0600196X.

Agoh:2007:CIL

[1153] Takashi Agoh and Karl Dilcher. Convolution identities and lacunary recurrences for Bernoulli numbers. Journal of Number Theory, 124(1): 105–122, May 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X06002046.

Koike:2007:IFP

[1154] Kenji Koike and Hironori Shiga. Isogeny formulas for the Picard modular form and a three terms arithmetic geometric mean. Journal of Number Theory, 124(1):123–141, May 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X06002058.

Chen:2007:IBT

[1155] Kwang-Wu Chen. Identities from the binomial transform. Journal of Number Theory, 124(1):142–150, May 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X0600223X.

Nathanson:2007:DSN

[1156] Melvyn B. Nathanson and Rohit Parikh. Density of sets of natural numbers and the Lévy group. Journal of Number Theory, 124(1):151– 158, May 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X06002241.

Bartholdi:2007:MUE

[1157] Laurent Bartholdi and Michael R. Bush. Maximal unramified 3extensions of imaginary quadratic fields and SL_{2}(Z_{3}). Journal of Number Theory, 124(1):159–166, May 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X06002253.

Chan:2007:PDB

[1158] Wai Kiu Chan and Anna Rokicki. Positive definite binary Hermitian forms with finitely many exceptions. Journal of Number Theory, 124(1):167–180, May 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X06002265.

Berend:2007:RPF

[1159] D. Berend and G. Kolesnik. Regularity of patterns in the factorization of n!. Journal of Number Theory, 124(1):181–192, May 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X06002289.

Konyagin:2007:TUE

[1160] S. Konyagin and K. Soundararajan. Two S-unit equations with many solutions. Journal of Number Theory, 124(1):193–199, May 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X06002290.

Kane:2007:AMN

[1161] Daniel M. Kane. Asymptotics of McKay numbers for S_{n}. Journal of Number Theory, 124(1):200–228, May 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X06002307.

Vu:2007:SNR

[1162] Van H. Vu. Some new results on subset sums. Journal of Number Theory, 124(1):229–233, May 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X06002319.

Eliahou:2007:SRM

[1163] Shalom Eliahou and Michel Kervaire. Some results on minimal sumset sizes in finite non-abelian groups. Journal of Number Theory, 124(1): 234–247, May 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X06002320.

Faber:2007:FFK

[1164] X. W. C. Faber. On the finite field Kakeya problem in two dimensions. Journal of Number Theory, 124(1):248–257, May 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X06002332.

Anonymous:2007:PJb

[1165] Anonymous. Pages 259–520 (June 2007). Journal of Number Theory, 124 (2):??, June 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2007:EBf

[1166] Anonymous. Editorial Board. Journal of Number Theory, 124(2): ifc, June 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X07000674.

Khan:2007:FMS

[1167] Rizwanur Khan. The first moment of the symmetric-square L-function. Journal of Number Theory, 124(2):259–266, June 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X06002344.

Surroca:2007:EDT

[1168] Andrea Surroca. Sur l’effectivité du théorème de Siegel et la conjecture abc. (French) [On the effectiveness of Siegel ’s theorem and the abc conjecture]. Journal of Number Theory, 124(2):267–290, June 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X06002356.

Tangedal:2007:CFS

[1169] Brett A. Tangedal. Continued fractions, special values of the double sine function, and Stark units over real quadratic fields. Journal of Number Theory, 124(2):291–313, June 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X06002368.

Hegyvari:2007:MSS

[1170] Norbert Hegyvári and François Hennecart. On monochromatic sums of squares and primes. Journal of Number Theory, 124(2):314–324, June 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X0600237X.

Xia:2007:NHC

[1171] Honggang Xia. On L^{∞} norms of holomorphic cusp forms. Journal of Number Theory, 124(2):325–327, June 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X06002381.

Gon:2007:MMM

[1172] Yasuro Gon, Nobushige Kurokawa, and Hideo Oyanagi. Multiple q-Mahler measures and zeta functions. Journal of Number Theory, 124(2): 328–345, June 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X06002393.

Cenkci:2007:DCD

[1173] Mehmet Cenkci, Mümün Can, and Veli Kurt. Degenerate and character Dedekind sums. Journal of Number Theory, 124(2):346–363, June 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X0600240X.

Dabrowski:2007:E

[1174] Andrzej Da̧browski and Małgorzata Wieczorek. On the equation y^{2} = x(x − 2^{m})(x + q − 2^{m}). Journal of Number Theory, 124(2):364– 379, June 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X06002411.

Mihailescu:2007:NBC

[1175] Preda Mihăilescu. New bounds and conditions for the equation of Nagell–Ljunggren. Journal of Number Theory, 124(2):380–395, June 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X06002435.

Pacetti:2007:SCL

[1176] Ariel Pacetti and Gonzalo Tornaría. Shimura correspondence for level p^{2} and the central values of L-series. Journal of Number Theory, 124(2): 396–414, June 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X06002447.

Xia:2007:ZCF

[1177] Honggang Xia. On zeros of cubic L-functions. Journal of Number Theory, 124(2):415–428, June 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X06002459.

Cohen:2007:PRQ

[1178] Joseph Cohen. Primitive roots in quadratic fields, II. Journal of Number Theory, 124(2):429–441, June 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X06002460.

Elsner:2007:STI

[1179] Carsten Elsner. On a sequence transformation with integral coefficients for Euler’s constant, II. Journal of Number Theory, 124(2):442– 453, June 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X06002472.

Ghosh:2007:MDA

[1180] Anish Ghosh. Metric Diophantine approximation over a local field of positive characteristic. Journal of Number Theory, 124(2):454– 469, June 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X06002484.

Klazar:2007:MON

[1181] Martin Klazar and Florian Luca. On the maximal order of numbers in the “factorisatio numerorum” problem. Journal of Number Theory, 124(2):470–490, June 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X06002496.

Alaca:2007:ECS

[1182] Şaban Alaca and Kenneth S. Williams. Evaluation of the convolution sums ∑ +l + 6m = nσ(l)σ(m) and ∑_{2l+3m=n} σ(l)σ(m). Journal of Number Theory, 124(2):491–510, June 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X06002502.

Bremner:2007:SLS

[1183] A. Bremner and N. Tzanakis. On squares in Lucas sequences. Journal of Number Theory, 124(2):511–520, June 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X06002514.

Anonymous:2007:CVc

[1184] Anonymous. Contents of volume 124. Journal of Number Theory, 124 (2):III–IV, June 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X07000728.

Anonymous:2007:PJc

[1185] Anonymous. Pages 1–266 (July 2007). Journal of Number Theory, 125 (1):??, July 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2007:EBg

[1186] Anonymous. Editorial Board. Journal of Number Theory, 125(1): ifc, July 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X07000777.

Kim:2007:SCT

[1187] Myung-Hwan Kim and Sung-Geun Lim. Square classes of totally positive units. Journal of Number Theory, 125(1):1–6, July 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X06000825.

Liu:2007:NUB

[1188] Huaning Liu. A note on the upper bound estimate of high-dimensional Cochrane sums. Journal of Number Theory, 125(1):7–13, July 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X06002423.

Chen:2007:IF

[1189] Yong-Gao Chen. On integers of the forms k ± 2^{n} and k2^{n} ± 1. Journal of Number Theory, 125(1):14–25, July 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X06002526.

Popescu:2007:TBG

[1190] Cristian D. Popescu, Jack Sonn, and Adrian R. Wadsworth. n-Torsion of Brauer groups as relative Brauer groups of abelian extensions. Journal of Number Theory, 125(1):26–38, July 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X06002721.

Elizalde:2007:PCP

[1191] Sergi Elizalde and Kevin Woods. The probability of choosing primitive sets. Journal of Number Theory, 125(1):39–49, July 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X06002733.

Omar:2007:CLH

[1192] Sami Omar and Kamel Mazhouda. Le critère de Li et l’hypothèse de Riemann pour la classe de Selberg. (French) [Li’s criterion and Riemann’s hypothesis for Selberg’s class]. Journal of Number Theory, 125(1):50–58, July 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X06002757. See corrigendum and ad-

dendum [1709].

Park:2007:CNO

[1193] Sun-Mi Park and Soun-Hi Kwon. Class number one problem for normal CM-fields. Journal of Number Theory, 125(1):59–84, July 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X06002769.

Ghioca:2007:TVC

[1194] Dragos Ghioca. The Tate–Voloch conjecture for Drinfeld modules. Journal of Number Theory, 125(1):85–94, July 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X06002770.

Desrousseaux:2007:VFP

[1195] P. A. Desrousseaux. Valeurs d’une fonction de Picard en des points algébriques. (French) [Values of a function of Picard at algebraic points]. Journal of Number Theory, 125(1):95–116, July 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X06002782.

Okazaki:2007:F

[1196] Takeo Okazaki. L-functions of S_{3}(Γ(4, 8)). Journal of Number Theory, 125(1):117–132, July 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X06002794.

Laubie:2007:SDL

[1197] François Laubie. Systèmes dynamiques et lois de groupe formel sur les corps finis. (French) [Dynamical systems and formal group laws on finite bodies]. Journal of Number Theory, 125(1):133–148, July 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X06002800.

Urtis:2007:SVF

[1198] Çetin Ürtiş. Special values of L-functions by a Siegel–Weil–Kudla–Rallis formula. Journal of Number Theory, 125(1):149–181, July 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X06002812.

Tachiya:2007:TCI

[1199] Yohei Tachiya. Transcendence of certain infinite products. Journal of Number Theory, 125(1):182–200, July 2007. CODEN JNUTA9.

ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X06002824.

Garaev:2007:NSD

[1200] M. Z. Garaev. On the number of solutions of a diophantine equation with symmetric entries. Journal of Number Theory, 125(1):201– 209, July 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X06002848.

An:2007:ESS

[1201] Ta Thi Hoai An and Julie Tzu-Yueh Wang. An effective Schmidt’s subspace theorem for non-linear forms over function fields. Journal of Number Theory, 125(1):210–228, July 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X0600285X.

Vojta:2007:NCR

[1202] Paul Vojta. On the Nochka–Chen–Ru–Wong proof of Cartan’s conjecture. Journal of Number Theory, 125(1):229–234, July 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X06002861.

Steurer:2007:GGC

[1203] Aliza Steurer. On the Galois groups of the 2-class towers of some imaginary quadratic fields. Journal of Number Theory, 125(1):235– 246, July 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X06002885.

Lamzouri:2007:NLF

[1204] Youness Lamzouri. On the number of linear forms in logarithms. Journal of Number Theory, 125(1):247–253, July 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X06002897.

Munshi:2007:DPR

[1205] Ritabrata Munshi. Density of positive rank fibers in elliptic fibrations. Journal of Number Theory, 125(1):254–266, July 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X06002903.

Anonymous:2007:PAb

[1206] Anonymous. Pages 267–528 (August 2007). Journal of Number Theory, 125(2):??, August 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2007:EBh

[1207] Anonymous. Editorial Board. Journal of Number Theory, 125(2): ifc, August 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X07000972.

Liu:2007:FTW

[1208] Chunlei Liu. The L-functions of twisted Witt extensions. Journal of Number Theory, 125(2):267–284, August 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X06001193.

Berenhaut:2007:NMC

[1209] K. S. Berenhaut and F. Saidak. A note on the maximal coefficients of squares of Newman polynomials. Journal of Number Theory, 125(2):285–288, August 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X06002745.

Cipra:2007:HCW

[1210] James Arthur Cipra, Todd Cochrane, and Christopher Pinner. Heilbronn’s conjecture on Waring’s number (modp). Journal of Number Theory, 125(2):289–297, August 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X06002915.

Murty:2007:TVD

[1211] M. Ram Murty and N. Saradha. Transcendental values of the digamma function. Journal of Number Theory, 125(2):298–318, August 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X06002927.

Knafo:2007:GSF

[1212] Emmanuel Knafo. On a generalization of the Selberg formula. Journal of Number Theory, 125(2):319–343, August 2007. CODEN JNUTA9.

ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X06002939.

Offen:2007:URD

[1213] Omer Offen and Eitan Sayag. On unitary representations of GL_{2n} distinguished by the symplectic group. Journal of Number Theory, 125(2):344–355, August 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X06002940.

Cipu:2007:NSS

[1214] Mihai Cipu and Maurice Mignotte. On the number of solutions to systems of Pell equations. Journal of Number Theory, 125(2):356– 392, August 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X06002952.

Farhi:2007:NLB

[1215] Bakir Farhi. Nontrivial lower bounds for the least common multiple of some finite sequences of integers. Journal of Number Theory, 125(2):393–411, August 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X06002964.

Pinter:2007:PVP

[1216] Ákos Pintér. On the power values of power sums. Journal of Number Theory, 125(2):412–423, August 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X06002976.

Nakamura:2007:ENE

[1217] Takashi Nakamura. The existence and the non-existence of joint tuniversality for Lerch zeta functions. Journal of Number Theory, 125(2):424–441, August 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X06002988.

Miyata:2007:GSI

[1218] Yoshimasa Miyata. On Galois structure of the integers in elementary abelian extensions of local number fields. Journal of Number Theory, 125(2):442–458, August 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X0600299X.

Luca:2007:AFL

[1219] Florian Luca and Igor E. Shparlinski. Arithmetic functions with linear recurrence sequences. Journal of Number Theory, 125(2):459–472, August 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X06003003.

Hayashi:2007:CWR

[1220] Tsukasa Hayashi. Computation of Weng’s rank 2 zeta function over an algebraic number field. Journal of Number Theory, 125(2):473– 527, August 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X06003039.

Anonymous:2007:CVd

[1221] Anonymous. Contents of volume 125. Journal of Number Theory, 125(2):III–IV, August 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X07001011.

Anonymous:2007:PS

[1222] Anonymous. Pages 1–154 (September 2007). Journal of Number Theory, 126(1):??, September 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2007:EBi

[1223] Anonymous. Editorial Board. Journal of Number Theory, 126(1):ifc, September 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X0700131X.

Grabitz:2007:SCP

[1224] Martin Grabitz. Simple characters for principal orders in M_{m}(D). Journal of Number Theory, 126(1):1–51, September 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X06001387.

Ionascu:2007:HTT

[1225] Eugen J. Ionascu, Florian Luca, and Pantelimon Stănică. Heron triangles with two fixed sides. Journal of Number Theory, 126(1):52–67, September 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X06003015.

Shparlinski:2007:BIM

[1226] Igor E. Shparlinski. Bounds of incomplete multiple Kloosterman sums. Journal of Number Theory, 126(1):68–73, September 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X06003027.

Hanke:2007:AGS

[1227] Timo Hanke. On absolute Galois splitting fields of central simple algebras. Journal of Number Theory, 126(1):74–86, September 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X06003040.

Murray:2007:EMR

[1228] Brian J. Murray. Explicit multiplicative relations between Gauss sums. Journal of Number Theory, 126(1):87–109, September 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X06003052.

Alon:2007:LSF

[1229] Noga Alon. Large sets in finite fields are sumsets. Journal of Number Theory, 126(1):110–118, September 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X0700011X.

Postelmans:2007:I

[1230] Kelly Postelmans and Walter Van Assche. Irrationality of ζ_{q}(1) and ζ_{q}(2). Journal of Number Theory, 126(1):119–154, September 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X07000121.

Anonymous:2007:PO

[1231] Anonymous. Pages 155–296 (October 2007). Journal of Number Theory, 126(2):??, October 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2007:EBj

[1232] Anonymous. Editorial Board. Journal of Number Theory, 126(2): ifc, October 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X07001527.

Liu:2007:PGR

[1233] Tong Liu. Potentially good reduction of Barsotti–Tate groups. Journal of Number Theory, 126(2):155–184, October 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X07000133.

Friesen:2007:RFF

[1234] Christian Friesen. Rational functions over finite fields having continued fraction expansions with linear partial quotients. Journal of Number Theory, 126(2):185–192, October 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X07000145.

Ahlgren:2007:CMF

[1235] Scott Ahlgren and Mugurel Barcau. Congruences for modular forms of weights two and four. Journal of Number Theory, 126(2):193–199, October 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X07000157.

Chamizo:2007:VLP

[1236] Fernando Chamizo, Elena Cristóbal, and Adrián Ubis. Visible lattice points in the sphere. Journal of Number Theory, 126(2):200–211, October 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X07000169.

Allen:2007:MLR

[1237] Patrick B. Allen. On the multiplicity of linear recurrence sequences. Journal of Number Theory, 126(2):212–216, October 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X07000170.

Bass:2007:IEG

[1238] Jared Bass. Improving the Erdős–Ginzburg–Ziv theorem for some non-abelian groups. Journal of Number Theory, 126(2):217–236, October 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X07000182.

Croot:2007:CMC

[1239] Ernie Croot. On a combinatorial method for counting smooth numbers in sets of integers. Journal of Number Theory, 126(2):237–253, October 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X07000297.

Constales:2007:HAF

[1240] D. Constales, R. S. Kraußhar, and John Ryan. k-Hypermonogenic automorphic forms. Journal of Number Theory, 126(2):254–271, October 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X07000303.

Tzermias:2007:EVT

[1241] Pavlos Tzermias. An explicit version of a theorem of Stickelberger. Journal of Number Theory, 126(2):272–286, October 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X07000315.

Sun:2007:CSB

[1242] Zhi-Wei Sun and Roberto Tauraso. Congruences for sums of binomial coefficients. Journal of Number Theory, 126(2):287–296, October 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X07000327.

Anonymous:2007:CVe

[1243] Anonymous. Contents of volume 126. Journal of Number Theory, 126 (2):III, October 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X07001564.

Anonymous:2007:PN

[1244] Anonymous. Pages 1–144 (November 2007). Journal of Number Theory, 127(1):??, November 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2007:EBk

[1245] Anonymous. Editorial Board. Journal of Number Theory, 127(1):ifc, November 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X07001709.

Liu:2007:EHE

[1246] Sheng-Chi Liu. Equidistribution of Hecke eigenforms on the Hilbert modular varieties. Journal of Number Theory, 127(1):1–9, November 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X07000339.

Rosen:2007:IHP

[1247] Michael Rosen and Joseph H. Silverman. On the independence of Heegner points associated to distinct quadratic imaginary fields. Journal of Number Theory, 127(1):10–36, November 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X07000431.

Languasco:2007:NMF

[1248] A. Languasco and A. Zaccagnini. A note on Mertens’ formula for arithmetic progressions. Journal of Number Theory, 127(1):37–46, November 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X07000443.

Sun:2007:HIT

[1249] Hae-Sang Sun. Homological interpretation of a theorem of L. Washington. Journal of Number Theory, 127(1):47–63, November 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X07000455.

Fan:2007:AMP

[1250] Ai-Hua Fan, Bao-Wei Wang, and Jun Wu. Arithmetic and metric properties of Oppenheim continued fraction expansions. Journal of Number Theory, 127(1):64–82, November 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X07000467.

Yang:2007:IML

[1251] Jae-Hyun Yang. Invariant metrics and Laplacians on Siegel–Jacobi space. Journal of Number Theory, 127(1):83–102, November 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X07000479.

Zaimi:2007:APB

[1252] Toufik Zaimi. Approximation by polynomials with bounded coefficients. Journal of Number Theory, 127(1):103–117, November 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X07000480.

Kaplan:2007:FCP

[1253] Nathan Kaplan. Flat cyclotomic polynomials of order three. Journal of Number Theory, 127(1):118–126, November 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X07000492.

Rodseth:2007:DPC

[1254] Øystein J. Rødseth. Distribution of points on the circle. Journal of Number Theory, 127(1):127–135, November 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X07000509. See

addendum [1391].

Burshtein:2007:EDO

[1255] Nechemia Burshtein. The equation ∑_{i=1}^{9} \frac{1}{x_{i}} = 1 in distinct odd integers has only the five known solutions. Journal of Number Theory, 127(1):136–144, November 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X07000510.

Anonymous:2007:PD

[1256] Anonymous. Pages 145–330 (December 2007). Journal of Number Theory, 127(2):??, December 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2007:EBl

[1257] Anonymous. Editorial Board. Journal of Number Theory, 127(2):ifc, December 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X0700193X.

Beiglbock:2007:SCS

[1258] Mathias Beiglböck. Strong characterizing sequences of countable groups. Journal of Number Theory, 127(2):145–152, December 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X07000923.

Berthe:2007:PPE

[1259] Valerie Berthé and Anne Siegel. Purely periodic β-expansions in the Pisot non-unit case. Journal of Number Theory, 127(2):153–172, December 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X07001643.

Bettner:2007:EFF

[1260] Stefan Bettner. An explicit factorization formula for the singular values of ϕ and the generalized principal ideal theorem. Journal of Number Theory, 127(2):173–183, December 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X05002350.

Laughlin:2007:SSC

[1261] J. Mc Laughlin. Symmetry and specializability in the continued fraction expansions of some infinite products. Journal of Number Theory, 127(2):184–219, December 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X07000935.

Gaudron:2007:EDC

[1262] Éric Gaudron. Étude du cas rationnel de la théorie des formes linéaires de logarithmes. (French) [Study of the rational case of the theory of linear forms of logarithms]. Journal of Number Theory, 127(2):220–261, December 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X07001667.

Gu:2007:OII

[1263] Haihua Gu, Dawu Gu, and Ya Liu. Ono invariants of imaginary quadratic fields with class number three. Journal of Number Theory, 127(2):262–271, December 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X07001370.

Davison:2007:QPC

[1264] J. L. Davison. Quasi-periodic continued fractions. Journal of Number Theory, 127(2):272–282, December 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X0700162X.

Zhang:2007:CEN

[1265] Wenpeng Zhang and Zhefeng Xu. On a conjecture of the Euler numbers. Journal of Number Theory, 127(2):283–291, December 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X07000947.

vanFrankenhuijsen:2007:AIR

[1266] Machiel van Frankenhuijsen. ABC implies the radicalized Vojta height inequality for curves. Journal of Number Theory, 127(2):292–300, December 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X07000753.

Datskovsky:2007:MVT

[1267] Boris A. Datskovsky and Behailu Mammo. A mean value theorem for discriminants of abelian extensions of a number field. Journal of Number Theory, 127(2):301–325, December 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X07001382.

Wagstaff:2007:CM

[1268] Samuel S. Wagstaff. Congruences for r_{s}(n) modulo 2s. Journal of Number Theory, 127(2):326–329, December 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X07001151.

Anonymous:2007:CVf

[1269] Anonymous. Contents of volume 127. Journal of Number Theory, 127(2):III–IV, December 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X07001977.

Anonymous:2008:PJa

[1270] Anonymous. Pages 1–234 (January 2008). Journal of Number Theory, 128(1):??, January 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2008:EBa

[1271] Anonymous. Editorial Board. Journal of Number Theory, 128(1): ifc, January 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X07002259.

Matala-aho:2008:IM

[1272] Tapani Matala-aho. On irrationality measures of ∑_{l=0}^{∞} d^{l}/ ∏_{j=1}^{l}(1+d^{j}r+ d^{2j}s). Journal of Number Theory, 128(1):1–16, January 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X07001680.

Buchholz:2008:CPR

[1273] Ralph H. Buchholz and James A. MacDougall. Cyclic polygons with rational sides and area. Journal of Number Theory, 128(1):17–48, January 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X07001126.

Mattarei:2008:LRR

[1274] Sandro Mattarei. Linear recurrence relations for binomial coefficients modulo a prime. Journal of Number Theory, 128(1):49–58, January 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X0700114X.

Filaseta:2008:GGP

[1275] M. Filaseta, F. Luca, P. Stănică, and R. G. Underwood. Galois groups of polynomials arising from circulant matrices. Journal of Number Theory, 128(1):59–70, January 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X07000959.

Mao:2008:GSC

[1276] Zhengyu Mao. A generalized Shimura correspondence for newforms. Journal of Number Theory, 128(1):71–95, January 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X07000613.

Levin:2008:ZDI

[1277] Aaron Levin. On the Zariski-density of integral points on a complement of hyperplanes in P^{n}. Journal of Number Theory, 128(1):96– 104, January 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X07001138.

Solomon:2008:TZF

[1278] D. Solomon. On twisted zeta-functions at s = 0 and partial zetafunctions at s = 1. Journal of Number Theory, 128(1):105–143, January 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X07000650.

Burger:2008:SPL

[1279] Edward B. Burger, Jesse Gell-Redman, Ross Kravitz, Daniel Walton, and Nicholas Yates. Shrinking the period lengths of continued fractions while still capturing convergents. Journal of Number Theory, 128(1):144–153, January 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X07000595.

Baier:2008:ILS

[1280] Stephan Baier and Liangyi Zhao. An improvement for the large sieve for square moduli. Journal of Number Theory, 128(1):154–174, January 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X07000625.

Baruah:2008:MRN

[1281] Nayandeep Deka Baruah and Jonali Bora. Modular relations for the nonic analogues of the Rogers–Ramanujan functions with applications to partitions. Journal of Number Theory, 128(1):175–206, January 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X07000583.

Yasaki:2008:ESP

[1282] Dan Yasaki. An explicit spine for the Picard modular group over the Gaussian integers. Journal of Number Theory, 128(1):207–234, January 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X0700056X.

Anonymous:2008:PF

[1283] Anonymous. Pages 235–452 (February 2008). Journal of Number Theory, 128(2):??, February 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2008:EBb

[1284] Anonymous. Editorial Board. Journal of Number Theory, 128(2):ifc, February 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X07002326.

Kozima:2008:GPF

[1285] Noritomo Kozima. Garrett’s pullback formula for vector valued Siegel modular forms. Journal of Number Theory, 128(2):235–250, February 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X07000285.

Chisholm:2008:RTE

[1286] C. Chisholm and J. A. MacDougall. Rational tetrahedra with edges in geometric progression. Journal of Number Theory, 128(2):251–262, February 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X07001631.

Uchida:2008:DBO

[1287] Yukihiro Uchida. The difference between the ordinary height and the canonical height on elliptic curves. Journal of Number Theory, 128(2):263–279, February 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X07002168.

Sun:2008:CIB

[1288] Zhi-Hong Sun. Congruences involving Bernoulli and Euler numbers. Journal of Number Theory, 128(2):280–312, February 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X07000649.

Shi:2008:EMV

[1289] Sanying Shi. On the equation n_{1}n_{2} = n_{3}n_{4} and mean value of character sums. Journal of Number Theory, 128(2):313–321, February 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X0700217X.

Fujita:2008:EDP

[1290] Yasutsugu Fujita. The extensibility of Diophantine pairs {k − 1, k + 1}. Journal of Number Theory, 128(2):322–353, February 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X07001114.

Kim:2008:LSM

[1291] Henry H. Kim. Langlands–Shahidi method and poles of automorphic L-functions III: Exceptional groups. Journal of Number Theory, 128(2):354–376, February 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X07001795.

Gonzalez:2008:CFI

[1292] Josep González. On cubic factors of j-invariants of quadratic Qcurves of prime degree. Journal of Number Theory, 128(2):377–389, February 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X07001497.

Benito:2008:MIF

[1293] Manuel Benito, Luis M. Navas, and Juan L. Varona. Möbius inversion formulas for flows of arithmetic semigroups. Journal of Number Theory, 128(2):390–412, February 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X07001205.

Cheon:2008:GHN

[1294] Gi-Sang Cheon and M. E. A. El-Mikkawy. Generalized harmonic numbers with Riordan arrays. Journal of Number Theory, 128(2):413–425, February 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X07002004.

Yamamoto:2008:KLF

[1295] Shuji Yamamoto. On Kronecker limit formulas for real quadratic fields. Journal of Number Theory, 128(2):426–450, February 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X07001485.

Jossey:2008:CGE

[1296] John Jossey. Corrigendum to “Galois 2-extensions unramified outside 2” [J. Number Theory 124 (2007) 42–56]. Journal of Number Theory, 128(2):451–452, February 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X07000819. See [1150].

Anonymous:2008:PMa

[1297] Anonymous. Pages 453–708 (March 2008). Journal of Number Theory, 128(3):??, March 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2008:EBc

[1298] Anonymous. Editorial Board. Journal of Number Theory, 128(3): ifc, March 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X08000115.

Rosales:2008:SSP

[1299] J. C. Rosales, P. A. García-Sánchez, and J. M. Urbano-Blanco. The set of solutions of a proportionally modular Diophantine inequality. Journal of Number Theory, 128(3):453–467, March 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X07002302.

Xu:2008:EPP

[1300] Kejian Xu and Yongliang Wang. On the elements of prime power order in K_{2}(Q). Journal of Number Theory, 128(3):468–474, March 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X07002156.

Helou:2008:WTC

[1301] Charles Helou and Guy Terjanian. On Wolstenholme’s theorem and its converse. Journal of Number Theory, 128(3):475–499, March 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X07001771.

Homma:2008:DUD

[1302] Kosuke Homma. On the discrepancy of uniformly distributed roots of quadratic congruences. Journal of Number Theory, 128(3):500– 508, March 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X07002223.

Ng:2008:LGB

[1303] Nathan Ng. Large gaps between the zeros of the Riemann zeta function. Journal of Number Theory, 128(3):509–556, March 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X07001059.

Koninck:2008:IDS

[1304] Jean-Marie De Koninck and Florian Luca. Integers divisible by sums of powers of their prime factors. Journal of Number Theory, 128(3):557– 563, March 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X07000522.

Pekker:2008:SMA

[1305] Alexander Pekker. On successive minima and the Absolute Siegel’s Lemma. Journal of Number Theory, 128(3):564–575, March 2008. CO-DEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL

http://www.sciencedirect.com/science/article/pii/S0022314X07001163.

Murabayashi:2008:CCC

[1306] Naoki Murabayashi. On construction of certain CM elliptic curves. Journal of Number Theory, 128(3):576–588, March 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X07001916.

Jarvis:2008:MSE

[1307] Frazer Jarvis and Jayanta Manoharmayum. On the modularity of supersingular elliptic curves over certain totally real number fields. Journal of Number Theory, 128(3):589–618, March 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X07002181.

Vauclair:2008:NTC

[1308] David Vauclair. Noyaux de Tate et capitulation. (French) [Tate kernels and capitulation]. Journal of Number Theory, 128(3):619–638, March 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X07000601.

Kane:2008:LBS

[1309] Daniel M. Kane. On lower bounds on the size of sums-of-squares formulas. Journal of Number Theory, 128(3):639–644, March 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X07002065.

Urban:2008:AND

[1310] Roman Urban. Algebraic numbers and density modulo 1. Journal of Number Theory, 128(3):645–661, March 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X07001503. See

corrigendum [1688].

Dokchitser:2008:PRE

[1311] Tim Dokchitser and Vladimir Dokchitser. Parity of ranks for elliptic curves with a cyclic isogeny. Journal of Number Theory, 128(3):662– 679, March 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X07000881.

Chan:2008:EST

[1312] Heng Huat Chan and Shaun Cooper. Eisenstein series and theta functions to the septic base. Journal of Number Theory, 128(3):680– 699, March 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X07002053.

Choi:2008:EFT

[1313] Dohoon Choi, Daeyeol Jeon, Soon-Yi Kang, and Chang Heon Kim. Exact formulas for traces of singular moduli of higher level modular functions. Journal of Number Theory, 128(3):700–707, March 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X07000911.

Anonymous:2008:PAa

[1314] Anonymous. Pages 709–1090 (April 2008). Journal of Number Theory, 128(4):??, April 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2008:EBd

[1315] Anonymous. Editorial Board. Journal of Number Theory, 128(4): ifc, April 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X08000322.

Elliott:2008:RSG

[1316] Jesse Elliott. Ring structures on groups of arithmetic functions. Journal of Number Theory, 128(4):709–730, April 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X07001886.

Hubbard:2008:DSE

[1317] David Hubbard. Dihedral side extensions and class groups. Journal of Number Theory, 128(4):731–737, April 2008. CODEN JNUTA9.

ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X07001175.

Young:2008:DBP

[1318] Paul Thomas Young. Degenerate Bernoulli polynomials, generalized factorial sums, and their applications. Journal of Number Theory, 128(4): 738–758, April 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X07000893.

Thillainatesan:2008:KAF

[1319] Meera Thillainatesan. A kernel for automorphic L-functions. Journal of Number Theory, 128(4):759–780, April 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X07001412.

Fukuhara:2008:DSP

[1320] Shinji Fukuhara. Dedekind symbols with plus reciprocity laws. Journal of Number Theory, 128(4):781–795, April 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X07001394.

Luca:2008:CGQ

[1321] Florian Luca and Allison M. Pacelli. Class groups of quadratic fields of 3-rank at least 2: effective bounds. Journal of Number Theory, 128(4):796–804, April 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X07001400.

Lu:2008:SPF

[1322] Guangshi Lü. On sums of a prime and four prime squares in short intervals. Journal of Number Theory, 128(4):805–819, April 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X0700090X.

Machide:2008:SPK

[1323] T. Machide. Sums of products of Kronecker’s double series. Journal of Number Theory, 128(4):820–834, April 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X0800005X.

Hiroe:2008:HSP

[1324] Kazuki Hiroe and Takayuki Oda. Hecke–Siegel’s pull-back formula for the Epstein zeta function with a harmonic polynomial. Journal of Number Theory, 128(4):835–857, April 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X07002077.

Ichimura:2008:HSN

[1325] Humio Ichimura. Hilbert–Speiser number fields for a prime p inside the p-cyclotomic field. Journal of Number Theory, 128(4):858– 864, April 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X0700176X.

Byeon:2008:CNP

[1326] Dongho Byeon and Jungyun Lee. Class number 2 problem for certain real quadratic fields of Richaud–Degert type. Journal of Number Theory, 128(4):865–883, April 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X0700087X.

Akhtari:2008:LCP

[1327] Shabnam Akhtari and Stephen K. Choi. On the Littlewood cyclotomic polynomials. Journal of Number Theory, 128(4):884–894, April 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X07000856.

Murabayashi:2008:MCE

[1328] Naoki Murabayashi. Modularity of CM elliptic curves over division fields. Journal of Number Theory, 128(4):895–897, April 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X07000534.

Mizuno:2008:FCE

[1329] Yoshinori Mizuno. On Fourier coefficients of Eisenstein series and Niebur Poincaré series of integral weight. Journal of Number Theory, 128(4):898–909, April 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X07002375.

Steele:2008:CNN

[1330] G. Ander Steele. Carmichael numbers in number rings. Journal of Number Theory, 128(4):910–917, April 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X07002089.

Wright:2008:QRC

[1331] Steve Wright. Quadratic residues and the combinatorics of sign multiplication. Journal of Number Theory, 128(4):918–925, April 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X07001618.

Madore:2008:ERH

[1332] David A. Madore. Équivalence rationnelle sur les hypersurfaces cubiques de mauvaise réduction. (French) [Rational equivalence on cubic hypersurfaces of poor reduction]. Journal of Number Theory, 128(4):926– 944, April 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X07000571.

Kim:2008:ECM

[1333] Daeyeoul Kim, Ja Kyung Koo, and Yoon Kyung Park. On the elliptic curves modulo p. Journal of Number Theory, 128(4):945– 953, April 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X07001448.

Bruche:2008:RGM

[1334] Clement Bruche and Bouchaïb Sodaïgui. On realizable Galois module classes and Steinitz classes of nonabelian extensions. Journal of Number Theory, 128(4):954–978, April 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X07001102.

Buium:2008:DE

[1335] Alexandru Buium. Differential eigenforms. Journal of Number Theory, 128(4):979–1010, April 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X07002211.

Liu:2008:BDH

[1336] H.-Q. Liu. Barban–Davenport–Halberstam average sum and exceptional zero of L-functions. Journal of Number Theory, 128(4):1011– 1043, April 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X07001679.

Luca:2008:PNZ

[1337] Florian Luca and József Sándor. On a problem of Nicol and Zhang. Journal of Number Theory, 128(4):1044–1059, April 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X07001060.

Machide:2008:EAG

[1338] Tomoya Machide. An elliptic analogue of generalized Dedekind– Rademacher sums. Journal of Number Theory, 128(4):1060–1073, April 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X07001047.

Policky:2008:ICU

[1339] Zdeněk Polický. On the index of circular units in the full group of units of a compositum of quadratic fields. Journal of Number Theory, 128(4):1074–1090, April 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X07001837.

Anonymous:2008:PMb

[1340] Anonymous. Pages 1091–1388 (May 2008). Journal of Number Theory, 128(5):??, May 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2008:EBe

[1341] Anonymous. Editorial Board. Journal of Number Theory, 128(5): ifc, May 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X08000395.

Hattori:2008:TCR

[1342] Shin Hattori. Tame characters and ramification of finite flat group schemes. Journal of Number Theory, 128(5):1091–1108, May 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X07001199.

Lasjaunias:2008:REC

[1343] Alain Lasjaunias. On Robbins’ example of a continued fraction expansion for a quartic power series over F_{13}. Journal of Number Theory, 128(5):1109–1115, May 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X07000546.

Wang:2008:ASM

[1344] Yonghui Wang. The analytic strong multiplicity one theorem for GL_{m}(A_{K}). Journal of Number Theory, 128(5):1116–1126, May 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X07001461.

Madritsch:2008:NNG

[1345] Manfred G. Madritsch, Jörg M. Thuswaldner, and Robert F. Tichy. Normality of numbers generated by the values of entire functions. Journal of Number Theory, 128(5):1127–1145, May 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X07001072.

Javaheri:2008:GDA

[1346] Mohammad Javaheri. A generalization of Dirichlet approximation theorem for the affine actions on real line. Journal of Number Theory, 128(5):1146–1156, May 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X07002016.

Stoll:2008:CDD

[1347] Thomas Stoll. Complete decomposition of Dickson-type polynomials and related Diophantine equations. Journal of Number Theory, 128(5):1157–1181, May 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X07000844.

Chan:2008:ARS

[1348] Tsz Ho Chan. Approximating reals by sums of two rationals. Journal of Number Theory, 128(5):1182–1194, May 2008. CODEN JNUTA9.

ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X07001783.

Matomaki:2008:BGP

[1349] Kaisa Matomäki. The binary Goldbach problem with one prime of the form p = k^{2} + l^{2} + 1. Journal of Number Theory, 128(5):1195– 1210, May 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X07000558.

Sunic:2008:FPD

[1350] Zoran Šunić. Frobenius Problem and dead ends in integers. Journal of Number Theory, 128(5):1211–1223, May 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X07001187.

Shparlinski:2008:NDB

[1351] Igor E. Shparlinski and Arne Winterhof. On the number of distances between the coordinates of points on modular hyperbolas. Journal of Number Theory, 128(5):1224–1230, May 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X07001898.

Lalin:2008:MMC

[1352] Matilde N. Lalín. Mahler measures and computations with regulators. Journal of Number Theory, 128(5):1231–1271, May 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X07000637.

Billerey:2008:FHD

[1353] Nicolas Billerey. Formes homogènes de degré 3 et puissances p-ièmes. (French) [Homogeneous forms of degree 3 and powers of p]. Journal of Number Theory, 128(5):1272–1294, May 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X08000036.

Sun:2008:QCQ

[1354] Zhi-Hong Sun. On the quadratic character of quadratic units. Journal of Number Theory, 128(5):1295–1335, May 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X07001849.

Cai:2008:CTI

[1355] Yingchun Cai. On Chen’s theorem (II). Journal of Number Theory, 128(5):1336–1357, May 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X07001217.

Brenner:2008:SLG

[1356] Eliot Brenner. Stability of the local gamma factor in the unitary case. Journal of Number Theory, 128(5):1358–1375, May 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X07002387.

Bartolo:2008:WVF

[1357] A. Di Bartolo and G. Falcone. Witt vectors and Fermat quotients. Journal of Number Theory, 128(5):1376–1387, May 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X07002363.

Anonymous:2008:PJb

[1358] Anonymous. Pages 1389–1892 (June 2008). Journal of Number Theory, 128(6):??, June 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2008:EBf

[1359] Anonymous. Editorial Board. Journal of Number Theory, 128(6): ifc, June 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X08000693.

Petho:2008:APP

[1360] A. Pethő and V. Ziegler. Arithmetic progressions on Pell equations. Journal of Number Theory, 128(6):1389–1409, June 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X08000474.

Jones:2008:OAF

[1361] John W. Jones and David P. Roberts. Octic 2-adic fields. Journal of Number Theory, 128(6):1410–1429, June 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X08000309.

Virdol:2008:TCF

[1362] Cristian Virdol. Tate classes and L-functions on a product of a quaternionic Shimura surface and a Picard modular surface. Journal of Number Theory, 128(6):1430–1447, June 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X08000176.

Lorenzini:2008:GPJ

[1363] Dino Lorenzini. Grothendieck’s pairing for Jacobians and base change. Journal of Number Theory, 128(6):1448–1457, June 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X0800019X.

Hsia:2008:RNT

[1364] Liang-Chung Hsia. On the reduction of a non-torsion point of a Drinfeld module. Journal of Number Theory, 128(6):1458–1484, June 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X08000188.

Yamauchi:2008:MMF

[1365] Takuya Yamauchi. Q-motives and modular forms. Journal of Number Theory, 128(6):1485–1505, June 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X08000516.

Zhong:2008:MPC

[1366] Ting Zhong. Metrical properties for a class of continued fractions with increasing digits. Journal of Number Theory, 128(6):1506–1515, June 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X07001096.

Conrey:2008:LOT

[1367] J. B. Conrey, D. W. Farmer, J. P. Keating, M. O. Rubinstein, and N. C. Snaith. Lower order terms in the full moment conjecture for the Riemann zeta function. Journal of Number Theory, 128(6):1516– 1554, June 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X0700145X.

Filipin:2008:TDE

[1368] Alan Filipin. There does not exist a D(4)-sextuple. Journal of Number Theory, 128(6):1555–1565, June 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X07001904.

Broughan:2008:OMN

[1369] Kevin A. Broughan and Qizhi Zhou. Odd multiperfect numbers of abundancy 4. Journal of Number Theory, 128(6):1566–1575, June 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X07000868.

Ranieri:2008:GAE

[1370] Gabriele Ranieri. Générateurs de l’anneau des entiers d’une extension cyclotomique. (French) [Generators of the ring of integers of a cyclotomic extension]. Journal of Number Theory, 128(6):1576– 1586, June 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X07001850.

Kontogeorgis:2008:AGG

[1371] A. Kontogeorgis. Actions of Galois groups on invariants of number fields. Journal of Number Theory, 128(6):1587–1601, June 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X07001242.

Corteel:2008:EOR

[1372] Sylvie Corteel, Jeremy Lovejoy, and Olivier Mallet. An extension to overpartitions of the Rogers–Ramanujan identities for even moduli. Journal of Number Theory, 128(6):1602–1621, June 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X07001084.

Namboothiri:2008:CET

[1373] K. Vishnu Namboothiri and Rajat Tandon. Completing an extension of Tunnell’s theorem. Journal of Number Theory, 128(6):1622– 1636, June 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X08000073.

Popa:2008:WNA

[1374] Alexandru A. Popa. Whittaker newforms for Archimedean representations of GL(2). Journal of Number Theory, 128(6):1637–1645, June 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X07001436.

Pan:2008:CHN

[1375] Hao Pan. On consecutive happy numbers. Journal of Number Theory, 128(6):1646–1654, June 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X08000097.

Vijay:2008:EED

[1376] Sujith Vijay. Eleven Euclidean distances are enough. Journal of Number Theory, 128(6):1655–1661, June 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X07002041.

Kazalicki:2008:ZCD

[1377] Matija Kazalicki. Zeros of certain Drinfeld modular functions. Journal of Number Theory, 128(6):1662–1669, June 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X07000832.

Muriefah:2008:DE

[1378] Fadwa S. Abu Muriefah. On the diophantine equation px^{2} + q^{2m} = y^{p}. Journal of Number Theory, 128(6):1670–1675, June 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X08000541. See note [1713].

Deleglise:2008:GVN

[1379] M. Deléglise, M. O. Hernane, and J.-L. Nicolas. Grandes valeurs et nombres champions de la fonction arithmétique de Kalmár. (French) [Large values and champion numbers of the arithmetic function of Kalmár]. Journal of Number Theory, 128(6):1676–1716, June 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X07001473.

Yin:2008:AQC

[1380] Linsheng Yin and Chong Zhang. Arithmetic of quasi-cyclotomic fields. Journal of Number Theory, 128(6):1717–1730, June 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X07001230.

Berkovich:2008:KSC

[1381] Alexander Berkovich and Frank G. Garvan. K. Saito’s conjecture for nonnegative eta products and analogous results for other infinite products. Journal of Number Theory, 128(6):1731–1748, June 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X07000820.

Buckingham:2008:CFG

[1382] Paul Buckingham. The canonical fractional Galois ideal at s = 0. Journal of Number Theory, 128(6):1749–1768, June 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X07002120.

Koyama:2008:MES

[1383] Shin ya Koyama and Nobushige Kurokawa. Multiple Eisenstein series and multiple cotangent functions. Journal of Number Theory, 128(6):1769–1774, June 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X07001369.

Kurokawa:2008:LVE

[1384] Nobushige Kurokawa. Limit values of Eisenstein series and multiple cotangent functions. Journal of Number Theory, 128(6):1775– 1783, June 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X07001357.

Thorne:2008:IDP

[1385] Frank Thorne. Irregularities in the distributions of primes in function fields. Journal of Number Theory, 128(6):1784–1794, June 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X07002028.

Ballet:2008:TRM

[1386] Stéphane Ballet. On the tensor rank of the multiplication in the finite fields. Journal of Number Theory, 128(6):1795–1806, June 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X07001813.

Amdeberhan:2008:APS

[1387] Tewodros Amdeberhan, Luis A. Medina, and Victor H. Moll. Arithmetical properties of a sequence arising from an arctangent sum. Journal of Number Theory, 128(6):1807–1846, June 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X07001424.

Brown:2008:ECM

[1388] Brittany Brown, Neil J. Calkin, Timothy B. Flowers, Kevin James, Ethan Smith, and Amy Stout. Elliptic curves, modular forms, and sums of Hurwitz class numbers. Journal of Number Theory, 128(6):1847– 1863, June 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X07002417.

Dahmen:2008:LBN

[1389] Sander R. Dahmen. Lower bounds for numbers of ABC-hits. Journal of Number Theory, 128(6):1864–1873, June 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X07001801.

Belding:2008:WPT

[1390] Juliana V. Belding. A Weil pairing on the p-torsion of ordinary elliptic curves over K[ϵ]. Journal of Number Theory, 128(6):1874– 1888, June 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X08000486.

Rodseth:2008:ADP

[1391] Øystein J. Rødseth. Addendum to “Distribution of points on the circle” [J. Number Theory 127 (1) (2007) 127–135]. Journal of Number Theory, 128(6):1889–1892, June 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X07002090. See [1254].

Anonymous:2008:PJc

[1392] Anonymous. Pages 1893–2198 (July 2008). Journal of Number Theory, 128(7):??, July 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2008:EBg

[1393] Anonymous. Editorial Board. Journal of Number Theory, 128(7): ifc, July 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X08000929.

Sinnott:2008:DSF

[1394] W. Sinnott. Dirichlet series in function fields. Journal of Number Theory, 128(7):1893–1899, July 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X08000528.

Underwood:2008:SHS

[1395] Robert G. Underwood. Sequences of Hopf–Swan subgroups. Journal of Number Theory, 128(7):1900–1915, July 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X08000589.

Filaseta:2008:PAS

[1396] Michael Filaseta, Carrie Finch, and Mark Kozek. On powers associated with Sierpiński numbers, Riesel numbers and Polignac’s conjecture. Journal of Number Theory, 128(7):1916–1940, July 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X08000462.

Li:2008:HOD

[1397] Wen-Ching Winnie Li and Yotsanan Meemark. Hecke operators on Drinfeld cusp forms. Journal of Number Theory, 128(7):1941– 1965, July 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X08000590.

Azaiez:2008:RQF

[1398] Najib Ouled Azaiez. The ring of quasimodular forms for a cocompact group. Journal of Number Theory, 128(7):1966–1988, July 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X07001862.

Kurth:2008:MFS

[1399] Chris A. Kurth and Ling Long. On modular forms for some noncongruence subgroups of SL_{2}(Z). Journal of Number Theory, 128(7):1989– 2009, July 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X07002399.

Chae:2008:IAG

[1400] Hi joon Chae and Dae San Kim. Identities arising from Gauss sums for finite classical groups. Journal of Number Theory, 128(7):2010– 2024, July 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X08000498.

Katsurada:2008:CDS

[1401] Hidenori Katsurada and Hisa aki Kawamura. A certain Dirichlet series of Rankin–Selberg type associated with the Ikeda lifting. Journal of Number Theory, 128(7):2025–2052, July 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X07001825.

Aubry:2008:SQR

[1402] Yves Aubry and Régis Blache. On some questions related to the Gauss conjecture for function fields. Journal of Number Theory, 128(7):2053– 2062, July 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X07002466.

Haessig:2008:AMF

[1403] C. Douglas Haessig. On the p-adic meromorphy of the function field height zeta function. Journal of Number Theory, 128(7):2063– 2069, July 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X07001229.

Lengler:2008:CIP

[1404] Johannes Lengler. A combinatorial interpretation of the probabilities of p-groups in the Cohen–Lenstra measure. Journal of Number Theory, 128(7):2070–2084, July 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X0700203X.

Chen:2008:DPD

[1405] Yong-Gao Chen and Ying Shi. Distribution of primes and dynamics of the w function. Journal of Number Theory, 128(7):2085– 2090, July 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X0800053X.

Beli:2008:SAQ

[1406] Constantin N. Beli and Wai Kiu Chan. Strong approximation of quadrics and representations of quadratic forms. Journal of Number Theory, 128(7):2091–2096, July 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X07001874.

Koike:2008:EGA

[1407] Kenji Koike and Hironori Shiga. An extended Gauss AGM and corresponding Picard modular forms. Journal of Number Theory, 128(7): 2097–2126, July 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X08000218.

Lee:2008:RTD

[1408] Yoonjin Lee. Reflection theorem for divisor class groups of relative quadratic function fields. Journal of Number Theory, 128(7):2127– 2137, July 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X07002132.

Chen:2008:AVF

[1409] Yen-Mei J. Chen. Average values of L-functions in characteristic two. Journal of Number Theory, 128(7):2138–2158, July 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X0800022X.

Vigni:2008:RCE

[1410] Stefano Vigni. On ring class eigenspaces of Mordell–Weil groups of elliptic curves over global function fields. Journal of Number Theory, 128(7):2159–2184, July 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X08000048.

Cluckers:2008:RSA

[1411] Raf Cluckers and Eva Leenknegt. Rectilinearization of semi-algebraic p-adic sets and Denef’s rationality of Poincaré series. Journal of Number Theory, 128(7):2185–2197, July 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X07002454.

Anonymous:2008:PAb

[1412] Anonymous. Pages 2199–2504 (August 2008). Journal of Number Theory, 128(8):??, August 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2008:EBh

[1413] Anonymous. Editorial Board. Journal of Number Theory, 128(8): ifc, August 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X08000991.

delaMaza:2008:HAN

[1414] Ana Cecilia de la Maza and Eduardo Friedman. Heights of algebraic numbers modulo multiplicative group actions. Journal of Number Theory, 128(8):2199–2213, August 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X08000619.

Farhi:2008:UBO

[1415] Bakir Farhi. Upper bounds for the order of an additive basis obtained by removing a finite subset of a given basis. Journal of Number Theory, 128(8):2214–2230, August 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X08000577.

Vandervelde:2008:MMP

[1416] Sam Vandervelde. The Mahler measure of parametrizable polynomials. Journal of Number Theory, 128(8):2231–2250, August 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X08000206.

Klosin:2008:ILI

[1417] Krzysztof Klosin. Ihara’s lemma for imaginary quadratic fields. Journal of Number Theory, 128(8):2251–2262, August 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X08000565.

Ash:2008:CCS

[1418] Avner Ash, Paul E. Gunnells, and Mark McConnell. Cohomology of congruence subgroups of SL(4, Z) II. Journal of Number Theory, 128(8):2263–2274, August 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X0700220X.

An:2008:NAA

[1419] Ta Thi Hoai An, Julie Tzu-Yueh Wang, and Pit-Mann Wong. Non-Archimedean analytic curves in the complements of hypersurface divisors. Journal of Number Theory, 128(8):2275–2281, August 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X07002442.

Chadozeau:2008:REP

[1420] Arnaud Chadozeau. Sur la répartition des entiers premiers à un entier sans petit facteur premier. (French) [On the distribution of prime integers to an integer without small prime factor]. Journal of Number Theory, 128(8):2282–2317, August 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X08000267.

Guo:2008:DAB

[1421] Li Guo and Bin Zhang. Differential algebraic Birkhoff decomposition and the renormalization of multiple zeta values. Journal of Number Theory, 128(8):2318–2339, August 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X07002144.

Asgari:2008:AFA

[1422] Mahdi Asgari and Ralf Schmidt. On the adjoint L-function of the p-adic GSp(4). Journal of Number Theory, 128(8):2340–2358, August 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X07002119.

Fukshansky:2008:DWR

[1423] Lenny Fukshansky. On distribution of well-rounded sublattices of Z^{2}. Journal of Number Theory, 128(8):2359–2393, August 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X08000085.

Wu:2008:RCF

[1424] Jun Wu. A remark on continued fractions with sequences of partial quotients. Journal of Number Theory, 128(8):2394–2397, August 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X08000231.

Denomme:2008:ECP

[1425] Robert Denomme and Gordan Savin. Elliptic curve primality tests for Fermat and related primes. Journal of Number Theory, 128(8):2398– 2412, August 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X08000243.

Long:2008:ASD

[1426] Ling Long. On Atkin and Swinnerton-Dyer congruence relations (3). Journal of Number Theory, 128(8):2413–2429, August 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X08000607.

Chenevier:2008:QCH

[1427] Gaëtan Chenevier. Quelques courbes de Hecke se plongent dans l’espace de Colmez. (French) [Some Hecke curves embed in Colmez space]. Journal of Number Theory, 128(8):2430–2449, August 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X07002107.

Kishi:2008:NFI

[1428] Yasuhiro Kishi. A new family of imaginary quadratic fields whose class number is divisible by five. Journal of Number Theory, 128 (8):2450–2458, August 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X08000668.

Zhang:2008:MVF

[1429] Wenpeng Zhang. On the mean value of L-functions with the weight of character sums. Journal of Number Theory, 128(8):2459–2466, August 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X08000061.

Lanphier:2008:CMS

[1430] Dominic Lanphier. Combinatorics of Maass–Shimura operators. Journal of Number Theory, 128(8):2467–2487, August 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X07002430.

Cilleruelo:2008:S

[1431] Javier Cilleruelo. Squares in (1^{2} + 1) ⋯ (n^{2} + 1). Journal of Number Theory, 128(8):2488–2491, August 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X07002235.

Virdol:2008:MPP

[1432] Cristian Virdol. Moduli problem and points on some twisted Shimura varieties of PEL type. Journal of Number Theory, 128(8):2492–2504, August 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X08000371.

Anonymous:2008:PS

[1433] Anonymous. Pages 2505–2766 (September 2008). Journal of Number Theory, 128(9):??, September 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2008:EBi

[1434] Anonymous. Editorial Board. Journal of Number Theory, 128(9):ifc, September 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X08001248.

Langmann:2008:FGR

[1435] Klaus Langmann. Fermat und die Güte Riemannscher Näherungen für ∫_{−1}^{1} ∫_{−1}^{1} ∫_{−1}^{1} Φ((ax + by + cz)R)Ψ(x, y, z) dx dy dz. (German) [Fermat and the goodness of Riemann approximations for ∫_{−1}^{1} ∫_{−1}^{1} ∫_{−1}^{1} Φ((ax + by + cz)R)Ψ(x, y, z) dx dy dz]. Journal of Number Theory, 128(9):2505–2519, September 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X08000796.

Garaev:2008:EFP

[1436] M. Z. Garaev and V. C. Garcia. The equation x_{1}x_{2} = x_{3}x_{4} + λ in fields of prime order and applications. Journal of Number Theory, 128 (9):2520–2537, September 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X08000802.

Muneta:2008:SEE

[1437] Shuichi Muneta. On some explicit evaluations of multiple zeta-star values. Journal of Number Theory, 128(9):2538–2548, September 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X08000851.

Vaananen:2008:LIV

[1438] Keijo Väänänen and Wadim Zudilin. Linear independence of values of Tschakaloff functions with different parameters. Journal of Number Theory, 128(9):2549–2558, September 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X08000814.

Chakraborty:2008:CNM

[1439] Kalyan Chakraborty, Florian Luca, and Anirban Mukhopadhyay. Class numbers with many prime factors. Journal of Number Theory, 128 (9):2559–2572, September 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X08000826.

Chen:2008:AET

[1440] Yong-Gao Chen. The analogue of Erdős–Turán conjecture in Z_{m}. Journal of Number Theory, 128(9):2573–2581, September 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X08000887.

Alnaser:2008:WN

[1441] Ala Alnaser and Todd Cochrane. Waring’s number modm. Journal of Number Theory, 128(9):2582–2590, September 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X08000899.

Anderson:2008:AFC

[1442] Jaclyn Anderson. An asymptotic formula for the t-core partition function and a conjecture of Stanton. Journal of Number Theory, 128 (9):2591–2615, September 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X07002193.

Batson:2008:NHF

[1443] Joshua Batson. Nathanson heights in finite vector spaces. Journal of Number Theory, 128(9):2616–2633, September 2008. CODEN JNUTA9.

ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X08000875.

Arnold:2008:CMP

[1444] Trevor Arnold. Complex multiplication and parity in Iwasawa theory. Journal of Number Theory, 128(9):2634–2654, September 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X08000553.

Hu:2008:NPR

[1445] Su Hu. A note on the d-th power residue symbol of F_{q}[t]. Journal of Number Theory, 128(9):2655–2662, September 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X07002429.

Kessebohmer:2008:FAS

[1446] Marc Kesseböhmer and Bernd O. Stratmann. Fractal analysis for sets of non-differentiability of Minkowski’s question mark function. Journal of Number Theory, 128(9):2663–2686, September 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X08000255.

Li:2008:GUS

[1447] Yan Li and Xianke Zhang. Global unit squares and local unit squares. Journal of Number Theory, 128(9):2687–2694, September 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X0800084X.

Shparlinski:2008:VPM

[1448] Igor E. Shparlinski and Arne Winterhof. Visible points on multidimensional modular hyperbolas. Journal of Number Theory, 128(9): 2695–2703, September 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X08000644.

Ki:2008:ZSR

[1449] Haseo Ki. On the zeros of sums of the Riemann zeta function. Journal of Number Theory, 128(9):2704–2755, September 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X08000620.

Hare:2008:BEI

[1450] Kevin G. Hare and David Tweedle. Beta-expansions for infinite families of Pisot and Salem numbers. Journal of Number Theory, 128 (9):2756–2765, September 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X08000632.

Anonymous:2008:PO

[1451] Anonymous. Pages 2767–2860 (October 2008). Journal of Number Theory, 128(10):??, October 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2008:EBj

[1452] Anonymous. Editorial Board. Journal of Number Theory, 128(10): ifc, October 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X08001340.

Fieker:2008:LRB

[1453] Claus Fieker and Michael E. Pohst. A lower regulator bound for number fields. Journal of Number Theory, 128(10):2767–2775, October 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X08001121.

Kueh:2008:SEE

[1454] Ka-Lam Kueh, King Fai Lai, and Ki-Seng Tan. Stickelberger elements for Z_{p}^{d}-extensions of function fields. Journal of Number Theory, 128 (10):2776–2783, October 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X0800111X.

Jin:2008:RSP

[1455] Yuan Jin, Zhi-Juan Lu, and Asmus L. Schmidt. On recurrences for sums of powers of binomial coefficients. Journal of Number Theory, 128(10):2784–2794, October 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X08000905.

Xiong:2008:PCR

[1456] Maosheng Xiong and Alexandru Zaharescu. Pair correlation of rationals with prime denominators. Journal of Number Theory, 128 (10):2795–2807, October 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X08000280.

Schein:2008:WSC

[1457] Michael M. Schein. Weights in Serre’s conjecture for GL_{n} via the Bernstein–Gelfand–Gelfand complex. Journal of Number Theory, 128 (10):2808–2822, October 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X08000978.

Malle:2008:CLH

[1458] Gunter Malle. Cohen–Lenstra heuristic and roots of unity. Journal of Number Theory, 128(10):2823–2835, October 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X08000504.

Hall:2008:LSB

[1459] Richard R. Hall. Large spaces between the zeros of the Riemann zetafunction and random matrix theory II. Journal of Number Theory, 128 (10):2836–2851, October 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X08000024.

Maclachlan:2008:OEQ

[1460] C. Maclachlan. Optimal embeddings in quaternion algebras. Journal of Number Theory, 128(10):2852–2860, October 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X08000279.

Anonymous:2008:PN

[1461] Anonymous. Pages 2861–2962 (November 2008). Journal of Number Theory, 128(11):??, November 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2008:EBk

[1462] Anonymous. Editorial Board. Journal of Number Theory, 128(11):ifc, November 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X08001662.

Nicolae:2008:SAF

[1463] Florin Nicolae. On the semigroup of Artin’s L-functions holomorphic at s_{0}. Journal of Number Theory, 128(11):2861–2864, November 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X08001133.

Agoh:2008:FPF

[1464] Takashi Agoh and Ladislav Skula. The fourth power of the Fermat quotient. Journal of Number Theory, 128(11):2865–2873, November 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X0800108X.

Nicole:2008:SAV

[1465] Marc-Hubert Nicole. Superspecial abelian varieties and the Eichler basis problem for Hilbert modular forms. Journal of Number Theory, 128 (11):2874–2889, November 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X08001108.

Bertin:2008:MMH

[1466] Marie José Bertin. Mesure de Mahler d’hypersurfaces K3. (French) [Mahler measurement of K3 hypersurfaces]. Journal of Number Theory, 128(11):2890–2913, November 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X08000292.

vandeWoestijne:2008:NNS

[1467] Christiaan van de Woestijne. Noncanonical number systems in the integers. Journal of Number Theory, 128(11):2914–2938, November 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X08000656.

Savitt:2008:BMR

[1468] David Savitt. Breuil modules for Raynaud schemes. Journal of Number Theory, 128(11):2939–2950, November 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X08001066.

Young:2008:AFB

[1469] Paul Thomas Young. A 2-adic formula for Bernoulli numbers of the second kind and for the Nörlund numbers. Journal of Number Theory, 128(11):2951–2962, November 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X0800067X.

Anonymous:2008:PD

[1470] Anonymous. Pages 2963–3072 (December 2008). Journal of Number Theory, 128(12):??, December 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2008:EBl

[1471] Anonymous. Editorial Board. Journal of Number Theory, 128(12):ifc, December 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X0800187X.

Gallardo:2008:SWP

[1472] Luis H. Gallardo and Leonid N. Vaserstein. The strict Waring problem for polynomial rings. Journal of Number Theory, 128(12):2963–2972, December 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X08001601.

Thunder:2008:CSG

[1473] Jeffrey Lin Thunder. Counting subspaces of given height defined over a function field. Journal of Number Theory, 128(12):2973–3004, December 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X08001595.

Sleziak:2008:LGD

[1474] Martin Sleziak and Miloš Ziman. Lévy group and density measures. Journal of Number Theory, 128(12):3005–3012, December 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X08001169.

Berger:2008:TSD

[1475] Lisa Berger. Towers of surfaces dominated by products of curves and elliptic curves of large rank over function fields. Journal of Number Theory, 128(12):3013–3030, December 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X08000863.

Im:2008:GTR

[1476] Bo-Hae Im and Michael Larsen. Generalizing a theorem of Richard Brauer. Journal of Number Theory, 128(12):3031–3036, December 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X08001091.

Achimescu:2008:HMF

[1477] Sever Achimescu and Abhishek Saha. Hilbert modular forms of weight 1/2 and theta functions. Journal of Number Theory, 128(12):3037–3062, December 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X08000838.

Liu:2008:CEN

[1478] Guodong Liu. On congruences of Euler numbers modulo powers of two. Journal of Number Theory, 128(12):3063–3071, December 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X08001078.

Anonymous:2009:PJa

[1479] Anonymous. Pages 1–246 (January 2009). Journal of Number Theory, 129(1):??, January 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2009:EBa

[1480] Anonymous. Editorial Board. Journal of Number Theory, 129(1): ifc, January 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X08002096.

Glass:2009:GSR

[1481] Darren Glass and Sonin Kwon. Galois structure and de Rham invariants of elliptic curves. Journal of Number Theory, 129(1):1–14, January 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X08001637.

Pande:2009:LCD

[1482] Aftab Pande. Local constancy of dimensions of Hecke eigenspaces of automorphic forms. Journal of Number Theory, 129(1):15–27, January 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X08001455.

Trnkova:2009:EGC

[1483] Veronika Trnková. Enlargement of the group of circular units of a bicyclic field. Journal of Number Theory, 129(1):28–35, January 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X08001820.

Gasbarri:2009:DTC

[1484] C. Gasbarri. Dyson’s theorem for curves. Journal of Number Theory, 129(1):36–58, January 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X08001844.

Breuer:2009:DMP

[1485] Florian Breuer and Hans-Georg Rück. Drinfeld modular polynomials in higher rank. Journal of Number Theory, 129(1):59–83, January 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X08001728.

Byott:2009:NNR

[1486] Nigel P. Byott and G. Griffith Elder. On the necessity of new ramification breaks. Journal of Number Theory, 129(1):84–101, January 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X08001613.

Draziotis:2009:SDE

[1487] Konstantinos Draziotis and Dimitrios Poulakis. Solving the Diophantine equation y^{2} = x(x^{2} − n^{2}). Journal of Number Theory, 129(1):102– 121, January 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X08001480. See corrigendum [1533].

Kim:2009:CSN

[1488] Sun Kim. Covering systems in number fields. Journal of Number Theory, 129(1):122–141, January 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X08001303.

Taelman:2009:AM

[1489] Lenny Taelman. Artin t-motifs. Journal of Number Theory, 129(1):142– 157, January 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X08001716.

Bruggeman:2009:ETO

[1490] R. W. Bruggeman and T. Mühlenbruch. Eigenfunctions of transfer operators and cohomology. Journal of Number Theory, 129(1):158– 181, January 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X08001777.

Ingram:2009:MIP

[1491] Patrick Ingram. Multiples of integral points on elliptic curves. Journal of Number Theory, 129(1):182–208, January 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X08001492.

Flatters:2009:PDS

[1492] Anthony Flatters. Primitive divisors of some Lehmer–Pierce sequences. Journal of Number Theory, 129(1):209–219, January 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X08001224.

Chen:2009:RQC

[1493] Alex Chen. On the reducible quintic complete base polynomials. Journal of Number Theory, 129(1):220–230, January 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X08001625.

Nover:2009:CGG

[1494] Harris Nover. Computation of Galois groups associated to the 2-class towers of some imaginary quadratic fields with 2-class group C_{2} × C_{2} × C_{2}. Journal of Number Theory, 129(1):231–245, January 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X0800139X.

Anonymous:2009:PF

[1495] Anonymous. Pages 247–498 (February 2009). Journal of Number Theory, 129(2):??, February 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2009:EBb

[1496] Anonymous. Editorial Board. Journal of Number Theory, 129(2):ifc, February 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X08002229.

Bornhofen:2009:PAM

[1497] Matthias Bornhofen and Urs Hartl. Pure Anderson motives over finite fields. Journal of Number Theory, 129(2):247–283, February 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X08001790.

Jetchev:2009:EHP

[1498] Dimitar Jetchev, Kristin Lauter, and William Stein. Explicit Heegner points: Kolyvagin’s conjecture and non-trivial elements in the Shafarevich–Tate group. Journal of Number Theory, 129(2):284–302, February 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X08001212.

Luca:2009:PML

[1499] Florian Luca and T. N. Shorey. Products of members of Lucas sequences with indices in an interval being a power. Journal of Number Theory, 129(2):303–315, February 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X08001753.

Chan:2009:ARS

[1500] Tsz Ho Chan. Approximating reals by sums of rationals. Journal of Number Theory, 129(2):316–324, February 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X08001807.

Laca:2009:BCT

[1501] Marcelo Laca, Nadia S. Larsen, and Sergey Neshveyev. On Bost– Connes type systems for number fields. Journal of Number Theory, 129(2):325–338, February 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X08001935.

Eliahou:2009:PVS

[1502] Shalom Eliahou, Michel Kervaire, and Cédric Lecouvey. On the product of vector spaces in a commutative field extension. Journal of Number Theory, 129(2):339–348, February 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X0800142X.

Moshchevitin:2009:SFP

[1503] N. G. Moshchevitin. On small fractional parts of polynomials. Journal of Number Theory, 129(2):349–357, February 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X08001546.

Komatsu:2009:SPL

[1504] Takao Komatsu. Shrinking the period length of quasi-periodic continued fractions. Journal of Number Theory, 129(2):358–366, February 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X08001789.

Garcia-Selfa:2009:TET

[1505] Irene García-Selfa and José M. Tornero. Thue equations and torsion groups of elliptic curves. Journal of Number Theory, 129(2):367–380, February 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X08001418.

Nguyen:2009:PSE

[1506] Lan Nguyen. Picard sequences with every orbit containing infinitely many perfect n-powers. Journal of Number Theory, 129(2):381–401, February 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X08001510.

Buyukboduk:2009:TDE

[1507] Kâzım Büyükboduk. Tamagawa defect of Euler systems. Journal of Number Theory, 129(2):402–417, February 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X08001467.

Hoelscher:2009:GER

[1508] Jing Long Hoelscher. Galois extensions ramified only at one prime. Journal of Number Theory, 129(2):418–427, February 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X08001431.

Prakash:2009:LSI

[1509] Gyan Prakash and D. S. Ramana. The large sieve inequality for integer polynomial amplitudes. Journal of Number Theory, 129(2):428– 433, February 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X08001200.

Guo:2009:VTM

[1510] Song Guo and Zhi-Wei Sun. A variant of Tao’s method with application to restricted sumsets. Journal of Number Theory, 129(2):434– 438, February 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X08001170.

Lyall:2009:PCD

[1511] Neil Lyall and Ákos Magyar. Polynomial configurations in difference sets. Journal of Number Theory, 129(2):439–450, February 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X08001157.

Hofer:2009:DPN

[1512] Roswitha Hofer. On the distribution properties of Niederreiter–Halton sequences. Journal of Number Theory, 129(2):451–463, February 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X08001315.

Brink:2009:FPT

[1513] David Brink. Five peculiar theorems on simultaneous representation of primes by quadratic forms. Journal of Number Theory, 129(2):464– 468, February 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X08001145.

Perucca:2009:PVO

[1514] Antonella Perucca. Prescribing valuations of the order of a point in the reductions of abelian varieties and tori. Journal of Number Theory, 129(2):469–476, February 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X08001509.

Lu:2009:AFA

[1515] Guangshi Lü. Additive functions on arithmetic progressions with large moduli. Journal of Number Theory, 129(2):477–487, February 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X08001558.

Lu:2009:AVF

[1516] Guangshi Lü. The average value of Fourier coefficients of cusp forms in arithmetic progressions. Journal of Number Theory, 129(2):488–494, February 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X0800156X.

Gras:2009:NRF

[1517] Georges Gras and Jean-François Jaulent. Note on 2-rational fields. Journal of Number Theory, 129(2):495–498, February 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X08001583.

Anonymous:2009:PMa

[1518] Anonymous. Pages 499–740 (March 2009). Journal of Number Theory, 129(3):??, March 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2009:EBc

[1519] Anonymous. Editorial Board. Journal of Number Theory, 129(3): ifc, March 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X09000109.

Sun:2009:QOR

[1520] Zhi-Hong Sun. Quartic, octic residues and Lucas sequences. Journal of Number Theory, 129(3):499–550, March 2009. CODEN JNUTA9.

ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X08002291.

Suzuki:2009:RHW

[1521] Masatoshi Suzuki. The Riemann hypothesis for Weng’s zeta function of Sp(4) over Q. Journal of Number Theory, 129(3):551–579, March 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X08001479.

Weng:2009:AZF

[1522] Lin Weng. Appendix: Zeta functions for Sp(2n). Journal of Number Theory, 129(3):567–579, March 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X08001534.

Kurlberg:2009:FNP

[1523] Pär Kurlberg and Zeév Rudnick. The fluctuations in the number of points on a hyperelliptic curve over a finite field. Journal of Number Theory, 129(3):580–587, March 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X08001765.

Nebe:2009:FWR

[1524] Gabriele Nebe. Finite Weil representations and associated Hecke algebras. Journal of Number Theory, 129(3):588–603, March 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X08002187.

Lee:2009:CDN

[1525] Jungyun Lee. The complete determination of narrow Richaud–Degert type which is not 5 modulo 8 with class number two. Journal of Number Theory, 129(3):604–620, March 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X08001996.

Benamar:2009:CBA

[1526] H. Benamar and A. Chandoul. Convergence of the Brun algorithm over the field of formal power series. Journal of Number Theory, 129(3):621–631, March 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X0800231X.

deShalit:2009:IVP

[1527] Ehud de Shalit and Eran Iceland. Integer valued polynomials and Lubin–Tate formal groups. Journal of Number Theory, 129(3):632– 639, March 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X08002394.

Nielsen:2009:CSW

[1528] Pace P. Nielsen. A covering system whose smallest modulus is 40. Journal of Number Theory, 129(3):640–666, March 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X08002199.

Moree:2009:ICP

[1529] Pieter Moree. Inverse cyclotomic polynomials. Journal of Number Theory, 129(3):667–680, March 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X0800200X.

Murty:2009:EEE

[1530] M. Ram Murty and Kaneenika Sinha. Effective equidistribution of eigenvalues of Hecke operators. Journal of Number Theory, 129(3):681– 714, March 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X08002369.

Baran:2009:MCL

[1531] Burcu Baran. A modular curve of level 9 and the class number one problem. Journal of Number Theory, 129(3):715–728, March 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X08002035.

Chang:2009:NRV

[1532] Chieh-Yu Chang. A note on a refined version of Anderson–Brownawell– Papanikolas criterion. Journal of Number Theory, 129(3):729–738, March 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X08002400.

Draziotis:2009:CSD

[1533] Konstantinos Draziotis and Dimitrios Poulakis. Corrigendum to “Solving the Diophantine equation y^{2} = x(x^{2} − n^{2})” [J. Number Theory 129 (1) (2009) 102–121]. Journal of Number Theory, 129(3):739– 740, March 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X08002412. See [1487].

Anonymous:2009:PAa

[1534] Anonymous. Pages 741–970 (April 2009). Journal of Number Theory, 129(4):??, April 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2009:EBd

[1535] Anonymous. Editorial Board. Journal of Number Theory, 129(4): ifc, April 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X09000304.

Sidorov:2009:ENI

[1536] Nikita Sidorov. Expansions in non-integer bases: Lower, middle and top orders. Journal of Number Theory, 129(4):741–754, April 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X09000079.

Kawashima:2009:CRA

[1537] Gaku Kawashima. A class of relations among multiple zeta values. Journal of Number Theory, 129(4):755–788, April 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X09000055.

Bars:2009:CPS

[1538] Francesc Bars and Ignazio Longhi. Coleman’s power series and Wiles’ reciprocity for rank 1 Drinfeld modules. Journal of Number Theory, 129(4):789–805, April 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X09000043.

Pixton:2009:SSS

[1539] Aaron Pixton. Sequences with small subsum sets. Journal of Number Theory, 129(4):806–817, April 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X09000031.

Ichikawa:2009:SMF

[1540] Takashi Ichikawa. Siegel modular forms of degree 2 over rings. Journal of Number Theory, 129(4):818–823, April 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X09000080.

Bensebaa:2009:GG

[1541] B. Bensebaa, A. Movahhedi, and A. Salinier. The Galois group of X^{p} + aX^{p−1} + a. Journal of Number Theory, 129(4):824–830, April 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X08002060.

Ren:2009:RSC

[1542] Tian Ren and Robert Sczech. A refinement of Stark’s conjecture over complex cubic number fields. Journal of Number Theory, 129(4):831– 857, April 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X08002382.

Ford:2009:SCS

[1543] Ben Ford, Hoàng Mai, and Lawrence Sze. Self-conjugate simultaneous p- and q-core partitions and blocks of A_{n}. Journal of Number Theory, 129(4):858–865, April 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X08002047.

Pink:2009:IGR

[1544] Richard Pink and Egon Rütsche. Image of the group ring of the Galois representation associated to Drinfeld modules. Journal of Number Theory, 129(4):866–881, April 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X08002436.

Pink:2009:AOD

[1545] Richard Pink and Egon Rütsche. Adelic openness for Drinfeld modules in generic characteristic. Journal of Number Theory, 129(4):882– 907, April 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X08002424.

Eie:2009:RSF

[1546] Minking Eie, Wen-Chin Liaw, and Yao Lin Ong. A restricted sum formula among multiple zeta values. Journal of Number Theory, 129(4): 908–921, April 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X08001741.

Cho:2009:ARG

[1547] Bumkyu Cho, Ja Kyung Koo, and Yoon Kyung Park. Arithmetic of the Ramanujan–Göllnitz–Gordon continued fraction. Journal of Number Theory, 129(4):922–947, April 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X08002205.

Zhang:2009:CSS

[1548] Tianping Zhang, Zhefeng Xu, and Zhanhu Li. On character sums over a short interval. Journal of Number Theory, 129(4):948–963, April 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X08001443.

Oh:2009:MSS

[1549] Byeong-Kweon Oh and Zhi-Wei Sun. Mixed sums of squares and triangular numbers (III). Journal of Number Theory, 129(4):964– 969, April 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X08001959.

Anonymous:2009:PMb

[1550] Anonymous. Pages 971–1204 (May 2009). Journal of Number Theory, 129(5):??, May 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2009:EBe

[1551] Anonymous. Editorial Board. Journal of Number Theory, 129(5): ifc, May 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X09000687.

Sun:2009:NRA

[1552] Zhi-Hong Sun. On the number of representations of n by ax(x − 1)/2 + by(y − 1)/2. Journal of Number Theory, 129(5):971–989, May 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X09000225.

Schmid:2009:RTS

[1553] Wolfgang A. Schmid. A realization theorem for sets of lengths. Journal of Number Theory, 129(5):990–999, May 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X09000237.

Hans-Gill:2009:USP

[1554] R. J. Hans-Gill, Madhu Raka, Ranjeet Sehmi, and Sucheta. A unified simple proof of a conjecture of Woods for n ≤ 6. Journal of Number Theory, 129(5):1000–1010, May 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X09000274.

Hans-Gill:2009:CMW

[1555] R. J. Hans-Gill, Madhu Raka, and Ranjeet Sehmi. On conjectures of Minkowski and Woods for n = 7. Journal of Number Theory, 129(5):1011–1033, May 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X09000262.

Gross:2009:CFC

[1556] Benedict H. Gross, Eriko Hironaka, and Curtis T. McMullen. Cyclotomic factors of Coxeter polynomials. Journal of Number Theory, 129(5):1034–1043, May 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X09000432.

Matala-aho:2009:DAR

[1557] Tapani Matala-aho and Ville Merilä. On Diophantine approximations of Ramanujan type q-continued fractions. Journal of Number Theory, 129(5):1044–1055, May 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X09000250.

Arias-de-Reyna:2009:TGR

[1558] Sara Arias de Reyna and Núria Vila. Tame Galois realizations of GL_{2}(F_{ℓ}) over Q. Journal of Number Theory, 129(5):1056–1065, May 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X09000390.

Ionascu:2009:CRT

[1559] Eugen J. Ionascu. A characterization of regular tetrahedra in Z^{3}. Journal of Number Theory, 129(5):1066–1074, May 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X09000249.

He:2009:IAW

[1560] Yuan He and Qunying Liao. On an identity associated with Weil’s estimate and its applications. Journal of Number Theory, 129(5):1075– 1089, May 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X09000286.

Schinzel:2009:PRL

[1561] A. Schinzel and W. M. Schmidt. Powers of roots in linear spaces. Journal of Number Theory, 129(5):1090–1098, May 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X09000420.

Ayala:2009:NLJ

[1562] E. J. Gómez Ayala. A new look at Jarvis’ distribution formula. Journal of Number Theory, 129(5):1099–1121, May 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X09000377.

Xie:2009:DUS

[1563] Bingyong Xie. On Drinfeld’s universal special formal module. Journal of Number Theory, 129(5):1122–1135, May 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X09000444.

Luca:2009:MNL

[1564] Florian Luca and Juan Luis Varona. Multiperfect numbers on lines of the Pascal triangle. Journal of Number Theory, 129(5):1136– 1148, May 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X08001984.

Kim:2009:SSP

[1565] Byoung Du Kim. The symmetric structure of the plus/minus Selmer groups of elliptic curves over totally real fields and the parity conjecture. Journal of Number Theory, 129(5):1149–1160, May 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X08001923.

Haessig:2009:ZFD

[1566] C. Douglas Haessig. On the zeta function of divisors for projective varieties with higher rank divisor class group. Journal of Number Theory, 129(5):1161–1177, May 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X08001182.

Manoharmayum:2009:LGR

[1567] Jayanta Manoharmayum. Lifting Galois representations of number fields. Journal of Number Theory, 129(5):1178–1190, May 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X08001947.

Stern:2009:NGG

[1568] Leonid Stern. On the norm groups of Galois 2n-extensions of algebraic number fields. Journal of Number Theory, 129(5):1191– 1204, May 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X08001297.

Anonymous:2009:PJb

[1569] Anonymous. Pages 1205–1622 (June 2009). Journal of Number Theory, 129(6):??, June 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2009:EBf

[1570] Anonymous. Editorial Board. Journal of Number Theory, 129(6): ifc, June 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X09000936.

Mouhib:2009:PCN

[1571] A. Mouhib. On the parity of the class number of multiquadratic number fields. Journal of Number Theory, 129(6):1205–1211, June 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X09000468.

Vergnaud:2009:MIL

[1572] Damien Vergnaud. Mesures d’indépendance linéaire de carrés de périodes et quasi-périodes de courbes elliptiques. (French) [Measures of linear independence of squares of periods and quasi-periods of elliptic curves]. Journal of Number Theory, 129(6):1212–1233, June 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X09000493.

Deckelbaum:2009:MPS

[1573] Alan Deckelbaum. Minimum product set sizes in nonabelian groups of order pq. Journal of Number Theory, 129(6):1234–1245, June 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X09000778.

Fourquaux:2009:ALC

[1574] Lionel Fourquaux. Applications Q_{p}-linéaires, continues et Galoiséquivariantes de C_{p} dans lui-même. (French) [Q_{p}-linear, continuous and Galois-equivariant applications of C_{p} in itself]. Journal of Number Theory, 129(6):1246–1255, June 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X09000523.

Yesilyurt:2009:GMI

[1575] Hamza Yesilyurt. A generalization of a modular identity of Rogers. Journal of Number Theory, 129(6):1256–1271, June 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X0900050X.

Pitale:2009:IRF

[1576] Ameya Pitale and Ralf Schmidt. Integral representation for Lfunctions for GSp_{4} × GL_{2}. Journal of Number Theory, 129(6):1272– 1324, June 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X09000626.

Prodinger:2009:DGF

[1577] Helmut Prodinger and Stephan Wagner. The digit generating function of a polynomial. Journal of Number Theory, 129(6):1325–1337, June 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X09000559.

Park:2009:CNO

[1578] S.-M. Park and G.-N. Lee. The class number one problem for some totally complex quartic number fields. Journal of Number Theory, 129(6):1338–1349, June 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X09000560.

Emmons:2009:SVA

[1579] Caleb J. Emmons and Cristian D. Popescu. Special values of Abelian L-functions at s = 0. Journal of Number Theory, 129(6):1350– 1365, June 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X08001571.

Berkovich:2009:RIS

[1580] Alexander Berkovich and Hamza Yesilyurt. On the representations of integers by the sextenary quadratic form x^{2}+y^{2}+z^{2}+7s^{2}+7t^{2}+7u^{2} and 7cores. Journal of Number Theory, 129(6):1366–1378, June 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X08001649.

Rhoades:2009:SGB

[1581] Robert C. Rhoades. 2-Selmer groups and the Birch–Swinnerton-Dyer Conjecture for the congruent number curves. Journal of Number Theory, 129(6):1379–1391, June 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X09000602.

Ghioca:2009:PPL

[1582] D. Ghioca and T. J. Tucker. Periodic points, linearizing maps, and the dynamical Mordell–Lang problem. Journal of Number Theory, 129(6):1392–1403, June 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X08002023.

Alaca:2009:SII

[1583] Ayşe Alaca, Şaban Alaca, and Kenneth S. Williams. Some identities involving theta functions. Journal of Number Theory, 129(6):1404– 1431, June 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X08002072.

Weingartner:2009:LDD

[1584] Andreas Weingartner. The limiting distribution of the divisor function. Journal of Number Theory, 129(6):1432–1442, June 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X09000766.

Hooley:2009:PFV

[1585] C. Hooley. On the power-free values of polynomials in two variables: II. Journal of Number Theory, 129(6):1443–1455, June 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X08002448.

Cornelissen:2009:TAF

[1586] Gunther Cornelissen and Oliver Lorscheid. Toroidal automorphic forms for some function fields. Journal of Number Theory, 129(6):1456– 1463, June 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X08001194.

Cheng:2009:SCS

[1587] Chuangxun Cheng, Brian Dietel, Mathilde Herblot, Jingjing Huang, Holly Krieger, Diego Marques, Jonathan Mason, Martin Mereb, and S. Robert Wilson. Some consequences of Schanuel’s conjecture. Journal of Number Theory, 129(6):1464–1467, June 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X09000201.

Gross:2009:CRQ

[1588] Benedict H. Gross and Mark W. Lucianovic. On cubic rings and quaternion rings. Journal of Number Theory, 129(6):1468–1478, June 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X08001406.

Barreira:2009:FDL

[1589] Luis Barreira and Godofredo Iommi. Frequency of digits in the Lüroth expansion. Journal of Number Theory, 129(6):1479–1490, June 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X08001327.

Jao:2009:EGB

[1590] David Jao, Stephen D. Miller, and Ramarathnam Venkatesan. Expander graphs based on GRH with an application to elliptic curve cryptography. Journal of Number Theory, 129(6):1491–1504, June 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X09000213.

Spencer:2009:MC

[1591] Craig V. Spencer. The Manin conjecture for x_{0}y_{0} + ⋯ + x_{s}y_{s} = 0. Journal of Number Theory, 129(6):1505–1521, June 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X08001960.

Wang:2009:ENC

[1592] Nianliang Wang, Junzhuang Li, and Duansen Liu. Euler numbers congruences and Dirichlet L-functions. Journal of Number Theory, 129(6):1522–1531, June 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X09000407.

Connes:2009:F

[1593] Alain Connes, Caterina Consani, and Matilde Marcolli. Fun with F_{1}. Journal of Number Theory, 129(6):1532–1561, June 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X08001856.

Goren:2009:DBS

[1594] Eyal Z. Goren and Kristin E. Lauter. The distance between superspecial abelian varieties with real multiplication. Journal of Number Theory, 129(6):1562–1578, June 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X08001522.

Heath-Brown:2009:SDT

[1595] D. R. Heath-Brown. Sums and differences of three k-th powers. Journal of Number Theory, 129(6):1579–1594, June 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X09000572.

Jancevskis:2009:CSS

[1596] Martin Jancevskis. Convergent sieve sequences in arithmetic progressions. Journal of Number Theory, 129(6):1595–1607, June 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X08002059.

Nathanson:2009:SRS

[1597] Melvyn B. Nathanson. Supersequences, rearrangements of sequences, and the spectrum of bases in additive number theory. Journal of Number Theory, 129(6):1608–1621, June 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X08001832.

Anonymous:2009:PJc

[1598] Anonymous. Pages 1623–1812 (July 2009). Journal of Number Theory, 129(7):??, July 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2009:EBg

[1599] Anonymous. Editorial Board. Journal of Number Theory, 129(7): ifc, July 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X09001000.

Manes:2009:MSF

[1600] Michelle Manes. Moduli spaces for families of rational maps on P^{1}. Journal of Number Theory, 129(7):1623–1663, July 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X09000857.

Pinter:2009:CDE

[1601] Á. Pintér. On a class of diophantine equations related to the numbers of cells in hyperplane arrangements. Journal of Number Theory, 129(7):1664–1668, July 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X08002357.

Maier:2009:CES

[1602] H. Maier and A. Sankaranarayanan. On certain exponential sums over primes. Journal of Number Theory, 129(7):1669–1677, July 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X09000638.

Fujita:2009:DQC

[1603] Yasutsugu Fujita. Any Diophantine quintuple contains a regular Diophantine quadruple. Journal of Number Theory, 129(7):1678– 1697, July 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X0900002X.

Fili:2009:NAM

[1604] Paul Fili and Charles L. Samuels. On the non–Archimedean metric Mahler measure. Journal of Number Theory, 129(7):1698–1708, July 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X09000419.

Walling:2009:RHS

[1605] Lynne H. Walling. Restricting Hecke–Siegel operators to Jacobi modular forms. Journal of Number Theory, 129(7):1709–1733, July 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/

pii/S0022314X09000845. See corrigendum [1710].

Hartl:2009:DBF

[1606] Urs Hartl. A dictionary between Fontaine-theory and its analogue in equal characteristic. Journal of Number Theory, 129(7):1734– 1757, July 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X09000195.

Bringmann:2009:RCM

[1607] Kathrin Bringmann, Jeremy Lovejoy, and Robert Osburn. Rank and crank moments for overpartitions. Journal of Number Theory, 129(7): 1758–1772, July 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X08002333.

Pickett:2009:ECS

[1608] Erik Jarl Pickett. Explicit construction of self-dual integral normal bases for the square-root of the inverse different. Journal of Number Theory, 129(7):1773–1785, July 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X09000870.

Chan:2009:RRC

[1609] Heng Huat Chan, Song Heng Chan, and Zhi-Guo Liu. The Rogers– Ramanujan continued fraction and a new Eisenstein series identity. Journal of Number Theory, 129(7):1786–1797, July 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X09000596.

Kim:2009:NEZ

[1610] Taekyun Kim. Note on the Euler q-zeta functions. Journal of Number Theory, 129(7):1798–1804, July 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X08002308.

Tripathi:2009:LSP

[1611] Amitabha Tripathi. On the largest size of a partition that is both score and t-core. Journal of Number Theory, 129(7):1805–1811, July 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X08001972.

Anonymous:2009:PAb

[1612] Anonymous. Pages 1813–2020 (August 2009). Journal of Number Theory, 129(8):??, August 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2009:EBh

[1613] Anonymous. Editorial Board. Journal of Number Theory, 129(8): ifc, August 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X09001164.

Cao:2009:NSC

[1614] Hui-Qin Cao and Hao Pan. Note on some congruences of Lehmer. Journal of Number Theory, 129(8):1813–1819, August 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X09000791.

Zhai:2009:PPF

[1615] Wenguang Zhai. On the prime power factorization of n!. Journal of Number Theory, 129(8):1820–1836, August 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X09000985.

Agoh:2009:HOR

[1616] Takashi Agoh and Karl Dilcher. Higher-order recurrences for Bernoulli numbers. Journal of Number Theory, 129(8):1837–1847, August 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X09000808.

Zudilin:2009:RTS

[1617] Wadim Zudilin. Ramanujan-type supercongruences. Journal of Number Theory, 129(8):1848–1857, August 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X09000584.

Gun:2009:LID

[1618] Sanoli Gun, M. Ram Murty, and Purusottam Rath. Linear independence of digamma function and a variant of a conjecture of Rohrlich. Journal of Number Theory, 129(8):1858–1873, August 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X0900081X.

Liu:2009:TAS

[1619] Jianya Liu and Yonghui Wang. A theorem on analytic strong multiplicity one. Journal of Number Theory, 129(8):1874–1882, August 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X08002345.

Perucca:2009:TVS

[1620] Antonella Perucca. Two variants of the support problem for products of abelian varieties and tori. Journal of Number Theory, 129 (8):1883–1892, August 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X09000365.

Sharma:2009:IIF

[1621] Aribam Chandrakant Sharma. Iwasawa invariants for the False–Tate extension and congruences between modular forms. Journal of Number Theory, 129(8):1893–1911, August 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X09000882.

Jeong:2009:QG

[1622] Sangtae Jeong. On a question of Goss. Journal of Number Theory, 129(8):1912–1918, August 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X08001819.

Kurokawa:2009:MEP

[1623] Nobushige Kurokawa and Hiroyuki Ochiai. A multivariable Euler product of Igusa type and its applications. Journal of Number Theory, 129(8):1919–1930, August 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X08002321.

Yau:2009:FDS

[1624] Stephen T. Yau and Letian Zhang. On formulas for Dedekind sums and the number of lattice points in tetrahedra. Journal of Number Theory, 129(8):1931–1955, August 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X09000912.

Fujii:2009:TRI

[1625] Satoshi Fujii. The torsion of p-ramified Iwasawa modules for Z_{p}^{2}- extensions. Journal of Number Theory, 129(8):1956–1964, August 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X09000614.

Cynk:2009:GKC

[1626] Sławomir Cynk and Matthias Schütt. Generalised Kummer constructions and Weil restrictions. Journal of Number Theory, 129(8):1965– 1975, August 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X08002011.

Zhang:2009:DEA

[1627] Yuqing Zhang. Diophantine exponents of affine subspaces: the simultaneous approximation case. Journal of Number Theory, 129(8):1976– 1989, August 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X09000833.

Kobayashi:2009:DSN

[1628] Mitsuo Kobayashi, Paul Pollack, and Carl Pomerance. On the distribution of sociable numbers. Journal of Number Theory, 129(8):1990– 2009, August 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X08002370.

Sengun:2009:WRB

[1629] Mehmet Haluk Şengün and Seyfi Turkelli. Weight reduction for modℓ Bianchi modular forms. Journal of Number Theory, 129(8):2010–2019, August 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X09000869.

Anonymous:2009:PS

[1630] Anonymous. Pages 2021–2226 (September 2009). Journal of Number Theory, 129(9):??, September 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2009:EBi

[1631] Anonymous. Editorial Board. Journal of Number Theory, 129(9):ifc, September 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X09001309.

Tanaka:2009:QDR

[1632] Tatsushi Tanaka. On the quasi-derivation relation for multiple zeta values. Journal of Number Theory, 129(9):2021–2034, September 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X09001103.

Spreafico:2009:BDZ

[1633] M. Spreafico. On the Barnes double zeta and Gamma functions. Journal of Number Theory, 129(9):2035–2063, September 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X09001127.

Chang:2009:CSB

[1634] Mei-Chu Chang. On character sums of binary quadratic forms. Journal of Number Theory, 129(9):2064–2071, September 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X09001073.

Xue:2009:MRA

[1635] Hui Xue. Minimal resolution of Atkin–Lehner quotients of X_{0}(N). Journal of Number Theory, 129(9):2072–2092, September 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X09001061.

Harris:2009:CID

[1636] Kelley Harris. On the classification of integers n that divide ϕ(n) + σ(n). Journal of Number Theory, 129(9):2093–2110.e10, September 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X09000547.

Chen:2009:EZA

[1637] William Y. C. Chen and Lisa H. Sun. Extended Zeilberger’s algorithm for identities on Bernoulli and Euler polynomials. Journal of Number Theory, 129(9):2111–2132, September 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X0900078X.

Fishman:2009:SGB

[1638] Lior Fishman. Schmidt’s game, badly approximable matrices and fractals. Journal of Number Theory, 129(9):2133–2153, September 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X09000742.

Gun:2009:TLG

[1639] Sanoli Gun, M. Ram Murty, and Purusottam Rath. Transcendence of the log gamma function and some discrete periods. Journal of Number Theory, 129(9):2154–2165, September 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X09000511. See

corrigendum [1711].

Kim:2009:ENP

[1640] Min-Soo Kim. On Euler numbers, polynomials and related p-adic integrals. Journal of Number Theory, 129(9):2166–2179, September 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X09000067.

Szpruch:2009:CLC

[1641] Dani Szpruch. Computation of the local coefficients for principal series representations of the metaplectic double cover of SL_{2}(F). Journal of Number Theory, 129(9):2180–2213, September 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X09000754.

Matomaki:2009:BVT

[1642] Kaisa Matomäki. A Bombieri–Vinogradov type exponential sum result with applications. Journal of Number Theory, 129(9):2214–2225, September 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X09000535.

Anonymous:2009:PO

[1643] Anonymous. Pages 2227–2646 (October 2009). Journal of Number Theory, 129(10):??, October 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2009:EBj

[1644] Anonymous. Editorial Board. Journal of Number Theory, 129(10): ifc, October 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X09001644.

Shlapentokh:2009:RSI

[1645] Alexandra Shlapentokh. Rational separability of the integral closure. Journal of Number Theory, 129(10):2227–2259, October 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X09001255.

Sander:2009:AUN

[1646] J. W. Sander. On the addition of units and nonunits modm. Journal of Number Theory, 129(10):2260–2266, October 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X09001449.

Cesaratto:2009:NEA

[1647] E. Cesaratto. A note on “Euclidean algorithms are Gaussian” by V. Baladi and B. Vallée. Journal of Number Theory, 129(10):2267–2273, October 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X09001139. See [784].

Gurel:2009:GSM

[1648] E. Gürel. Galois structure of modular forms of even weight. Journal of Number Theory, 129(10):2274–2288, October 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X09001413.

Louboutin:2009:OII

[1649] Stéphane R. Louboutin. On the Ono invariants of imaginary quadratic number fields. Journal of Number Theory, 129(10):2289–2294, October 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X09001437.

Paugam:2009:GAG

[1650] Frédéric Paugam. Global analytic geometry. Journal of Number Theory, 129(10):2295–2327, October 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X0900105X.

Li:2009:PF

[1651] Hongze Li and Hao Pan. Primes of the form ⌊αp + β⌋. Journal of Number Theory, 129(10):2328–2334, October 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X09001358.

Adhikari:2009:VP

[1652] Sukumar Das Adhikari and Andrew Granville. Visibility in the plane. Journal of Number Theory, 129(10):2335–2345, October 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X09001140.

Petras:2009:FED

[1653] Oliver Petras. Functional equations of the dilogarithm in Motivic cohomology. Journal of Number Theory, 129(10):2346–2368, October 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X09001383.

Yamasaki:2009:EMD

[1654] Yoshinori Yamasaki. Evaluations of multiple Dirichlet L-values via symmetric functions. Journal of Number Theory, 129(10):2369–2386, October 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X09001450.

Tsuzuki:2009:SSM

[1655] Masao Tsuzuki. Spectral square means for period integrals of wave functions on real hyperbolic spaces. Journal of Number Theory, 129 (10):2387–2438, October 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X09001395.

Zahid:2009:ZAF

[1656] Jahan Zahid. Zeros of p-adic forms. Journal of Number Theory, 129 (10):2439–2456, October 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X09001371.

Bureau:2009:IDT

[1657] Jean Bureau and Jorge Morales. Isospectral definite ternary F_{q}[t]- lattices. Journal of Number Theory, 129(10):2457–2473, October 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X09001280.

Hattori:2009:RBT

[1658] Shin Hattori. On a ramification bound of torsion semi-stable representations over a local field. Journal of Number Theory, 129(10):2474– 2503, October 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X09001498.

Meng:2009:LEP

[1659] Xianmeng Meng. On linear equations with prime variables of special type. Journal of Number Theory, 129(10):2504–2518, October 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X09001085.

Zhao:2009:CFD

[1660] Yusheng Zhao, Lianrong Ma, and Xianke Zhang. Calculating formula and divisibility for relative class numbers of abelian function fields. Journal of Number Theory, 129(10):2519–2525, October 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X09001462.

Dani:2009:OPI

[1661] S. G. Dani and Arnaldo Nogueira. On SL(n, Z)_{+}-orbits on R^{n} and positive integral solutions of linear inequalities. Journal of Number Theory, 129(10):2526–2529, October 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X09000456.

Fukshansky:2009:SCW

[1662] Lenny Fukshansky. On similarity classes of well-rounded sublattices of Z^{2}. Journal of Number Theory, 129(10):2530–2556, October 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X09000730.

Brown:2009:DPP

[1663] Jim L. Brown and Yingkun Li. Distribution of powers of the partition function modulo ℓ^{j}. Journal of Number Theory, 129(10):2557– 2568, October 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X09000353.

Kerz:2009:CDH

[1664] Moritz Kerz and Alexander Schmidt. Covering data and higher dimensional global class field theory. Journal of Number Theory, 129 (10):2569–2599, October 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X09001231.

Lafforgue:2009:VSF

[1665] Vincent Lafforgue. Valeurs spéciales des fonctions L en caractéristique p. (French) [Special values of L-functions in characteristic p]. Journal of Number Theory, 129(10):2600–2634, October 2009. CODEN JNUTA9.

ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X09000900.

Wittmann:2009:DCR

[1666] Christian Wittmann. Densities for 4-class ranks of quadratic function fields. Journal of Number Theory, 129(10):2635–2645, October 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X09001115.

Anonymous:2009:PN

[1667] Anonymous. Pages 2647–2882 (November 2009). Journal of Number Theory, 129(11):??, November 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2009:EBk

[1668] Anonymous. Editorial Board. Journal of Number Theory, 129(11):ifc, November 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X09001887.

deRoton:2009:ASH

[1669] Anne de Roton. Une approche séquentielle de l’hypothèse de Riemann généralisée. (French) [A sequential approach to the generalized Riemann hypothesis]. Journal of Number Theory, 129(11):2647–2658, November 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X0900153X.

Chamberland:2009:BSR

[1670] Marc Chamberland and Karl Dilcher. A binomial sum related to Wolstenholme’s theorem. Journal of Number Theory, 129(11):2659–2672, November 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X09001528.

Kaplan:2009:BMH

[1671] Nathan Kaplan. Bounds for the maximal height of divisors of x^{n} − 1. Journal of Number Theory, 129(11):2673–2688, November 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X09001504.

Chen:2009:PNN

[1672] Yong-Gao Chen and Min Tang. Partitions of natural numbers with the same representation functions. Journal of Number Theory, 129 (11):2689–2695, November 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X09001553.

Fu:2009:SIB

[1673] Amy M. Fu, Hao Pan, and Iris F. Zhang. Symmetric identities on Bernoulli polynomials. Journal of Number Theory, 129(11):2696–2701, November 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X09001711.

Harada:2009:SFG

[1674] Shinya Harada and Toshiro Hiranouchi. Smallness of fundamental groups for arithmetic schemes. Journal of Number Theory, 129(11):2702–2712, November 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X09001516.

Akatsuka:2009:ZMM

[1675] Hirotaka Akatsuka. Zeta Mahler measures. Journal of Number Theory, 129(11):2713–2734, November 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X09001607.

Smith:2009:GBD

[1676] Ethan Smith. A generalization of the Barban–Davenport–Halberstam Theorem to number fields. Journal of Number Theory, 129(11):2735– 2742, November 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X09001577.

Ki:2009:ZDS

[1677] Haseo Ki and Yoonbok Lee. Zeros of derivatives of sum of Dirichlet L-functions. Journal of Number Theory, 129(11):2743–2746, November 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X09001590.

Guo:2009:WSF

[1678] Li Guo and Bingyong Xie. Weighted sum formula for multiple zeta values. Journal of Number Theory, 129(11):2747–2765, November 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X09001486.

Freeze:2009:CNF

[1679] Michael Freeze, Weidong Gao, and Alfred Geroldinger. The critical number of finite abelian groups. Journal of Number Theory, 129 (11):2766–2777, November 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X09001735. See corrigendum [1716].

Shchur:2009:LDF

[1680] V. Shchur, Ya. Sinai, and A. Ustinov. Limiting distribution of Frobenius numbers for n = 3. Journal of Number Theory, 129(11):2778–2789, November 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X09001541.

Lu:2009:SEM

[1681] Guangshi Lü. The sixth and eighth moments of Fourier coefficients of cusp forms. Journal of Number Theory, 129(11):2790–2800, November 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X0900064X.

Tran:2009:CA

[1682] Linh Tran, Van Vu, and Philip Matchett Wood. On a conjecture of Alon. Journal of Number Theory, 129(11):2801–2807, November 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X0900047X.

McLeman:2009:GSE

[1683] Cam McLeman. A Golod–Shafarevich equality and p-tower groups. Journal of Number Theory, 129(11):2808–2819, November 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X09001760.

Dartyge:2009:DSM

[1684] Cécile Dartyge, Florian Luca, and Pantelimon Stănică. On digit sums of multiples of an integer. Journal of Number Theory, 129(11):2820–2830, November 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X09001279.

Akbary:2009:POM

[1685] Amir Akbary and Dragos Ghioca. Periods of orbits modulo primes. Journal of Number Theory, 129(11):2831–2842, November 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X09001267.

Gras:2009:RSC

[1686] Georges Gras. On the T -ramified, S-split p-class field towers over an extension of degree prime to p. Journal of Number Theory, 129 (11):2843–2852, November 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X09000821.

Jutila:2009:AFH

[1687] Matti Jutila. Atkinson’s formula for Hardy’s function. Journal of Number Theory, 129(11):2853–2878, November 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X09000894.

Urban:2009:CAN

[1688] Roman Urban. Corrigendum to “Algebraic numbers and density modulo 1” [J. Number Theory 128 (3) (2008) 645–661]. Journal of Number Theory, 129(11):2879–2881, November 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X09001772. See [1688].

Anonymous:2009:PD

[1689] Anonymous. Pages 2883–3114 (December 2009). Journal of Number Theory, 129(12):??, December 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2009:EBl

[1690] Anonymous. Editorial Board. Journal of Number Theory, 129(12):ifc, December 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X09002170.

Huynh:2009:LOT

[1691] D. K. Huynh, J. P. Keating, and N. C. Snaith. Lower order terms for the one-level density of elliptic curve L-functions. Journal of Number Theory, 129(12):2883–2902, December 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X09000389.

Rubin:2009:PCR

[1692] K. Rubin and A. Silverberg. Point counting on reductions of CM elliptic curves. Journal of Number Theory, 129(12):2903–2923, December 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X09000663.

Gong:2009:PRC

[1693] Sherry Gong. On a problem regarding coefficients of cyclotomic polynomials. Journal of Number Theory, 129(12):2924–2932, December 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X0900136X.

Johnson:2009:LCS

[1694] Keith Johnson. Limits of characteristic sequences of integer-valued polynomials on homogeneous sets. Journal of Number Theory, 129 (12):2933–2942, December 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X09001401.

Nakano:2009:FQC

[1695] Shin Nakano. A family of quintic cyclic fields with even class number parameterized by rational points on an elliptic curve. Journal of Number Theory, 129(12):2943–2951, December 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X09001425.

Bosser:2009:CFD

[1696] Vincent Bosser and Federico Pellarin. On certain families of Drinfeld quasi-modular forms. Journal of Number Theory, 129(12):2952–2990, December 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X09001474.

Hare:2009:IBS

[1697] Kevin G. Hare. Infinite Barker series. Journal of Number Theory, 129 (12):2991–2999, December 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X09001620.

Kim:2009:DFQ

[1698] Henry H. Kim and Yichao Zhang. Divisor function for quaternion algebras and application to fourth moments of L-functions. Journal of Number Theory, 129(12):3000–3019, December 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X09001619.

Tseng:2009:BAA

[1699] Jimmy Tseng. Badly approximable affine forms and Schmidt games. Journal of Number Theory, 129(12):3020–3025, December 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X09001589.

Yuan:2009:NRD

[1700] Pingzhi Yuan and Xiangneng Zeng. A new result on Davenport constant. Journal of Number Theory, 129(12):3026–3028, December 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X09001565.

Lu:2009:AIC

[1701] Jun Lu, Mao Sheng, and Kang Zuo. An Arakelov inequality in characteristic p and upper bound of p-rank zero locus. Journal of Number Theory, 129(12):3029–3045, December 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X09001759.

Koo:2009:IIA

[1702] Koopa Tak-Lun Koo. Iwasawa invariants of Λ-adic representation. Journal of Number Theory, 129(12):3046–3051, December 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X09001747.

Hegarty:2009:ATQ

[1703] Peter Hegarty. Answers to two questions posed by Farhi concerning additive bases. Journal of Number Theory, 129(12):3052–3058, December 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X09001851.

Guerzhoy:2009:ALS

[1704] P. Guerzhoy and Z. Kent. p-Adic liftings of the supersingular j-invariants and j-zeros of certain Eisenstein series. Journal of Number Theory, 129 (12):3059–3068, December 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X0900184X.

Kawamura:2009:CFE

[1705] Katsunori Kawamura, Yoshiki Hayashi, and Dan Lascu. Continued fraction expansions and permutative representations of the Cuntz algebra O_{∞}. Journal of Number Theory, 129(12):3069–3080, December 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X09001838.

Tanaka:2009:CAI

[1706] Taka aki Tanaka. Conditions for the algebraic independence of certain series involving continued fractions and generated by linear recurrences. Journal of Number Theory, 129(12):3081–3093, December 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X09001802.

Scarpello:2009:HI

[1707] Giovanni Mingari Scarpello and Daniele Ritelli. The hyperelliptic integrals and π. Journal of Number Theory, 129(12):3094–3108, December 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X09001796.

Virdol:2009:PME

[1708] Cristian Virdol. Potential modularity for elliptic curves and some applications. Journal of Number Theory, 129(12):3109–3114, December 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X09001991.

Omar:2010:CAF

[1709] Sami Omar and Kamel Mazhouda. Corrigendum et addendum à ^{≪} Le critère de Li et l’hypothèse de Riemann pour la classe de Selberg ^{≫} [J. Number Theory 125 (1) (2007) 50–58]. Journal of Number Theory, 130(4):1109–1114, April 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X09002868. See [1192].

Walling:2010:CRH

[1710] Lynne H. Walling. Corrigendum to “Restricting Hecke–Siegel operators to Jacobi modular forms” [J. Number Theory 129 (7) (2009) 1709–1733]. Journal of Number Theory, 130(4):1115, April 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X09002765. See

[1605].

Gun:2010:CTL

[1711] Sanoli Gun, M. Ram Murty, and Purusottam Rath. Corrigendum to “Transcendence of the log gamma function and some discrete periods” [J. Number Theory 129 (9) (2009) 2154–2165]. Journal of Number Theory, 130(5):1251, May 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X09002777. See [1639].

Kojima:2010:CFC

[1712] Hisashi Kojima. Corrigendum to “On the Fourier coefficients of Hilbert– Maass wave forms of half integral weight over arbitrary algebraic number fields” [J. Number Theory 107 (1) (2004) 25–62]. Journal of Number Theory, 130(5):1252–1253, May 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X10000223. See [690].

Goss:2010:NDE

[1713] David Goss. Note on “On the diophantine equation px^{2} + q^{2m} = y^{p}” [J. Number Theory 128 (6) (2008) 1670–1675]: Fadwa S. Abu Muriefah Riyadh University for Women, Mathematics Department, Saudi Arabia. Journal of Number Theory, 130(10):2393, October 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X0800214X. See [1378].

Borisov:2014:CQI

[1714] Alexander Borisov. Corrigendum to “Quantum integers and cyclotomy” [J. Number Theory 109 (1) (2004) 120–135]. Journal of Number Theory, 145(??):632–634, December 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X14001942. See [746].

Togbe:2015:CPF

[1715] Alain Togbé. Corrigendum to “A parametric family of cubic Thue equations” [J. Number Theory 107 (1) (2004) 63–79]. Journal of Number Theory, 150(??):206–207, May 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X14003850. See [691].

Freeze:2015:CCN

[1716] Michael Freeze, Weidong Gao, and Alfred Geroldinger. Corrigendum to “The critical number of finite abelian groups” [J. Number Theory 129 (11) (2009) 2766–2777]. Journal of Number Theory, 152(??):205– 207, July 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X15000219. See [1679].